A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (2000–2004)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 March 2019 Version 1.02

Title word cross-reference

(0 < p ≤ 1) [DJK01b]. (C, 1) [LM02b]. (I + S_{max}) [KHMN02]. (n − 1) [Che01]. (r, s) [Nis03b]. −∞ < x < ∞ [Tas00a]. 0 < x < ∞ [Tas04]. 1 [WKM04]. 12 [LPSSP00]. 2 [Age02, AM03, KP03a, Liu02a, Rat00, Wün01a, Wün03a]. 3 [Aco01, AD02, DN02a, GLZ03, HKD04, JO03, LF03, LF04, MS03b, Rat00, Wan04a, vdH00]. 4 [LLG04]. 5 [NS04a]. 9 [KZ03, Van00a]. [0, 1] [DS02a]. [0, ∞) [DS02a]. [J_{n+1}(z)]^{2} − J_{n}(z)J_{n+2}(z) [KV01a]. 2F_{1}(a, b; c; 4) [VD03b]. _{2}F_{1}(a, b; c; z) [BS00b]. 2F_{2} [Mil03b]. _{3} [Bru00]. _{3}F_{2} [DL01b, RKS04]. _{4}F_{3} [Lie04]. _{8}ϕ_{7} [Sch03c]. A [WW01]. A_{T,S}^{(2)} [WW01, WZ03]. α [Zub04].

B [WLYL04, YYT02]. C^{1} [DL01a, DN00, Far02, Lai00, LH03]. C^{2} [SA00]. C_{r} [Sch03c]. D [SC04, dAMR03, CZ03, DS02a, Hag01]. d^{(1)} [Sid00]. D_{q} [FR03]. Δ^{2} [Wen00]. E [Osa00]. ϵ [GL02]. f_{4}(O_{s}, −) [BG03]. F_{D} [FL03]. F_{p,μ} [GKS00, GKS00]. g [Dmy04]. G^{1} [SYW02]. G^{2} [MW02, WM02, WMA03]. Γ [Dav02]. GREP^{(1)} [Sid00]. H [HK00, KNSmG00, ONU03, Ste00b]. hp [Ste00b, Sur01]. I [DSD04]. j [Van00a]. J_{n}(z) ± iJ_{n+1}(z) [KV01a]. K [Pom01, ELW04]. K_{ϕ} [PP02b]. L [BdSR03, CMRS00, IN03]. L^{1} [ADL^{+}02, DFC02, HLY04]. L^{2} [BS02b, BGP02, Lau00]. L_{n}^{−k} [ELW04]. L_{p} [DJK01b, Gao01, LM02b, Sun01]. L_{q} [PS01].

LDL^{t} [Par00b]. M [AP04b, jGwHsZ04, GS04b, Arg01b, BBW04, Eve04, GSG03, Sim04, ZH04]. }↕(∈) [SS03a]. ∫⊓(∞, ∞) [LV03, Van00a]. ∫⊓(∈) [LV03]. N [Din02, Kan00, Miy02, Aga03, BV02, CK01, DKK03, EM04b, NS04a]. n + 1 [SW02a]. p [BR02, CG02, DS02b, He03, Koh01, Koh02, LZ04, LS00b, SS03a, Son01, Ste00b, Sur01]. Q [dAMR03, ÁNM01, ÁNAA03, AGMMB00a, AGRZ01, DS02a, FR01, IS03, KS01b, Lew03a, Lew03b, MÁNM01, OP03, SS03a, SC04, SJ03b, Van02, VYZ01, VZ01, dMPL03]. q ≥ 4 [DN00]. QR [CKR02, Wat00a]. r [FR03]. R^{d} [DDGH03b, Yoo03]. R^{n} [SW02a]. S′_{n} [CG03b]. σ [MS02]. sin Θ [Ips00]. su(1, 1) [Lie04]. SU_{q}(1, 1) [SC04]. SU_{q}(2) [SC04]. T [Par00b]. T (r) [Pet02a]. θ [Kot03]. u′(t) = au(t) + a_{0}u([t]) [LSY04]. U_{r,s}(}↕_{△}) [Nis03b]. V [Pet01]. W [BG02, HLY04, JPW04]. X − AXB = C [TML03]. X_{0} [WS02b]. ξ [Cof04]. y′′ + g(y)y′ + f(y)y = 0 [KW03]. y′′ = f(x, y) [CD00]. y′′ = F (x, y, y′) [dRM03]. Z_{2} [WS02b]. ζ(2n + 1) [ST00b]. ζ(3) [Cof03]. ζ(4) [Cof03]. ζ(6) [Cof03].

-1 [BDGV01]. -algorithm [CKR02, Osa00]. -analogue [AGRZ01, Van02]. -analogues [SJ03b]. -Appel [VZ01]. -approximation [Sun01]. -Askey [ÁNM01]. -band [jGwHsZ04]. -Based [HLY04]. -Bernstein [OP03]. -breaking [WS02b]. -classical [FR01, Lew03a, Lew03b, ÁNM01]. -coefficient [Eve04]. -colour [Aga03]. -Columns [dMPL03]. -Convergence [Dav02]. -curve [CMRS00]. -cyclic [LZ04, Son01]. -D [GLZ03, LF03, Liu02a]. -differences [AGMMB00a]. -dimensional [CK01, NS04a, Kan00]. -divergence [PP02b]. -extremal [PS01]. -FEM [Mel02]. -fold [Hag01, LPSSP00]. -form [WLYL04]. -fractional [Miy02]. -fractions [Dmy04]. -Fréchet [Arg01b]. -function [BBW04, Sim04]. -functions [SC04]. -hypergeometric [Nis03b]. -Krall [VYZ01]. -Krawtchouk [SC04]. -Laguerre [FR03, AGMMB00a]. -Laplacian [BR02, He03]. -like [Wat00a]. -matrices [GS04b, HLY04]. -matrix [HK00, KNSmG00, ONU03]. -Meixner [SC04]. -method [BG02]. -methods [Kot03, JPW04]. -norm [LS00b, Gao01]. -norms [Koh01, Koh02, BS02b]. -orthogonal [CZ03, KS01b, MS02, BdSR03]. -parameter [Age02]. -point [GSG03, KZ03, LLG04]. -polynomials [ÁNAA03, MÁNM01]. -preinvex [YYT02]. -process [Pet01, Pet02a]. -representations [SS03a]. -series [Lie04]. -shadowing [Far02]. -sparse [BGP02]. -spectra [GL02]. -spline [SA00]. -splines [Zub04]. -subdivision [JO03]. -symmetric [Che01]. -symmetry [Nis03b, WS02b]. -Taylor [IS03]. -th [EM04b, FR03]. -transformation [Sid00]. -type [DS02b]. -version [CG02]. -versions [Ste00b]. -widths [Din02].

107 [KO01]. 117 [BGM01]. 129 [QZ03]. 133 [Van01]. 149 [PGG03]. 157 [Win04a]. 15th [Ano02y].

2 [BMN01]. 2000 [BMPV00, BGPW02]. 2001 [ERV04]. 2002 [GVW04]. 20th [Bre00a, But00, GvdV00, SvdV00, WW00c]. 2nd [Ano02w, Ano02x].

3 [AVG^{+}04].

89 [Die02].

=f [Kob00].

Abel [Won02a, Won02b, Won04]. ABS [SXZ00]. abscissae [Not01, Not03]. Absolute [HW03, CJ03]. absolutely [GW04]. absorbing [MT03]. absorption [LCL01]. abstract [AO00b, Mas03, STW00]. academic [Sch04a]. accelerated [Lui01, Pal02a]. accelerating [NHMS04]. acceleration [Bre00a, EES00, Tam03, Yam02]. according [Sch04b]. accuracy [GM00b, Lan04, MM02a, Min04, Sto01, Sty04, YH03]. Accurate [Tas00a, Tas04, BM01a, DKL04, SU04]. accurately [Wri02]. acoustic [ARV00, MO00b, PDVS04, Pot00]. acoustics [Ava00]. activator [HL04a]. activator-inhibitor [HL04a]. active [FD00]. actuator [FD00]. actuator/sensor [FD00]. Acyclic [Par00a]. Adams [UH00]. Adaptive [BM04a, BHB04, CELM00, DFI02, Hop04, HR01, LRS02, Ran01, SW00a, ZK04, AD00, CACK04, CL03b, Lam03, PP00c, Ran04, Shi04b, VALM00, WKM04, Zeg04]. adaptivity [LS02a, LRvSS04]. Addendum [PGG03]. Addition [Rad00]. Additive [BJ02, Bai03a, ZLF01]. adhesion [CFHS03, CFSS03]. ADI [TML03]. adiabatic [LJ03]. adjacencies [PR02]. Adjoint [CLP02, BM00c, Rui02]. admitting [SB01]. adsorption [dCCSR00, Rem04]. Advanced [YSNM02]. Advances [Coo02, XZ03]. advection [ALM03, BS00c, CVB04, EW01, HC03b, MT04, WF02, vdH00]. advection-dispersion [MT04]. advection-dominated [EW01]. advection-reaction [ALM03, vdH00]. aerodynamic [SW01]. aerospace [WPS02]. affine [ZH04, Zhu03a]. after [BPT02]. after-effect [BPT02]. aggregation [Mar03]. aggregation/disaggregation [Mar03]. air [BV03]. Airy [EM04a, GST03]. Airy-type [GST03]. Aitken [Wen00]. ALA’01 [ERV04]. Algebra [Ano00x, ERV04, DE00, Fab02, Nac04, O’L00]. Algebraic [HHC01, RSvR03, Sha01, Wal03, BPT02, BJK04a, CLP02, IZ00, LVH04, LP00, Lor00, MAK01, PS02, Sch04a, Sch03d, Stü01, Wan00a, WL02a, WZ04b, Win03, Win04a, Win04b, Wol00a]. algorithm [BTFY02, CKR02, CCL03, CZ04, CR03, Cui02, Dal04, DvM01, DS02a, Din03, DSW^{+}03, Dra04, EM04b, GS04a, GMRS00, GLZ03, Han03b, HKD04, Ji04, KyGUY02, Lam03, LR04a, LZ02b, MKS^{+}01, MK01, MO00b, NMST02, Osa00, PT03, QSZ02, RDA04, RCZV04, Sch03a, SS04, SC01, Ste02, Swa00, Swa02, Thu02, Tru04, VB04, Wen00, Zhu00, Zhu03a, Zhu03b, dAMR03]. Algorithms [FLR03, FGJ00, HL02, Nak01, AZPF04, Bar02, BBBCD00, Bog04a, Bou04, BRZS00, CP04, CLP02, CMKM03, DSVB04, DL00, Din04, DE00, FQR00, Gao01, GHR02, GL00, HLY04, IRVD01, IVD02, KL04a, KP03b, LP00, Min03, NSP04, Noo01, Sch02b, SSV04, SXZ00, VMV04, VALM00, WPS02, Wat00a, WJ01, YH02, ZD02]. allocation [Dra02]. alloys [VV00]. Almost [FW03, Beh02, Miy02]. alternating [DZN00, HWT04, WH01]. Alternative [UH00]. aluminium [BQ00]. analog [Chi03]. Analogs [Sim04]. analogue [Aga03, AGRZ01, CJ04, Van02]. analogues [MN02a, SJ03b]. analogy [Hat03]. analyses [LWY02]. Analysis [AP00, Ano00x, Ano01s, BMPV00, Bre00b, EES00, Fre03, KZ03, LCL01, Miu02, Tid02, WBBF00, WW00c, Yan02, AS04a, AM03, AM00, ARV00, Arg03a, ATG04, Bac04, BQ00, BTSHK04, BC04, BO04b, BS02b, BZPFB04, Bri04, Buc00, BM00c, CLP02, CFHS03, CFSS03, Che03a, CMV01, DLM^{+}00, DLTS02, De 01, DND04, ELR00, ELW02, ELW04, Fuj02, Gar04, Gov00, Gru03, GH03b, HSS01, Hau02, HRK04, JL00, JS00b, KyGUY02, Kui01a, Lac03, LR04a, LP00, LHHW04, Löt00, Lou02, LLN04, Mar00a, MRT00, MS04, MS01b, NY03, Pet01, PW02, Pet02a, PDVS04, QST00, Rih03, SKSV03, SIM02, Sla00, SVV01, SV04, Str00, TL01, Tho01, TWV02, Tsu01, TYI03, WL02b, WZ04a, WKS^{+}03, lXsQsJ03, Yam00, YSNM02, ZV04, ZD02]. Analytic [BE02b, MF01, SW00b, SZ04, AT00, AJ01, BS00b, BL01c, BEM00b, DSZ02, EP02b, GL04a, Göt01, Kaz02, KSSV03, MS04, Pet02c, SS02a, SKSV03, SV04, SS03d, Wal03]. analytic-numerical [AJ01]. Analytical [HR03, Vel01, BH03, FFX04]. analyzing [Gro02]. angular [AR03]. Anisotropic [Ran04, Aco01, CCH02, FCP02, Häh00, IY03, JT03]. any [ABC03]. AOR [EMT01]. Appel [VZ01]. Appell [Ism03b, CS03a, Ell01, HR02a]. Appl [BGM01, Die02, KO01, PGG03, QZ03, Van01, Win04a]. Application [AT00, CL00, KL04a, MAK01, PDVS04, SGG^{+}04, VWE04, DEL04, Dmy04, DDPT00, EM04a, Fab02, FKM02, HWT04, IN03, KSSV03, Li02, LCZ04, Oou01, Oou03, PCR04, Pen00b, Sch01, Sha01, SW02b]. Applications [IS03, JKVV03, Sla00, AM00, Arg01b, BRS00, BB02, BJK04b, BS02b, DSVB04, Dat00, DRC02, Doh02, DMS00, Eli00, Fre01, Her03, Hil02b, HR01, Kal00, KT01, Koh01, Koh02, Koh03b, KyGUY02, KLT04, LN03, LLL03, ML03, MO00b, Müh00, Nie00, RSvR03, Ric00, SRLAD03, Sch04a, Sid00, SS01, Ste02, SVV01, TO03, TYI03, WO01, Win03, Win04a, Win04b, xZqZX02, JS00a]. Applied [Ano00w, MP02b, eMEM00, BD04b, FTY02, FL04, LZ04, LWW01, RDA04, TQ03, UC03a]. applying [Ism00]. approach [eMKM02, BO04a, BC04, BG03, DKM^{+}01, DSW^{+}03, GH01a, GDD04, HRS02, HR01, Hui04, LF03, LQQ01, Mar03, MFMGZ02, PR00, Sai04, Sza02, WL01, WG02, YH02]. approaches [BRZS00, MOTT03, Sch04c, SdSP01]. appropriate [NOI02]. approximant [Gu04]. Approximants [BL01b, AB03, DMGVO01, GPTT02, HR00, Kha02, RLS01, Zho01]. Approximate [BL01c, vdHS01, ADL^{+}02, BV03, BT00b, CN00]. approximating [Fer03, Swa02]. Approximation [AsKS04, DVM02, Fas02, KP00, KP02, Wat00b, WW00c, AGC00, AAD02, AB04a, BM01a, BEM00a, BRS03b, BM04a, BJG02, BD00b, BVV04, BW03, CDV03, CHW02, CS04c, CS03b, DL01a, Dar01, DV00, Dem04, DFC02, DS02c, Fre03, GS01b, GK03, GR01, GH00, GSS03, HJ02, Ham00, HLS^{+}03, HR00, IM01, Ing03, KS00, KL01a, Lee02, Lev00, LYF03, LL03b, LJ03, MGL01, MRT00, MO02, MY00, MW04a, MSKS04, Mil01b, MN01, MN00, NKK03, Pet02c, SW02b, Slo04, Sun01, UY04, Wal00, WdZ04, WW00b, Wri04, YTI02, YY02, Yoo03, dACCS03, dSPPT01]. Approximations [CM04, Bai01, Dav02, DLZ02, EGV03, Gau02, GDD04, Has04, KO99, KO01, KOH03a, KO03, Kor01, LMO01, Mai03, MT04, Pie00, Pré00b, Shi04a, SB04, Swa04, Yan02]. arbitrary [AS04c, BS00b, DvM01, FQR00, LNLL02, PP00b, WW00a]. arc [Che03b, MW04a]. arc-length [Che03b]. arcs [AsKS04, Dom03]. area [PP00b, UY04]. area-preserving [UY04]. argument [BW03]. arguments [CK02b, Jan03a, KR02, MY03b]. arises [NS03]. arising [BE04b, CN02, IY03, MRVM00, PS04, PP00c, SW00a]. Arithmetic [ERV04, HK00, Nak01, SK01]. ARKN [Fra03]. Árpád [LMS03]. array [OASO03]. art [SD00]. Asian [AP04a]. Askey [ÁNM01, AGRZ01, SRD01, TL01]. Aspects [Sch03a, ADG03, CACK04, Gor04, Joh02, KLyD^{+}03, MF01, MD04, Tem00]. asperities [AB04a]. assessing [VBL^{+}04]. assessment [DFT04]. assigned [ZH04]. assisted [BMZ00, Nag04]. associated [AB04b, BE04a, CL03a, Che03a, CS00b, CS03a, CV01, CV03, Dac03, Dun03, Fab00, FR03, Kui01a, KV01b, Kza00, Let01, LL03a, Lew03b, Man00b, Miy03, Mou03, Nak01, NR01, Yak02, vD03a]. associative [WZ04a]. assumptions [HS00]. astronomy [Nie00]. Asymptotic [AVV00, And00, BZPFB04, CP00b, FL03, GR01, KRT02, Kra03, Kui01a, KV01b, Kza00, KMV03, LL01a, LP01, LMMD03, Mac02a, PP02b, Sto05, Wal00, AM02b, BD04a, BW04b, CCL03, CG00, Čer02, DKM^{+}01, GST03, Kar01, Lar02, LB01, MFMGO01, Nay02, Par02a, Par02b, Par03b, Par03a, Tem00, TL01, Zha03]. Asymptotic-numerical [BZPFB04]. asymptotical [CD03a]. Asymptotics [AMBP^{+}01, CV03, FP02, Gat02, LL03a, Pan02, RLS01, Ron01, AGMMB00b, BC03]. Asynchronous [LLP03, FS00]. atmosphere [ZFYY04]. atmospheres [ADL^{+}02, KSCI00]. atmospheric [BV00, BKB04]. attainable [MIB03]. attractivity [LCC04]. augmented [SD03]. Author [Ano00a, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano02i, Ano02j, Ano02l, Ano02o, Ano02k, Ano02m, Ano02n, Ano02g, Ano02h, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano03a, Ano03g, Ano03m, Ano03b, Ano03c, Ano03h, Ano03d, Ano03i, Ano03j, Ano03e, Ano03f, Ano03k, Ano03n, Ano03l, Ano04a, Ano04b, Ano04k, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04j, Ano04h, Ano04i]. Authors [Ano01f, Ano01g, Ano01h, Ano01l, Ano01i, Ano01j, Ano01k]. automata [LS04a]. Automatic [BBBCD00, KR03a]. automation [BC04]. autonomous [DKST04, FPS00]. average [Aco01]. Averages [FK01]. axis [GS01b]. axisymmetric [Abd00, RdAR04].

B [Bru00, BM03b, BM04c, jGwHsZ04, HM00, KED04, MVDB04, NS04b, SYW02, SW02a]. B-spline [jGwHsZ04, HM00, KED04, SYW02]. B-splines [BM03b, BM04c, MVDB04, NS04b, SW02a]. Bacharach [LCZ04, WZ04b]. back [Dal04, Zhu03b]. back-tracking [Zhu03b]. Bäcklund [KCMK02]. backward [Cas00, Dra04, HY03, Les02, Liu02a, Vec00, YLC04, ZV04, xZqZX02]. backward-facing [YLC04]. balanced [Lie04]. balancing [MO00b]. ball [NS04a]. balls [Pet04, Xu01b]. Bampton [Rix04]. Banach [ABC02, Arg01a, BL01a, DX02, Din03, Hau02, Pre00a]. band [jGwHsZ04, ZH04]. bank [FQR00]. banks [BS02a]. Bargmann [SV04]. Barnes [MP03a]. barotropic [SM04a]. Based [HLY04, ACGR01, AK01c, BLSS03, BZS02, BHB04, CO02, CW04, CP03, DLM^{+}00, DDN01, DM02a, Die98, Die02, DNS00, DHK02, Dra02, GVV04, Ham01, HRK04, Hug01, qJkSlS04, KO99, KO01, KOH03a, KO03, LS02a, LLG04, Mel02, MN01, PR02, PDVS04, SSV04, VV01, Wan04b, Yu02, Zeg04]. bases [Peñ00a, SVV01, vG00]. basic [AT00, CS03a, DSS03, Lew00]. basins [Mal04a]. basis [BVV04, CS01, Far00, FK01, LM02a, LH03, LS00b, NMST02, Sch00a, Sza02, Yoo03]. Baum [MO00b]. Bayesian [Iqb00]. BDF [XLFC00]. be [WFV01]. beam [AVV00, AMS00, Fer03, LL00]. beams [SAA01]. behavior [CP00b, DKST04, Fre03, LB01]. behaviour [Aas02, AB01, BW04b, GR01, MGL01, SB04]. Bell [Col01, Zha03]. Bellman [HWT04]. below [BMZ00]. BEM [MG01]. benefits [Dul04]. Benney [Öze04]. Bernoulli [EST03, SAA01]. Bernstein [CS03b, Far00, GL03, LS00b, OP03, Sza02, VB01]. Bernstein-type [GL03]. Bessel [Sto05, AJDG02, BW03, CV03, Dac01, EL01, Elb01, Fab02, GKK00, GST03, GKS00, KT00, KRT02, LL01a, MFMGO01, PD03, PSS03, Vel01, Yak00]. Bessel-type [GKS00]. Best [AKTvD00, DFC02, El 04, Hei03a, FK01]. beta [KR03a, TO03]. better [CP04]. between [BDGV01, CO02, CFMV03, Che03a, O’L00, Par03b, VD03b, WH04]. beyond [SRD01]. Bézier [AsKS04, CO01, CS03b, LPSSP00, OP03, Rab03, WMA03]. BFGS [AB01, LF01]. Bi [AS04a, Dav02, LY02]. bi-cubic [LY02]. Bi-factorial [AS04a]. bi-Laplacian [Dav02]. biaxially [FCP02]. bibliography [Ano02a]. BiCGstab [IN03]. bidimensional [CM01]. bidirectional [MN02a, WZ04a]. Biedenharn [Lie04]. Bifurcation [JKN00, TWV02, BBMS04, Cai02, ELR00, FW00, Gov00, WZ04a, WS02b, WF00]. bifurcations [CHW02, CL01b]. biharmonic [lCHcH00, Chr01, Jeo00, Jeo01, LY02, Moh00, Wan04b]. bilateral [DSS03, MC03]. BiM [BM04b]. Binary [PP02a, Nei02]. binomial [TO03, lXsQsJ03]. biodegradation [WKS^{+}03]. bioremediation [CL00]. biorthogonal [VZ01]. Biorthogonality [Zhe04]. biosciences [BR00]. biquartic [SYW02]. Birkhoff [CG00, dBS01]. Birth [vD03a, Mar03]. Birth-death [vD03a]. bispectral [GI03]. bistable [DKST04]. Bivariate [Alf00, CG00, Liu02b, CS03a, CJ00, DL01a, DSS00, DFI02, Lai00, LCZ04, NZ00a]. blade [HM00]. blading [Ega00]. Blended [Bru00]. blending [TW04, TT02]. Block [AH04, CN02, CCM01, Bru00, CW04, GS04b, KLT04, LGL03, Min03, SZ01, Sch03a, Sch02b, Son01]. Block-circulant [CN02]. blocks [KT01]. blow [CY00, IY03, YL03]. blow-up [CY00, IY03, YL03]. blowing [GR01]. Blumenthal [Chi03]. board [Ano02w]. Bochner [Ism03a]. bodies [Kai01, Mar04a]. body [CF04, Iwa02]. Bogoliubov [Ism00]. Bohnenblust [Ste02]. Boltzmann [AM03, LWW01]. bone [MOTT03]. Borg [BPW02]. boson [BTFY02]. boson-fermion [BTFY02]. both [Zhu00]. bottom [SM04a]. bound [Bat04, Cha04, GJS03, GS01b, LZ02b, MS01a, PP02c, Rum03]. boundaries [DKL04, PP00b]. Boundary [DDG04, HZZ03, LL00, ADG03, ADG04, AZ04, ABO02b, AD02, ASN02, ADR02, And02, And00, AG01a, AH03, AG01b, AG02a, AA02c, AK01c, AK02, BM00a, BL01c, Bha02, BBW02, BD04b, BW04a, CL04, CACK04, CGM02, CMSV03, Cha02, CK04, Chi01a, CH02b, CL02, Dav02, DH00, Deh02, DB01, EP02b, EP03, EGV03, FTY02, FHM^{+}04, Faz02, Fuj02, GS01a, Gan04, GN03, Gru03, Gug00, GSG03, GT04, GJHY00, GGDL04, HP00, HC03a, HOO03, HY02, HK01, HSW00, Jan02, Jan03a, Jeo00, Jeo01, JW00, JFW01, JW04, Joh02, JE01, JE02, KWW01, Kum02, KK03, Law02, LCL01, Lin00, LY00, LW02, LLN04, MMOS01, Mar00a, MSKS04, NS03, NJVA03, NRL03, Pal02b, PS04, Pao01, Pre00a, QST00, Rac00, RSS04, SGG^{+}04, Ste00b, Sug02a, Swa00, Tay00, VV01, Wat03, Waz01, WWA^{+}04]. boundary [WA00a, Won02a, WY02, ZW02]. Boundary-type [HZZ03]. boundary-value [ADG04, AK02, LW02]. Bounded [AK00, Din02, Dra04, MO01, OOAS03, Pre00a]. Boundedness [EF02, BH04, LM02b]. Bounding [BS02b, Xu03b]. Bounds [Kub01, AR04, AJ01, Bar02, BS00d, DHP02, Die98, Die02, Gat02, Hei03a, Her03, Neh03, Nie03, Par02a, Tas00a, Tas04, Won02b, xZqZqT04]. box [Kál02]. bracket [Cas04]. branch [LZ02b]. branching [CR01]. breaking [WS02b]. Brezinski [Wen00]. brine [MP02a]. Brown [BE02a, Mil03a]. Brownian [Dai02]. buckling [BZPFB04, CCM01]. buffers [Dai02]. Building [KT01, Sch02b]. built [Rad00]. Burgers [dGK01, eMEM00, KED04, LZ01, WG02]. Burgers’-type [dGK01]. BVM [JkSlS04]. BVMS [Bru00]. BVPs [Faz02].

C [MW02]. C-shaped [MW02]. CAGD [GM00a]. Cahn [Ye03]. calculation [DMN01, aH02]. calculations [EM04a, LLN04]. calculus [GKS00, Kir00, Koh01, Koh02, Koh03b, LNBI02, Miy02]. Calderón [GHK03]. calligraphy [LHYaC03]. Can [YEWE03, WFV01]. Cancellation [CES02, WS02a]. canonical [Ava00, Tor03]. cardiac [PS02]. cardinality [DDN01]. cardinality-based [DDN01]. Carlo [Nie03, Wan00b, Wan01b]. cascade [GL00]. cascades [BD00a]. case [BC03, BG03, DKST04, Gao01, GPTT02, GK03, Hom04, JO03, Mil00, MSKS04, PW02, RR01, Ron01, Sey01, Xu02, Xu03a]. cases [LGL03, Tem03]. casting [BQ00, LLP03]. categorized [NOI02]. Cauchy [Bad01, CC03a, JM00, JS00b, Jun04, KC00b, KC00a, KY02, Mas03, Müh00, Tak03, YH02, Yan03]. caustics [BLSS03, BLSS04]. cavity [TWV02]. Cayley [LCZ04, WZ04b]. cell [Miu02, Swa00, Swa02]. cells [Bog03]. cellular [LWY02, Sla00]. censored [RCZV04]. censoring [Age02]. center [AR03, CGG00]. centers [CGG00, Sun01, Yoo03]. Central [BBK04, AG03b, KL01a]. Century [WW00c, Bre00a, But00, GvdV00, SvdV00]. Certain [CS00b, CS04a, DSS03, DMGVO01, Grü01, JV03, Li03b, Mac02a, Mal04b, Mil00, Smi03, ST00b, WW03, Weg01a]. Chain [BDR02, Pöt02, And02, AA02c, EPM00, LS04b]. chains [AP02b, CH02b, Elo02, Hil02a, Pöt02, Won02a, Won02b]. change [Čer02, WWA^{+}04]. changing [AO00c, CDN01, GT04]. chaotic [KN02, YB03]. characteristic [BG03, Büh00]. characteristics [Ban04, ZL03]. Characterization [FHMS04, VZ04]. characterizations [MC03]. characters [LHYaC03]. charge [OOA03, OOAS03]. Charlier [HHR00, Let01]. charring [TVV04]. Chebyshev [BGM01, BGM00, BE03, HS00, KSCI00, Not01, Pie00, VAR03, Vou01, dFN00]. chemistry [BV00]. Chihara [AIV01]. childhood [MG03]. Chinese [Tan02]. Cholesky [Sch03a]. CIP [YOT^{+}02]. CIR [MA04]. circle [BCM01, BDGV01, CFMV03, DGVM02, Dik03, MN02b, RLS01, Sim04]. circles [LMYL01]. circuit [DHW04, Fre00, LVH04, Pen00b, Win03, Win04a, Win04b]. circuits [MT02]. Circulant [JkSlS04, Ng03, CN02]. Circular [CS03b, AsKS04, SK01, Weg01a]. class [ABC02, AW03, AK01c, BRS00, BC02, CFSP04, CHW02, DDN01, DM02a, DX02, DP02b, ENE04, FR03, GHMY00, Ifa01, IK03, JNS04, JKY02, Kum02, LR02, Liu01, Mlo01, PS04, PJ03, SD03, SZ04, SR04, ST00b, Suá01, VALM00, WL03, Weg01a, WJ01, WF00, XLFC00, XL02, Zha02, ZW02]. Classes [GVSJ01, AKM03, CS00b, Her03, KMCK02, Mal04b, MM02a, NZ00b]. Classical [RZAG00, AH04, BM03a, CS01, ELW00, FR01, FKR04, Her01, HHR00, JV03, KS01b, Lew03a, Lew03b, NR01, Rad00, VC01, VZ04, ÁNM01]. classification [SS02c, TL04]. Classifications [LZ02a]. Clausenian [Kar00b, Van02]. Clenshaw [Bar02]. Closed [BD00a, GDD04, KO99, KO01, Cha02, KOH03a]. Closed-form [BD00a, GDD04, KO99, KO01]. clothoid [MW04a]. clothoids [MW04b]. cluster [KATS02]. clusters [Neu03]. CNN [SZ03a]. co [FR01, FR03, Let01, NR01]. co-recursive [FR01, FR03, Let01, NR01]. coarse [DNS00]. coarse-space [DNS00]. coatings [HJ02]. Coaxing [Mee02]. code [AVG^{+}04, BM04b]. codes [FLR03, UH00]. coding [CW04]. coefficient [AJ01, AMS00, CW00, Eve04, Van00a, Wri02, Wri04]. coefficients [Bar04, Beh02, BBK04, CS04a, CCL03, CG03a, CL03a, Dim01b, Doh02, GZ02, KS02a, KS02b, KV01b, Kza00, Lew03a, LHHW04, LZ01, Neh03, PP00a, Ron01, TC01, WT01, dAMR03]. Coercive [Fab00]. coherent [AMBP^{+}01, Fuj02, MdB02]. collocation [AK01a, Ant02, BB01, BF01, CD00, Dom03, FGJ00, Hei03b, HR02b, JM00, KSCI00, Kou03, LS02a, MIB03, Sug02a, SU04, VVV03b, Ye03]. colloidal [Iqb00]. colour [Aga03]. Columns [dMPL03]. combination [CO01]. combinations [KLM01]. Combinatorial [Hof00, Aga03, Rad00]. Combined [NSP04, Tan02, AB01, AG03b]. combustion [AR04]. Comment [Dim01a]. Comments [Pet02b, Sto05, Van01, BCJW00]. Committee [Ano02y]. common [PSS03]. commutative [Röß04]. commutators [GHK03]. compact [BS02a, DN01, DF00, Elo02, GZ02, Hu01, KZ03, YD03, ZL04]. compactional [sY01]. Compactly [Han03a]. compactness [BH04]. Comparison [Bha02, Kou03, Oht02, WH04, Arg04a, BV00, CK04, CS01, FW02, JS00a, KP01, KHMN02, LRS02, Mor00, WKM04]. comparisons [FLR03, TO03]. Compensating [BBCH00, BBCH02]. competition [DKST04]. complement [MRT00]. Complementarity [BM00b, LQQ01, LZ02b, ZLF01]. complements [FF02]. Complete [CY00, AQ04, yCjC01]. completely [BP01a, DX02, LCL01]. Complex [Bec01, MWL04, AB03, BPW02, DSZ02, Ips00, Koh02, LCI01, RSvR03, SK01, Par04b]. complex-valued [Koh02]. complexity [HK00, KV04]. compliance [HSS01]. comply [Mee02]. component [Gra04, LGL03, VV00]. componentwise [Hei03a]. Composed [Dar01, WMA03]. Composite [TT02, BT01]. compositions [Aga03, GKS00]. Compressible [BSKP04, KK03]. compression [HSV04, LL03b]. compressor [Ega00]. compressors [VVD04]. Comput [BGM01, Die02, KO01, PGG03, QZ03, Van01, Win04a]. computability [Liu01]. computable [Hei03a, Tas00b]. Computation [Fre01, GST03, HH04, KT00, Lau01, NR01, Yu02, AG03a, CKR02, DEL04, Fre02, GZ02, GS04a, GvdV00, Gug00, Hop04, KTIS03, KOZ03, Mac02b, Mai03, MO00a, Sed04, SDMV04, Sid00, WZ04a]. computation/simulation [Fre02]. Computational [Ano00w, BG03, BBC00, Kon04, LLN04, WKS^{+}03, ADG03, BBCH02, CMKM03, Gor04, Hui04, HC04, KP01, PP00c, Wat02, WO01, WS02a, YSNM02]. computations [BLSS03, BLSS04, KATS02, Mar00a, MT03, TNW02, ZK04]. Computer [Koz00, BMZ00, Fab02, Nag04]. computer-assisted [Nag04]. computers [DB01]. Computing [AR03, BA00, BR02, DKK03, DvM01, Gem00, KV04, Lu03, Pie00, RO03, WT01, Duf00, KSSV03, Man00b, MZ02, PNV01, SKSV03]. concave [Dem04, KS01a]. concerning [Arg03b]. Condition [Lau00, WZ03, ADR02, BQ00, BC03, Cui02, DH02, Fan03b, LS00b, MS03a, Mel02, PP02c, Sen01]. conditions [AR03, And00, AH03, Arg03a, AG01b, BS00a, BBW02, BS00d, CG03a, CG00, CL02, DJK01a, Deh02, DHZ04, EGV03, EH00, Fuj02, GN03, HP00, Her01, Jan02, Joh02, KP02, KP04, Lin00, Pao01, RSS04, SYW02, SW01, Wat03, dSPT00]. conduction [Liu02a]. conductor [FMRW04, YOO02]. Conference [Ani02, Ano00b, Ano00c, Ano02p, Ano02q, ERV04, JKVV03, Ano01r, Ano02y]. confluent [Cam01, Gau02, Ses04, Yan03]. conformal [OOA03, OOAS03, Weg01b, dG01]. conforming [WL02b]. congruences [CL03a]. conjecture [Bak03, EL01, ELS01, Sam01a, Szy03]. Conjugate [WG02, CS02, KRW00, WA00a]. connected [Las02, MS02, OOAS03, Weg01a, Weg01b]. Connection [DV01, Dim01b, LZ01, AGRZ01, BDGV01, CFMV03, Par03b, SRD01]. consecutive [CES02]. consequence [CG03a]. Conservation [MTA^{+}03, BBK04, Coc01, Fur01, LYC02, Mon01, ZL03]. conservative [EHS00, LMMD03, Mat03, YOT^{+}02]. Conserving [Sch00b]. consistent [Li03a, Yam02]. constant [AR03, AMS00, Beh02, LHHW04]. constants [Alz02, GS04a, WT01]. Constrained [CD01, CZ04, Dem02, DSW^{+}03, NW00, SD03, WZ03, Zhu00, Zhu01, Zhu03b]. Constraint [Tas00b, GG01]. Constraint-selected [Tas00b]. constraints [BM00c, Dul04, Gra04, KYF04, LZ02b, MM00, Zhu03a]. Constructing [LLL03, JFW01]. Construction [BM03b, CJ03, HH03a, JPW04, KP04, Lew03a, dSPT00, FHMS04, VBL^{+}04]. constructions [ABG03]. Constructive [Weg01a]. contact [Abd03, AZPF04, BQ00, CSRB04, CFHS03, CFSS03, DHK02, FI03, FC04, HS01, HKD04, RDA04]. containing [Sam01b, Sri02]. Contaminant [MD00]. content [Dia03]. Contents [Ano01m, Ano01n, Ano01o]. Contiguous [Vid03]. continuation [AT00, BS00b, BEM00b, BE02b, CR00, DKK03, DSVB04, Pre00a, Rhe00]. Continued [BSS03, GPTT02, Lup01, BS03, BL01b, Bea01, BL04, TJ04, xZqZX02, ZZ03a, xZqZqT04, dAMR03]. Continuity [HR00, Kza00, TW04]. Continuous [Enr00, PD03, eMKM04, CZW04, CL03c, GKMN00, GW04, Gra04, LLP03, LLC01, LWY02, Oou01, Oou03, SYW02, Won04]. continuous-time [CZW04]. contracts [WFV01]. control [Bor03, BD00b, BHB04, BM00c, CL00, EK01, Enr00, FD00, GJL^{+}00, HW03, Jun04, KW00, Leu00, LY00, MRVM00, MX00, MM00, SAA01, Sar00, SBSA03, SZ03b, TW04, VID01]. controllability [Tay00]. controls [Gug00, Kas00, Mar00a]. convection [eMKM04, AH03, AG03b, Ban04, BE03, BJG02, Bog01, CELM00, CN02, CJL03, Fan03a, FHM^{+}04, HRK04, JY01, KZ03, KL01a, KK03, Len02, Rem04, RZ03, Shi04a, ST02, Tid02, Tid03, WL02b]. convection-dominated [Bog01, KK03, ST02, Tid02, Tid03]. convective [VMD04]. Convergence [Bre00a, CR01, Dav02, Fan03a, LWT04, LWW01, LCF04, Löt00, MO03, SL03, UY04, WW00a, Yam02, YB03, YK04, AA02a, AVMRVM02, Arg01a, BW02, Bai03a, Bai03b, BJK04a, BM03a, CL04, Cao02, CS02, Che03b, Cui02, DJK01c, DJK01a, DJK01b, Fre03, FS03, Gar04, HS00, Hom03, Hom04, KYF04, KY02, KL01a, LF01, Li03a, Lui01, NHMS04, PS04, PR04, QST00, SS02a, SHS02, Tam03, WH01, WX02b, XLFC00, Yam00, LHHW04]. Convergent [Chi01c, CJL03, CMKM03, HKT^{+}03, Ji04, PT02, ST00b, VV01]. Convex [Lai00, CCL03, CO01, Dem04, DF00, KYF04, KS01a, Mar04a, MN00, YTI02]. Convexity [Dim03, Dem02, LPSSP00, Mil03a, MWL04, TW04]. convexity-preservation [TW04]. convolution [FW02, Kan00]. convolutions [EST03]. coolant [BTSHK04]. cooled [Las04]. coordinate [LS02a, Zeg04]. coordinates [Hui04]. cope [dAMR03]. core [DDG04]. Corporation [Ani02]. Correction [GJHY00, AZPF04, And00, CMSV03, CELM00, Chi01a, VV01]. corrections [GGM01, Xu01a]. corrector [PS03b, UH00]. correlation [Iqb00]. corresponding [dAMR03]. Corrigendum [KO01]. cosine [Cof03]. cost [Par03c, SBSA03]. Cotes [KS03a]. Coulomb [DHK02, HKD04, YEWE03]. Coulomb-like [YEWE03]. counter [Bak03]. counter-examples [Bak03]. Counting [MZ02]. Coupled [APT02, CDV03, CFSP04, DHW04, LVH04, LV03, Tom02]. Coupling [FMRW04, De 01, Dul04, HL04a, SKD04, SRRS04]. Cowell [vdHMS00]. CP [Ixa00]. crack [DND04, KL03a, Oht02]. Craig [Rix04]. Cramer [MAK01]. Creating [Kál02]. criteria [AO00b, CS04a, EP02a, HHC01, Oht02]. criterion [BB02, DS02b, GK02, ONU03, Par01, TC01]. critical [BH03, Mil00]. crossflow [BTSHK04]. crystal [KT02]. crystalline [IY03, UY04]. Cubature [CMS01, Pet04, CK01, MM02a, NS04a, Sto01, Wan01a, Xu01b]. cube [CK01]. Cubic [Coo03, Sya03, WdZ04, BS03, CO01, DM02b, DDPT00, DWZ02, Hom03, Hom04, Lai00, LY02, LH03, Miy03, WM01, WMA03, WA02]. cuboid [BH03]. Current [Hof00, IYO03, KTIS03, YOO02]. currents [FMRW04, Moo02]. Curvature [WM01, Xu04]. curve [AW03, CMRS00, JKS04, Mee02, WM02, WL02a, YY02]. curved [OOA03]. curves [AsKS04, FHMS04, KP02, MW02, Miy03, OP03, Sya03, WM01, WMA03, WZ04b, WLYL04, dSPT00]. curvilinear [Zhu01]. cutting [FLW01, KyGUY02]. cutting-plane [FLW01]. cyclic [CES02, DKK03, DSD04, LZ04, Son01, ZKO02]. cyclic- [DKK03]. cylinder [Tem00]. cylindrical [BH03, MS03b].

D [Aco01, AD02, AM03, DN02a, GLZ03, HKD04, KP03a, LF03, LF04, Liu02a, MS03b, Rat00, WKM04, Wan04a, Wün01a, Wün03a, vdH00]. DAE [BBKS00, BCJW00]. DAEs [Cas00]. d’Alembert [PR00]. damage [HSS01]. damped [BD00a, JKN00, Nay02, SW03]. damper [Tru04]. damping [Fab00, Miy02, WW03]. Darboux [Grü01, Koe03]. Darcy [MRT00]. Data [Sch04a, AKTvD00, AD00, AW03, BBW04, BW04a, CO01, yCjC01, Dem02, Dem04, DFT04, DFI02, Ing03, Isk03, KLY04, Lai00, LM02a, MO00b, OS04, RCZV04, Wol00b, dSPPT01]. data-dependent [AKTvD00]. Daubechies [Tas00b]. DEA [NOI02]. death [Mar03, vD03a]. decay [BE02b, Han03a]. decaying [BEM00b, BE02c, DN02b, Kui01a, Kui01b, Oou01]. decomposition [BS00c, Bog01, BD04b, CB00, Chi01a, DS04, DNS00, HY03, Han03b, Hop04, HC03b, HSW00, HLY04, LLP03, Les02, Leu00, LMO01, Lui00, Swa00, Waz01, ZZ02a, ZZ03b]. Deconvolution [Iqb03, MMK02]. Defect [Chi01a, CELM00, Enr00]. deferred [CMSV03, VV01]. defined [DLTS02, FGJ00, Sya03]. defining [BS02a]. Definite [Psa03, BW02, ELW00, ELW02, KyGUY02, MZ02]. deformation [APT02]. deformed [DKL04]. degenerate [CY00, Slo04]. degree [AsKS04, CGG00, DN00, GS02, KO03, NS04a, Pom01, Rab03, Sto01, Swa04, Sza02]. Delay [Zou02, BGZ00, BB03b, BB03a, BR00, Buc00, CS00a, CS04a, CZY03, CP00b, ELR00, FW00, GH03b, HW03, HHC01, JT02b, Kai03, Kot02, Kot03, KS02b, LWY02, LCF04, MS03a, NW00, Par01, Pau00, QM01, SBSA03, WF00, ZV04]. delay-dependent [Par01]. delay-integro-differential [ZV04]. delayed [EK01, HL04b, Jan03a, McC01, PJ03, WL04]. delays [BB00, Čer02, FW03, LG04, LS01, Par01, WZ04a]. delta [Tél00]. denoising [FK01, LL03b]. denominator [BM04a, ZZ02b]. densities [GDD04, HH04, IZ01, Tak03]. Density [sY01, Ing03, Kui01a]. Density-driven [sY01]. dependence [Hat03, HSV04, RZAG00]. dependent [AKTvD00, Bai01, BBW02, EW01, HS01, Len02, Les02, MD00, Par01, WL04, ZL03, vdHS01]. derivation [BJK03]. derivative [ABC03, Arg01a, EH04, Has04, Suá03, Tse04]. derivatives [Cof04, KO99, KO01, KO03, KCI02, LYF03, Shi04a, Sid00]. derived [ZD02]. descent [SHS02]. design [BBKS00, Fre02, GL04b, HM00, HCT^{+}02, KL04a, McC00, MP02b, RRWT00, WL01, YSNM02]. Designing [Pau00]. designs [TO03]. Detecting [Chr01]. Detection [CL01a, KATS02, dSPT02]. Determinacy [BGVHN01a]. determinant [KV04]. Determinantal [KN02, Kal00, KG00, KP01]. determinants [EM04b, PW02]. determination [Lee02, Pet02b, VID01]. determining [Alf00, KCB02]. Deterministic [WX02b]. detonation [aYtZ03]. developing [KLyD^{+}03]. Development [PCR04, Koh03b]. Developments [NZ00a, Bav01, Dah01, PRT00, SM04b, Yam00]. Devising [Coc01]. DFP [PT03]. diagonal [EM04b, MO00b, VMV04]. diagonal-plus-semiseparable [VMV04]. diagrams [CRSL00, OO02]. diameter [JT02a, YLD02]. Dickson [DRC02]. dielectrics [Bri04]. Difference [FR03, RLZ03, ALM03, AG03b, AK01c, BJK03, BM01a, CL04, CD03a, CG04a, CG04b, CP00b, CL03c, DN01, DN02a, EF02, FTY02, Fan03a, FR01, Fur01, GZ02, He03, HR04a, JT02b, KO99, KO00, KO01, KOH03a, KO03, KL01a, Kra03, Kum02, Let01, LYF03, LL01b, LB01, MT04, Moh00, Mou03, PS04, SS00, Sty04, TC01, WA00b, XCZ02, Yam02, aYtZ03, ZKO02]. differences [AGMMB00a, BS04, Dem02, Tho01, VMD04]. different [Che03a]. differentiability [EH00, Law02]. differentiable [Arg01b]. Differential [Ano01s, BMPV00, BPT02, Bav00a, Bav01, Bav03, BCM01, BC02, HR02a, Ism03b, Jan02, KK00, Koh01, Koh02, LW02, AM02a, AK00, AB04b, AGRZ03, AG01b, Bad01, Bak00b, BL01a, Bar04, BJK04a, Bav00b, Beh02, BF02, BGZ00, BB03b, BF01, BR00, BW04b, BH04, Buc00, BP01b, Büh00, BBM00, BT01, BHB04, But00, CS00a, CS04a, Cam01, CLP02, Čer02, Che03a, CH02b, Daa04, DZN00, DV01, DHP02, DS02b, Doh02, EFS02, EK01, ELR00, EPM00, ELW00, EKLW01, ELW02, EP03, EW01, Far02, FW00, FKR04, GZ02, GT04, GPS03, HP00, HY02, HG03, HR02b, HK01, JPW04, Jan03a, Jan03b, JW00, Kai03, Kha02, Koh03b, KW00, Kot02, Kot03, KS02b, KL03c, LV00, Lam03, Lee02, LP00, LZ02a, LHHW04, LCF04, LS01, MS03a, MFMGZ02, Mas03, MY03b, Mou03]. differential [MIB03, NRL03, NW00, Par03c, Pau00, QM01, Ran01, Röß04, Sch04a, Sch04b, Sch03d, SW00b, SZ04, Slo04, SAE04, SB04, SS01, SS00, TYZL04, Tho01, TA02, Tom02, Tse04, TA03, Vec00, VAR03, Vil03, WZW03, WW03, Win03, Win04a, Win04b, WF00, ZV04, dFN00, vdHS01]. differential-algebraic [BJK04a, CLP02]. differential-difference [Mou03]. differential-equation [Lee02]. differential-functional [Jan03b]. differentiation [BBBCD00, Cas00, KO00, Vec00, ZV04]. diffusion [eMKM02, eMBH02, AK01a, AH03, AG03b, Ban04, BM01a, BE03, BS00c, BJG02, Bog04b, CH02a, CVB04, CFSP04, CELM00, CZY03, CN02, CL01b, CJL03, Fan03a, FHM^{+}04, GVSJ01, GDD04, GLM00, GLZ03, HGI04, HC03b, JY01, JB04, KZ03, KL01a, LCL01, Len02, MMOS01, MOS02, MP03b, Pao01, PS02, Rem04, RZ03, SB01, Shi04a, ST02, Tid02, WL02b, lXsQsJ03, YL03, Yeh04, ZK04, Zou02]. diffusion-wave [GLM00]. diffusions [Ell01]. digital [Liu02b, Sch02b]. dilation [Han03a]. dilatonic [BTFY02]. dimension [DSW^{+}03]. dimensional [eMBH02, AA02a, AG01a, CG02, CN02, CK01, CR01, DN01, GZ02, GW04, Glo03, Gug00, GJ01, HW02, HC03a, Ji04, JS00b, Kan00, LR02, Moh00, MN00, NBS04, NS03, Nis03a, NS04a, PS03a, Qui02, SS02b, SW01, Ver01, Xu02, Xu03a, ZD02]. dimensions [KMCK02]. Diophantine [Pré00b]. dipoles [IYO03, YOO02]. Dirac [Car03, Sch02a, Tél00]. Dirac-orthogonality [Car03]. Direct [Ano01o, Ano01p, HOO03, LTT00, PGG02, PGG03, BBCH00, Bet00, BBCH02, LNBI02]. direction [DZN00, HWT04, Oht02, SW02a]. directional [CJ00, Gu04, NS04b]. directions [Hof00]. directly [ZZ02b]. Dirichlet [HC03b, LR04b, Mar04c, MY00, RSS04]. Dirichlet/Neumann [RSS04]. Dirichlet/Robin [HC03b]. disaggregation [Mar03]. disc [BS02a, UC03a, UC03b]. Disclosure [DFT04]. disconjugacy [GK02]. discontinuities [Wri04]. discontinuous [BBK04, CH02a, Coc01, Hei03b, RdAR04, Wri04, dSPPT01]. Discrete [HHR00, MN02a, MSM04, Pre00a, RS00, Shi04a, hWqX03, AGC00, ACV03, BVV04, CVB04, CMRS00, CZ03, DSD04, EM04a, FL04, FW02, Gan04, GHM01, HW02, Ing03, KZ03, KP00, KP04, LCL01, Lor03a, MSI04, MG01, ML03, PD03, RR01, SBSA03, Sun01, TRTG03, WA00a, Won04]. Discrete-time [MN02a, RS00, MSI04]. discretisation [GGDL04]. Discretization [RZ02, CN02, EHS00, Leu00, MP02a, MN02b, Sch00b, Swa00, Swa02, ZV04]. discretizations [Sch03d]. discretizing [Dul04]. discriminant [KyGUY02]. diseases [MG03]. disentangling [SJ03b]. Disjoint [Rhe02]. disk [Ari03, BC03, KL04b, dG01]. dispersal [LCC04]. dispersion [MD00, MT04]. dispersive [GN03, SWSZ04]. Displacement [DWQ04, YH02, BRS03b, KTIS03, LS02b, MKS^{+}01, Yan03]. dissimilarity [Hat03]. Dissipative [VAR03, Mat03, dFN00]. dissolution [VV00]. Distance [Rab03, CO02, RO03]. distances [AW03]. distinct [LGL03]. distributed [BA00, BW04a, Bor03, Dik03, Lou02, dBS01]. distribution [Age02, AM02b, AL03, BGP02, GS02, IZ01, Kra03, MFMGO01, MFMGZ02, Rem00]. distributional [MÁNM01, Tél00]. distributions [CO02, Car03, DMGVO01, EST03, FP02, Iqb00, Koz00, Mar03, RO03]. divergence [BL04, LR04b, PP00a, PP02b]. divided [Dem02]. divisibility [Tak03]. division [Dal01]. do [FW00]. Domain [Bog01, HSW00, Lui00, AGC00, BRS03b, BS00c, Chi01a, DNS00, HY03, Hop04, HC03b, JKN00, LLP03, Leu00, LMO01, NKK03, Swa00, UC03a, UC03b, VVD04, WWA^{+}04, XCZ02, ZZ02a, ZZ03b]. domains [BH03, CCL03, CDN01, De 01, Din02, LHHW04, Mar04c, Nis03a, OOA03, OOAS03, Sur01, YH03, ZL04]. dominated [BE03, Bog01, CELM00, EW01, KK03, ST02, Tid02, Tid03]. Double [MM03, Sug02a, BM01b, Cas04, FKM02, GKK00, Kar00b, KOZ03, Min04, MS01b, Ped03, Van02, WS02a, WS02b]. double-bracket [Cas04]. double-exponential [MS01b]. double-index [GKK00]. doublets [GK03]. drainage [Mal04a]. draining [NBS04]. Drazin [DWQ04, WW00b]. driven [DPM04, TWV02, sY01]. drug [eMKM02]. DS [IN03]. dual [DHK02, Dul04, GH01a, QSZ02, Rix04, VZ04, YF03]. dual-dual [GH01a]. Duality [Bri04, DNS00, LL02]. Duality-based [DNS00]. Dubins [Dal01]. due [AD01, RKS04]. Durfee [Mut02]. during [Bre00a]. Dynamic [ABOP02, AP02b, BD00a, CB00, CFHS03, DSVB04, DH02, Dra02, Elo02, EP02a, GK02, Rih03, Rix04, Sie02, Yan00]. Dynamical [DKST04, GJL^{+}00, LS04b, San03, SV00]. Dynamics [NW00, BO04b, CMKM03, FC04, Hui04, RDA04, RCZV04, SZ03a, SAE04, SS01, WO01].

Earth [DDG04]. Eastham [BE02e]. economic [Cai02]. economical [Bru00]. economy [STW00]. ECT [BM03b]. ECT-systems [BM03b]. ed [Ano02w]. eddy [BI03, FMRW04, YZCL04]. edge [Dul04]. Editorial [AO00a, Ano00d, Ano00e, Ano00f, Ano00g]. editors [Ano02u, Ano02v, Ano02x, Ano03-34, Ano03-35, Ano02r, Ano02s, Ano02t, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28]. EEG [HR04a]. Effect [ZFYY04, BPT02, DKST04]. Effects [HL04a]. efficiency [ZD02]. Efficient [BJ04, PS02, RZ04, SFK^{+}04, SEKW01, vdV02, BK02, FKM02, GLZ03, HS02, HGI04, HC03a, Mac02b, Pau00, PDVS04, Tru04]. eigenfrequencies [DMN01]. eigenfunctions [Bav00a, Bav01, Bav03]. eigenparameter [BBW02]. eigenproblems [GGM04]. Eigenvalue [And02, GvdV00, Mar04c, WA00a, And00, Bac04, CKR02, Cap04, DV00, De 01, DHZ04, KLT04, KSSV03, LZ04, Lui00, SS03b, ST00a, Wat00a]. Eigenvalues [HT01, Sch02a, AM02b, BMZ00, BR02, CP00a, IP01, KWZ04, MZ02, Mor00, TNW02]. eighth [LW02]. eighth-order [LW02]. eikonal [CR03]. elastic [Fab00, Leu00, Qui02, TQ03]. elasticity [ADG03, MG01]. elastico [Ari03]. elastico-viscous [Ari03]. elastoacoustic [BRS03b]. elastodynamics [PS03a]. Elbert [LMS03]. electric [Bri04, IYO03, YOO02]. electrical [BG02]. electroencephalography [HR04a]. Electromagnetic [ATG04, ADEL00, BLSS03, BLSS04, DHW04, HJ02, HH04, Hop04, NGGZ04, SIM02]. electromagnetism [Fre03]. electronic [aH02]. Electrostatic [Grü01]. electrotechnic [ADEL00]. element [AD02, AP02a, ARV00, Bai01, BC04, BO02, BE03, BRS03b, CG02, CL03b, lCHcH00, dCCSR00, DZN00, De 01, FTY02, FF02, GH01a, GGDL04, HGI04, HC03a, HZZ03, HLS^{+}03, Hug01, JY01, JW04, JE01, JE02, JL00, KTIS03, KT02, KRW00, KED04, KK03, LR04a, LCL01, LY02, LMO01, MD04, MP02a, NKK03, PCR04, Ran01, RdAR04, Rui02, SGG^{+}04, Sai04, SW00a, Shi02, Ste00b, Sur01, Tid02, Tid03, Tsu01, TYI03, WL02b, Wan04b, YH03, Yan00]. element-alternating [DZN00]. element-boundary [GGDL04, SGG^{+}04]. element/finite [Tid02, Tid03]. elemental [CB00]. elementary [JNS04]. elements [Aco01, CDV03, CS04b, Dul04, NGGZ04, PRK04, Ran04, Tho01]. elevated [MD00]. elevation [AD01]. Elimination [GM00a, AH04, BCM03]. Elliot [Lie04]. ellipsoid [PGG02, PGG03]. ellipsoidal [El 04]. elliptic [AQ04, BW04a, BH03, CCL03, CF00, CW00, CDN01, ENE04, FFX04, GHMY00, GZ02, JW04, Kno01, KW00, LMZ02, LR04b, MM00, Rui02, Sim00, SS03c, SdSP01, Sty04, TNW02, YL03, ZL04]. elliptical [CO02]. embedded [Fra02]. embedding [CO02]. Enclosing [Neu03]. endogenous [Cai02]. Energy [PP00b, AG01a, BLSS04, CD01, Fur01, WA02]. energy-minimization [WA02]. engine [HCT^{+}02]. engineering [WPS02]. enhance [TL04]. Enhanced [SZ01, DPM04, dFN00]. enhancement [BM03a]. enriched [DND04]. enthalpy [GN02]. entire [Wün01b]. Entropic [CFT02, SRD00]. entropies [DMFSR01]. entropy [Lar02, LL03a]. environment [HCT^{+}02, Yeh04]. environmental [ZFYY04]. environments [SC00]. epsilon [GMRS00, Thu02, Wen00]. equality [DSW^{+}03, PP00a, Zhu01, Zhu03a]. equality-constrained [DSW^{+}03]. equally [Wol00b]. equation [eMEM00, eMBH02, Abd02, Abd03, ADL^{+}02, AM03, AK01a, ADR02, AG01a, AMS00, Bad01, BRS03a, Ban04, BJK03, Bav00b, BO02, BB00, BB03b, BB03a, BJG02, BK02, Cam01, CS01, lCHcH00, CN02, CS04c, CR03, Daa04, DN01, DN02a, DHP02, DS02c, DH02, EM04a, EK01, EPM00, ELW00, Fer03, FR01, Fu04, Fur01, GS01a, GNPB01, Gla00, GLM00, GT04, Ham00, HY03, HC03a, HOO03, HR02a, HV02, IM01, Ism03b, Ixa00, Jeo00, Jeo01, JKN00, KMS03, Kan00, Kan04, KZ03, KCB02, Kob00, KS03b, KW00, Kou03, KL03a, KED04, KL03c, Lee02, Len02, Let01, LYF03, LD02, LZ01, LWW01, Liu02a, LCF04, LAT04, LSY04, McC01, Min04, Miy02, Pal02a, PRK04, PR00, RSS04, Röß04, San00, SIM02, Shi04a, SW00b, SZ03a, Slo04, Sty04, TML03, Tru04, WG02]. equation [Ye03, ZZ03b]. Equations [Ano01s, BMPV00, WBBF00, AS04a, Abd00, ABOP02, AM02a, AG00a, AJ01, ABG03, ALM03, Ant02, AGRZ03, Arg01b, Arg04b, AG01b, AP02b, Bac01, Bai03a, Bak00a, Bak00b, BPT02, BRS00, BL01a, BO04a, BGZ00, BB02, BS00c, BF01, BR00, DFM04, BW04b, Buc00, BP01b, BH03, Büh00, BBM00, BT01, BHB04, But00, CN00, CS00a, CS04a, CLP02, CF04, Čer02, CD03a, CY00, CHW02, CL03b, CH03, Chr01, CH02b, Cla03, CMKM03, CD03b, CDN01, CM04, CL02, CL03c, DZN00, DS02b, Doh02, Dom03, DMN01, EF02, EFS02, Elo02, EG00, ELR00, EP02a, EKLW01, EW01, Far02, FW02, FW03, FW00, FR03, FKR04, FLM03, GVV04, GKMN00, Gan04, GHMY00, GZ02, GN03, GLZ03, GJ01, GPS03, GK02, HW02, HGI04, Hau02, HY02, He03, HG03, HRE00, HR04b, HR02b]. equations [HK01, HSW00, HWT04, HC04, HSV04, Ism00, JPW04, Jan02, Jan03a, JKY02, JW00, JT02b, Joh02, JL00, JM00, JS00b, Jun04, Kai03, KYF04, Kaz02, KCMK02, KMCK02, Koe03, KK00, Kot02, Kot03, Kra03, KS02b, LV00, Lam03, Lau00, LF03, LF04, LZ02a, LY02, LR02, LC04, LHHW04, LLC01, Lin00, LL01b, LW02, LMP00, LRvSS04, LMO01, LB01, LS01, Mal04b, MS03a, MFMGZ02, Mas03, MOS02, MT04, Min01, MY03b, MAK01, MN02a, MIB03, NY03, NRL03, Nis03a, NW00, Öze04, Pao01, Pau00, PC00, PH01, Pie00, PT02, Ran01, RV00, Sch04a, Sch04b, Sha01, SW00a, SZ04, Sie02, SEKW01, SAE04, SB04, SXZ00, SS01, SS00, Swa02, Swa04, TRTG03, TC01, TYZL04, Tho01, TA02, Tom02, TLQ02, Tse04, Tsu01, Ues04, Vec00, VAR03, Vil03, WL03]. equations [WZW03, WW03, Wan04b, Win03, Win04a, Win04b, Wol00a, WA00b, WF00, XCZ02, YF03, Yan00, ZX01, ZV04, Zou02, dFN00, dG01, dGK01, dMPL03, vdH00, vdHS01]. equations-numerical [Joh02]. equiconvergence [VB01]. equidistribution [BM01a, Che03b, QST00]. equilibria [CZW04]. equilibrium [LY01, Par00a, STW00, Son02]. equistage [CP04]. Erdélyi [JV03]. Erdos [DJK01c]. Erratum [BGM01, Die02, QZ03, Win04a]. Error [Die98, Die02, KL01b, MS04, Nie03, Par02a, Tid03, UC03a, YH03, AJ01, ARV00, Arg03a, Bar02, BD04a, CMRS01, CC03b, DLTS02, GGM04, GM00b, Hug01, JL00, KSSV03, LY02, LY00, MRT00, Ran04, RSS04, SKSV03, Smi03, SS03d, TYI03, Won02b, Yan00, xZqZqT04]. errors [AP00, GPTT02, Hei03a, RCZV04]. essential [BMZ00, Hin02]. estimate [GM00b, ZLF01]. estimates [ARV00, And00, BH03, FK01, FPO01, HK00, Hei03a, KLT04, KL01b, LY02, LY00, NSZ04, PP00b, Pot00, Tid03, UC03a, YL03]. estimating [Age02]. Estimation [San03, BCM03, CC03b, GGM04, Hug01, KSSV03, LMYL01, Li03b, MO00b, PH01, RSS04, TO03, YH03, Yan00]. estimator [Ran04]. estimators [CMRS01]. Euclidean [QSZ02]. Euler [Aga03, BT01, CL03a, Kar01, KW03, LCF04, OO01, Oou01, Oou03, SAA01]. evaluating [EM04b, JE01, JE02, KY02]. evaluation [Bar02, BMS03, CRSL00, GHK03, IVD02, Par04a, Par04b, Vel01]. evaluations [BS03]. even [CG03a, KT01]. evidence [Sch02a]. Evolution [SD00, dG01, Cha02, Hau02, KCMK02, KMCK02, MG00, Mal04b, NY03]. evolutionary [JB04, WPS02]. Exact [LS02c, SW01, KCMK02, Mak04, PT03, Rem00]. examples [Bak03]. excitable [Miu02]. excitation [PS02]. exciting [Fre02]. exclusion [Geo03]. exhibiting [CFSP04, NJVA03]. Existence [AO00b, AR04, AA02c, CZY03, CZW04, DH00, Din03, EPM00, GHMY00, HP00, Hit02b, Jan03a, LV00, LG04, MP01, Rac00, AG01b, BMZ00, BH03, DDGH03a, DS02b, He03, Kaz02, KS00, LZ02a, LJ03, Mig01, Min01, Sim00]. expanded [GH01a]. Expansion [Lie04, RR01, ADG04, CCL03, GST03, Kza00, LL01a, Lar02, LMMD03, Lov00, Nay02, Par02a, Par02b, Par03a, Sto05, Zha03]. Expansions [SRLAD03, Cho03, Doh02, FL03, KR02, KL01b, KMV03, Lew03a, LP01, LM02b, Mac02a, Par04a, Par04b, Smi03, TVA03, Wal00, dAMR03]. experience [BBCH02]. experimental [BBKS00]. experiments [GCL02, Jeo00]. Explicit [DM02a, Her03, BS03, BE03, CH02a, Fra02, Fra03, Fra04, KK03, LH03, NSP04, TYZL04, XCZ02, dSPT02]. Explicitly [YYT02]. exploitation [MKS^{+}01]. exponent [Büh00]. Exponential [McC01, Qui02, VVV03b, DJK01a, Ham01, Han03a, IRVD01, IVD02, KOZ03, Kub01, KL03b, LL03a, Lu03, LM02b, Mac02b, MN01, MS01b, MM03, Par03b, PJ03, SJ03b, Sug02a, TYZL04, TO03, VAR03, WdZ04]. exponential-fitted [VAR03]. Exponentially [Fra04, VDVV00, BEM00b, BE02c, Fra02, JY01, VID01, WL02b]. exponentially-fitted [VID01]. exponentiated [Fas02]. exponents [BH03, DEL04]. expression [ELW02]. expressions [KO99, KO01, KOH03a, Mac02c]. Extended [BE04b, Cam01, Cas00, GG01, Ram03, XL02]. Extending [XLFC00]. Extension [Koh03b, BC03, Oht02, Sam01a, dAMR03]. Extensions [JV03, Ehr02, GKMN00, MÁN01]. exterior [HC03a, Mar04c, MG01, Nis03a, UC03a, UC03b]. external [Dav02]. Exton [RKS04]. extra [AB01]. extra-updating [AB01]. extra-updating/self-scaling [AB01]. Extrapolation [GJ01, Bre00b, GM00a, GGM04, GJHY00, HW02, JS00a, LD02, Sid00]. extrema [WM01]. extremal [JFW01, Li03b, PS01]. Extremality [CH02a]. extreme [Dim03, Gao01].

f [Kob00]. Faber [MN01]. fabrics

[ZFYY04]. face [CG02]. facing [YLC04]. FACR [HV02]. factorial [AS04a]. factorials [Sam01b, Sri02]. Factorization [FKR04, BV03, BX02, Gem00, HR02a, Tas00b, YK04, vdHS01]. factorized [XKC01]. factors [CF00, SS02a]. failure [KyGUY02, VBL^{+}04]. fair [Dal01, MW02, WM02]. Fairing [LLG04, Xu04]. fairness [KP02]. families [ÁNAA03, Miy03, Tas00b]. family [Bru00, CS02, GG01, Jun04, Kal00, KG00, KP01, TO03, Zhu01]. far [SW01]. far-field [SW01]. Fast [HC04, KP03b, SIM02, Weg01b, AP00, Aya03, EM04b, GHK03, IY03, MKS^{+}01, SGG^{+}04, VMV04, YH02]. Fatigue [DND04]. fault [dSPT02]. Favard [MÁN01]. FDEM [SA03]. FDM [SA03]. feasible [Xu03b]. featuring [Xu04]. Feedback [Kas00]. feedforward [LWT04]. Fehlberg [FKM02]. Fejér [DJK01c, VB01]. Fejér-type [VB01]. Fekete [SdSP01]. FEM [MG01, Mel02, SA03]. Fériet [CS03a]. fermion [BTFY02]. fertility [Cai02]. Feynman [CRSL00, OO02]. Fibers [Bog03]. fictitious [NKK03, VVD04]. Field [DHW04, GGDL04, KTIS03, LVH04, RdAR04, SDMV04, SIM02, SW01]. Field-circuit [DHW04, LVH04]. fields [AR03, BLSS03, BLSS04, BE04a, Bri04, Hop04, Miy03]. fifth [CGG00, Waz01]. fifth-order [Waz01]. filled [Liu01]. Filler [Ano01o, Ano01p]. film [DN02a, NBS04]. films [BCM03]. Filter [FQR00, BS02a, WG02, ZH04]. Filtered [BP04]. filtering [Isk03, MSI04]. filters [Tas00b]. Finding [Sia01a, SdSP01, Sya04, Van01, YF03, Bou04, KP01, MW04b, RZ02, SS02a, SC01, SK01, ZZ02b]. fingering [CFSP04]. Finite [BRS03b, CDV03, De 01, DLZ02, FTY02, Fur01, HLS^{+}03, KTIS03, KT02, MT04, Tsu01, aYtZ03, AA02a, AP02a, ARV00, ALM03, AK01c, Bai01, BJK03, BM01a, BC04, BO02, BE03, CG02, CS04b, CL03b, lCHcH00, CG04a, CG04b, dCCSR00, DZN00, DN01, Dai02, DN02a, DKL04, Dra04, EGV03, Fan03a, FF02, GH01a, GGDL04, HGI04, Has04, HR04a, Hug01, IM01, JT03, JY01, JW04, JL00, KO99, KO00, KO01, KOH03a, KO03, KRW00, Kum02, KED04, KK03, LR04a, LCL01, LY02, LMZ02, LMO01, MD04, MP02a, MS03b, Moh00, NKK03, NGGZ04, PS04, PCR04, PRK04, Ran04, Ran01, RdAR04, Rui02, SGG^{+}04, Sai04, SW00a, Shi02, SS00, Sty04, Sur01, Tho01, Tid02, Tid03, TYI03, Vel01, Vou01, WL02b, Wan04b, XCZ02, Yam02, YH03, Yan00, ZKO02, dGK01]. Finite-difference [Fur01, AK01c, DN01, XCZ02]. first [CK02a, CV03, EP02b, EP03, HY02, IRVD01, KP01, KO99, KO01, KCI02, Lew03b, Moh00, Sch00b, Tse04, dRM03]. first-order [EP02b, EP03, HY02, IRVD01]. Fisher [AK01a, San00, Zou02]. fit [PP02b]. fitted [Fra02, Fra04, JY01, PS03b, VDVV00, VID01, VVV03b, VAR03, WL02b]. fitting [AW03, Dem02, IRVD01, IVD02, JKS04, McC01, Sch04a, TYZL04]. FitzHugh [SZ03a]. five [CL01a]. Fixed [CN00, TO03, WA00b, BEM00a, DHK02, KS00, LV00, Lan04, Sim00, STW00, VVV03b]. fixed-point [Sim00]. Fixed-sign [WA00b]. flat [MSKS04]. flexible [CW04, SA03]. Flow [VVD04, eMKM04, AA02a, Ari03, Cas04, CHW02, DV01, Hat03, HC04, KCB02, KLY04, Las04, Mal04a, MT04, MSKS04, MN02b, OASO03, RdAR04, SEKW01, SRRS04, TWV02, WWA^{+}04, sY01, YLC04]. flows [AB04a, CLMR00, CM04, DPM04, DKL04, FF02, Fuj02, Ji04, Nag04, SW01, VWE04, Wat03]. Fluid [MMPZ04, Ari03, CMKM03, FF02, HLS^{+}03, Hui04, KCB02, Las04, Nag04, PGG02, PGG03, Qui02, Wat03, WO01]. fluids [LJ03]. Flux [SKD04, CC03b, SU04, ZL03]. FM [Ano02u, Ano02v, Ano03-34, Ano03-35]. Fokker [LAT04]. fold [BLSS03, BLSS04, CHW02, Hag01, HC04, LPSSP00]. followed [AH04]. following [MdB02]. force [HH04]. Forced [FPS00, FW03, HRK04]. forcing [HGI04]. Ford [Osa00]. form [BD00a, BH04, GDD04, GN02, KO99, KO01, KOH03a, LR04b, WLYL04]. formal [BZS02]. formally [Fre01]. forming [NSP04, PCR04]. forms [AG03a, AFGG03, Bar04, Fab00, Öze04, Riv03, Yu02]. formula [BD04a, CGMB03, Ham01, KS01b, SJ03b, Van02]. formulae [Cas00, CGM02, CMS01, DGVM02, DV01, FKM02, KS03a, Kar00b, Lew00, MM02a, MS02, NS04a, Not01, Not03, Pet04, Sto01, Wan01a, Xu01b]. formulas [AGRZ01, BDGV01, CK01, DD01, Ehr01b, Ehr02, Eli00, Jeo00, Jeo01, Kal00, KO00, Lau01, Lie04, MM03, Rad00, Vec00, YH02, Yan03]. formulation [Abd00, ADG03, BG02, BRS03b, GH01a, Yan02]. formulations [Dul04]. Forsythe [Bar02]. forward [HY03, HR04a, Les02, Liu02a]. forward-backward [HY03]. foundation [Peñ00a]. four [CJ00, DM02a, KS03b, NS04b, Sty04]. four-directional [CJ00]. four-step [KS03b]. Fourier [ADG04, AAD02, AD00, BS02b, CL03a, Fu04, Ing03, KP03b, Kza00, Lov00, Oou01, VB01, Wri02, Wri04]. Fourier-style [Lov00]. Fourth [FR01, Let01, ASN02, AK01c, DM02a, FKR04, Moh00, SK01, XCZ02]. Fourth-order [FR01, Let01, AK01c, FKR04, Moh00, XCZ02]. fractals [dACCS03]. fraction [BS03, dAMR03]. fractional [AB04b, EFS02, GKS00, KT00, KRT02, Kir00, LL02, LAT04, MG00, MT04, MWLS03, Miy02, Tor03]. fractions [BSS03, BL01b, Bea01, BL04, Dmy04, GPTT02, Lup01, TJ04, xZqZX02, ZZ03a, xZqZqT04]. Frame [DDGH03b]. frames [Han03a, LW03a, ZH04]. framework [Man00a, WFV01, WA02]. Fréchet [ABC03, Arg01b, Arg01a]. Fréchet-derivative [Arg01a]. Fredholm [Abd03, DFM04, FLM03, GKMN00, GJ01, HW02]. free [eMKM04, AD01, BE02b, EH04, Faz02, LCL01, Miy02, NBS04, NS03]. free-convection [eMKM04]. free-surface [AD01]. Frequency [IVD02, VID01, AGC00, AD00, BLSS03, BLSS04, Pet01, PW02, Pet02a]. Freud [KL01b]. friction [DHK02, Fuj02, HSS01, HKD04, Joh02, SM04a]. frictional [Gor04]. frictionless [CFHS03]. fringes [Dal04]. Froissart [Gil01, GK03]. front [CR03]. fronts [Zou02]. Fucik [BR02, HR03]. Fucik-spectrum [BR02]. Fully [Gan04, Gen03, MG01]. function [Abd00, Alz03, BPV02, BKR03, Bar04, BS00b, BP01a, BLM04, BBW04, BBC00, CS00b, Cof04, CK02b, DSZ02, FLR03, FL03, GST03, GL01, HKT^{+}03, JS03, Kar01, KRT02, LL01a, Lu03, Man00b, Mil03b, MY03a, Oou01, Par02a, Par02b, Par03a, Ped03, Ram03, RKS04, Sim04, Sto05, Tem03, Vel01, WX02a, WdZ04, Wün01b, Yak00, YTI02, Zho01]. functional [CL02, DS02b, EK01, Far02, HY02, HL04b, Jan03b, JW00, Kaz02, Kob00, LZ02a, LS01, Man00a, Mas03, MY03b, NRL03, SW00b, SZ04, SS01, Thu02, WW03, WL04]. functionals [Che01, Ehr01b]. functionology [HRS02]. Functions [JKVV03, AJDG02, ABG03, AHP04, AG02a, AG03b, Bac01, BPV02, BKR03, Bar03, BL01c, BE04b, BGVHN01b, BGVHN03, BVV04, BW03, CRSL00, CC03a, CH03, CS00b, Cho03, CS03a, Cla03, Coo03, CV01, CV03, Dac03, DV00, Din02, Dun03, EL01, Elb01, Eli00, Fab02, Fas02, FK04, GKK00, GJS03, Gau01, Gau02, Gla00, GLM00, Göt01, IYO03, JV03, KG00, Kál02, Kar00a, Kar00b, KPS01, KNS03, KT00, Kir00, Koh02, KR02, KS02a, KSSV03, Kub01, KL03c, Kza00, LL01a, LM02a, LW03b, Lew00, Liu01, LLL03, Lor03b, Lov00, Mac02a, Mac02c, MG00, MDSR04, MGL01, MO04, Mel02, Mil01a, MS04, MC05, Nay02, Nei02, OA04, Pan02, PP02a, PKA03, Ped03, PW02, Pet02c, PSS03, PNV01, Pom01, SS02a, SKSV03, Sch00a, Seg03]. functions [Ses04, SC04, SL03, Sto05, SS03d, TRTG03, Tem00, VB03, VB04, Wal03, Wri02, Wri04, XL02, Yak02, YYT02, Yoo03, dG01, dSPPT01]. fundamental [Koh01, OASO03, Sch02b, UC03a, UC03b]. Further [WL03, Koh03b]. fusion [aH02]. Future [Oya02, Hof00]. fuzzy [Bou04, DP02b, Hu01, Hu02, LV00].

Gabor [LW03a]. Galerkin [AD02, Coc01, HW02, HC03a, HL02, LLC01, Ran01, SB01]. Gamma [BP01a, RO03, Alz03, CS00b, GL01, Kar01, Par02a, Par02b, Par03a, Ped03]. Gammel [Bak03]. gaps [Sch02a]. gas [AS04b, HCT^{+}02, SFK^{+}04, YOT^{+}02, YLC04, ZFYY04]. gas-particle [YLC04]. GasTurbnLab [HCT^{+}02]. Gauss [AH04, CK04, Ehr02, Gau02, GG01, Hag01, JV03, Lau01, Li03a, Mel02, MS04, NHMS04, Tem03, XZ01]. Gauss-type [Lau01]. Gaussian [BS00b, DGVM02, Ehr01b, LMYL01, MO04, Smi03]. Gaussians [SVV01]. Gautschi [Alz03]. GCD [Sed04]. Gegenbauer [Bav00a, Bel01, Dim03, KNS03, KV01b, Lar02, Not03]. Gelfand [KR03b, Min04]. general [BJK04a, BS03, BL04, CCL03, CH00, DHP02, DL00, Fu04, HK00, Han03a, HW03, KL03b, LM02b, Man00a, Moh00, Noo01, Öze04, Ric00, RCZV04, SZ01, SRD00, SR04, VB01, YSNM02, ZL04]. general-purpose [YSNM02]. generalisation [AP04b, JE01, WM04a]. generalised [Bac04, JT03, Mac02b, Par03a]. Generalization [JS04, Kal00, Dmy04, Goh02, Ism03a, Lag01, Mil00, Oou03]. generalizations [Wün03b]. Generalized [AJDG02, Bel01, Bou04, CRSL00, Dat00, Hu02, Kai01, KY00, Yan03, ASB04, ABC02, Arg03a, Bav00b, BBW04, BDR02, BM03b, CT00, CG04b, CR03, DX02, DP02b, Din04, GKK00, Ham01, HV02, HHR00, Kir00, KK00, KSSV03, Liu01, MFMGO01, Mil03b, MWLS03, MWL04, NP03, PR00, PT02, Ron01, SS03b, ST00a, Sid00, STW00, Son02, Tam03, Tan02, TML03, Tse04, Wan01a, WW01, WZ03, XZ01, Yan02]. generated [BM03b, CL01b, EP03]. generating [Dun03, HRS02, KP03a, LHYaC03, NMST02, Pan02]. generation [KG00, KR03a, YOT^{+}02]. generator [KS03b, SZ03b, Tan02]. genetic [KL04a]. geodesy [Nie00, SW02b]. Geometric [ABG03, Iwa02, LW03b, LN03, TQ03, WO01, BLSS04, BP01b, CF04, Koz00, LLG04, Mar03, Nak01, PR00, Wün03b]. geometrical [BLSS03]. geometrically [Lou02]. geometry [DZN00, KCB02, MS03b, Sch02b, Wan00a]. Gibbs [JS04, Pas04, SJ03a]. Gilbert [CS04c]. Ginzburg [DH00]. given [AK00, Wri02, Wri04]. Global [BM03a, CS02, LJ03, LCC04, NS03, RCZV04, SHS02, Dul04, HL04a, Ji04, Leu00, LF01, PRT00, Xu02, Xu03a]. Globally [CMKM03, PT02]. GMRES [Aya03, CCM01]. Gompertz [JKS04]. Gontscharoff [Won02a, Won02b, Won04]. Good [SW02b]. goodness [PP02b]. goodness-of-fit [PP02b]. governed [KW00]. governing [RCZV04]. GPS [AS04a]. graded [JNT02]. gradient [CS02, KRW00, LWT04, SS04, VS01, VALM00, WX02b]. gradient-related [SS04]. Graev [KR03b]. granular [FC04, RDA04]. GRAPE [Mak02]. graphs [AA02b, DWZ02, YLD02]. Gravity [Moo02, AD01, BBK04, BTFY02, DPM04]. Green [AG02a, AG03b]. Grid [MSKS04, BM01a, CACK04, GL04b, TVV04]. grids [FMRW04, Glo03, Zeg04]. Ground [LLXS04]. groundwater [KY04, Mal04a, MP02a]. Group [eMEM00, eMKM02, eMBH02, CO02, DEL04, Mar04b, Öze04, Van02]. groups [BS02a, Dun03, LMP00, SC04]. growth [AHP04, Cai02, DND04, KT02]. GSMAC [KT02]. Guaranteed [Par03c, Min04, PR04, SBSA03, YH03].

Hadamard [Lag01, Par04a, Par04b]. Hahn [FR03, KY00]. half [MO02]. Halley [EH00, EH04]. Halley-type [EH04]. Hamilton [HWT04, LC04]. Hamiltonian [AG00b, CG04a, Kno02]. Hamiltonians [VASF03]. Hammerstein [HRE00]. hand [LR04b]. handling [Kál02]. Hans [CP00a]. harmonic [Alz03, Che03a, CJ04, CD03b, GGDL04, IYO03, Nak01, PP02a, RLZ03]. harvest [LCC04]. having [Bav00a, Bav01, Bav03]. head [KLY04]. heat [AG01a, CC03b, DN01, DN02a, FL04, Ham00, HY03, Kou03, Las04, Liu02a, MSKS04, VMD04, ZFYY04]. heated [BTSHK04]. heaters [SDMV04]. heavy [Ism00, MP02a]. Heine [CGMB03]. Heine-type [CGMB03]. held [ERV04]. helical [FHMS04]. Helmholtz [HC03a]. hemivariational [Mig01]. Hermite [AMBP^{+}01, CG00, CGMB03, DLM^{+}00, DJK01c, Ehr02, ELW00, Gru03, Hag01, KP03a, LS02a, LY02, Lor00, MW02, SV00, WdZ04, Wün01a, Wün03a]. Hermitian [BW02, Psa03]. Heuristic [Dra02]. hexagonal [Liu02b]. hierarchical [Glo03]. High [AZPF04, Dom03, GN03, Kha02, LC04, Mat03, SSV04, VV01, AsKS04, AL03, AB03, BLSS03, BLSS04, CP04, CM04, CM03, Duf00, GZ02, GLZ03, HS02, SIM02, Swa04, Ver01, WL01, WZW03]. high-degree [Swa04]. high-dimensional [Ver01]. high-frequency [BLSS04]. High-order [AZPF04, Dom03, GN03, Kha02, LC04, Mat03, VV01, AL03, GLZ03, WL01, WZW03]. High-performance [SSV04, Duf00]. Higher [AM02b, BI03, Daa04, GM00c, KO99, KO01, KO03, LZ02a, SS00, Sty04, TQ03, WA00b]. higher-degree [KO03]. higher-order [BI03, Daa04, SS00, TQ03]. highly [CCL03, CMSV03, Eva04, JPW04, JT03]. Hilbert [DVM02, DKM^{+}01, DD01, Has04, Sch00a, SVV01, Vou01, Weg01a]. Hilliard [Ye03]. Historical [Yam00]. history [Eve04, GS00, Nie00, Tho01]. HJB [ZZ03b]. HLCP [XZ01]. HODIE [CG04b]. hodograph [WM02, WM04a]. Holling [WL04]. holomorphic [Din02, KPS01, Sai04]. homeomorphism [Lup01]. homogeneous [AFGG03, CGG00]. Homogenization [Las04]. homotopies [Wat02]. homotopy [DKK03, DvM01]. homotopy-like [DvM01]. Hopf [BBMS04, FW00, WZ04a, WS02b, WF00]. Hopfield [CZW04]. horizontal [ATG04]. host [MP01, MP03b]. host-parasite [MP01, MP03b]. hot [APT02]. hp [Mel02, AP02a]. hp-MITC [AP02a]. Hurwitz [BE04b, Gem00, HKT^{+}03]. Hybrid [LV02, SS02c, CS04b, CH00, DZN00, FF02, GGDL04, KS03b, MRT00, SGG^{+}04, Tid02, Tid03, Xu02, Xu03a]. hydrodynamic [SKD04]. Hydrodynamical [Gro02]. hydrodynamics [Ben02, LLL03]. hygroscopic [ZFYY04]. hyper [Gen03]. hyper-spherical [Gen03]. hyperasymptotic [Par04a, Par04b]. Hyperbolic [ZL03, BSKP04, Coc01, DB01, GH01b, Kan00, Koe03, LYC02, LRvSS04, Min01, Mon01, NS03, Tay00, Tid02, Tid03]. hypercone [Tél00]. Hyperelliptic [FK04]. Hypergeometric [AGRZ03, Gla00, Van00a, AT00, BPV02, BS00b, BE04b, CRSL00, Cam01, CS03a, DSS03, DL01b, DM02b, FL03, Gau02, JV03, Kar00a, KT00, KRT02, KR03a, KR03b, LL01a, Lew00, Mil01a, Mil03b, Nis03b, RKS04, SRD00, Ses04, Sto05, TL01, Tem03, VD03b, Van02, Vid03]. hypergeometric-Bessel [KRT02]. Hypergeometric-type [AGRZ03]. hypergraph [KL02]. hyperoctahedron [Sto01]. Hypersingular [AD02, GS01a, Ham00]. hypersonic [VWE04]. hypersurfaces [PP00b]. hypotheses [Arg01a, Mil03a]. hypothesis [Arg04b, Lac03].

IA [QM01]. ICCAM [BGPW02, GVW04]. ICCAM-2000 [BGPW02]. ICCAM-2002 [GVW04]. ice [CDV03]. ICSF2002 [JKVV03]. ideal [BKR03]. ideal-function-like [BKR03]. Identification [AMS00, FI03, IZ01, BX02, Faz02, HOO03, IYO03, Kno01, Mal04a, MKS^{+}01, SSV04, YOO02]. identities [CL03a, Dat00, KR03a]. identity [Aga03, KR03b, Lie04, VD03b]. IFC [Ano02w, Ano02x]. II [Beh02, BLSS04, BBCH02, BBW02, Bre00b, Din02, Kar00a, MN00, Par04b, Sch00a, dBS01]. III [Ano00x, LP01]. IIIC [QM01]. ill [CMRS00]. ill-posed [CMRS00]. ILU [GS04b, SZ01]. Image [LL03b, CMV01, Liu02b, MS03b]. images [PS01, Peh03]. imaginary [GST03]. imaging [LLXS04, PP00c]. immunity [MG03]. impact [Duf00, PCR04]. impedance [KL03a]. impermeable [LS02c, Mak04]. Implementation [DHK02, GS04a, Aya03, BBM00, FFX04, Joh02, MRV04]. implementations [FGJ00]. implications [Goh02]. implicit [BvdV01, BV03, Bru00, BJ04, CH02a, CGM02, DZN00, DSVB04, DP02b, FT03, LCF04, LRvSS04, MS03b, NSP04, RZ02]. implicit/explicit [NSP04]. implicitly [Sya03]. Improved [Neh03, Yan00, Arg03a, Liu01, Thu02, Zhu01]. Improvement [SS02a, IN03]. Improving [Wan00b]. Impulsive [SZ03b, BB03a, HY02, HG03, HK01, LS01, SS01]. including [PJ03]. Inclusion [Koe03, Pet03]. inclusions [AO00b, DL00, DX02, DP02b, HP00, Hei03a]. incomplete [Par02a, Par02b, Par03a, YK04, dB02]. incompressible [BJK03, BSKP04, DKL04, LJ03, Nag04, RdAR04, SEKW01, TWV02, Wat03]. inconsistent [Yam02]. incorporating [YTI02]. incorrectly [Iqb03]. increasing [GM00b, Gro02]. indefinite [BLM04, MM03]. independent [Sug02b]. indeterminate [Chi01d, GHM01]. Index [Ano00a, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano00v, Ano01a, Ano01f, Ano01g, Ano01h, Ano01l, Ano01i, Ano01j, Ano01k, Ano01b, Ano01c, Ano01d, Ano01e, Ano01t, Ano02i, Ano02j, Ano02l, Ano02o, Ano02k, Ano02m, Ano02n, Ano02g, Ano02h, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano03a, Ano03g, Ano03m, Ano03b, Ano03c, Ano03h, Ano03d, Ano03i, Ano03j, Ano03e, Ano03f, Ano03k, Ano03n, Ano03l, Ano03-36, Ano04a, Ano04b, Ano04k, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04j, Ano04h, Ano04i, Yak02, Ano03-37, Ano04-29, DRC02, GKK00, Sch03d, WW00a, Win03, Win04a, Win04b, Yak00]. indexes [CHW02]. indirect [Jeo01]. induced [DWZ02, Zou02]. Inductance [Kar00a]. induction [Dra04, SDMV04]. industrial [Sch04a]. Inequalities [Alz02, BE04a, GL01, QZ03, ASB04, AQ04, BS04, CFT02, Dev02, Din03, Din04, DNS00, GL03, HS01, Hu01, Hu02, ML03, Mig01, Noo01, SR04, XZ03]. inequality [Alz03, CT00, Goh02, Han03b, Hil02b, SQ01, Zhu03a]. inertia [LL00]. Inertial [HL02, MO03]. inexact [CG02, Cui02, PT02, SS04]. infinite [ABO02b, BL01a, EST03, FLW01, Faz02, LS01, OO02, Tak03, MN00]. Infinite-dimensional [MN00]. inflow [KK03]. influence [BCM03]. Information [BS02a, DMFSR01, KATS02, Wri02, Wri04]. inherit [Fur01]. inhibitor [HL04a]. inhibitory [GH03a]. inhomogeneous [Cap04, Daa04, aH02, Yeh04]. Initial [SC00, AD01, BS00d, Cas00, DB01, Gru03, SA00, Wol00b]. initial-boundary-value [Gru03]. initializers [CP04]. injected [BTSHK04]. injection [MMPZ04]. Inner [AGMMB00a, BS00a, Gu04, YTI02]. inputs [LWT04]. insoluble [NBS04]. instability [AVG^{+}04, Ben02, BBMS04, DDG04]. instruments [HT01]. integer [AdMR02, CK02b, DvM01, KV04, Mut02, SD00]. integers [ELW04]. Integrability [VASF03, Sch04b]. Integrable [Lor03a, BSS03, MN02b, Nak01]. Integral [Abd02, BMPV00, CD03b, CK02b, KL03a, Abd03, ADG03, AG01a, BRS00, BRS03a, BO04a, DFM04, BMN01, BW04b, BS04, Chr01, Dom03, DMN01, EG00, FCP02, FLM03, GKMN00, GS01a, Gan04, GKS00, GDD04, GJ01, HW02, HRE00, HR02b, HSW00, HSV04, Jan02, Jeo00, Jeo01, JM00, JS00b, Jun04, KCB02, KR03a, LL01a, Lau00, LR02, LLC01, LD02, Mil00, Pie00, Riv03, Sto05, TVV04, Yak00]. integrals [Abd02, AQ04, AGRZ03, CRSL00, CC03a, Cof03, Die98, Die02, Doh02, Gen03, Hor01, JE01, JE02, JV03, KT00, KRT02, KC00b, KC00a, KY02, LL03a, LP01, Mac02b, MP03a, MO02, MO00a, NMST02, NS04a, Par04a, Par04b, SRD00, Sch00b, Vel01]. integrands [CF00, KCI02, Ver01]. integrated [Doh02]. integration [AD02, BV00, BV03, BP01b, CLMR00, Cas04, CF04, Col01, Coo02, Ehr01a, Fra03, KS03a, LN03, MM03, Nie03, NSP04, Pet02c, SS03b, SA00, Sch00b, Sug02b, SS03d, Sza01, TQ03, WX04, vdH00]. integrators [HW00, KMS03, Lu03]. Integro [Bad01, AB04b, HR02b, Kot02, Kot03, MIB03, SB04, SS00, Vec00, ZV04]. Integro-differential [Bad01, AB04b, HR02b, Kot02, Kot03, MIB03, SB04, SS00, Vec00]. integrodifferential [MN02a]. interaction [MP03b]. interactions [CL01b]. Interest [MA04]. interesting [LN03]. interface [HL04a, JW04]. interfaces [LGL03]. Interior [PW00, BM03a, CACK04, FHM^{+}04, GG01, MM00, Shi04b, Zhu03a, Zhu03b]. Interior-point [PW00]. internal [MT02, WY02]. International [JKVV03]. interpolant [CMSV03, DLTS02]. interpolants [CM03, SL03, WG04, Zhe04, ZZ02b]. Interpolating [CVB04]. Interpolation [Bre00b, DSS00, DN00, Li02, Müh00, MN00, NS04b, NZ00b, TW04, Aco01, CG00, CO01, DLM^{+}00, DJK01c, DJK01a, DJK01b, Dik03, DDPT00, GS00, Gru03, Ham01, JT03, Kal00, KL03b, KL01b, Lai00, LM02a, LCZ04, Lor00, MO01, MW02, NZ00a, RV00, Sya03, TT02, VB04, WW04, Wan04a, WA02, Won02b, Won04, dBS01, dB02, dB03]. interpolational [MM02a]. interpolations [WW04, Yan03]. Interpolatory [Not01, Gen03, KC00b, KY02, LLG04, MN01, Not03]. interpolatory-type [KY02]. interpretation [Grü01]. interproximation [CM01]. intersection [Kai01]. Interval [AM00, Wol00a, AKM03, BDGV01, FPO01, GG01, Gra04, LZ01, PFO03, Sch03a, SK01, VB03, dGK01]. intervals [ABO02b, And02, Chi03, Faz02, FLM03, Mar04a, MO01]. Introduction [Buc00, Ter03, Thu02]. invariance [Pöt02, Van02]. Invariant [Sto01, BP01b, CK01, GLM00, Ips00, SB01, TA03, Wan01a]. invariants [IZ00]. Inverse [Peh03, PS03a, BK02, BT00b, CK04, DHZ04, DWQ04, ENE04, JS04, Kno02, Nac04, Pas04, PS01, Pot00, RS04, WH00, WW00b, WW01, WZ03, dMPL03]. inversion [AGRZ01, Bar03, KRW00, MK01, Sey01, YH02, Yan03]. inverted [SW03, WS02a]. invertible [Kaz02]. investigate [HRK04]. investigation [CDN01, WS02a]. investor [WFV01]. involving [AGMMB00a, Arg01b, BS04, Dav02, HKT^{+}03, Tse04]. IRK [BCJW00]. irreducible [Bar04, SS03a]. irregular [Büh00, DZN00, LW03a]. Isochronous [CGG00]. isolation [RZ04]. isomorphic [LGL03]. isothermal [KSCI00]. Isotropic [Rem00]. Issue [ERV04, Ano02w, Ano02x, QZ03]. issues [BBM00]. Itô [TVA03]. iterated [HW02, Man00b, Wen00]. iteration [BW02, Bai03b, CRSS00, Cha04, FGJ00, HC03b, KG00, Lui03, Oht03, PR04, ZLF01]. iteration-by-subdomain [HC03b]. iterations [Bai03a, EH04, FS00]. Iterative [CR00, CP03, Din04, Kob00, SvdV00, ASB04, AH04, ABG03, ABC03, BJK04a, BJG02, Bog04a, Bog04b, CMRS01, Cao02, CL03b, CS04c, Dal04, EMT01, HC03a, KZ03, KNSmG00, KHMN02, LR04a, LZ02a, LZ04, Löt00, Mai03, SS02a, SW00b, SZ04, TLQ02, WH04, YK04, vdV02]. Ito [Vil03]. IV [WBBF00]. IVPs [PS03b]. ix [QZ03].

J [BGM01, Die02, KO01, PGG03, QZ03, Sto05, Van01, Win04a]. Jacobi [AdMR02, Bav00b, Bav03, Bec01, BM01b, JNS04, KK00, Las02, Li03b, LC04, Man00b, WH04, Wün03a]. Jacobin [HWT04]. jets [BTSHK04]. Jordan [Jea04]. jump [MSI04, lXsQsJ03]. jump-diffusion [lXsQsJ03]. jumps [Kra04].

K. [VVV03a]. Kalman [MSI04]. Kampé [CS03a]. Kantorovich [Arg01b, Arg03b, Arg04a, Arg04b, HS00, Her01]. KdV [FW02]. KdV-type [FW02]. Kepler [Pal02a]. kernel [AD02, Bad01, Dom03, HSV04]. kernels [EP02b, GKMN00]. kind [Abd03, AGRZ03, CV03, GKMN00, GST03, LLC01, Moh00, ZZ02a]. Kinetic [BJK03]. Kirk [Akk02]. Klein [DD01]. knot [CS03b, VVV03b]. knots [AS04c]. Koekoeks [Bav00b]. Kolmogorov [Nag04]. Kontorovich [Nay02]. Korobov [Mai03]. Korobov-like [Mai03]. KPP [Zou02]. Krall [VYZ01]. Kramer [EP02b, GHM01]. Krawtchouk [SC04]. Kronecker [HV02, LS04a, Van00b]. Krylov [Fre00, Ism00, WW00a, Zha02]. Krylov-subspace [Fre00, Zha02]. Krzyz [Szy03, Sam01a]. Kummer [Mil03b]. Kummer-type [Mil03b]. Kutta [BE03, BM02, Fra04, VV01, BM02, BJ02, CLMR00, CGM02, CJ03, DM02a, Fra02, FKM02, IZ00, Kot02, LSY04, Röß04, TA02, TS02, VDVV00, VID01, VVV03b].

L [Sto05]. L. [Akk02, Arg03b]. ladder [RLZ03]. lag [Las02]. Lagrange [Aco01, DJK01a, DJK01b, KL03b, KL01b, LCZ04, MO01, Zhe04]. Lagrangian [BE03, Öze04, SD03, YOT^{+}02]. lags [Rih03]. Laguerre [FR03, AGMMB00a, Bav01, Dac03, DVM02, Dim03, Ehr02, ELW04, Gat02, KMV03, MFMGO01, MS01a, NP03, Pet03, SRLAD03, Sch01, Szy03, Wün01a, Wün03a]. Laguerre-like [KMV03, Pet03]. Laguerre-type [Sch01]. laminar [AB04a]. Lanczos [BRZS00, BZS02, CRSS00]. Lanczos-based [BZS02]. Lanczos-type [BRZS00]. Landau [Sto05, Alz02, CS04c, DH00, KS02a]. Landweber [CK04]. Laplace [Bar03, BW04b, BW03, Fab02, HOO03, LP01, Par04a, Par04b]. Laplace-type [BW04b, LP01, Par04a, Par04b]. Laplacian [BR02, Dav02, He03, HR03]. Lardy [SS03c]. Large [Aas02, Tem03, YZCL04, BI03, BT00a, BZPFB04, CR00, CMRS00, CJ03, Dev02, GJL^{+}00, LM02a, LP00, RS00, Sch04c, VMD04, YSNM02]. large-scale [BT00a, GJL^{+}00, LP00, YSNM02]. Largest [YLD02]. lattice [IN03, Liu02b, Lor03a, hWqX03]. lattices [AB03, Dra04, Peh01, RLZ03]. Laudatum [BE02e, BEE04]. Laurent [VZ01]. Lauricella [FL03]. law [MRT00]. Laws [Dev02, BBK04, Coc01, DM02a, LYC02, Mon01, ZL03]. layer [DDG04, HJ02, Las04, MSKS04, Shi04b, WY02, YZCL04]. layered [ATG04, ZL04]. layers [AS04b, BD04b, CACK04, FHM^{+}04, MMOS01, NJVA03]. leading [Jea04]. learning [CF04]. Least [Hei03b, JKS04, Bai01, BVV04, CP03, Dem02, Hin02, HR00, JS03, KED04, LS00a, Nie00, XKC01]. least-square [CP03]. Least-squares [Hei03b, JKS04, BVV04, HR00, JS03, KED04, XKC01]. Lebedev [Nay02]. Lebesgue [Alz02, Kub01]. Leffler [Kir00, MG00]. left [ELW00, ELW02, MZ02]. left-definite [ELW00, ELW02, MZ02]. Legendre [ELW02, Far00, Ing03, Lov00, PFO03, Ye03]. legs [BV02]. Leja [CVB04, CD01]. length [Che03b, VID01, YY02]. Leonard [Ter03]. less [BV02]. level [FI03, Mar03, SZ01]. level-geometric [Mar03]. level-set [FI03]. Levenberg [KYF04]. Levin [Hom00]. Levin-type [Hom00]. Lévy [AP04a]. Lewy [CP00a]. Liapunov [GS04a]. Lie [BS02a, Hil02a, KP03a, Yu02]. Lie-theoretic [KP03a]. Liénard [FW03]. Liénard-type [FW03]. life [IM01]. life-span [IM01]. Lifshitz [CS04c]. like [Arg03a, BKR03, DvM01, DL00, Din03, HRE00, KMV03, Mai03, Pet03, TJ04, Wat00a, Yam00, YEWE03, ZD02]. likelihood [MO00b]. Limit [EGV03, GS02, IP01, BB02, Bri04]. Limitations [Shi04b]. limiters [SKD04]. limits [BSKP04, Sid00]. line [DFM04, CFMV03, CS02, MO02, MO00a, Mou03, PT03, SS04, Zhu00, WJ01]. line-SOR [WJ01]. Linear [AGC00, Ano00x, Tor03, AH04, AKM03, AG00a, AP02b, AG02b, BW02, Bai03a, BB01, Bar04, BMS03, BS00a, BW04b, BH03, Büh00, BJ02, BD00b, CCH02, CRSS00, Cao02, CW00, CGG00, Che01, CFT02, Coc01, Daa04, DSS00, De 00, DE00, Duf00, EFS02, EMT01, EKLW01, FQR00, FFX04, GZ02, GSS03, HJ02, HHC01, IN03, Kob00, KLM01, Li03a, LZ02b, LCF04, LMMD03, LRvSS04, MFMGZ02, Min03, MAK01, Nac04, O’L00, Riv03, RLZ03, SvdV00, SZ01, SIM02, SSV04, Sim00, SW01, Son01, SXZ00, Tam03, TC01, Tid02, Tid03, Tru04, Tse04, VMV04, Wat00b, WW00a, WZ03, Woź01, XL02, Yan02, ZLF01, Zha02, Zhu03a, Zhu03b, vdV02, ERV04]. linearization [AGRZ01, PRK04, SRD01]. Linearized [BB03b, BB03a, NSZ04, Nis03a, SEKW01]. linearly [BJ04]. lines [SWSZ04]. linesearch [BM03a]. Liouville [AM02b, BBW02, BLM04, BBW04, BMZ00, BEM00b, BE02b, BE02c, CS01, DH02, GGM01, GGM04, GM00c, KWZ04, Mar04b, MZ02, Tas00a, Tas04]. Lipschitz [MS03a]. liquid [YOT^{+}02]. List [Ano01q, Ano01r]. little [AGMMB00a]. LMF [qJkSlS04]. LMF-based [qJkSlS04]. load [VBL^{+}04]. loading [DSD04]. Lobatto [Mel02, MS04, QM01]. Lobatto-based [Mel02]. Local [JY01, Wri04, AG03b, BS00a, BPW02, BM03b, Hug01, KYF04, LH03, MS03a]. localized [Fre03]. locally [Glo03, Kaz02]. Locating [DSZ02, PNV01, HC04]. location [BE02c, FD00, Hin02, Wri04]. locking [SR04]. lofted [HM00]. log [Cof03]. logarithmic [Dom03, LG04]. logistic [BB00, JS03]. long [KS03a]. long-time [KS03a]. longest [Gro02]. look [LHYaC03]. loop [CRSL00]. Lorentz [DEL04]. lost [BYY03]. Lotka [DKST04, IM01]. low [Bet00, GHR02, Kza00, Wri04]. low-order [Wri04]. low-thrust [Bet00]. Lower [GJS03, DHP02, GS01b, ZW02]. Löwner [RV00]. LP [KLyD^{+}03]. LPABO [KLyD^{+}03]. LPAKO [KLyD^{+}03]. lubrication [AS04b, GNPB01, KL04a]. lumped [Sai04]. Lur’e [HW03]. Luswili [Sto05]. Lyapunov [DEL04, RS00, SS01, Tru04, WT01].

M [Akk02]. Macdonald [CV01, FLR03]. machines [HCT^{+}02, TL04]. Machining [MD04]. macro [APT02]. macro-micro [APT02]. Magnetic [SDMV04, BE04a, GGDL04, Hit02b, KTIS03]. magnetodynamic [Dul04]. magnetohydrodynamics [BBMS04]. magnetostatic [KSCI00, KCMK02]. MAH [AVG^{+}04]. MAH-3 [AVG^{+}04]. major [KATS02]. make [SA03]. management [RRWT00]. manifolds [FPS00]. many [Iwa02]. many-body [Iwa02]. map [Bog03]. mapping [OOA03, Weg01b]. mappings [DP02b, LV00, OOAS03, PS01, Peh03]. maps [BEM00a, Mar04c]. Marchenko [BPW02]. marching [CR03]. Markov [DLZ02, MO04]. Markovian [MSI04]. Marquardt [KYF04]. mass [BD00a, KRW00, Sai04]. mass-spring [BD00a]. Master [BM01b]. matching [MSM04]. matchings [DWZ02]. Material [Doi02]. Materials [Fre02, aH02, TVV04]. Math [BGM01, Die02, KO01, PGG03, QZ03, Van01, Win04a]. Mathematical [Gor04, KOH03a, MOTT03, LHYaC03, Lor03b, MG03, Sla00, SVV01, VV00]. Mathematics [Ano00w, Wol00a]. matrices [Bar03, Bec01, BS00d, BKB04, BGVHN03, CKR02, DHZ04, GM00a, GS04b, HLY04, Ips00, JNS04, KRW00, KLT04, LZ04, Man00b, MO00b, Psa03, RSvR03, RS00, Sad01, Son01, YH02, Yan03]. Matrix [Gu04, Hil02a, KATS02, SV00, BS02b, BV03, BRZS00, BW03, CFMV03, Cha04, Dal04, DHP02, DS02a, EP02a, Fan03b, Fre01, HK00, Han03a, HSV04, Jea04, KV04, Koh01, Koh03b, KNSmG00, Kui01a, LMP00, Lu03, MK01, MN01, ONU03, OA04, PP02c, Psa03, Sza01, TML03, WW01]. matrix-valued [Fre01]. max [Bac04, ZLF01]. max-norm [ZLF01]. max-plus [Bac04]. maxima [LB01]. maximal [YLD02]. Maximovic [Sch04b]. Maximum [DWZ02, Dam03, GS02, MO00b]. maximum-likelihood [MO00b]. Maxwell [CD03b, HH04, LF03, LF04, Sha01, XCZ02]. May [ERV04]. Mazur [BD04a]. McKendrick [IM01]. MCPs [BM03a]. mean [Alz03, DJK01c, DJK01a, DJK01b, Xu04]. means [Age02, Fas02, LM02b, Nak01]. measure [And02, AP02b, AA02c, CH02b, Dia03, Elo02, EPM00, Hil02a, Ifa01, LS04b, LR04b, Pöt02, Sia01a, Van01, Won02a, Won02b]. measurement [Hei03a]. measurements [Hei03a, Iqb00]. measures [DDN01, IK03, Mlo01]. mechanics [CSRB04, FC04, Iwa02, YSNM02]. Mechanisms [SRRS04]. media [AA02a, CCH02, FCP02, JT03, NGGZ04, RDA04, sY01]. medium [yCjC01, MT03]. Mehler [CGMB03]. Meixner [Ara04, AGMMB00b, KPS01, Let01, Ron01, SC04]. Mellin [ÁNAA03, EG00, MP03a, Mil01a]. members [KP01]. membrane [eMKM02]. memorial [LMS03]. memory [HGI04, WZ04a]. memory-efficient [HGI04]. Menten [HL04b]. meromorphic [Zho01]. mesh [AK01c, BBCH00, BBCH02, CK03, CJL03, CJ00, HR01, LNLL02, NS04b, Ran04, SW02a, Shi04b, Wan04b, ZK04]. meshes [CG04b, FL04, GCL02, JNT02, LC04, PR02, VWE04, WF02]. Meshkov [AVG^{+}04]. meshless [DND04]. Meso [aH02]. Meso-scale [aH02]. metal [NSP04, PCR04]. metal-forming [PCR04]. Method [SWSZ04, eMBH02, ADG04, AD02, AP02a, AB01, AK01a, ASN02, ASB04, AH04, AP00, AG01a, Ant02, Arg01a, Arg03b, AG03b, Aya03, AK01c, AK02, BB01, Ban04, BG02, BO02, BLSS03, BLSS04, BE03, Bet00, BK02, BD04b, BZPFB04, BX02, Bri04, BT01, CCH02, CRSS00, CMRS01, CK04, CFT02, CCM01, dCCSR00, CJ00, DZN00, DSVB04, DKL04, DHK02, DND04, Elo02, ENE04, EMT01, EH00, Fab02, FL04, FS03, FFX04, Gal00, GH01a, GH04, GJL^{+}00, GHK03, GN02, Gug00, GJHY00, Ham00, HY03, HGI04, HR02a, HZZ03, HR04a, HV02, HS00, Her01, HR04b, Hom03, Hom04, HC03b, HWT04, aH02, Ism00, JFW01, qJkSlS04, JW04, JE01, JS04, JM00, Kal00, KG00, Kan04, KCB02, KT02, KHMN02, KR03a, KSSV03, Kum02, KED04, KK03, LCL01, Lee02]. method [Les02, LNBI02, LF01, LHYaC03, LWT04, LZ04, LQQ01, LCF04, LMO01, MMOS01, Mal04b, MRV04, MOS02, MG01, Min04, Moh00, MO03, NKK03, NJVA03, NHMS04, Nis03a, NOI02, OASO03, OOA03, OOAS03, Pal02a, PS04, Pen00b, Pet01, Pet03, PR04, Pre00a, RKS04, Rix04, RZ03, SGG^{+}04, Sai04, SS02a, SS03b, ST00a, SB01, Sha01, SD03, Son01, SS01, Ste00b, SS00, Sty04, Sug02a, SK01, Sun01, Sur01, SU04, SS03d, TJ04, TML03, TWV02, TLQ02, Tse04, UC03a, UC03b, VMD04, VYZ01, VVD04, WX04, Wan00b, WL02b, Wan04b, WKS^{+}03, Waz01, Wol00b, Woź01, Wri02, WX02b, XCZ02, XZ01, Xu01a, Xu02, Xu03a, YTI02, YH03, YF03, YOO02, YEWE03, Ye03, aYtZ03, YB03, ZZ02a, ZW02, ZZ03b, ZZ02b, dFN00]. methods [eMEM00, Abd03, AG03a, AG00a, ABC02, ABC03, AG00b, Arg03a, AW03, BP04, BW02, Bai03b, BJK04a, BBKS00, BM03a, BS00c, BBCH00, BBCH02, BJG02, BF01, BCJW00, BM02, BV00, Bog04b, DFM04, Bor03, BZS02, Bru00, BJ02, BJ04, But00, CR00, CLMR00, Cao02, CG02, Cas04, CELM00, CW00, CS02, CL03b, CH03, lCHcH00, CJ03, CP03, Coc01, CD00, Dah01, DKK03, DM02a, DS02c, Die98, Die02, Dom03, DB01, EES00, EG00, Enr00, Eva04, EW01, FTY02, Fan03a, FI03, FF02, FW00, FT03, Fra02, Fra03, Fra04, FK04, Fre00, Fu04, Gan04, Gar04, GCL02, GGM04, GHR02, GD04, GH01b, GM00c, Gro02, Had00, HRE00, Hop04, HR02b, HSW00, HC04, IZ00, Ixa00, JR02, JPW04, JNS04, Jan03b, JS00a, JkSlS04, JB04, JNT02, Kai03, KP01]. methods [Kál02, KYF04, KZ03, KSCI00, KS03b, Kon04, Kot02, Kot03, Kou03, KL03a, LLP03, LF04, LTT00, LY02, Li03a, LLC01, LD02, LSY04, LLXS04, LMP00, Löt00, Lui00, Lui01, Lui03, LS00a, LLN04, Mar00b, MX00, Mil03a, MM00, MIB03, Nas00, Ng03, Oht03, Pas04, PH01, Pet02b, PW00, PS03b, PT02, QM01, Ran01, Rat00, Rhe00, Röß04, RdAR04, Rui02, Rum03, SKSV03, SA00, SFK^{+}04, Shi02, ST02, Sla00, Ste00a, Ste00b, SM04b, SHS02, Tam03, Tid02, Tid03, TA02, VV01, VDVV00, VID01, VVV03b, VAR03, Wan00b, WH01, Wan01b, WH04, WW00a, XLFC00, XKC01, XZ03, lXsQsJ03, Yam00, Yan00, YK04, ZX01, Zha02, Zhu01, vdHMS00, vdV02, dMPL03, BM00c]. metric [KS00, LS00a]. Mhaskar [Dam03]. Michael [BE02e]. Michaelis [HL04b]. micro [APT02]. microscale [DN01, DN02a]. mild [EH00]. Milne [YEWE03]. Mindlin [AP02a]. Minimal [De 00, AM03, Alf00, EES00, Jea04]. minimax [LL02, MWLS03, MWL04]. minimization [LF01, ST00a, VALM00, WA02]. minimizing [AW03, Oht03, PP00b, QSZ02]. minimum [GH01a, LS02b]. Minkowski [LK01]. Miranda [Arg04a]. MITC [AP02a]. Mitropolski [Ism00]. Mittag [Kir00, MG00]. Mittag-Leffler [Kir00]. Mittag-Leffler-type [MG00]. Mixed [AJ01, CS04b, CD00, MP02a, PRK04, Abd00, Abd02, ASB04, CS04a, CL04, lCHcH00, CM01, Dia03, Din03, Din04, FF02, GH01a, GD04, Ham01, MRT00, NKK03, Noo01, RSS04, Swa00, VS01, Wan04b]. mixed-boundary [Swa00]. Mixed-finite [MP02a]. Mixed-hybrid [CS04b, FF02, MRT00]. mixed-type [NKK03]. mixing [YZCL04]. MMPDE [HR01]. MO [NMST02]. Möbius [Man00b]. mode [CL01b, DSD04]. model [Age02, Ava00, BI03, Cai02, CDV03, Dai02, DSD04, GVV04, GH03b, LY00, LWW01, LG04, MMPZ04, MP01, MP03b, MG03, NBS04, RRWT00, SZ03a, TVV04, VV00, WL04, XL02, lXsQsJ03]. modeling [APT02, BBK04, BV03, Doi02, Fre00, Gor04, LLG04]. Modelisation [DBDPF04]. Modelling [dCCSR00, NGGZ04, SAE04, AVV00, ADEL00, BR00, BTFY02, GGDL04, KY04, MSKS04, MN02a, SS02b]. models [AAD02, BKB04, CFSP04, DHW04, KCMK02, Kon04, Kui01a, LS02b, LCC04, Miu02, SGG^{+}04, SKD04]. modes [DBDPF04]. Modification [HS00, Pet01, EH00, Her01, HR04b]. Modified [Cas00, FS03, MGL01, Noo01, CV03, Fab02, GST03, GKS00, Hom03, Hom04, LF01, Li03a, Rui02, SD03, WFV01, WJ01, ZX01]. modifying [Kon04]. modular [Sed04]. modulated [DBDPF04]. modulus [yCjC01, Dam03]. molds [MMPZ04]. molecular [aH02, NMST02, SRLAD03]. molten [WWA^{+}04]. moment [BGVHN01a, Chi01d, GHM01, Mil01b]. moment-preserving [Mil01b]. Moments [EST03, BA00]. Mono [CGM02]. Mono-implicit [CGM02]. monomials [Nei02]. Monotone [Bog04a, CL03b, Kor01, Pal02b, BP01a, Bog04b, JFW01, Lui03, MO03, San03]. monotonic [Dra04, WA02]. Monotonicity [AQ04, Dem02, GL01, NP03]. Monte [Nie03, Wan00b, Wan01b]. Mordell [Miy03]. Morocco [ERV04]. morphology [LHYaC03]. motion [HLS^{+}03, IY03, Ism00, PGG02, PGG03, UY04]. motional [GGDL04]. motions [Xu04]. Motor [Ani02]. movement [HR01]. Moving [Bai01, eMKM04, dCCSR00, FMRW04, Shi04b, TVV04, TA03, ZK04]. Multi [GGDL04, RdAR04, eMBH02, DRC02, EFS02, FT03, GPS03, Las04, Law02, LGL03, SZ01, Sya04, VV00, Xu03a]. multi-component [LGL03, VV00]. multi-dimensional [eMBH02, Xu03a]. Multi-field [RdAR04]. Multi-harmonic [GGDL04]. multi-index [DRC02]. multi-layer [Las04]. multi-level [SZ01]. multi-point [Law02]. multi-resultant [Sya04]. multi-step [FT03]. multi-term [EFS02]. multi-variate [GPS03]. multiclass [Dai02]. multicomponent [dCCSR00]. multidimensional [BJ04, Coo02, Dmy04, MMK02]. multidimensions [Bai01]. multidisciplinary [HCT^{+}02, RRWT00]. multidomain [BD04b]. Multigrid [Bor03, DMS00, CW00, GCL02, Hop04, KT01, RSvR03, Stü01, VWE04, VMD04, WO01]. multiindex [Kir00]. multiindices [PKA03]. multilayers [DZN00]. Multilevel [GS04b, Ste00b, FGJ00, JNT02, LVH04, Sha01]. multimaps [LY01]. Multiple [CL01b, Eva04, FKM02, Kir00, LS04b, Pet02a, ACV03, BW04b, Cai02, CS00b, CV01, CV03, FS03, Gen03, HW03, Hor01, KLY04, LG04, Mil01b, NOI02, Par01, Rum03, TNW02, VC01]. Multiplication [GKK00]. multiplicative [Bai03a]. Multiplicity [KWZ04, DH00, EPM00, FPS00, JKN00]. multiply [OOAS03, Weg01a, Weg01b]. multipoint [MSM04, WA00a]. multipole [AP00, SGG^{+}04]. multirate [BG02]. Multiresolution [DLM^{+}00]. Multiscale [CCL03]. multisplitting [BW02, Bai03b].

Multistage [LW01]. multistep [IRVD01, IVD02, Sch03d, TYZL04]. multivariable [SSV04]. multivariables [PKA03]. Multivariate [GH00, Lor00, Wan00a, WX02a, Zho01, GS00, jGwHsZ04, Hom04, HR00, LM02a, LW04, LMMD03, LS00b, RR01, TT02, Wal00, WLYL04, dSPPT01]. Müntz [MC05, MDSR04]. musical [HT01]. Muskhelishvili [DS02c].

N [Sto05]. Nagumo [SZ03a]. Nash

[FGJ00, Par00a]. Native [Sch00a]. natural [And00, DNS00, GL04b, VMD04]. Navier [Abd00, BJK03, CM04, Joh02, SEKW01, Swa04]. near [Ari03, BLSS03, BLSS04, LS02a, WS02b]. near-singular [LS02a]. nearly [Bat04, BvdV01, TNW02]. Necessary [CL02, DJK01a, BC03, DH02]. needles [Rhe02]. negative [AdMR02]. nets [LPSSP00]. network [CZW04, GH03a, HCT^{+}02, Ji04]. networks [BG02, BO04b, Dai02, Hig03, JT02a, LS04a, Leu00, Li02, LWT04, LQQ01, LWY02, MN02a, Sla00, SAE04, VBL^{+}04, WZ04a, WX02b]. Neumann [Deh02, Mar04c, Par00a, RSS04]. Neural [CF04, VBL^{+}04, BO04b, CZW04, Li02, LWT04, LQQ01, LWY02, MN02a, Sla00, SAE04, WZ04a, WX02b]. neuron [GH03a, GH03b]. neurons [BJK04b]. neutral [BGZ00, CL02, HHC01, LG04, LB01, Par01, Par03c, PC00, QM01, WL03]. Nevai [Chi03]. Neville [TJ04]. Neville-like [TJ04]. Newsletter [Ano00w]. Newton [ABC02, Arg01b, Arg01a, Arg03a, Arg03b, Arg04a, Arg04b, CK04, FT03, FS03, Gal00, Gug00, Her01, HR04b, Hom03, Hom04, Kal00, KG00, KS03a, LNBI02, LMP00, Mar00b, Mil03a, Nas00, NR01, Pal02a, PT02, TWV02, XKC01, XZ01, Yam00, ZX01, ZD02]. Newton-like [Arg03a, Yam00]. Newton-type [LMP00]. Newtonian [PGG03, PGG02]. next [YOT^{+}02]. NIG [AP04a]. Nikiforov [ÁNM01]. nilpotent [SZZ01]. nine [BJG02, KP01]. nine-point [BJG02]. ninth [Sto01]. nodal [Dul04]. nodes [Dik03, MO01, Mil01b, NS04a]. noise [DKST04, FD00, GK03, Röß04]. non [AK01c, Chr01, Coc01, DKST04, HY03, HGI04, MG03, OS04, RDA04, RCZV04, Str00, Swa02, Yoo03]. non-autonomous [DKST04]. non-censored [RCZV04]. non-linear [Coc01]. non-overlap [HY03]. non-permanent [MG03]. non-smooth [HGI04, OS04]. non-stationary [Swa02, Yoo03]. non-uniform [AK01c, Str00]. non-uniqueness [Chr01]. nonautonomous [BB03b, CZY03, LCC04, San00, WL04]. nonclassical [CS01]. noncommutative [GI03]. noncompact [CT00]. Nonconforming [Shi02]. nonconvex [CMV01, El 04, LF01, YH03]. nondefinite [AM02b]. Nondifferentiable [MWLS03, HR04b, LL02]. nonequilibrium [Rem04]. Nonexistence [YL03, EPM00]. nonhomogeneous [HCT^{+}02, LL00]. nonidentically [BA00]. Nonlinear [AD01, Ben02, CK03, SB01, WBBF00, Aas02, ASB04, AG00a, ABG03, AG01b, ADEL00, BRS03a, BM00a, BGZ00, BB03b, BB03a, Bha02, BJK04b, BT00a, Bog04a, Bog04b, BW04a, BJ04, CN00, CFSP04, CGM02, CMSV03, CHW02, CH03, Cla03, CMKM03, DP02b, Din03, Din04, ENE04, Fre03, Fur01, GS01a, Gao01, GVSJ01, GJ01, HJ02, HP00, HG03, HRE00, JKY02, JT02b, JB04, JM00, Jun04, KYF04, Kaz02, KSCI00, KCMK02, KMCK02, KWW01, LZ02a, LS04b, LQQ01, Lin00, LL01b, LMP00, Lui03, LS00a, LLN04, Mal04b, Mar00a, Mar00b, ML03, Min01, Moh00, PJ03, Rac00, SWSZ04, Slo04, SAE04, SC01, Tsu01, XLFC00, XKC01, YF03, ZW02, Zhu03a, dMPL03, QZ03]. nonlinearities [AO00c, GT04, HW03]. nonlinearity [KS01a]. nonlocal [Pao01]. nonmatching [FMRW04]. nonmonotone [CZ04, SHS02, Zhu01]. Nonmonotonic [Zhu00, Zhu03b]. nonnegative [Fas02, KS01a]. nonoscillating [MMK02]. nonoscillation [Ues04]. nonoverlapping [LMO01, Lui01]. nonperiodic [Wri02]. Nonreflecting [ADR02, GN03]. nonselfadjoint [BB02]. nonseparable [LL03b]. nonsingular [DKK03]. nonsmooth [CW00, CMKM03, PT02, SQ01]. nonstandard [DV00, KW03]. nonstationary [BW02, Bai03b]. nonsymmetric [AVV00, GS04b, Woź01, Zha02]. nonterminating [Sch03c]. Nonuniform [GL00, CJL03, Gor04]. nonuniformly [Dik03, dBS01]. nonuniquely [LD02]. nonvanishing [KPS01, KS02a]. norm [BD04a, DFC02, Gao01, LS00b, PP02c, ZLF01]. normal [AG03a, AFGG03, HSS01, Sad01, YD03, Yu02]. normality [KV01b, PP02b]. normalized [MVDB04, SU04]. normed [Tak03, Wat00b]. Norms [vG00, BS02b, Koh01, Koh02, Koh03b, QSZ02]. Note [Cho03, vdH00, Bav00b, Cha04, DJK01b, DRC02, DS02a, Ehr01a, Fan03b, HSV04, JR02, LZ04, MW04b, Sad01, San00, WZW03, YD03]. notebook [BYY03]. notion [Sie02]. Novel [Lee02, Age02]. nuclear [SZ03b]. Number [Zub04, CM04, Fan03b, HSV04, LS00b, NS04a, PP02c, Tan02, WZ03]. numbers [CK02a, Dev02, Dia03, ELW02, Lau00, Mel02, Sen01]. Numerical [AK01a, AB04a, Ano00x, Ano01s, AH03, AVG^{+}04, Bac04, BQ00, BBKS00, BTSHK04, BGZ00, BR00, DFM04, Bre00b, BH03, BBM00, But00, Cas04, CMV01, DE00, DMN01, EK01, EG00, ELR00, ENE04, Fer03, Gao01, Gov00, GM00c, Hau02, JB04, JS00b, KLyD^{+}03, KOZ03, KED04, LK01, LD02, LR04b, Liu02a, LAT04, MS03a, Mas03, MO02, MT03, MX00, Min01, Min04, Nac04, NY03, Nis03a, OOAS03, Pao01, QST00, Rhe00, RS04, RN03, Str00, SS03d, Tem00, WBBF00, WH00, WF00, WW00c, lXsQsJ03, Xu03b, YLC04, AG03a, AM03, AK01b, AJ01, AVMRVM02, Bak00a, BO04a, Bar03, BMS03, BLSS03, BBMS04, BS00d, BZPFB04, BTFY02, BM04b, Buc00, CLMR00, CDV03, Cas00, CGM02, Cha02, CFHS03, CFSS03, CD01, CDN01, DV00, DKL04, EFS02, Enr00, EW01, Far02]. numerical [Faz02, FW00, FLM03, Fra03, FFX04, GKMN00, GS01a, GCL02, Gau01, GGM04, GHK03, GR01, HSS01, HKD04, HOO03, HR03, Iqb03, JS00a, Jeo00, Joh02, KMS03, Kál02, KO00, KS03b, LS02a, LRS02, LJ03, LSY04, Mar00a, MO00a, MOTT03, MOS02, MS01b, MM03, NJVA03, NSP04, OOA03, Oou03, Pen00b, SS03b, Sch02a, SFK^{+}04, SSV04, SB04, Ste00a, SdSP01, SM04b, Sug02b, Tas00a, Tas04, Tho01, TA02, TNW02, TQ03, VS01, VAR03, WZW03, Waz01, WWA^{+}04, ZX01, BMPV00]. Numerov [And00, Tse04]. NURBS [LW04]. Nystrom [GJ01, BM02, Fra04, VV01].

O.D.E. [San00]. object [PCR04]. object-oriented [PCR04]. objective [XL02]. objects [PS03a]. oblique [dSPT02]. observations [Ava00, BS00a, EP03]. obstacle [AZ04, ASN02, Häh00, MC03, RN03]. obstacles [OASO03]. ocean [SM04a]. odd [CG03a, CK02b, KT01]. odd-even [KT01]. odd-integer [CK02b]. ODE [JkSlS04, qJkSlS04]. ODES [Bru00, BM04b, Cas00, Enr00, Gov00, IRVD01, IVD02, RZ02, SC00]. offset [AsKS04]. Oldroyd [NSZ04]. One [Sch03d, AA02a, AM02b, CRSL00, DN01, DSZ02, DFC02, GW04, Gug00, JS00b, Moh00, NS03, Qui02, Tom02, Wat02, XL02, Xu02, ZD02]. One- [Sch03d]. one-dimensional [AA02a, DN01, GW04, Gug00, JS00b, Moh00, NS03, Qui02, Xu02, ZD02]. one-loop [CRSL00]. one-sided [DFC02]. ones [KV01b]. online [LWT04, WX02b]. only [Mar00a]. onto [vG00]. Open [Mul01, Mul03, Dom03]. operational [Dat00]. operator [AO00b, BE02a, BH04, CN00, Ham00, Kan00, KLT04, LR04b, MTA^{+}03, Mou03, Slo04]. operators [AP04b, AB04b, Arg01b, Bav00a, Bav01, Bav03, Beh02, BGP02, Che03a, DN02b, DLZ02, EP03, GW04, GKS00, IP01, LMMD03, Löt00, MO03, RLZ03, Sch02a, SJ03b, TNW02, YD03]. optical [JT02a]. optics [BLSS03, BLSS04, DRC02, PP00c]. optics-based [BLSS03]. Optimal [FD00, Göt01, Jun04, KW00, MO01, MM02a, SAA01, Sar00, SBSA03, SVV01, Won02b, Bat04, Bor03, BM00c, CL00, FP02, GJL^{+}00, Gug00, Leu00, LS02c, Mak04, McC00, MP02b, xZqZqT04]. Optimal-by-accuracy [MM02a]. optimal-by-order [MM02a]. optimality [Leu00]. optimisation [FT03]. Optimization [BCM03, WBBF00, AG00a, Bet00, BCJW00, BT00a, CB00, CZ04, Dal01, DSW^{+}03, DMS00, Ega00, Hof00, LS00a, O’L00, PRT00, RRWT00, Sch03d, Sch04c, SXZ00, SHS02, Tru04, WPS02, Wol00a, XL02, Xu02, Xu03a, ZZ02a, ZD02, Zhu00, Zhu01, Zhu03a, Zhu03b]. Optimized [MMK02, TS02, BM04a, SDMV04, Tas00b]. optimum [BBKS00]. option [AP04a]. options [WFV01]. Order [HW00, AZPF04, AM02b, ASN02, AL03, ABC03, AB03, AA02c, AK01c, AK02, BA00, BI03, BBCH00, BBCH02, BH04, Büh00, CS04a, CP04, CCL03, CGM02, Cha02, Chi01a, Daa04, DV00, Deh02, DM02a, Dom03, EM04b, EP02a, EP02b, EP03, FR01, FKR04, FKM02, Fre00, FS03, GZ02, GN03, GHR02, GM00c, GLZ03, GSG03, GT04, GK02, HP00, HY02, HG03, IRVD01, JW00, JFW01, Kha02, KWW01, KL01a, KL03c, Len02, Let01, LZ02a, LYC02, LC04, LL01b, LW02, LB01, Lup01, MMOS01, MFMGZ02, Mat03, MM02a, Moh00, MSM04, MIB03, SA00, SW00b, Slo04, SS00, Sty04, SK01, TYZL04, Tor03, Tse04, TQ03, VWE04, VV01, VAR03, WL01, WL03, WZW03, WW03, Waz01, WA00b, Wri04, XCZ02, Yan00, ZW02, dRM03]. order-preserving [Lup01]. orders [Che03a, GST03]. Ordinary [BMPV00, CS03a, BP01b, But00, Daa04, EKLW01, HG03, JPW04, SA00, TYZL04]. Organizing [Ano02y]. oriented [PCR04]. Orlicz [BD04a]. Orthogonal [Bar03, BF01, BGVHN03, BVV04, EKLW01, KL03c, Lor03b, PS01, VB03, Xu01b, Abd02, AMBP^{+}01, Ara04, AGMMB00b, ACV03, AIV01, AW03, Bar02, Bel01, BC02, BDR02, BdSR03, BGVHN01b, CLMR00, CFMV03, CGMB03, CZ03, Chi01c, Chi01b, CMS01, CV01, CV03, DMFSR01, DKM^{+}01, DS02a, Dim01b, Elb01, FR01, FR03, FKR04, Fre01, HHR00, Ifa01, KS01b, KLM01, KMV03, LCI01, Lew03a, Lew03b, LLC01, Lor03a, MF01, MdB02, MS02, MN02b, NR01, Peh01, Peh03, RR01, RLS01, RZAG00, SRD00, Sim04, Suá01, TL01, VC01, VB04, VZ04]. Orthogonality [AdMR02, KLM01, MDSR04, MC05, BCM01, BZS02, Car03, ELW04, Mil01b, Sia01a, Suá03, Van01]. orthonormal [Han03a, KL01b, LM02b]. oscillating [PS03b, TC01, TS02]. Oscillation [BLM04, CL03c, EP02a, JKY02, JT02b, KS02b, LL01b, PC00, Ues04, WW03, BB00, BB03b, BB03a, CL02, DH02, TC01, WL03]. Oscillations [HG03, MY03b, CMV01, FPS00]. oscillator [RLZ03]. oscillators [Fra03]. Oscillatory [LB01, vdHMS00, CCL03, CC03a, Ehr01a, Eva04, HSV04, KCI02, Nay02]. osculatory [WG04]. Oseen [LMO01, Nis03a]. other [WH04]. outer [DDG04]. outlier [KATS02]. output [FKM02]. overconvergence [dB03]. overlap [HY03]. Overlapping [AS04b, De 01, HC03b, MMOS01]. overrelaxation [Had00]. overview [De 00, Ips00]. oxidation [CS04b].

Padé [AB03, Dar01, DMGVO01, GPTT02, GK03, Gu04, GH00, HR00, Pré00b, RLS01, SV00, WdZ04, Zho01]. Padé-type [Dar01, Gu04]. page [BYY03]. Painlevé [Cla03, TRTG03]. pair [Fra02, Fra03]. pairs [AMBP^{+}01, MdB02, Ter03, TS02]. pairwise [Sug02b]. Pál [Dik03, dB02]. Pál-type [Dik03, dB02]. Paley [AB04b]. pantograph [MIB03]. parabolic [Ant02, Bog01, Bog04a, Bor03, BJ02, BJ04, CY00, CK04, CJL03, CL02, DZN00, EM04a, Fu04, GR01, JKY02, Kas00, KS02b, Lin00, Lui03, Mig01, MY03b, NY03, OS04, PC00, Rui02, Shi04a, Slo04, Tem00, VS01, WKM04, Yan00]. parabolic-gradient [VS01]. parachute [LS02c, Mak04]. Parallel [ADEL00, GM00b, ST02, WPS02, Bai03b, BvdV01, DB01, NMST02, PNV01, RDA04, SAA01, YB03]. parallelism [KP04, dSPT00]. Parameter [Kno01, Mal04a, NJVA03, Age02, CS02, Hei03a, MMOS01, MSI04, Tem03]. parameter-uniform [MMOS01]. Parameterization [ZH04]. parameterized [VV01]. parameters [AdMR02, BS00b, BCM03, BMN01, BW04a, CZ04, CL01a, Fer03, Hat03, KLY04, KY04, MFMGO01, RZAG00, Yu02]. parametric [Man00a, Par03c, WM01]. parametrized [Jun04, Tsu01]. parasite [MP01, MP03b]. Part [BLSS03, BLSS04, BMN01, LYF03, LHHW04, Sza02]. Partial [AM02a, Ano01s, BF01, BS02b, BP01b, CL03c, EW01, GZ02, KW00, LHHW04, LL01b, QH04, Ran01, Sch04a, SAE04, Tho01, dFN00, vdHS01]. participants [Ano01q, Ano02p]. particle [LLL03, Wol00b, YLC04]. particles [FC04, VV00]. Particular [LHHW04, AKM03]. Partition [CL03a, Mut02]. partitioned [BM02]. partitions [BYY03, PR02]. past [CM04]. path [Bou04, GDD04]. path-finding [Bou04]. paths [Cai02, Zhu01]. PATRICIA [Dev02]. PCG [ZD02]. PCG-like [ZD02]. PDE [GVSJ01, SS03c]. PDEs [AG03a, Bai01, Dah01, FFX04, Lui03, MRV04, SA01, WKM04]. peeling [NS03]. Penalized [MO00b]. penalty [CZ04, YTI02]. pendulum [SW03, TQ03, WS02a]. penetrable [Häh00, PS03a]. Penetrating [LLXS04]. penetration [Joh02]. perform [FW00]. Performance [AG00b, CHW02, Duf00, SIM02, SSV04, Woź01, ZX01]. Periodic [GH03a, HY02, HL04b, Ism00, WL04, AG01b, BL01c, CZY03, FW03, Jan03b, JFW01, JKN00, KWW01, LG04, MGL01, NGGZ04, NRL03, OASO03, Rac00, Sch02a]. periodically [FPS00, JNS04]. periods [Bog03]. permanence [WL04]. permanent [MG03]. permutation [DP02a]. personal [Eve04]. perspective [Bak00a, GPS03, Rhe00]. Perturbation [TL04, CFT02, Hei03b, HRK04, KSSV03, WW01, Yu02]. perturbations [Jea04, LW03a, MR01, PJ03]. Perturbed [DL00, KCB02, AG03b, AK02, BM01a, Bog01, Bog04a, Bog04b, BD04b, CL04, CGG00, Che03b, FHM^{+}04, FKR04, Fra03, FPS00, HR02b, JNS04, JY01, Len02, LHHW04, MOS02, MSKS04, Min04, NJVA03, OS04, QST00, Sch02a, Shi04a, Shi04b, SS03c, Yeh04, ZL04]. Phase [Rem00, AA02a, KL02, WWA^{+}04]. phase-change [WWA^{+}04]. phenomena [Daa04, MT03, MOTT03, NS03, SJ03a]. phenomenon [GW04, JS04, Pas04, SR04]. photo [Ano02q]. photon [Iqb00]. physical [EGV03]. physics [Lor03b]. physiology [PS04]. Picard [TWV02]. Pick [Ped03]. Piecewise [AG00a, WL02a, AMS00, Dem04, FQR00, Kob00, Pet02c, WZ04b, Yan02]. Pincherle [MP03a]. pioneer [MP03a]. pivot [AG02b]. pivoting [PN00]. place [BQ00]. Planar [MW02, Rem00, WM02, WMA03, Mee02, TA03, WM01, YLD02]. Plancherel [AGMMB00b]. Planck [LAT04]. plane [ADG03, ADG04, Dam03, FLW01, KyGUY02, LCI01, MG01, NS04b]. plate [AP02a, CCM01, MSKS04]. plates [Las04]. platform [Doi02]. plus [Bac04, VMV04]. Poincaré [GS04a]. Point [PH01, Rem00, AM02b, And02, AK01c, BM03a, BB02, BM04a, BJG02, CGM02, CMSV03, CH03, CD03b, Cui02, DvM01, DL00, DHK02, FTY02, GPTT02, GSG03, GJHY00, Hin02, KZ03, KS00, Kum02, Law02, LLG04, Min04, MM00, NJVA03, Nie03, PW00, QST00, RZ03, Sim00, STW00, Sug02a, Tom02, WS02b, Xu03b, ZW02, Zhu03a, Zhu03b]. points [BEM00a, Büh00, CN00, CD01, GL04b, HC04, IP01, LV00, Seg03, SdSP01, Tas00a, Tas04, Tsu01, VVV03b, hWqX03, Wol00b]. Pointwise [Ava00, Che03b, Hug01, KL01a, SAA01]. Poisson [HV02, LYF03, RSS04, Wol00b]. pole [WS02a]. pole-zero [WS02a]. poles [BM04c]. Pollaczek [Ara04]. pollution [BV03]. poly [DSS03]. poly-basic [DSS03]. Polya [Goh02]. polydisperse [BBK04]. polyethylene [WKS^{+}03]. polygonal [YH03]. polyhedral [DKK03]. Polynomial [ACGR01, Bar02, BRZS00, CG03a, CG00, CZ03, CL01a, DKK03, EKLW01, FHMS04, GS00, Gem00, GK03, JS04, KLM01, PS01, Peh03, PH01, Pet02b, Pet03, Pie00, Pom01, RZ04, SS02a, Swa04, Sya04, VASF03, WM01, Wal00]. polynomials [Abd02, AMBP^{+}01, AdMR02, ÁNM01, ÁNAA03, ACGR01, Ano02a, Ara04, AGMMB00a, AGMMB00b, ACV03, AIV01, Bar03, Bav00a, Bav00b, Bav01, Bav03, BGM00, BGM01, Bel01, BC02, BM01b, BDR02, BdSR03, CFMV03, CGMB03, CGG00, Chi01c, Chi01d, Chi01b, Col01, CMS01, CV01, CV03, Dam03, Dat00, DRC02, DMFSR01, DKM^{+}01, DFC02, DS02a, Dim01b, Dim03, Doh02, DL01b, DD01, DM02b, ELS01, Elb01, ELW02, ELW04, Fas02, FPO01, FR01, FR03, FKR04, Fre01, GJS03, Gat02, Gil01, GL03, Grü01, HR02a, Her03, HHR00, Ifa01, Ism03b, KP03a, KPS01, KNS03, KK00, KS01b, Kra03, KL04b, KV01b, KMV03, LCI01, Lar02, Las02, Let01, Lev00, Lew03a, Lew03b, Li02, Li03b, Lor03a, Lor03b, Lor00, MDSR04, MF01, MFMGO01, MÁNM01, MdB02]. polynomials [MS01a, MS02, MC05, MN01, MN02b, NR01, NP03, Neu03, OP03, PFO03, PS01, Peh01, Peh03, Pet01, Psa03, Rad00, RR01, RLS01, RZAG00, Rum03, SRD00, SRD01, SRLAD03, Sch01, Sim04, SC04, Suá01, Sza02, Sza01, Szy03, TL01, VC01, VYZ01, VZ01, VZ04, Vou01, Wün01a, Wün03a, Xu01b, Zha03, dG01, vD03a]. polytope [hWqX03]. population [LG04, LCC04, SS01]. porous [AA02a, HRK04, JT03, Qui02, sY01]. portfolio [XL02]. posed [CMRS00, Iqb03]. Positive [BM00a, GSG03, KWW01, AG01b, AG02a, AA02c, BW02, Bar03, CZY03, CL01a, DH00, DS02b, EPM00, ELW04, GHMY00, GM00a, GS01b, GT04, He03, Her03, LZ02a, LS04b, LG04, Pal02b, Rac00, Sta03, WF02, YL03]. positivity [Cof04]. possible [Hei03a]. post [CCL03]. post-processing [CCL03]. posteriori [ARV00, Hug01, JL00, LY00, NSZ04, Ran04, RSS04]. postprocessor [HS02]. potentials [BGP02, BEM00b, BE02b, BPW02, BE02c, DN02b, FMRW04, SRLAD03, VASF03, YEWE03]. poverty [Cai02]. Powell [LPSSP00, MVDB04, WVDB03]. Power [KR02, BE02b, CHW02, DL01a, HC04, Mil01b]. powerful [Bri04]. powers [BS00d, CG03a, hWqX03]. Practical [BM02, Mar00b, CHW02]. pre [HS02, ZH04]. pre-assigned [ZH04]. pre/postprocessor [HS02]. precise [HS02, TYI03, WX04]. precise/efficient [HS02]. precision [Lan04]. Preconditioned [BJG02, qJkSlS04, LRvSS04, EMT01, JkSlS04, KRW00]. preconditioner [CG02, GH04, GS04b, KHMN02, LVH04]. preconditioners [CN02, GSS03, HC03a, NHMS04]. Preconditioning [Mor00, DKL04, Glo03, SZ01, SEKW01]. predator [CZY03, HL04b, WL04]. Prediction [Wen00, AZPF04, PDVS04]. predictor [PS03b, UH00]. predictors [Sya03]. Preface [ABO02a, ACY03, BMPV00, BE02d, BE04c, BR01, DHV04, GMR01, Hit02a, HH03b, LR03, LB00, MM02b, NSK03, QZ01, QYZ02, Sia01b, SVA03, WW02, WT04, WM04b, QZ03]. preinvex [YYT02]. prescribed [BM04c]. presentation [PKA03]. presented [Ano01r]. preservation [TW04]. Preserving [IZ00, CM03, DM02a, Lai00, Lev00, Lup01, Mil01b, UY04]. pressure [Ava00, Yan02]. pressures [RdAR04]. prewavelets [FQR00, HH03a]. prey [CZY03, HL04b, WL04]. pricing [AP04a, WFV01]. primal [QSZ02]. primal-dual [QSZ02]. principal [CC03a]. principle [HH04, Pöt02, Ric00, TYI03]. principles [KLT04]. Pringsheim [BL01b, xZqZqT04]. priori [AJ01, FPO01, Hei03a]. priors [TO03]. probabilistic [Lou02]. probabilities [Sch03b]. Probability [Wat02, IK03, Ing03, Mlo01]. Probability-one [Wat02]. problem [eMKM02, AZ04, AP02a, AR04, AS04b, AVMRVM02, And02, AH03, AA02c, BQ00, BM01a, BL01c, Bha02, BK02, Bog01, Bog04a, Bog04b, BD04b, Bou04, BMZ00, BGVHN01a, Cap04, Cha02, CK04, CFHS03, CFSS03, Che03b, CL01b, CMV01, CP00a, CM01, CD01, Dal01, DH00, De 01, Deh02, DHZ04, El 04, Fre03, Gao01, GR01, GI03, HL04a, HSS01, HC03a, HY02, HLS^{+}03, HR04a, HRS02, HCT^{+}02, HK01, JW04, JS03, Kno02, KWW01, KL01a, KS02a, LLP03, LK01, Law02, LCL01, Li03b, LR04b, Liu02a, LS02c, Mak04, MRVM00, MSKS04, NSZ04, NS03, QST00, Qui02, Rac00, Ran04, RZ03, RS04, ST00a, SC00, SV00, Str00, Sun01, UC03b, WL02b, WY02, YTI02, Yan02, YY02, Yeh04, ZL03, ZD02, dB03]. Problems [Kui01b, Abd03, ADG04, AO00c, ABO02b, AZPF04, AM02b, ASN02, AJ01, AP00, AGRZ01, AG02a, AG03b, AK01c, AK02, Bac04, BM00a, BE03, BM00b, BBW02, BLM04, BBW04, BCM03, BV00, BS00d, Bor03, BJ02, BJ04, BM00c, CH02a, CL04, CR00, CMRS00, CCL03, Cas00, CGM02, CMSV03, CELM00, Chi01a, Chi01d, CH02b, CJL03, CP03, CD03b, CL02, Cui02, Dav02, DV00, DHK02, Duf00, DB01, ENE04, EP02b, EP03, EGV03, EHS00, FD00, FLW01, FTY02, Fan03a, FHM^{+}04, FL04, GHM01, GGM01, GH01b, GM00c, GN02, Gru03, GSG03, GT04, GJHY00, GGDL04, HS01, Han03b, HKD04, Hei03b, HC03b, Iqb03, Jan03a, JW00, JY01, JFW01, JB04, KTIS03, KS01a, Kno01, Koe03, KyGUY02, KSSV03, Kum02, LVH04, LS02a, Les02, Leu00, LYF03, LMZ02, LQQ01, LY01, LY00, Lui00]. problems [MMOS01, Mal04a, MZ02, Mar04c, Mas03, MY00, MC03, MM00, Moh00, Mul01, Mul03, Nac04, NKK03, NJVA03, NRL03, OS04, OASO03, Oht03, Pal02b, PS04, PP00c, PDVS04, Pre00a, Rat00, Rem04, RZ02, Rui02, RN03, SS03b, SA00, SRD01, Sch04c, SC00, SD03, SW00a, Shi04b, Sim00, ST02, Son02, Sug02a, SQ01, SHS02, Sur01, Swa00, Tas00a, Tas04, Tid02, Tid03, TS02, UC03a, VV01, WPS02, Wat00a, Waz01, WH00, Weg01a, WA00a, Won02a, XLFC00, XKC01, XL02, ZLF01, ZZ02a, ZW02, ZL04, dB02]. procedure [GGM04, PRK04, dACCS03]. procedures [CSRB04, Sch00b]. Proceedings [ERV04, GVW04, BGPW02]. process [CB00, dCCSR00, Grü01, KT02, MRVM00, PS02, Pet01, Pet02a, Sid00, Wen00].

Processes [Rem00, AP04a, MG00, Mar03, MD04, vD03a]. processing [CCL03, Liu02b, MS03b, Sch02b]. product [CG03b, Gu04, HV02, LS04a, LV03, Mil01a, Tak03, Tél00, Van00b, Zeg04, Zha02]. product-type [Zha02]. products [AGMMB00a, BS00a, CL03a, RR01, Yak02]. Prof. [BE02e]. Profile [Sey01, CDV03]. profile-velocity-temperature [CDV03]. program [PCR04]. Programming [QZ03, BT00a, Dra02, FLW01, KyGUY02, LL02, LW01, MWLS03, MWL04, SD00, YTI02]. programs [LZ02b]. progress [FLR03]. Progressive [Isk03]. project [Mak02]. projected [Zhu00]. projection [ADL^{+}02, GVV04, Kai01, SS03b, XZ03]. projection-based [GVV04]. projection-type [XZ03]. projectors [vG00]. proof [BMZ00, KOH03a, Min01, Nag04]. propagation [JR02]. propagators [CVB04]. proper [CP03]. Properties [BJK04b, ZX01, Aga03, Bak03, BE02a, BF02, BGM00, BGM01, BCM01, BC02, BKB04, CG03b, CV01, GL01, JKY02, KYF04, KS01a, LW03b, MTA^{+}03, Öze04, Ram03, WW00a, Wen00]. property [Bat04, BO04b, Fur01, ONU03, Tay00, Ues04, WW04]. proportions [TO03]. protection [DFT04]. proximal [DL00, Han03b, MO03]. pseudodifference [Löt00]. pseudodifferential [Ant02]. Pseudorandom [DP02a, CK02a]. pseudospectral [JR02, dFN00]. pulse [MT03]. pump [KL04a]. pumps [VVD04]. purification [MRVM00]. purpose [YSNM02]. purposes [NOI02]. PVM [PNV01]. pyrolysis [TVV04]. Pythagorean [BV02, WM02, WM04a].

QCD [IN03]. QMC [Kai03].

quadrangulation [LH03]. Quadratic [AS04c, DM02b, BRS00, BT00a, CF00, DLM^{+}00, DL01a, DM02a, FLW01, KED04, LW01, LZ02b, LMP00, LPSSP00, XL02]. Quadrature [BGVHN01b, KCI02, KL03b, MS02, Rat00, BDGV01, CC03a, DGVM02, Die98, Die02, Dom03, Ehr01b, Ehr02, Eva04, Gau01, Gau02, Göt01, Hag01, Ham01, HZZ03, KC00b, KC00a, KY02, KSSV03, Lau01, LW02, Not01, Not03, Oou03, SKSV03, Smi03, VB03]. Quadratures [Mil01b, DMGVO01, GG01, MS04, Sch04b]. quadrilateral [LNLL02]. Qualitative [SB04]. Qualitatively [ALM03]. Qualocation [Slo00]. quantiles [Age02]. quantizers [FP02]. Quantum [DMFSR01, Ell01, Vou01, DRC02, Fre02, SC04]. Quartic [ASN02, LW04]. Quasi [AFGG03, Mar03, WW04, ASB04, ABC02, CT00, DL00, DX02, FT03, GGM04, Mar00b, Nie03, Sim00, Wan00b, Wan01b, XKC01, ZX01]. Quasi-birth-and-death [Mar03]. quasi-extrapolation [GGM04]. Quasi-homogeneous [AFGG03]. Quasi-interpolations [WW04]. quasi-linear [Sim00]. quasi-Monte [Nie03, Wan00b, Wan01b]. quasi-Newton [ABC02, FT03, Mar00b, XKC01, ZX01]. quasi-variational [ASB04, CT00, DX02]. quasi-variational-like [DL00]. Quasilinear [Lin00, GHMY00, GT04, HP00, Koe03, KL01a, Mar04a, YL03, ZL04]. Quasilinearization [HK01, Elo02]. Quasioptimality [GD04]. quasistatic [CFSS03, HSS01]. quasivariational [DP02b]. questions [DKM^{+}01]. queueing [Dai02]. Quintic [SA00, WM02, WM04a].

Rabat [ERV04]. Radar [LLXS04]. Radau [MS04, QM01]. Radial [LM02a, AHP04, Kan04, Sch00a, Tor03, Yoo03]. radially [BR02, Min04]. radiation [ADL^{+}02, NKK03, PDVS04, SFK^{+}04, SKD04]. radiative [Arg01b, FL04]. radii [GL02]. radius [Cha04]. raising [Rab03]. Rakhmanov [Dam03]. Ramanujan [BS03, BYY03, BM01b, Kar01]. Random [CD03a, Sch03b, Age02, Bar03, Dev02, DWZ02, GS02, KL02, Kui01a, LMYL01, LS02b, Lou02, MT02, Mut02, Rhe02, Sug02b, Tan02, WT01]. Range [CO01, Ega00]. rank [Büh00, DWQ04]. ranking [NOI02]. Rao [RO03, VVV03a]. rapidly [ST00b]. Rate [CW04, IY03, MA04, NHMS04]. Ratio [AGMMB00b, WL04]. ratio-dependent [WL04]. Rational [BD00b, GSS03, Mac02c, RV00, WLYL04, BB01, Bar04, BM04a, BM04c, BGVHN01a, BGVHN01b, BGVHN03, BVV04, DDN01, DDPT00, Gau01, GS01b, MDSR04, MC05, SL03, TT02, VB03, VB04, VB01, hWqX03, WG04, ZZ02b, dG01]. rationally [BBW02]. Ravenhall [BE02a]. Rayleigh [AVG^{+}04]. Razumikhin [SS01]. reacting [VWE04]. reaction [AK01a, ALM03, BM01a, BE03, Bog04b, CFSP04, CL01b, GLZ03, HGI04, JB04, MMOS01, MOS02, Pao01, PS02, SB01, YL03, Yeh04, ZK04, Zou02, vdH00]. Real [Sza02, DFM04, BM00c, CFMV03, DVM02, Dia03, Fab02, GS01b, KNS03, KS02a, Kra04, MO00a, Mou03, Par04a, RZ04, Sya04]. real-time [BM00c]. realization [De 00, HKD04]. Realizations [LV03]. reciprocal [Suá03]. reciprocity [MY03a]. Recognition [TA03]. Reconstruction [yCjC01, QH04, HLY04, JS04, Nac04, VWE04]. reconstructions [Pot00]. records [NN04]. Recovery [BBW04, KLY04, KY04]. recruitment [McC01]. recruitment/renewal [McC01]. rectangular [CCM01, LHHW04, Wan04b, WW01]. Recurrence [LNLL02, Pom01, Ron01, WLYL04, xZqZX02]. recurrences [BMS03, Lew03a, WT01]. recurrent [Man00b, MK01]. Recursion [Lew00]. recursions [CR01, Nei02, PP02a]. recursive [FR01, FR03, GS02, Let01, MT02, NR01, RV00]. redistribution [CK03]. reduced [CF04, DSW^{+}03, Fre00, KT01, UC03a, UC03b]. reduced-order [Fre00].

Reducibility [AP02b]. Reducing [dRM03]. Reduction [CF00, GVV04, MRT00, MSM04, Rab03, Sed04, Sza02, Wan01b, ZKO02]. refined [ADL^{+}02, Glo03, GL04b, LPSSP00]. refinement [BBCH00, BBCH02, Shi04b, Zub04]. Refining [BT00b]. reflection [Dun03]. reformation [MOTT03]. reformulation [LQQ01]. reformulations [SQ01]. region [CZ04, DSW^{+}03, RRWT00, Zhu00, Zhu03b]. regions [BE02b, BH03, CJ03, FI03, Glo03, SB01, WLYL04, Weg01a, Weg01b]. registered [Ano01q]. regression [RCZV04]. regressions [Kon04]. regular [ADG04, BD04b, JO03, PR02, Tom02]. Regularity [dB02, MC03]. regularization [AD00, CMRS00, Fu04, Iqb03, SS03c]. regularizing [CRSS00]. regulator [FD00]. Reissner [AP02a]. rejection [Wan00b]. related [AGRZ03, BP01a, CACK04, CMV01, Cof03, DMGVO01, FCP02, GMRS00, Had00, Kan00, Kui01b, OO01, OO02, Ped03, SKSV03, San00, SS04, SC04, Van00a, Vel01, WZ03]. relation [Hu02, MY03a, RKS04, xZqZX02]. Relations [Che03a, Cof04, BdSR03, KP03a, Kir00, Pom01, Vid03]. relationship [LZ04]. Relative [BH04, Ips00]. relaxation [BSKP04, yCjC01, El 04, Gar04, Jan03b, LYC02, YB03]. relaxed [FLW01]. relevance [BS00d]. reliable [IYO03, vdV02]. remainder [Neh03, Tan02]. Remarks [Ism03b, MC05, NRL03, Sri02, EP03, Mak04, Tam03]. removal [CS03b]. Removing [SM04a]. renewal [McC01]. Rényi [PP00a]. repetitive [DBDPF04]. replace [Par00b]. representation [Dra02, Kar01, SV04, WW00b]. Representations [Lew03b, CK02b, DV01, KRT02, LV03, SS03a, ST00b]. representing [Wri02]. rescaling [Yu02]. residual [AM03, EES00, GH01a, Hug01]. residual-based [Hug01]. resins [MMPZ04]. resistance [Joh02]. resolution [AS04a, BJ02, JS04, Pas04, SJ03a]. Resolvent [BS00d, Noo01]. resonance [BE02b, RS04, TQ03]. resonance-free [BE02b]. Resonances [BEM00b, BE02c, Hit02b]. Resonant [Bac01]. response [BD00a, HL04b, MSM04, WL04]. restarted [CKR02, EES00]. restricted [CO01]. Restrictions [CG03a]. result [Akk02, CD03a, Far02, KSSV03]. resultant [Sya04]. results [AO00c, BBMS04, BLM04, BE04b, CH02a, CJ04, Cof04, Elb01, FPS00, HR03, JKN00, KS00, Law02, LS01, NP03, Sim00, WH01, WL03, WH04]. Retakh [KR03b]. Retarded [Bak00b, DS02b, Far02, Mas03]. retrieval [KATS02]. return [CK02a]. reverse [YTI02]. Reversing [BS00a]. Review [FF02, BZS02, Stü01]. revised [PT03]. revisited [Dam03, MF01, Ses04]. revolution [QH04]. Reynolds [CM04, GNPB01]. Rhine [SS02b]. Riccati [DHP02]. Richards [BK02, PRK04]. Richardson [HW02, Sid00]. Richtmyer [AVG^{+}04]. ridge [Li02]. Riemann [BL01c, BBC00, Cof04, CK02b, DKM^{+}01, Lac03, Weg01a, ZL03]. Riesz [vG00]. right [Age02, LR04b]. right-hand [LR04b]. rigid [CF04]. rigid-body [CF04]. risk [DFT04, Oht03]. Ritz [CJ04]. river [SS02b, SRRS04]. RKN [GHR02]. Robin [AH03, CL02, HC03b]. Robust [CW00, MSI04, Sug02b, Wri02]. role [Col01, Fre02, Wün03a]. Rook [PN00]. root [Bat04, BM01b, CF00, Her03, KP01, PR04, RZ02, Sch03c]. root-bound [Bat04]. root-finding [KP01, RZ02]. rootfinding [Hom03]. roots [Ano02a, CL01a, FS03, Kra04, PNV01, RZ04, Rum03, dBS01]. Rotach [AGMMB00b]. rotary [VVD04]. rotating [Ari03]. Rotation [CK01]. rotations [BZPFB04]. rotatory [LL00]. Rothe [GN02]. rough [AB04a]. Rounding [Bar02, MRT00]. routing [JT02a]. row [BTSHK04]. Ruijgrok [LWW01]. rule [Ham01, Hor01, KC00b, KC00a, MO04, MAK01, Pöt02, Ses04]. rules [BS00a, Ehr01a, Gen03, KY02, KCI02, NR01, Smi03]. Runge [BE03, BM02, BJ02, CLMR00, CGM02, CJ03, DM02a, Fra02, Fra04, FKM02, IZ00, Kot02, LSY04, Röß04, TA02, TS02, VV01, VDVV00, VID01, VVV03b]. Runs [Lou02].

Saad [CJ04]. Sabin [LPSSP00, MVDB04, WVDB03]. saddle [CH03, CD03b, Cui02]. Saff [Dam03]. Saliga [Akk02]. Salvatore [MP03a]. Sampling [AB04b, CCH02, GHM01, GL04a, QH04, Str00, Wan00b]. satisfied [FR01]. satisfying [Kra03, KL03c]. saturated [HRK04]. saturation [Qui02]. Scalable [KRW00, CW04]. Scalar [Hom00, Jeo00, CRSL00, FMRW04, FP02, Jeo01]. scale [ABO02b, BI03, BT00a, BP01b, DHP02, EP02a, GJL^{+}00, GLM00, aH02, LP00, Won04, YSNM02]. scale-invariant [BP01b, GLM00]. scales [ABOP02, AM02a, AG02a, Bha02, DH02, GK02, Hil02b, LV02, Sie02]. scaling [AB01, ZH04, Zhu03a]. scattered [CO01, DFI02, Isk03, Lai00, LM02a, Yoo03, dSPPT01]. scattering [AP00, HJ02, Hit02b, KL03a, PS03a, Pot00, Sey01, SW02b]. Schafheitlin [Mil00]. scheme [AGRZ01, BJK03, BvdV01, CMSV03, CS04c, CJL03, DN01, DN02a, EHS00, FLW01, Fer03, KZ03, KW03, Len02, LMZ02, LWW01, LLG04, MS03b, Mon01, SRD01, Slo04, TL01, VWE04, WF02, Yoo03, ZLF01]. schemes [AA02a, ALM03, BBK04, CL04, CG04b, Deh02, DF00, FW02, Fur01, GZ02, JY01, JO03, LRS02, LYC02, LC04, Man00a, Mat03, TT02, VV01, WL01, Yam02, ZKO02]. Schrödinger [BGP02, DN02b, GW04, Ixa00, KMS03, Kan04, KS03b, RLZ03]. Schur [MRT00, MN02b]. Schwarz [Lui01, Lui03, MMOS01, ST02, ZLF01]. Science [Ano01p, Wat02]. scratched [LHYaC03]. SDEs [LRS02, Pen00b]. SDEs-a [LRS02]. search [CS02, LTT00, Nei02, PP02a, PT03, Tas00b, Zhu00]. search-optimized [Tas00b]. searches [SS04]. Secant [HRE00, Zhu01]. Secant-like [HRE00]. Second [Büh00, Deh02, Abd03, AGRZ03, Arg01a, AA02c, AK02, BH04, CS04a, CCL03, CGM02, Cha02, Chi01a, DV00, EP02a, EH04, GKMN00, GSG03, GT04, GK02, HP00, HG03, JW00, JFW01, KL01a, KL03c, Len02, LLC01, LL01b, LB01, MMOS01, MFMGZ02, SA00, SW00b, Slo04, Tse04, VV01, VAR03, WL03, WW03, Yan00, ZW02]. Second-order [Büh00, Deh02, AA02c, AK02, BH04, CS04a, CCL03, Chi01a, DV00, EP02a, GSG03, GT04, HG03, JW00, JFW01, KL01a, KL03c, LL01b, LB01, MMOS01, SA00, SW00b, Slo04, Tse04, VV01, VAR03, WL03, WW03, Yan00]. seemingly [Kon04]. segments [WMA03]. Seidel [Li03a, NHMS04]. Selberg [MY03a, Riv03]. selected [Tas00b]. Selecting [FC04]. selection [XL02]. selective [CMV01]. Self [Pet02c, Rui02]. self-adjoint [Rui02]. self-scaling [AB01]. Self-validating [Pet02c]. selfadjoint [LMZ02]. Semi [MS03b, BE03, BH03, FLW01, HHR00, KyGUY02, Leu00, LCF04, TA03, YOT^{+}02]. semi-classical [HHR00]. semi-definite [KyGUY02]. semi-differential [TA03]. semi-discretization [Leu00]. Semi-implicit [MS03b, LCF04]. semi-infinite [FLW01]. semi-Lagrangian [BE03, YOT^{+}02]. semi-linear [BH03]. semiaxis [DVM02]. semiclassical [Suá01, VZ04]. semiconductor [CL03b, FI03]. semiconductors [CS04b]. Semiconvergence [Son01]. semigroup [Sai04]. semilinear [CY00, Hau02, KS01a, Kas00, NY03, ST02, Ues04]. semilocal [Arg01a, HS00]. semirings [Bou04]. semiseparable [VMV04]. semismooth [TLQ02]. sense [BL04]. Sensitivity [MR01, Rih03, BM00c, CLP02, LP00]. sensor [FD00]. separable [AAD02, AG00b]. separated [AJ01]. separation [BBK04, SM04a]. sequence [BL01a, Hom00, Suá03, Wün03b]. sequences [BDR02, CVB04, Chi01c, KLM01, MDSR04, MC05, Rad00, Sid00]. Sequential [BT00a, TO03]. Series [Tom02, AJDG02, AT00, AAD02, Bar02, BS02b, CS00b, DSS03, HKT^{+}03, KO99, KO01, KOH03a, KO03, KR02, Lie04, Lov00, Mou03, Neh03, Nis03b, OO01, OO02, SRLAD03, ST00b, VB01, VD03b, Van00a, Van02, Vid03, Wün03b, dAMR03]. set [ASB04, AK00, Dia03, FI03, Hu01]. set-valued [ASB04]. sets [AKM03, Alf00, CZ03, CS01, DDGH03a, DDGH03b, DF00, LM02a, MN00, NMST02, Nie03, Pre00a, SdSP01, Xu03b]. settling [PGG02, PGG03]. Several [Tam03, BB00, Hat03, KS02b, Mac02c, XL02]. Shadowing [BO04b, Far02]. shallow [CDV03, GN03]. Shanks [Sen01]. Shape [CM03, Lev00, Häh00, Mel02, MP02b]. Shape-preserving [CM03, Lev00]. shaped [MW02]. shapes [TA03]. sharp [AB04a]. shearing [LJ03]. sheet [CDV03, NSP04]. shell [BZPFB04, Chi01d]. shells [Mar00a]. shielding [SDMV04]. shifts [BM04a]. Shishkin [CG04b, GCL02]. shooting [GM00b]. short [Tho01]. shortest [Bou04]. Shortley [LYF03, MY00]. Shout [WFV01]. side [LR04b, YLC04]. sided [DFC02]. sideways [Fu04]. Sidi [Osa00]. Siegel [CO02]. sign [AO00c, CDN01, GT04, KV04, WA00b]. sign-changing [CDN01]. signal [SVV01, SV04]. Signs [Dem02]. similar [Sch03b]. Similarity [eMKM04, DDN01, Öze04]. simple [AS04c, PNV01]. simplest [Yu02]. simplex [DvM01, YF03]. simplices [Pet04, Xu01b]. simplicial [DvM01]. simplicity [BG03]. Simplified [Fu04]. Simpson [Hor01]. simulating [JT03]. simulation [BI03, CS04b, DHW04, Fre02, Fre00, KT02, Koz00, MD04, NSP04, OOA03, OOAS03, PGG02, PGG03, Pen00b, RDA04, San03, SRRS04, Win03, Win04a, Win04b, YZCL04, YLC04]. simulations [AVG^{+}04, Hat03, MOTT03, MP02a, PCR04, VVD04]. Simultaneous [Sch04c, PH01, Pet02b, SK01]. Sinc [AK01a, AK01b, Sug02a, Ste00a, SM04b]. Sinc-collocation [Sug02a]. single [BTSHK04, GPS03, GH03b, HHC01, KT02, YLC04, Zha03]. single-side [YLC04]. singular [ADG04, AO00c, AK01c, BO04a, BM00a, Büh00, Cao02, CG04a, DS04, Die98, Die02, Fan03a, FW02, GKMN00, GHMY00, GJHY00, Hei03b, JW00, JW04, JE01, JE02, JM00, JS00b, Jun04, KWW01, Kum02, LS02a, Lau00, MO02, MO04, Pal02b, PS04, Pot00, Rac00, RZ02, Son01, SZZ01, TW04, Tas00a, Tas04, Tom02, WW00a, WS02b, ZW02]. singularities [Pan02]. singularity [SM04a]. Singularly [FHM^{+}04, OS04, AG03b, AK02, BM01a, Bog01, Bog04a, Bog04b, BD04b, CL04, Che03b, HR02b, JY01, Len02, LHHW04, MOS02, MSKS04, NJVA03, QST00, Shi04a, Shi04b, SS03c, Yeh04, ZL04]. sintering [APT02]. Sivashinsky [BO02]. size [HL02, Iqb00, Mut02, Sch00b]. skeleton [PR02]. skeleton-regular [PR02]. skew [Ng03]. skew-circulant [Ng03]. slackness [TW04]. slender [Gan04]. Śleszyński [BL01b, xZqZqT04]. Slip [Joh02, Gor04, Wat03]. slit [OOA03]. Slow [dGK01]. slowly [DN02b, HKT^{+}03, Kui01a, Kui01b, Oou01]. small [Hig03, KR02, Pan02]. Smooth [DEL04, Cha02, HGI04, OS04, RDA04, Wan04a, WT01, ZL04]. smoothed [LLL03]. Smoothing [GN02, AD00, CSRB04, CMV01, FGJ00, LLL03, SQ01, Tay00, XZ01]. smoothing-nonsmooth [SQ01]. smoothness [CM03, LYF03]. Sobolev [AMBP^{+}01, AGMMB00a, AGMMB00b, Bav00a, Bav01, Bav03, BCM01, BC02, BC03, BH03, CGMB03, ELW04, LCI01, MF01, MdB02]. Sobolev-type [Bav00a, Bav01, Bav03]. Software [LP00, CLP02, DE00, Pau00, SSV04, WKM04]. softwares [KLyD^{+}03]. soil [ATG04]. soils [Qui02]. solar [KCMK02]. solenoidal [Swa04]. solid [CM04, Ism00, YOT^{+}02]. solitary [AK01b, KOZ03]. solitons [AK01b]. Solution [Ban04, FL04, JS03, LR02, Rem04, ADL^{+}02, AK01b, AH04, AKM03, AH03, BO04a, BL01c, BM04b, CDV03, Cas00, CGM02, Cha02, CS01, CZY03, CL03b, CD01, CR03, Dal04, DHP02, Duf00, EFS02, EK01, Fer03, FLM03, GKMN00, GS01a, GJ01, HW02, HC03a, HL04b, Kai03, KMS03, KWW01, KS03b, LK01, LS02a, LRS02, LYF03, LD02, Liu02a, LAT04, LSY04, LRvSS04, LS02c, Mak04, Mas03, MC03, Min04, MAK01, OASO03, PS02, QST00, RV00, SvdV00, Sch04c, SW00a, SdSP01, Swa02, TA02, UC03a, UC03b, VAR03, WZW03, Waz01, WH00, Weg01a, Xu03b, ZW02]. Solutions [EM04a, Aas02, eMKM04, AJ01, AK00, AGRZ03, AG01b, AG02a, AA02c, BM00a, BW04b, BBM00, CN00, Cam01, Chr01, CH02b, CDN01, DKK03, DH00, DS02b, Din03, EF02, EPM00, EKLW01, Fan03a, FW03, GHMY00, GVSJ01, GNPB01, GLM00, GR01, GH03a, GSG03, GT04, HP00, He03, Iqb03, IY03, Ism00, Jan03a, JKY02, JFW01, JKN00, KS01a, Kaz02, KCMK02, KMCK02, Koe03, KN02, Kor01, KED04, Lee02, LZ02a, LS04b, LHHW04, LR04b, LZ01, LW02, LG04, MS03a, MFMGZ02, Mig01, MP01, Min01, MY03b, NS03, Pal02b, Pao01, PS03b, Rac00, RN03, Shi04a, SW00b, SZ04, SZ03a, SC01, Swa04, TRTG03, Tom02, TS02, Vil03, WL04, WA00b, YF03, YL03, ZL04]. Solvability [BL01a, BRS03a, AZ04, DHZ04]. solvable [KN02, LD02]. solve [eMBH02, ABG03, AG02b, El 04, HR04a, SA01]. solver [AM03, CG02, YOT^{+}02]. solvers [Mor00, SIM02]. Solving [AGRZ01, FLW01, Hu01, LLC01, LQQ01, MM00, Sun01, SQ01, eMKM02, ADG04, ASN02, AG00b, Arg04b, BK02, BM00c, Cao02, CFT02, DZN00, DN01, DN02a, Din04, Doh02, ENE04, GLZ03, HV02, HRE00, HCT^{+}02, KYF04, LZ02b, LMP00, Mal04b, Mar00b, Min03, Pau00, PH01, Pie00, SA00, SC00, SS00, TLQ02, VMV04, WG02, Woź01, ZLF01, Zha02, dMPL03]. Some [AO00c, ACV03, Bak03, BS03, BGM00, BGM01, BCJW00, BKB04, CACK04, CG03b, CZ03, lCHcH00, CV01, Elb01, EP03, GCL02, JNT02, Kar00b, KS00, LF04, Mak04, Min03, Miy03, NP03, OO02, PP00c, Ram03, SRD01, Sim00, Smi03, Sta03, VC01, XZ03, ÁNAA03, DFM04, BdSR03, Cam01, CH00, Che01, Cof03, Dah01, DV00, Dom03, EGV03, EHS00, FGJ00, FP02, Goh02, GD04, HKT^{+}03, IZ01, JW04, KCMK02, KMCK02, Koh01, LN03, LY01, LY00, Mac02b, MDSR04, MO00a, MW02, ML03, MS04, MC05, Mor00, Pet02b, Ric00, SRLAD03, Sch04a, Sla00, Vel01, WLYL04, ZL04, dB02]. Sonin [MO04]. Sonine [Vel01]. SOR [Had00, Cha04, CK03, Son01, WJ01, Woź01]. sound [SW02b]. source [IZ01, MD00, Mon01, RZ03]. sources [Pot00]. space [Arg01a, BG02, Car03, De 00, DNS00, Dra02, FHMS04, Gu04, LAT04, MKS^{+}01, SW00a, SW02a, SVV01, Vou01, WX02a, ZL03]. space-dependent [ZL03]. space-time [SW00a]. spaced [Wol00b]. Spaces [BEM00a, AO00b, Alf00, ABC02, AB04b, BPV02, BL01a, BM03b, DX02, Din02, Din03, GKS00, Hau02, Lau00, LH03, Mai03, Mar04a, NZ00b, Peñ00a, Pre00a, Sch00a, Wat00b, vG00]. span [IM01]. Spanier [Dal01]. Sparse [AA02b, BGP02, GS04b, GSS03, HK00, MAK01, SZ01, Ver01]. spatial [Kou03, WF02]. Special [Dun03, ERV04, JKVV03, TRTG03, Bac01, DFM04, Cla03, GNPB01, Lor03b, Mac02a, Mac02c, Min03, PNV01, QZ03, Seg03, VD03b, dB03]. specific [MD04]. specification [AG01a]. spectra [GL02]. Spectral [BF02, DB01, GGM01, GH01b, JNS04, Lac03, LMYL01, VZ01, BE02a, BBW04, BGP02, BE04b, Cha04, CS01, ELW00, ELW02, ELW04, FW02, FK04, Gan04, GD04, Hei03b, Ifa01, KSCI00, LS02b, PP02c, Sie02, Tas00a, Tas04, Tas00b]. Spectrally [DKL04]. spectrum [Beh02, BMZ00, BR02, GW04, Hin02, HR03]. speed [AR03]. sphere [AS04c, SW02b]. spheres [LMYL01, LN03, Pet04, Xu01b]. spherical [Gen03, KP03b, Mar00a]. spherically [YOO02]. spheroidal [KR02]. spheroids [Gan04]. Spike [ZL04]. Spike-layered [ZL04]. spin [SZ03b]. spiral [WMA03, WM04a]. Spitzer [Ste02]. Spline [DF00, SW02a, dSPPT01, ASN02, Alf00, AS04c, Ant02, AK02, BF01, BM03b, CS03b, Die98, Die02, DDPT00, jGwHsZ04, HR02b, HM00, KED04, LH03, MW04a, Mil01b, NZ00a, NZ00b, Peñ00a, SA00, SYW02, SU04, Sya03, Wan00a, WX02a, WVDB03, WA02, Wri04]. spline-based [Die98, Die02]. splines [BM03b, BM04c, CJ00, CM01, DL01a, DSS00, DN00, GL04b, KP00, KP02, KP04, Lai00, Liu02b, LPSSP00, MVDB04, NS04b, SW02a, Zub04]. split [LF03]. splitting [Bai03a, DHK02, LWW01, MTA^{+}03, MT03, Mon01, MO03, Ng03, Noo01]. splittings [CP03, WW01]. Sponsored [Ani02]. sport [NN04]. spreading [Kas00]. spring [BD00a]. SQP [Bet00, BM00c, GJL^{+}00, MM00]. SQP-methods [BM00c]. square [BM01b, CF00, CP03, Mut02, PR04]. square-root [PR04]. squares [AW03, Bai01, BVV04, Dem02, Hei03b, HR00, JS03, JKS04, KED04, LS00a, Nie00, XKC01]. Srinivasa [VVV03a]. Stability [BW04a, CS04a, GH03b, HJ02, HL02, Kot02, Kot03, LWY02, LSY04, LS01, Pot00, QM01, Vec00, ZV04, BGZ00, BS00d, BBM00, CS00a, CZW04, CJ03, DS02c, DMN01, EF02, GL02, Hat03, HW03, HHC01, JR02, KS01a, LCF04, LLN04, MVDB04, Nag04, PJ03, Qui02, RS00]. stabilization [LL00, Mar00a, Par03c]. stabilized [LMO01]. Stable [XKC01, AT00, ALM03, CMSV03, CB00, DN02a, JPW04, Koz00, LF04]. stacking [KL04a]. stage [BW02, DM02a, Gar04, GHR02, LCC04, YK04]. stage-structured [LCC04]. staggered [XCZ02]. standard [CP04]. starlike [LW03b]. stars [BTFY02, HW00]. Starting [GHR02, CP04, DvM01]. state [BG02, BM00c, De 00, Dra02, aH02, MKS^{+}01, MM00, SD00]. state-of-the-art [SD00]. state-space [BG02, De 00, Dra02, MKS^{+}01]. states [Kui01a]. static [Fre03]. stationary [KZ03, Mal04a, SW01, Swa02, Swa04, Yoo03]. statistical [DFT04, Goh02]. statistics [BA00]. steady [Bai01]. steel [WWA^{+}04]. Stefan [GN02]. Steffensen [ABC02]. stellar [ADL^{+}02]. Step [YY02, CP04, CJ03, FT03, HL02, JPW04, KS03b, LF03, MSM04, Sch00b, VID01, YLC04]. step-length [VID01]. stepping [BvdV01, LF04]. stepsize [BHB04, UH00, VALM00]. Stewart [CJ04]. Stieltjes [BRS00, GPTT02, IK03, SL03]. stiff [BT01, Cas00, HW00, XLFC00, YB03]. Stirling [ELW02]. Stochastic [Win03, Win04a, Win04b, Buc00, BBM00, BT01, BHB04, Hau02, Lam03, LS04a, LW01, LWT04, LCF04, MS03a, Röß04, TA02, Vil03, XL02]. stochastics [Sch01]. Stokes [Abd00, BJK03, CM04, Daa04, DV01, Fuj02, GW04, Joh02, JL00, KK03, Ran04, SEKW01, Swa02, Swa04, Yan02]. Stopping [Ehr01b, Oht03]. Störmer [vdHMS00]. strategies [BBC00, EES00, FKM02, Iqb00, JT03, SZ01, UH00]. strategy [PN00]. stratified [Ehr02]. Stratonovich [Röß04]. stream [Abd00]. stream-function [Abd00]. streamline [RZ03]. streamline-diffusion [RZ03]. stress [HH04, Yan02]. strings [Leu00]. strip [Ara04]. strong [DMGVO01, KYF04]. strongly [Die98, Die02, XLFC00]. structural [Ava00, Bak03]. structure [BKB04, Jea04, KS01b, MT02, MKS^{+}01, MR01, Ver01, YH02]. structured [CKR02, LCC04]. structures [BZPFB04, DBDPF04, McC00, Yan03]. study [AK01a, AVMRVM02, Ehr01a, MÁNM01, MG03, Mon01, SWSZ04, VS01, VYZ01, WWA^{+}04]. studying [SZ03a]. Sturm [AM02b, BBW02, BLM04, BBW04, BMZ00, BEM00b, BE02b, BE02c, CS01, DH02, GGM01, GGM04, GM00c, KWZ04, Mar04b, MZ02, Tas00a, Tas04]. style [Lov00]. sub [DN02a]. sub-microscale [DN02a]. subclasses [LW03b]. Subdivision [Sch02b, BS00a, CJ00, DLTS02, DF00, jGwHsZ04, JO03, LLG04]. subdomain [HC03b]. subfilter [BI03]. subfilter-scale [BI03]. Subject [Ano01t, Ano03-37, Ano03-36, Ano04-29, AG01a, Zhu03a]. subproper [WW01]. subsequences [Gro02]. subsets [Dra04]. subspace [DDGH03a, Fre00, MKS^{+}01, SSV04, WW00a, Xu01a, Zha02]. subspace-based [SSV04]. subspaces [Ips00]. substitution [Dal04]. substrate [BCM03]. substructuring [BS00c, Rix04]. subtracted [Pan02]. successes [Hof00]. Successive [Had00]. suction [MSKS04]. sufficient [CL02, Cui02, DH02]. sum [AW03, NR01, QSZ02, Sam01b, Ses04, Sri02]. summability [Tam03]. Summary [Ste00a, EW01]. summation [Sch03c]. sums [BS02b, KL03b, Par03b, Sta03]. super [Bar04, EH00]. super-Halley [EH00]. super-irreducible [Bar04]. supercomputing [Oya02]. Superconvergence [LYF03, MY00]. superlinear [CDN01]. supplementary [Ano02a]. support [SS02c, TL04, ZL04]. supported [Han03a]. supports [SRD00]. supreme [BD04a]. Surface [GL04b, Xu04, eMKM04, AD01, AW03, CC03b, JT03, MMPZ04, Wan04a]. surfaces [Gen03, HM00, KP04, LW04, LNLL02, QH04, Rab03, SYW02, WLYL04, YY02, dSPT02]. surfactant [NBS04]. survey [ABOP02, Faz02, Gat02, KV04, Lor00, MO01, Nas00, NHMS04, PH01]. survival [NN04]. suspensions [BBK04]. SVD [Chr01, DEL04]. swaps [MA04]. swelling [Qui02]. Sylvester [GVV04]. Symbiosis [O’L00]. Symbolic [AG03a, GZ02]. Symbolic-numerical [AG03a]. Symmetric [OA04, Rui02, BDR02, BR02, Che01, CDN01, Din02, EST03, Gen03, KS03b, LMZ02, MO00b, Min04, RSvR03, ST00a, YOO02]. symmetrizable [DHZ04]. symmetry [Nis03b, Öze04, WS02b]. Symplectic [KMS03, BM02]. symptotic [Abd03]. system [Aas02, Ara04, BF02, CFT02, CZY03, De 00, Fab02, HL04a, HL04b, Kaz02, KW00, KK03, MKS^{+}01, MOS02, MG01, MSM04, QM01, RCZV04, SAA01, Sch03c, SSV04, SC01, SS00, Tay00, Wol00b, WA00b, YL03, YSNM02]. systems [AH04, AKM03, AL03, AK00, AG00b, ADEL00, AG02b, BW02, Bai03a, BL01a, BSKP04, Bar04, BJK04a, BBKS00, BSS03, BBK04, BW04a, BX02, BD00b, CS00a, CES02, CRSS00, Cao02, CMSV03, CW00, CN02, CG04a, CP00b, Dal04, DSVB04, DKST04, Duf00, EFS02, Ell01, EK01, EMT01, Fab00, FC04, FPS00, GJL^{+}00, GLZ03, Gug00, GSS03, HGI04, HW03, HHC01, IN03, Iwa02, Jan03b, JkSlS04, qJkSlS04, KT01, Kas00, Kno02, KN02, LV02, Leu00, LP00, Li03a, LS04b, Lor03a, MSI04, Man00b, Mar00b, MO01, Mat03, MR01, Min03, MAK01, Müh00, Nak01, Ng03, Par01, PJ03, Par03c, PS02, PP00c, Rih03, Röß04, RV00, SvdV00, SZ01, San03, Sch03d, SB01, SBSA03, Son01, SV00, SZZ01, SRRS04, Sya04, TYZL04, Tru04, VMV04, VS01, Vou01, WW00a, WZ03, Woź01, YB03, ZK04]. systems [Zha02, Zub04, dMPL03, vdV02, BM03b]. Szego [ACGR01, BC03, BdSR03, Pet01, RLS01].

t [Kob00]. tableaus [ÁNM01]. tail [Dev02]. taking [BQ00]. talks [Ano01r]. tangent [KP02]. tangential [Yan03]. tanh [Mal04b]. Tarski [Sim00]. Tau [GH04, MRV04]. Taylor [KO01, AVG^{+}04, IS03, Kal00, KO99, KOH03a, KO03, MGL01, Mou03, Neh03, TVA03]. Tchebycheff [KL04b]. TE [Sey01]. tearing [BBMS04]. technique [DHK02, PDVS04, RS04, Yam02, Zhu01]. techniques [GM00a, JT03, Shi04b, VWE04, Wan01b]. technology [BS02a, CACK04, CW04]. Ted [AIV01]. telegraph [DKST04]. Telles [JE01]. Temme [Ses04]. temperature [CDV03, VMD04]. tempered [Car03]. Ten [Rum03]. Tension [HR02b, MMPZ04, Man00a]. Tensor [Zeg04, HH04, LV03]. Tensor-product [Zeg04]. term [Bog01, EFS02, RKS04, Tse04, xZqZX02]. terminating [Lie04]. terms [JB04, Mon01, Pom01]. tessellating [YY02]. test [CK02a, Geo03, KNSmG00, PP00a]. tests [Joh02, PP02b]. th [EM04b, FR03, Pom01]. Their [JKVV03, BRS00, Dat00, DRC02, Dmy04, Doh02, ML03, Sch03d, SC04, VZ04, ZV04]. them [Peh03]. Theorem [Bea01, Lag01, Arg01b, Arg01a, Arg03b, BD04a, Bha02, BL04, BPW02, Chi03, Dam03, Dem04, GHM01, Ism03a, Kal00, KHMN02, KY00, MÁN01, Sim00, STW00, Tan02, ZZ03a, xZqZqT04, BM01b, LCZ04, WZ04b]. Theorems [BYY03, AQ04, Arg04a, BS03, CF00, GKK00, Ips00, IS03, LL02, LL01b, Par00a, VB01, WA00a]. theoretic [eMEM00, eMKM02, eMBH02, KP03a]. Theoretical [AM03]. theories [Ben02]. Theory [WW00c, ADG03, AR04, AM00, AP04b, BB03b, BB03a, BBMS04, BCJW00, BGP02, BE04b, BHB04, Čer02, CL00, CG04a, Dal01, De 00, ELW00, Gal00, GL04a, Hil02a, Hug01, JNS04, KS00, LV00, Sch00a, Sey01, Sim04, Wün03a, ZL03]. thermal [CS04b, MT03]. Thermally [DPM04]. thermoelastic [Aas02]. theta [Cho03, Coo03, FK04, PW02, Wen00]. theta-functions [FK04]. Thiele [WG04]. thin [eMKM02, BCM03, DN02a, HJ02, JKN00, MMPZ04, Sur01]. thinning [DFI02]. Third [ABC03, FS03, GST03, KWW01, LYC02, VWE04]. Third-order [ABC03, FS03, KWW01, LYC02, VWE04]. three [AG01a, AA02c, BMN01, Büh00, CG02, CN02, GZ02, Glo03, HC03a, Kum02, Nis03a, SW01, VD03b, ZW02, xZqZX02]. three-dimensional [AG01a, CN02, Glo03, HC03a, Nis03a, SW01]. three-point [Kum02, ZW02]. three-term [xZqZX02]. thresholding [BP04]. thrust [Bet00]. tight [Han03a]. Tikhonov [CMRS00, Fu04]. Time [ABO02b, HS01, Aas02, ABOP02, AM02a, AG02a, Bai01, BRS03b, Bha02, BvdV01, BV03, BM00c, Cha02, CZY03, CZW04, CK02a, CD03b, DHP02, DH02, EK01, EP02a, EW01, FW03, Gug00, GK02, HL02, Hil02b, KS03a, Kou03, LV02, LF04, Len02, Les02, LWY02, MSI04, MD00, MN02a, NSP04, Rih03, RS00, SW00a, SBSA03, Sie02, Won04, XCZ02, ZV04, vdH00, vdHS01]. time-delay [SBSA03]. time-delayed [EK01]. Time-dependent [HS01, Bai01, EW01, Len02, Les02, MD00, vdHS01]. time-domain [BRS03b, XCZ02]. time-harmonic [CD03b]. time-integration [NSP04]. time-lags [Rih03]. time-optimal [Gug00]. time-stepping [BvdV01]. times [NN04]. timestepping [Lam03]. tip [LL00]. Titchmarsh [AP04b, CG04a, Yak00]. Toda [AB03, Peh01]. Toeplitz [Min03, Ng03, PW02]. toll [Nei02, PP02a]. tool [Mal04b]. top [DDG04]. topics [GMRS00]. topography [AD01]. total [NOI02]. totally [GM00a]. Toyota [Ano02y, Ani02]. trace [ST00a]. tracer [WF02]. Tracing [CCM01]. tracking [FD00, Zhu03b]. training [MO00b]. trajectory [Bet00, BCJW00]. transcription [BBCH00, BBCH02]. transfer [ADL^{+}02, Arg01b, FL04, Las04, MSKS04, VMD04, ZFYY04]. Transform [DVM02, Bar03, GKS00, Has04, KOZ03, Mil01a, Nay02, Oou01, Yu02]. Transformation [Van02, Mil03b, MS01b, Oou01, Oou03, Sen01, Sid00, Sug02a, Yak00]. Transformations [JE02, AL03, CS03a, DSS03, DM02b, Far00, GKMN00, Hom00, KCMK02, LS02a, VZ01, Zeg04]. transforming [dAMR03]. transforms [ÁNAA03, CH00, Dac01, Dac03, IK03, PD03, Pie00, Tor03, Yak02]. transient [AD01, KLY04]. transistors [FI03]. transition [KL02, Sch03b, WM02, WMA03]. Transmission [CP00a]. transonic [SW01]. transpiration [Las04]. transport [AM03, BV00, DN01, DN02a, JT03, KY04, WF02]. transport-chemistry [BV00]. trap [Cai02]. traveling [Zou02]. Travelling [KMCK02, SZ03a, dGK01]. treatment [AVMRVM02, Bak00a, SR04]. tree [GS02, Nei02, PP02a, lXsQsJ03]. Trees [Kra04, Dev02]. Trefftz [ADG04]. trends [Duf00, PRT00]. tri [EM04b]. tri-diagonal [EM04b]. triangles [BV02, CO01, DLM^{+}00, TA03]. Triangular [JO03, FL04, LC04, LS00b, PR02, Rab03]. triangulated [LH03]. triangulations [AKTvD00, DN00, FQR00, HH03a, LPSSP00, NZ00b, Wan04a]. tribute [LMS03]. tridiagonal [BGVHN03, Dal04, IP01, Min03]. tridiagonals [Par00b]. tries [Dev02]. trigonometric [Cho03, Fas02]. Trigonometrically [PS03b]. trimmed [YY02]. trimming [YY02]. trinomial [Gla00]. triple [SW03]. triples [Psa03]. truncated [Faz02, IP01, Nas00, hWqX03]. truncated-Newton [Nas00]. truncation [AP00]. Trust [RRWT00, CZ04, DSW^{+}03, Zhu00, Zhu03b]. trust-region [CZ04, DSW^{+}03]. tube [HRK04]. tubular [McC00]. Turán [MS04]. Turbine [HM00, Ega00, HCT^{+}02]. turbines [SFK^{+}04]. turbulent [WWA^{+}04]. turning [AM02b, Min04, NJVA03, Seg03, Tsu01]. tutorial [Nie00]. Twin [CH02b]. Two [GT04, SS02b, VMV04, AA02a, And02, AG00b, ATG04, AK01c, BW02, Bak03, BG03, Büh00, CF00, CGM02, CMSV03, Che03a, CS02, CJ03, CR01, FTY02, Gar04, GPTT02, GH03a, GJHY00, GJ01, HW02, JPW04, JT02a, Ji04, Kum02, KLM01, LR02, Mar00a, Mil01a, Par03b, PS03a, QST00, Rac00, RKS04, SKSV03, SdSP01, Sug02a, YLD02, YK04]. Two-dimensional [SS02b, CR01, HW02, Ji04, LR02, PS03a]. two-layered [ATG04]. two-neuron [GH03a]. two-parameter [CS02]. two-phase [AA02a]. two-point [And02, AK01c, CGM02, CMSV03, FTY02, GPTT02, GJHY00, Kum02, QST00, Sug02a]. two-stage [BW02, Gar04, YK04]. two-step [CJ03, JPW04]. two-term [RKS04]. type [Abd02, AGRZ03, BRS03a, Bav00a, Bav01, Bav03, BGZ00, BW04b, BRZS00, CGMB03, Dar01, DRC02, DS02b, Dik03, DHK02, EG00, EH04, FW02, FW03, Fuj02, GST03, GL03, GKS00, Gu04, Ham01, HZZ03, HRE00, Hil02b, Hom00, HL04b, IZ01, KT00, KRT02, KC00b, KY02, Lau01, LP01, LMP00, MG00, Mil03b, MY03b, NY03, NKK03, Par04a, Par04b, PC00, Rem00, Sch01, VB01, VVV03b, WL04, WG04, XZ03, Zha02, Zou02, dB02, dGK01]. types [DV00]. typically [KNS03].

ubiquitous [Van00b]. ultra [Kan00]. ultra-hyperbolic [Kan00]. ultraspherical [Doh02, DD01, ELS01, GL03]. unbounded [Čer02, Chi03, CP00b, FLM03, JNS04, KLT04]. uncertain [SBSA03]. uncertainty [MSI04, Par03c]. unconditionally [DN02a, LF04]. unconstrained [FT03, LS00a, SHS02, VALM00]. uncontrollable [SW03]. underdetermined [CRSS00]. undergoing [FW00]. unified [Hui04, Sch00a, Sza02]. Uniform [CL04, Che03b, Has04, KL01a, Mar00a, WVDB03, AK01c, BD04b, GST03, MMOS01, NJVA03, Nie03, Par02a, Par02b, Str00, Zha03]. uniformly [BM01a, CJL03, SU04]. unilateral [AZ04, AZPF04]. uniqueness [AR04, Chr01, Häh00, Kaz02, Min01]. unit [BCM01, BDGV01, CFMV03, CK01, DGVM02, Dik03, MN02b, RLS01, Sim04, dG01]. unitary [RLZ03]. unity [dBS01]. univalent [AHP04]. univariate [RR01]. universal [AA02b, Lee02]. univexity [MWLS03]. unknown [KCB02]. unorthodox [Wün01b]. Unravelling [Hig03]. unrelated [Kon04]. unstable [Kor01]. unsteady [eMKM04]. unstructured [Glo03, VWE04, WF02]. update [Ano02a]. updates [FT03]. Updating [DS04, dMPL03]. updating/self [AB01]. Upper [DHP02, Cha04, PP02c, ZW02]. Urysohn [BRS00]. USAOR [LZ04]. Use [PP00a, Gau01, JT03, Par04a, Par04b, Sch03d]. used [Kál02, NHMS04]. Using [eMBH02, ABC03, AVG^{+}04, Arg01a, BM01a, Bet00, BR00, CO01, Che03b, DKK03, DVM02, DMN01, Eva04, FLW01, Gru03, GGDL04, IYO03, Iqb00, KCI02, KT02, Lai00, LL01a, LS02a, LHYaC03, LD02, Mar00a, Miy02, MSM04, NKK03, NOI02, Pet01, Pie00, PNV01, Pot00, Pré00b, QST00, RLZ03, SS03b, SAE04, Sto05, Swa00, Swa02, Sya04, TW04, Tam03, TO03, Tru04, VWE04, Wol00b, YF03, YK04]. Uvarov [ÁNM01]. Uzawa [Cui02].

V [Arg03b]. validated [Neh03]. validating [Pet02c]. Validation [TVV04]. Value [BGP02, Oht03, ADG04, ABO02b, ASN02, And02, AG02a, AA02c, AK01c, AK02, BM00a, BL01c, Bha02, BS00d, CC03a, Cas00, CGM02, CMSV03, CK04, Chi01a, CH02b, DS04, Dav02, DH00, Deh02, DB01, EP02b, EP03, FTY02, Gru03, GSG03, GT04, GJHY00, HY02, HK01, Jan03a, JW00, JFW01, JW04, KV04, KWW01, Kum02, Law02, LW02, MSKS04, NRL03, Pal02b, PS04, Pre00a, QST00, Rac00, SA00, SC00, Sug02a, Swa00, VV01, Waz01, WA00a, Won02a, ZW02]. valued [ASB04, DV00, DS02a, Fab02, Fre01, Kál02, Koh02, WG04, xZqZX02, ZZ03a, xZqZqT04]. values [HOO03, Riv03, Wal03, ZZ02b]. Vandermonde [Müh00, OA04, RV00, YH02, Yan03]. Variable [GKMN00, LS00a, AJ01, BW04a, CP04, CG03a, DSZ02, GZ02, HL02, Lor03a, MR01, Par04a, Par04b, Sch00b, VVV03b, Zha03]. variables [BA00, BM00c, Čer02, CL03c, Lou02, Öze04, Sug02b]. Variance [Wan01b, Pom01]. variant [ADG04, Arg04a, HV02]. variate [GPS03]. variation [BBCH00, BBCH02, PP00a]. Variational [CFHS03, CFSS03, HSS01, KLT04, QZ03, ASB04, CT00, DL00, DX02, Din03, Din04, DNS00, HS01, Han03b, Hu01, Hu02, KP00, KP04, Noo01, Ric00, SR04, SQ01, XZ03]. variational-like [Din03]. variations [LNBI02]. variety [Law02]. various [TYZL04, WH00]. varying [MFMGO01, SRLAD03]. Vector [JS00a, WG04, AR03, DV00, Dra04, FMRW04, Koh02, KRW00, MY03b, Son02, SS02c, TL04, xZqZX02, ZZ03a, xZqZqT04]. vectorial [MTA^{+}03]. vectors [CJ04, LV03]. velocity [CDV03, Yan02]. velocity-stress-pressure [Yan02]. Verification [CH03, RN03]. Verified [TNW02]. version [CG02, GI03, Hil02b, Hor01, RDA04, Sch03a]. versions [Par00a, Ste00b]. Vertical [dSPT02, eMKM04, BKB04, NBS04]. Very [Bet00]. VI [BMPV00]. via [AGRZ01, CD03b, DSW^{+}03, FGJ00, GH01a, HR02a, HSW00, Kal00, KT00, LV00, LL03b, MKS^{+}01, Sim00, SQ01, TJ04, Tan02, Zhu01]. vibrating [Gug00, Tru04]. vibration [DBDPF04, Miy02]. vibrations [BRS03b]. vibro [PDVS04]. vibro-acoustic [PDVS04]. video [CW04]. view [Chi01b]. viewed [TL01]. VII [Ano01s]. virtual [HH04]. viscoelastic [CFSS03, yCjC01, HS01, HSS01, HLS^{+}03]. viscoelasticity [SW00a]. viscoplastic [BQ00]. viscous [Ari03, KK03, Nag04, OASO03, Wat03]. Vlasov [Wol00b]. VMs [Bru00]. Voigt [PKA03]. Vol [Ano00x, Ano01s, BMPV00, Bre00b, WBBF00, WW00c]. Volterra [Abd03, Bak00a, BRS03a, BO04a, BX02, CP00b, DKST04, GKMN00, HR02b, Kan04, LD02, SW00a, Slo04, SB04, Vec00, ZV04]. Volume [Ano00a, Ano01a, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01b, Ano01c, Ano01d, Ano01e, Ano02i, Ano02j, Ano02l, Ano02k, Ano02m, Ano02n, Ano03h, Ano03k, Ano03-36, Ano04i, AA02a, Ano02g, Ano02h, Ano03g, Ano03b, Ano03c, Ano03d, Ano03i, Ano03j, Ano03e, Ano03f, Ano03l, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Bai01, DKL04, EGV03, FTY02, JT03, LMZ02, MP02a, MS03b, Rui02, Tid02, Tid03, Wan04b]. Volumes [Ano01l, Ano01t, Ano02o, Ano03n, Ano03m, Ano03-37, Ano04k, Ano04j, Ano04-29]. volumetric [VVD04]. vorticity [Abd00]. vs [HH04].

W. [Akk02]. walks [Sch03b]. wall [CM04]. walls [AB04a]. waste [MRVM00]. wastewater [AVMRVM02]. water [GN03, KOZ03, MRVM00, ZFYY04]. wave [ADR02, AJ01, Fur01, GLM00, HSV04, JR02, JKN00, KMCK02, KR02, Mal04b, PDVS04, SZ03a, Ues04, UC03a, UC03b, Zou02]. wave-based [PDVS04]. Waveform [Jan03b, Gar04, YB03]. waveforms [SS02c]. Wavelength [JT02a]. Wavelet [CC03b, Dah01, LLXS04, BP04, CW04, DDGH03a, DDGH03b, FK01, Han03a, HSV04, LZ01, LL03b, PD03, SS02c, Tas00b]. wavelet-support [SS02c]. wavelets [AS04c, Han03a, HH03a, LLC01, WVDB03]. waves [AD01, AK01b, HJ02, KOZ03, MT03, NGGZ04, SWSZ04, aYtZ03, dGK01]. weak [Arg04a, FHM^{+}04, WX02a]. weakly [BO04a, BJK04b, Dra04, GKMN00, JE01, MO02]. Weber [Mil00]. Weibull [Age02, BA00]. weight [BLM04, Kui01a]. Weighted [DDPT00, ZLF01, BS02b, DJK01a, Dam03, KC00a, Kub01, Lau00]. Weights [IRVD01, DJK01c, DJK01a, Kub01, KL03b, Kui01b, KL01b, LL03a, LM02b]. Welch [MO00b]. well [YEWE03]. Weller [LYF03, MY00]. Werner [WG04]. Werner-type [WG04]. Weyl [AP04b, CG04a, Dac01, Dac03]. Which [AG02b, NS03, WFV01, dG01]. Whittaker [Yak02]. widths [Din02]. Wiener [AB04b]. Wigner [AL03, Ram03]. Wilf [HRS02]. Wills [Bak03]. windows [Las02]. wings [Chi01b]. wire [ATG04]. Wirtinger [Hil02b]. Wirtinger-type [Hil02b]. within [AGRZ01, ZFYY04]. without [ABC03, CS02, MMPZ04, PT03]. WKB [MFMGZ02]. work [AIV01, HH04, VVV03a, YEWE03].

Working [Ega00]. world [Hig03]. Worpitzky [ZZ03a, Bea01]. WR [JkSlS04, qJkSlS04]. Wright [GLM00]. Wronskian [Li03b]. Wu [LWW01]. Wynn [Wen00].

x [Kob00].

Yamamoto [TYI03]. years [Chi01b]. yield [Dem02].

Zassenhaus [SJ03b]. zero [MFMGO01, MFMGZ02, WS02a, dAMR03]. zerofinders [ACGR01]. Zeros [DL01b, DD01, Gil01, MdB02, CG03a, DVM02, DSZ02, Dim01b, Dim03, DM02b, EL01, ELS01, Elb01, FPO01, Gat02, Grü01, Kra03, KV01a, KSSV03, Mac02a, Mac02c, NP03, Neu03, Pet02a, Pet02b, Pet03, PSS03, SKSV03, Seg03, SK01, SS03d, Sya04]. Zeta [CS00b, Eli00, BBC00, CK02b, HKT^{+}03, Riv03]. zeta-function [HKT^{+}03].

References

Afif:2002:CFV

[AA02a] M. Afif and B. Amaziane. On convergence of finite volume schemes for one-dimensional two-phase flow in porous media. Journal of Computational and Applied Mathematics, 145(1):31–48, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005349.

Alon:2002:SUG

[AA02b] Noga Alon and Vera Asodi.

Sparse universal graphs. Journal of Computational and Applied Mathematics, 142(1): 1–11, May 1, 2002. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042701004551.

Avery:2002:ETP

[AA02c] Richard I. Avery and Douglas R. Anderson. Existence of three positive solutions to a second-order boundary value problem on a measure chain. Journal of Computational and Applied Mathematics, 141(1–2):65–73, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004368.

Amato:2002:FSA

[AAD02] U. Amato, A. Antoniadis, and I. De Feis. Fourier series approximation of separable models. Journal of Computational and Applied Mathematics, 146 (2):459–479, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003989.

Aassila:2002:LTB

[Aas02] Mohammed Aassila. Large time behaviour of solutions for a nonlinear thermoelastic system. Journal of Computational and Applied Mathematics, 146 (2):481–491, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003990.

Al-Baali:2001:BCE

[AB01] M. Al-Baali. On the behaviour of a combined extra-updating/ self-scaling BFGS method. Journal of Computational and Applied Mathematics, 134 (1–2):269–281, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005549.

Aptekarev:2003:PAC

[AB03] A. I. Aptekarev and A. Branquinho. Padé approximants and complex high order Toda lattices. Journal of Computational and Applied Mathematics, 155(2):231–237, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200866X.

Amirat:2004:NAL

[AB04a] Youcef Amirat and Olivier Bodart. Numerical approximation of laminar flows over rough walls with sharp asperities. Journal of Computational and Applied Mathematics, 164–165(??):25–38, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006393.

Annaby:2004:SPW

[AB04b] Mahmoud Annaby and Paul Butzer. Sampling in Paley– Wiener spaces associated with fractional integro-differential operators. Journal of Computational and Applied Mathematics, 171(1–2):39–57, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042704000470.

Amat:2002:CQN

[ABC02] Sergio Amat, Sonia Busquier, and Vicente Candela. A class of quasi-Newton generalized Steffensen methods on Banach spaces. Journal of Computational and Applied Mathematics, 149(2):397–406, December 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004843.

Amat:2003:TOI

[ABC03] Sergio Amat, Sonia Busquier, and Vicente Candela. Thirdorder iterative methods without using any Fréchet derivative. Journal of Computational and Applied Mathematics, 158(1):11–18, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004606.

Abdellatif:2000:MSF

[Abd00] Nehla Abdellatif. A mixed stream-function and vorticity formulation for axisymmetric Navier–Stokes equations. Journal of Computational and Applied Mathematics, 117(1):61–83, May 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003337.

Abdou:2002:IEM

[Abd02] M. A. Abdou. Integral equation of mixed type and integrals of orthogonal polynomials. Journal of Computational and Applied Mathematics, 138(2):273–285, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003776.

Abdou:2003:SMF

[Abd03] M. A. Abdou. On a symptotic methods for Fredholm– Volterra integral equation of the second kind in contact problems. Journal of Computational and Applied Mathematics, 154(2):431–446, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008622.

Amat:2003:GCI

[ABG03] S. Amat, S. Busquier, and J. M. Gutiérrez. Geometric constructions of iterative functions to solve nonlinear equations. Journal of Computational and Applied Mathematics, 157(1):197–205, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004205.

Agarwal:2002:P

[ABO02a] Ravi P. Agarwal, Martin Bohner, and Donal O’Regan. Preface. Journal of Computational and Applied Mathematics, 141(1–2):ix–x, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004538.

Agarwal:2002:TSB

[ABO02b] Ravi P. Agarwal, Martin Bohner, and Donal O’Regan. Time scale boundary value problems on infinite intervals. Journal of Computational and Applied Mathematics, 141(1–2):27–34, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004332.

Agarwal:2002:DET

[ABOP02] Ravi Agarwal, Martin Bohner, Donal O’Regan, and Allan Peterson. Dynamic equations on time scales: a survey. Journal of Computational and Applied Mathematics, 141(1–2):1–26, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004320.

Ammar:2001:PZB

[ACGR01] G. S. Ammar, D. Calvetti, W. B. Gragg, and L. Reichel. Polynomial zerofinders based on Szegő polynomials. Journal of Computational and Applied Mathematics, 127(1–2):1–16, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000491X.

Acosta:2001:LAI

[Aco01] Gabriel Acosta. Lagrange and average interpolation over 3D anisotropic elements. Journal of Computational and Applied Mathematics, 135(1):91–109, October 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005641.

Arvesu:2003:SDM

[ACV03] J. Arvesú, J. Coussement, and W. Van Assche. Some discrete multiple orthogonal polynomials. Journal of Computational and Applied Mathematics, 153(1–2):19–45, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005976.

Alefeld:2003:P

[ACY03] Götz Alefeld, Xiaojun Chen, and Tetsuro Yamamoto. Preface. Journal of Computational and Applied Mathematics, 152(1–2):xi–xii, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006921.

Angelini:2000:FFA

[AD00] Claudia Angelini and Daniela De Canditiis. Fourier frequency adaptive regularization for smoothing data. Journal of Computational and Applied Mathematics, 115(1–2):35–50, March 1, 2000. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001144.

Abou-Dina:2001:NTG

[AD01] M. S. Abou-Dina. Nonlinear transient gravity waves due to an initial free-surface elevation over a topography. Journal of Computational and Applied Mathematics, 130(1–2):173–195, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003842.

Aimi:2002:HKI

[AD02] A. Aimi and M. Diligenti. Hypersingular kernel integration in 3D Galerkin boundary element method. Journal of Computational and Applied Mathematics, 138(1):51–72, January 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003636.

Azizi:2000:PEM

[ADEL00] M. Azizi, E. M. Daoudi, R. El Hani, and A. Lakhouaja. Parallel electromagnetic modelling for the nonlinear electrotechnic systems. Journal of Computational and Applied Mathematics, 115(1–2):51–61, March 1, 2000. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003076.

Abou-Dina:2003:BIF

[ADG03] M. S. Abou-Dina and A. F. Ghaleb. On the boundary integral formulation of the plane theory of elasticity (computational aspects). Journal of Computational and Applied Mathematics, 159(2):285–317, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005351.

Abou-Dina:2004:VTM

[ADG04] M. S. Abou-Dina and A. F. Ghaleb. A variant of Trefftz’s method by boundary Fourier expansion for solving regular and singular plane boundary-value problems. Journal of Computational and Applied Mathematics, 167(2):363–387, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009269.

Ahues:2002:RPA

[ADL^{+}02] M. Ahues, F. D’Almeida, A. Largillier, O. Titaud, and P. Vasconcelos. An L^{1} refined projection approximate solution of the radiation transfer equation in stellar atmospheres. Journal of Computational and Applied Mathematics, 140(1–2):13–26, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004034.

Alfaro:2002:OJP

[AdMR02] Manuel Alfaro, María Álvarez de Morales, and M. Luisa Rezola. Orthogonality of the Jacobi polynomials with negative integer parameters. Journal of Computational and Applied Mathematics, 145(2):379–386, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005891.

Aladl:2002:NBC

[ADR02] Usaf E. Aladl, A. S. Deakin, and H. Rasmussen. Nonreflecting boundary condition for the wave equation. Journal of Computational and Applied Mathematics, 138(2):309–323, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003727.

Algaba:2003:QHN

[AFGG03] A. Algaba, E. Freire, E. Gamero, and C. García. Quasihomogeneous normal forms. Journal of Computational and Applied Mathematics, 150(1):193–216, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200660X.

Allgower:2000:PLM

[AG00a] Eugene L. Allgower and Kurt Georg. Piecewise linear methods for nonlinear equations and optimization. Journal of Computational and Applied Mathematics, 124 (1–2):245–261, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004271.

Antohe:2000:PTM

[AG00b] V. Antohe and I. Gladwell. Performance of two methods for solving separable Hamiltonian systems. Journal of Computational and Applied Mathematics, 125 (1–2):83–92, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000460X.

Ang:2001:BIM

[AG01a] W. T. Ang and A. B.

Gumel. A boundary integral method for the threedimensional heat equation subject to specification of energy. Journal of Computational and Applied Mathematics, 135(2):303–311, October 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005896.

Atici:2001:EPS

[AG01b] F. Merdivenci Atici and G. Sh.

Guseinov. On the existence of positive solutions for nonlinear differential equations with periodic boundary conditions. Journal of Computational and Applied Mathematics, 132(2):341–356, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004386.

Atici:2002:GFP

[AG02a] F. Merdivenci Atici and G. Sh.

Guseinov. On Green’s functions and positive solutions for boundary value problems on time scales. Journal of Computational and Applied Mathematics, 141(1–2):75–99, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270100437X.

Azulay:2002:WPS

[AG02b] David-Olivier Azulay and Ian Gambini. Which pivot to solve linear systems? Journal of Computational and Applied Mathematics, 147 (2):471–484, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200482X.

Ahamadi:2003:SNM

[AG03a] Malidi Ahamadi and Jean-Jacques Gervais. Symbolicnumerical methods for the computation of normal forms of PDEs. Journal of Computational and Applied Mathematics, 158(2):443–472, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042703004825.

Axelsson:2003:CML

[AG03b] O. Axelsson and S. V. Gololobov. A combined method of local Green’s functions and central difference method for singularly perturbed convection– diffusion problems. Journal of Computational and Applied Mathematics, 161 (2):245–257, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300671X.

Agarwal:2003:AEI

[Aga03] A. K. Agarwal. An analogue of Euler’s identity and new combinatorial properties of n-colour compositions. Journal of Computational and Applied Mathematics, 160(1–2):9–15, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006095.

Alonso:2000:LAD

[AGC00] Sergio Alonso and Concepción González-Concepción. Linear approximation in the discrete frequency domain. Journal of Computational and Applied Mathematics, 126 (1–2):255–267, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900357X.

Ageel:2002:NME

[Age02] M. I. Ageel. A novel means of estimating quantiles for 2-parameter Weibull distribution under the right random censoring model. Journal of Computational and Applied Mathematics, 149 (2):373–380, December 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004788.

Area:2000:IPI

[AGMMB00a] I. Area, E. Godoy, F. Marcellán, and J. J. Moreno-Balcázar. Inner products involving q-differences: the little q-Laguerre–Sobolev polynomials. Journal of Computational and Applied Mathematics, 118(1–2):1–22, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002788.

Area:2000:RPR

[AGMMB00b] I. Area, E. Godoy, F. Marcellán, and J. J. Moreno-Balcázar. Ratio and Plancherel– Rotach asymptotics for Meixner– Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 116(1):63–75, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002812.

Area:2001:SCL

[AGRZ01] I. Area, E. Godoy, A. Ronveaux, and A. Zarzo. Solving connection and linearization problems within the Askey scheme and its qanalogue via inversion formulas. Journal of Computational and Applied Mathematics, 133(1–2):151–162, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006403.

Area:2003:HTD

[AGRZ03] I. Area, E. Godoy, A. Ronveaux, and A. Zarzo. Hypergeometrictype differential equations: second kind solutions and related integrals. Journal of Computational and Applied Mathematics, 157(1): 93–106, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703003777.

Hyodo:2002:MSF

[aH02] Shi aki Hyodo. Meso-scale fusion: a method for molecular electronic state calculation in inhomogeneous materials. Journal of Computational and Applied Mathematics, 149(1):101–118, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200523X.

Ansari:2003:NSC

[AH03] Ali R. Ansari and Alan F. Hegarty. Numerical solution of a convection diffusion problem with Robin boundary conditions. Journal of Computational and Applied Mathematics, 156(1):221–238, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702009135.

Alanelli:2004:BGE

[AH04] Maria Alanelli and Apostolos Hadjidimos. Block Gauss elimination followed by a classical iterative method for the solution of linear systems. Journal of Computational and Applied Mathematics, 163(2):381–400, February 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007623.

Anderson:2004:RGU

[AHP04] J. M. Anderson, W. K. Hayman, and Ch. Pommerenke. The radial growth of univalent functions. Journal of Computational and Applied Mathematics, 171(1– 2):27–37, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704000524.

Askey:2001:TCH

[AIV01] R. Askey, M. E. H. Ismail, and W. Van Assche. Ted chihara and his work on orthogonal polynomials. Journal of Computational and Applied Mathematics, 133(1–2):1–11, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006312.

Almenar:2001:MPS

[AJ01] Pedro Almenar and Lucas Jódar. Mixed problems for separated variable coefficient wave equations: analyticnumerical solutions with a priori error bounds. Journal of Computational and Applied Mathematics, 134 (1–2):301–323, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005562.

Al-Jarrah:2002:GSB

[AJDG02] A. Al-Jarrah, K. M. Dempsey, and M. L. Glasser. Generalized series of Bessel functions. Journal of Computational and Applied Mathematics, 143(1):1–8, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005052.

Andres:2000:BSG

[AK00] Jan Andres and Bohumil Krajc. Bounded solutions in a given set of differential systems. Journal of Computational and Applied Mathematics, 113(1–2):73–82, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002459.

Al-Khaled:2001:NSF

[AK01a] Kamel Al-Khaled. Numerical study of Fisher’s reaction– diffusion equation by the sinc collocation method. Journal of Computational and Applied Mathematics, 137 (2):245–255, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003569.

Al-Khaled:2001:SNS

[AK01b]  Kamel Al-Khaled.  Sinc

numerical solution for solitons and solitary waves. Journal of Computational and Applied Mathematics, 130(1–2):283–292, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042799003763.

Aziz:2001:FOF

[AK01c] Tariq Aziz and Manoj Kumar. A fourth-order finitedifference method based on non-uniform mesh for a class of singular two-point boundary value problems. Journal of Computational and Applied Mathematics, 136 (1–2):337–342, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006245.

Aziz:2002:SMS

[AK02] Tariq Aziz and Arshad Khan.

A spline method for secondorder singularly perturbed boundary-value problems. Journal of Computational and Applied Mathematics, 147 (2):445–452, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200479X.

Akkouchi:2002:RWK

[Akk02] Mohamed Akkouchi. On a result of W. A. Kirk and L. M. Saliga. Journal of Computational and Applied Mathematics, 142(2):445–448, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005441.  See

[KS00].

Alefeld:2003:SSP

[AKM03] Götz Alefeld, Vladik Kreinovich, and Günter Mayer. On the solution sets of particular classes of linear interval systems. Journal of Computational and Applied Mathematics, 152(1–2):1–15, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006933.

Alboul:2000:BDD

[AKTvD00] Lyuba Alboul, Gertjan Kloosterman, Cornelis Traas, and Ruud van Damme. Best data-dependent triangulations. Journal of Computational and Applied Mathematics, 119 (1–2):1–12, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270000368X.

Almeida:2003:WDT

[AL03] José B. Almeida and Vasudevan Lakshminarayanan. Wigner distribution transformations in high-order systems. Journal of Computational and Applied Mathematics, 160(1–2):17–26, November 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006101.

Alfeld:2000:BSS

[Alf00] Peter Alfeld. Bivariate spline spaces and minimal determining sets. Journal of Computational and Applied Mathematics, 119(1–2):13–27, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003691.

Anguelov:2003:QSF

[ALM03] Roumen Anguelov, Jean M.- S. Lubuma, and Simon K. Mahudu. Qualitatively stable finite difference schemes for advection-reaction equations. Journal of Computational and Applied Mathematics, 158(1):19–30, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004680.

Alzer:2002:ICL

[Alz02]  Horst Alzer.  Inequalities

for the constants of Landau and Lebesgue. Journal of Computational and Applied Mathematics, 139 (2):215–230, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004265.

Alzer:2003:GHM

[Alz03] Horst Alzer. On Gautschi’s harmonic mean inequality for the gamma function. Journal of Computational and Applied Mathematics, 157(1):243–249, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004564.

Alefeld:2000:IAT

[AM00] Götz Alefeld and Günter Mayer. Interval analysis: theory and applications. Journal of Computational and Applied Mathematics, 121 (1–2):421–464, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003423.

Ahlbrandt:2002:PDE

[AM02a] Calvin D. Ahlbrandt and Christina Morian. Partial differential equations on time scales. Journal of Computational and Applied Mathematics, 141(1–2):35–55, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042701004344.

Akbarfam:2002:HOA

[AM02b] A. Jodayree Akbarfam and Angelo B. Mingarelli. Higher order asymptotic distribution of the eigenvalues of nondefinite Sturm–Liouville problems with one turning point. Journal of Computational and Applied Mathematics, 149(2):423–437, December 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004879.

Akesbi:2003:TNA

[AM03] S. Akesbi and E. Maitre. Theoretical and numerical analysis of a minimal residual solver for 2D Boltzmann transport equation. Journal of Computational and Applied Mathematics, 150(2):357–374, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006702.

Alfaro:2001:ASO

[AMBP^{+}01] Manuel Alfaro, Juan J.

Moreno-Balcázar, Teresa E. Pérez, Miguel A. Piñar, and M. Luisa Rezola. Asymptotics of Sobolev orthogonal polynomials for Hermite coherent pairs. Journal of Computational and Applied Mathematics, 133(1–2):141–150, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006397.

Avdonin:2000:IPC

[AMS00] S. A. Avdonin, N. G. Medhin, and T. L. Sheronova. Identification of a piecewise constant coefficient in the beam equation. Journal of Computational and Applied Mathematics, 114(1):11–21, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704279900285X.

Alvarez-Nodarse:2003:MTS

[ÁNAA03] Renato Álvarez-Nodarse, M. K. Atakishiyeva, and N. M. Atakishiyev. Mellin transforms for some families of qpolynomials. Journal of Computational and Applied Mathematics, 153(1–2):9–18, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006386.

Andrew:2000:ACN

[And00] Alan L. Andrew. Asymptotic correction of Numerov’s eigenvalue estimates with natural boundary conditions. Journal of Computational and Applied Mathematics, 125 (1–2):359–366, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004799.

Anderson:2002:EIT

[And02] Douglas R. Anderson. Eigenvalue intervals for a twopoint boundary value problem on a measure chain. Journal of Computational and Applied Mathematics, 141 (1–2):57–64, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004356.

Anil:2002:TCS

[Ani02]  M. Anil.  Toyota confer-

ence: Sponsored by Toyota Motor Corporation. Journal of Computational and Applied Mathematics, 149(1): xv–xvi, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702006118.

Alvarez-Nodarse:2001:CPA

[ÁNM01] R. Álvarez-Nodarse and J. C.

Medem. q-Classical polynomials and the q-Askey and Nikiforov–Uvarov tableaus. Journal of Computational and Applied Mathematics, 135(2):197–223, October 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005859.

Anonymous:2000:AIV

[Ano00a] Anonymous. Author index to volume 113 (2000). Journal of Computational and Applied Mathematics, 113 (1–2):411–412, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003519.

Anonymous:2000:Ca

[Ano00b]  Anonymous.  Conference.

Journal of Computational and Applied Mathematics, 116(2):N1–N18, April 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003149.

Anonymous:2000:Cb

[Ano00c]  Anonymous.  Conference.

Journal of Computational and Applied Mathematics, 126 (1–2):N1–N9, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006919.

Anonymous:2000:Ea

[Ano00d] Anonymous. Editorial. Journal of Computational and Applied Mathematics, 114 (1):ix–x, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799002836.

Anonymous:2000:Eb

[Ano00e] Anonymous. Editorial. Journal of Computational and Applied Mathematics, 115 (1–2):xi, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799003118.

Anonymous:2000:Ec

[Ano00f] Anonymous. Editorial. Journal of Computational and Applied Mathematics, 118 (1–2):ix–x, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700002764.

Anonymous:2000:Ed

[Ano00g] Anonymous. Editorial. Journal of Computational and Applied Mathematics, 120 (1–2):ix–x, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700002995.

Anonymous:2000:Ia

[Ano00h] Anonymous. Index. Journal of Computational and Applied Mathematics, 114 (2):411–412, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003672.

Anonymous:2000:Ib

[Ano00i] Anonymous. Index. Journal of Computational and Applied Mathematics, 115(1– 2):617–618, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042799003957.

Anonymous:2000:Ic

[Ano00j] Anonymous. Index. Journal of Computational and Applied Mathematics, 116(2): 279–280, April 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700003137.

Anonymous:2000:Id

[Ano00k] Anonymous. Index. Journal of Computational and Applied Mathematics, 117(2):183, May 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042700003514.

Anonymous:2000:Ie

[Ano00l] Anonymous. Index. Journal of Computational and Applied Mathematics, 118(1– 2):363–364, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700004155.

Anonymous:2000:If

[Ano00m] Anonymous. Index. Journal of Computational and Applied Mathematics, 119(1– 2):391–392, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700004891.

Anonymous:2000:Ig

[Ano00n] Anonymous. Index. Journal of Computational and Applied Mathematics, 120 (1–2):215, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005240.

Anonymous:2000:Ih

[Ano00o] Anonymous. Index. Journal of Computational and Applied Mathematics, 120 (1–2):217–233, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005252.

Anonymous:2000:Ii

[Ano00p] Anonymous. Index. Journal of Computational and Applied Mathematics, 120 (1–2):235–253, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005264.

Anonymous:2000:Ij

[Ano00q] Anonymous. Index. Journal of Computational and Applied Mathematics, 121(1– 2):465, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700005665.

Anonymous:2000:Ik

[Ano00r] Anonymous. Index. Journal of Computational and Applied Mathematics, 122 (1–2):357, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700005677.

Anonymous:2000:Il

[Ano00s] Anonymous. Index. Journal of Computational and Applied Mathematics, 123(1– 2):531, November 1, 2000.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005689.

Anonymous:2000:Im

[Ano00t] Anonymous. Index. Journal of Computational and Applied Mathematics, 124(1– 2):373, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700006063.

Anonymous:2000:In

[Ano00u] Anonymous. Index. Journal of Computational and Applied Mathematics, 125 (1–2):539–540, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006087.

Anonymous:2000:Io

[Ano00v] Anonymous. Index. Journal of Computational and Applied Mathematics, 126 (1–2):479–480, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006907.

Anonymous:2000:NCA

[Ano00w]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 113 (1–2):N1–N14, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003520.

Anonymous:2000:NAV

[Ano00x] Anonymous. Numerical analysis 2000 vol. III: Linear algebra. Journal of Computational and Applied Mathematics, 123(1–2):ix–xii, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004532.

Anonymous:2001:AIVa

[Ano01a] Anonymous. Author index volume 127 (2001). Journal of Computational and Applied Mathematics, 127(1– 2):369, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700006932.

Anonymous:2001:AIVi

[Ano01b] Anonymous. Author index volume 134 (2001). Journal of Computational and Applied Mathematics, 134 (1–2):369–370, September 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004915.

Anonymous:2001:AIVj

[Ano01c] Anonymous. Author index volume 135 (2001). Journal of Computational and Applied Mathematics, 135 (2):335, October 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701005131.

Anonymous:2001:AIVk

[Ano01d] Anonymous. Author index volume 136 (2001). Journal of Computational and Applied Mathematics, 136 (1–2):405–406, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005143.

Anonymous:2001:AIVl

[Ano01e] Anonymous. Author index volume 137 (2001). Journal of Computational and Applied Mathematics, 137 (2):399–400, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005453.

Anonymous:2001:AIVb

[Ano01f]  Anonymous.  Authors in-

dex to volume 128 (2001). Journal of Computational and Applied Mathematics, 128 (1–2):467, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701003521.

Anonymous:2001:AIVc

[Ano01g] Anonymous. Authors index volume 129 (2001). Journal of Computational and Applied Mathematics, 129 (1–2):209, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701003740.

Anonymous:2001:AIVd

[Ano01h] Anonymous. Authors index volume 130 (2001). Journal of Computational and Applied Mathematics, 130(1– 2):387–388, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701003958.

Anonymous:2001:AIVf

[Ano01i] Anonymous. Authors index volume 131 (2001). Journal of Computational and Applied Mathematics, 131(1– 2):507–508, June 1, 2001.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004137.

Anonymous:2001:AIVg

[Ano01j] Anonymous. Authors index volume 132 (2001). Journal of Computational and Applied Mathematics, 132 (2):483–484, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004319.

Anonymous:2001:AIVh

[Ano01k] Anonymous. Authors index volume 133 (2001). Journal of Computational and Applied Mathematics, 133 (1–2):703–705, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004903.

Anonymous:2001:AIVe

[Ano01l]  Anonymous.  Authors in-

dex volumes 121–130 (2001). Journal of Computational and Applied Mathematics, 130(1–2):389–402, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100396X.

Anonymous:2001:Ca

[Ano01m] Anonymous. Contents. Journal of Computational and Applied Mathematics, 127 (1–2):vii–viii, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006920.

Anonymous:2001:Cb

[Ano01n] Anonymous. Contents. Journal of Computational and Applied Mathematics, 128(1– 2):vii–viii, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270100351X.

Anonymous:2001:FCD

[Ano01o] Anonymous. Filler contents direct. Journal of Computational and Applied Mathematics, 127(1–2):II, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006956.

Anonymous:2001:FSD

[Ano01p] Anonymous. Filler science direct. Journal of Computational and Applied Mathematics, 127(1–2):I, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006944.

Anonymous:2001:LRP

[Ano01q]  Anonymous.  List of reg-

istered participants. Journal of Computational and Applied Mathematics, 133 (1–2):xix–xxvii, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004897.

Anonymous:2001:LTP

[Ano01r] Anonymous. List of talks presented at the conference. Journal of Computational and Applied Mathematics, 133(1–2):xv–xviii, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004885.

Anonymous:2001:NAV

[Ano01s] Anonymous. Numerical analysis 2000: Vol. VII: Partial differential equations. Journal of Computational and Applied Mathematics, 128 (1–2):ix–xi, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005082.

Anonymous:2001:SIV

[Ano01t]  Anonymous.  Subject in-

dex volumes 121–130 (2001). Journal of Computational and Applied Mathematics, 130(1–2):403–420, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003971.

Anonymous:2002:USB

[Ano02a] Anonymous. A 2002 update of the supplementary bibliography on roots of polynomials. Journal of Computational and Applied Mathematics, 142(2):433–434, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005465.

Anonymous:2002:AIa

[Ano02b]  Anonymous.  Author in-

dex. Journal of Computational and Applied Mathematics, 143(2):311, June 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004715.

Anonymous:2002:AIb

[Ano02c]  Anonymous.  Author in-

dex. Journal of Computational and Applied Mathematics, 144(1–2):375–376, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004983.

Anonymous:2002:AIc

[Ano02d]  Anonymous.  Author in-

dex. Journal of Computational and Applied Mathematics, 145(2):535–536, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005496.

Anonymous:2002:AId

[Ano02e]  Anonymous.  Author in-

dex. Journal of Computational and Applied Mathematics, 146 (2):493–494, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006520.

Anonymous:2002:AIe

[Ano02f]  Anonymous.  Author in-

dex. Journal of Computational and Applied Mathematics, 149(2):489–490, December 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200804X.

Anonymous:2002:AIVh

[Ano02g] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 147(2):517–518, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006891.

Anonymous:2002:AIVi

[Ano02h] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 148(2):463–464, November 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007768.

Anonymous:2002:AIVa

[Ano02i] Anonymous. Author index volume 138. Journal of Computational and Applied Mathematics, 138(2):367–368, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270100574X.

Anonymous:2002:AIVb

[Ano02j] Anonymous. Author index volume 139. Journal of Computational and Applied Mathematics, 139(2):369–370, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005933.

Anonymous:2002:AIVe

[Ano02k] Anonymous. Author index volume 140. Journal of Computational and Applied Mathematics, 140(1–2):905–931, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003679.

Anonymous:2002:AIVc

[Ano02l] Anonymous. Author index volume 140 (2002). Journal of Computational and Applied Mathematics, 140(1– 2):877–879, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702003655.

Anonymous:2002:AIVf

[Ano02m] Anonymous. Author index volume 141. Journal of Computational and Applied Mathematics, 141(1–2):283, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003813.

Anonymous:2002:AIVg

[Ano02n] Anonymous. Author index volume 142. Journal of Computational and Applied Mathematics, 142(2):451–452, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004491.

Anonymous:2002:AIVd

[Ano02o] Anonymous. Author index volumes 131–140. Journal of Computational and Applied Mathematics, 140(1– 2):881–903, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702003667.

Anonymous:2002:CPa

[Ano02p] Anonymous. Conference participants. Journal of Computational and Applied Mathematics, 148(1):x, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007550.

Anonymous:2002:CPb

[Ano02q]  Anonymous.  Conference

photo. Journal of Computational and Applied Mathematics, 149(1):xi, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006106.

Anonymous:2002:Ea

[Ano02r] Anonymous. Editors. Journal of Computational and Applied Mathematics, 148 (2):ifc, November 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702007720.

Anonymous:2002:Eb

[Ano02s] Anonymous. Editors. Journal of Computational and Applied Mathematics, 149 (1):ii, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702007847.

Anonymous:2002:Ec

[Ano02t] Anonymous. Editors. Journal of Computational and Applied Mathematics, 149 (2):ifc, December 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702008002.

Anonymous:2002:FEa

[Ano02u]  Anonymous.  FM editors.

Journal of Computational and Applied Mathematics, 147(1):ii, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702006787.

Anonymous:2002:FEb

[Ano02v]  Anonymous.  FM editors.

Journal of Computational and Applied Mathematics, 148 (1):ii, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702007604.

Anonymous:2002:IEB

[Ano02w] Anonymous. IFC ed. board 2nd issue. Journal of Computational and Applied Mathematics, 146(2):ifc, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006477.

Anonymous:2002:IEI

[Ano02x] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 147(2):ifc, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006854.

Anonymous:2002:OCT

[Ano02y] Anonymous. Organizing Committee of the 15th Toyota Conference. Journal of Computational and Applied Mathematics, 149(1):xiii–xiv, December 1, 2002. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200612X.

Anonymous:2003:AI

[Ano03a]  Anonymous.  Author in-

dex. Journal of Computational and Applied Mathematics, 151(2):477–478, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300356X.

Anonymous:2003:AIVc

[Ano03b] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 152(1–2):613–614, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003698.

Anonymous:2003:AIVd

[Ano03c] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 153(1–2):549–550, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004035.

Anonymous:2003:AIVf

[Ano03d] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 155(2):461–462, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300517X.

Anonymous:2003:AIVi

[Ano03e]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 158(2):507–508, September 15, 2003. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042703006770.

Anonymous:2003:AIVj

[Ano03f] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 159(2):473–474, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007489.

Anonymous:2003:AIVa

[Ano03g] Anonymous. Author index volume 150. Journal of Computational and Applied Mathematics, 150(2):397, January 15, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008427.

Anonymous:2003:AIVe

[Ano03h] Anonymous. Author index volume 154 (2003). Journal of Computational and Applied Mathematics, 154 (2):477–478, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703004448.

Anonymous:2003:AIVg

[Ano03i] Anonymous. Author index volume 156. Journal of Computational and Applied Mathematics, 156(2):501–502, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005806.

Anonymous:2003:AIVh

[Ano03j] Anonymous. Author index volume 157. Journal of Computational and Applied Mathematics, 157(2):511–512, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006022.

Anonymous:2003:AIVk

[Ano03k]  Anonymous.  Author in-

dex volume 160. Journal of Computational and Applied Mathematics, 160 (1–2):335–336, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007817.

Anonymous:2003:AIVm

[Ano03l] Anonymous. Author index volume 161. Journal of Computational and Applied Mathematics, 161(2):497–498, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008872.

Anonymous:2003:AIVb

[Ano03m]  Anonymous.  Author in-

dex volumes 141–150. Journal of Computational and Applied Mathematics, 150 (2):399–418, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008439.

Anonymous:2003:AIVl

[Ano03n]  Anonymous.  Author in-

dex volumes 151–160. Journal of Computational and Applied Mathematics, 160 (1–2):337–360, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007829.

Anonymous:2003:Ea

[Ano03o] Anonymous. Editors. Journal of Computational and Applied Mathematics, 150 (1):ii, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702008312.

Anonymous:2003:Eb

[Ano03p] Anonymous. Editors. Journal of Computational and Applied Mathematics, 150 (2):ifc, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702008385.

Anonymous:2003:Ec

[Ano03q] Anonymous. Editors. Journal of Computational and Applied Mathematics, 151 (1):ii, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702008506.

Anonymous:2003:Ed

[Ano03r] Anonymous. Editors. Journal of Computational and Applied Mathematics, 151 (2):ifc, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703003522.

Anonymous:2003:Ee

[Ano03s] Anonymous. Editors. Journal of Computational and Applied Mathematics, 152 (1–2):ii, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703003649.

Anonymous:2003:Ef

[Ano03t] Anonymous. Editors. Journal of Computational and Applied Mathematics, 154(2):ifc, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004400.

Anonymous:2003:Eg

[Ano03u] Anonymous. Editors. Journal of Computational and Applied Mathematics, 155(1):ii, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005041.

Anonymous:2003:Eh

[Ano03v] Anonymous. Editors. Journal of Computational and Applied Mathematics, 155(2):ifc, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005144.

Anonymous:2003:Ei

[Ano03w] Anonymous. Editors. Journal of Computational and Applied Mathematics, 156(1):ii, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005260.

Anonymous:2003:Ej

[Ano03x] Anonymous. Editors. Journal of Computational and Applied Mathematics, 156(2):ifc, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005764.

Anonymous:2003:Ek

[Ano03y] Anonymous. Editors. Journal of Computational and Applied Mathematics, 157(1): ii, August 1, 2003. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042703005909.

Anonymous:2003:El

[Ano03z] Anonymous. Editors. Journal of Computational and Applied Mathematics, 157 (2):ifc, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703005983.

Anonymous:2003:Em

[Ano03-27] Anonymous. Editors. Journal of Computational and Applied Mathematics, 158 (1):ii, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703006575.

Anonymous:2003:En

[Ano03-28] Anonymous. Editors. Journal of Computational and Applied Mathematics, 158 (2):ifc, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703006733.

Anonymous:2003:Eo

[Ano03-29] Anonymous. Editors. Journal of Computational and Applied Mathematics, 159 (1):ii, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703006836.

Anonymous:2003:Ep

[Ano03-30] Anonymous. Editors. Journal of Computational and Applied Mathematics, 159 (2):ifc, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703007441.

Anonymous:2003:Eq

[Ano03-31] Anonymous. Editors. Journal of Computational and Applied Mathematics, 160 (1–2):ii, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703007763.

Anonymous:2003:Er

[Ano03-32] Anonymous. Editors. Journal of Computational and Applied Mathematics, 161 (1):ii, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703008422.

Anonymous:2003:Es

[Ano03-33] Anonymous. Editors. Journal of Computational and Applied Mathematics, 161 (2):ifc, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703008835.

Anonymous:2003:FEa

[Ano03-34] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 153(1–2):ii, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703003984.

Anonymous:2003:FEb

[Ano03-35] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 154(1): ii, May 1, 2003. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042703004266.

Anonymous:2003:SIVb

[Ano03-36] Anonymous. Subject index volume 151–160. Journal of Computational and Applied Mathematics, 160 (1–2):361–387, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007830.

Anonymous:2003:SIVa

[Ano03-37] Anonymous. Subject index volumes 141–150. Journal of Computational and Applied Mathematics, 150 (2):419–440, January 15, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008440.

Anonymous:2004:AIVa

[Ano04a] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 162(2):497–498, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009191.

Anonymous:2004:AIVb

[Ano04b] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 163(2):465–466, February 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000251.

Anonymous:2004:AIVd

[Ano04c] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 166(2):597–598, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001062.

Anonymous:2004:AIVe

[Ano04d] Anonymous. Author index volume 167. Journal of Computational and Applied Mathematics, 167(2):499–500, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001980.

Anonymous:2004:AIVf

[Ano04e] Anonymous. Author index volume 168. Journal of Computational and Applied Mathematics, 168(1–2):529–531, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704002110.

Anonymous:2004:AIVg

[Ano04f] Anonymous. Author index volume 169. Journal of Computational and Applied Mathematics, 169(2):475–476, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704002729.

Anonymous:2004:AIVh

[Ano04g]  Anonymous.  Author in-

dex volume 170. Journal of Computational and Applied Mathematics, 170 (2):467–468, September 15, 2004. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704002936.

Anonymous:2004:AIVj

[Ano04h] Anonymous. Author index volume 171. Journal of Computational and Applied Mathematics, 171(1–2):425–426, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704003061.

Anonymous:2004:AIVk

[Ano04i] Anonymous. Author index volume 172 (2004). Journal of Computational and Applied Mathematics, 172 (2):409–410, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704004157.

Anonymous:2004:AIVi

[Ano04j]  Anonymous.  Author in-

dex volumes 161–170. Journal of Computational and Applied Mathematics, 170 (2):469–493, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704002948.

Anonymous:2004:AIVc

[Ano04k]  Anonymous.  Author in-

dex volumes 164–165. Journal of Computational and Applied Mathematics, 164– 165(??):815–817, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000378.

Anonymous:2004:Ea

[Ano04l] Anonymous. Editors. Journal of Computational and Applied Mathematics, 162 (1):ii, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270300904X.

Anonymous:2004:Eb

[Ano04m] Anonymous. Editors. Journal of Computational and Applied Mathematics, 162 (2):ifc, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703009154.

Anonymous:2004:Ec

[Ano04n] Anonymous. Editors. Journal of Computational and Applied Mathematics, 163 (1):ii, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703010124.

Anonymous:2004:Ed

[Ano04o] Anonymous. Editors. Journal of Computational and Applied Mathematics, 163 (2):ifc, February 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704000214.

Anonymous:2004:Ee

[Ano04p] Anonymous. Editors. Journal of Computational and Applied Mathematics, 164– 165(??):ii, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704000329.

Anonymous:2004:Ef

[Ano04q] Anonymous. Editors. Journal of Computational and Applied Mathematics, 166(1):ii, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000780.

Anonymous:2004:Eg

[Ano04r] Anonymous. Editors. Journal of Computational and Applied Mathematics, 166(2):ifc, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001025.

Anonymous:2004:Eh

[Ano04s] Anonymous. Editors. Journal of Computational and Applied Mathematics, 167(1):ii, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001785.

Anonymous:2004:Ei

[Ano04t] Anonymous. Editors. Journal of Computational and Applied Mathematics, 167(2):ifc, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001943.

Anonymous:2004:Ej

[Ano04u] Anonymous. Editors. Journal of Computational and Applied Mathematics, 168(1–2):ii, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704002067.

Anonymous:2004:Ek

[Ano04v] Anonymous. Editors. Journal of Computational and Applied Mathematics, 169(1): ii, August 1, 2004. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042704002596.

Anonymous:2004:El

[Ano04w] Anonymous. Editors. Journal of Computational and Applied Mathematics, 169 (2):ifc, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002687.

Anonymous:2004:Em

[Ano04x] Anonymous. Editors. Journal of Computational and Applied Mathematics, 170 (1):ii, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002808.

Anonymous:2004:En

[Ano04y] Anonymous. Editors. Journal of Computational and Applied Mathematics, 170 (2):ifc, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002894.

Anonymous:2004:Eo

[Ano04z] Anonymous. Editors. Journal of Computational and Applied Mathematics, 171 (1–2):ii, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003012.

Anonymous:2004:Eq

[Ano04-27] Anonymous. Editors. Journal of Computational and Applied Mathematics, 172 (1):ii, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400353X.

Anonymous:2004:Er

[Ano04-28] Anonymous. Editors. Journal of Computational and Applied Mathematics, 172 (2):ifc, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400411X.

Anonymous:2004:SIV

[Ano04-29] Anonymous. Subject index volumes 161–170. Journal of Computational and Applied Mathematics, 170 (2):495–522, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270400295X.

Anttila:2002:SCM

[Ant02] Juha Anttila. A spline collocation method for parabolic pseudodifferential equations. Journal of Computational and Applied Mathematics, 140(1–2):41–61, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004010.

Agarwal:2000:E

[AO00a] Ravi P. Agarwal and Donal O’Regan. Editorial. Journal of Computational and Applied Mathematics, 113 (1–2):ix, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799002381.

Agarwal:2000:ECO

[AO00b] Ravi P. Agarwal and Donal O’Regan. Existence criteria for operator inclusions in abstract spaces. Journal of Computational and Applied Mathematics, 113(1–2):183–193, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002526.

Agarwal:2000:SNR

[AO00c] Ravi P. Agarwal and Donal O’Regan. Some new results for singular problems with sign changing nonlinearities. Journal of Computational and Applied Mathematics, 113(1–2):1–15, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002393.

Amini:2000:ATE

[AP00] Sia Amini and Anthony Profit. Analysis of the truncation errors in the fast multipole method for scattering problems. Journal of Computational and Applied Mathematics, 115(1–2):23–33, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001752.

Ainsworth:2002:HMF

[AP02a] Mark Ainsworth and Katia Pinchedez. The hp-MITC finite element method for the Reissner–Mindlin plate problem. Journal of Computational and Applied Mathematics, 148(2):429–462, November 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005605.

Aulbach:2002:RLD

[AP02b] Bernd Aulbach and Christian Pötzsche. Reducibility of linear dynamic equations on measure chains. Journal of Computational and Applied Mathematics, 141(1–2):101–115, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004381.

Albrecher:2004:AOP

[AP04a] Hansjörg Albrecher and Martin Predota. On Asian option pricing for NIG Lévy processes. Journal of Computational and Applied Mathematics, 172(1):153–168, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270400113X.

Amrein:2004:OGW

[AP04b] W. O. Amrein and D. B. Pearson. M operators: a generalisation of Weyl–Titchmarsh theory. Journal of Computational and Applied Mathematics, 171(1–2):1–26, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000603.

Aizawa:2002:CMM

[APT02] Tatstuhiko Aizawa, Yunan Prawoto, and Fujio Tsumori. Coupled, macro-micro modeling for hot deformation and sintering. Journal of Computational and Applied Mathematics, 149(1):307–324, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005381.

Alzer:2004:MTI

[AQ04] Horst Alzer and Song-Liang Qiu. Monotonicity theorems and inequalities for the complete elliptic integrals. Journal of Computational and Applied Mathematics, 172(2):289–312, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001207.

Algaba:2003:CCC

[AR03] A. Algaba and M. Reyes. Computing center conditions for vector fields with constant angular speed. Journal of Computational and Applied Mathematics, 154(1):143–159, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200818X.

Al-Refai:2004:EUB

[AR04] Mohammed Al-Refai. Existence, uniqueness and bounds for a problem in combustion theory. Journal of Computational and Applied Mathematics, 167(2):255–269, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009130.

Araaya:2004:MPP

[Ara04] Tsehaye K. Araaya. The Meixner–Pollaczek polynomials and a system of orthogonal polynomials in a strip. Journal of Computational and Applied Mathematics, 170 (2):241–254, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000688.

Argyros:2001:NSC

[Arg01a] I. K. Argyros. A new semilocal convergence theorem for Newton’s method in Banach space using hypotheses on the second Fréchet-derivative. Journal of Computational and Applied Mathematics, 130(1–2):369–373, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003307.

Argyros:2001:NKT

[Arg01b]  Ioannis K. Argyros.  A

Newton–Kantorovich theorem for equations involving m-Fréchet differentiable operators and applications in radiative transfer. Journal of Computational and Applied Mathematics, 131(1– 2):149–159, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700003174.

Argyros:2003:IEA

[Arg03a]  Ioannis K. Argyros.  An

improved error analysis for Newton-like methods under generalized conditions. Journal of Computational and Applied Mathematics, 157(1):169–185, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300390X.

Argyros:2003:TVK

[Arg03b]  Ioannis K. Argyros.  On

a theorem of L. V. Kantorovich concerning Newton’s method. Journal of Computational and Applied Mathematics, 155(2):223–230, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008658.

Argyros:2004:CWV

[Arg04a] Ioannis K. Argyros. On the comparison of a weak variant of the Newton–Kantorovich and Miranda theorems. Journal of Computational and Applied Mathematics, 166(2): 585–589, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703009233.

Argyros:2004:NKH

[Arg04b]  Ioannis K. Argyros.  On

the Newton–Kantorovich hypothesis for solving equations. Journal of Computational and Applied Mathematics, 169(2):315–332, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000937.

Ariel:2003:FEV

[Ari03] P. Donald Ariel. On the flow of an elastico-viscous fluid near a rotating disk. Journal of Computational and Applied Mathematics, 154(1):1–25, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007446.

Alonso:2000:PEE

[ARV00] A. Alonso, A. Dello Russo, and V. Vampa. A posteriori error estimates in finite element acoustic analysis. Journal of Computational and Applied Mathematics, 117(2):105–119, May 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003350.

Abad:2004:BFA

[AS04a] Pilar Abad and José P. Suárez. Bi-factorial analysis for resolution of GPS equations. Journal of Computational and Applied Mathematics, 164–165(??):1–10, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009476.

Aliu:2004:OLG

[AS04b] E. Aliu and J. J. Shepherd. Overlapping layers in a gas lubrication problem. Journal of Computational and Applied Mathematics, 166(1):1–12, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008574.

Ameur:2004:QSW

[AS04c] El Bachir Ameur and Driss Sbibih. Quadratic spline wavelets with arbitrary simple knots on the sphere. Journal of Computational and Applied Mathematics, 162(1):273–286, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007076.

Al-Shemas:2004:IMG

[ASB04] Eman Al-Shemas and Stephen C.

Billups. An iterative method for generalized set-valued nonlinear mixed quasi-variational inequalities. Journal of Computational and Applied Mathematics, 170(2):423–432, September 15, 2004. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042704000913.

Ahn:2004:ACA

[AsKS04] Young Joon Ahn, Yeon soo Kim, and Youngsuk Shin. Approximation of circular arcs and offset curves by Bézier curves of high degree. Journal of Computational and Applied Mathematics, 167(2):405–416, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009294.

Al-Said:2002:QSM

[ASN02] Eisa A. Al-Said and Muhammad Aslam Noor. Quartic spline method for solving fourth order obstacle boundary value problems. Journal of Computational and Applied Mathematics, 143 (1):107–116, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701004976.

Ali:2000:ABH

[AT00] Ismail Ali and Vu Kim Tuan. Application of basic hypergeometric series to stable analytic continuation. Journal of Computational and Applied Mathematics, 118(1–2):193–202, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002776.

Arnautovski-Toseva:2004:EAH

[ATG04] Vesna Arnautovski-Toseva and Leonid Grcev. Electromagnetic analysis of horizontal wire in two-layered soil. Journal of Computational and Applied Mathematics, 168(1–2):21–29, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009609.

Avalos:2000:PPO

[Ava00]  George Avalos.  Pointwise

pressure observations of a canonical structural acoustics model. Journal of Computational and Applied Mathematics, 114(1):121–135, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002927.

Anuchina:2004:NSR

[AVG^{+}04] N. N. Anuchina, V. I. Volkov, V. A. Gordeychuk, N. S. Es’kov, O. S. Ilyutina, and O. M. Kozyrev. Numerical simulations of Rayleigh–Taylor and Richtmyer–Meshkov instability using MAH-3 code. Journal of Computational and Applied Mathematics, 168 (1–2):11–20, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703009592.

Alvarez-Vazquez:2002:WTP

[AVMRVM02] L. J. Alvarez-Vázquez, A. Martínez, C. Rodríguez, and M. E. Vázquez-Méndez. A wastewater treatment problem: study of the numerical convergence. Journal of Computational and Applied Mathematics, 140(1–2):27–39, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003995.

Alvarez-Vazquez:2000:AMN

[AVV00] L. J. Alvarez-Vázquez and J. M. Viaño. Asymptotic modelling of a nonsymmetric beam. Journal of Computational and Applied Mathematics, 126 (1–2):433–447, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003702.

Atieg:2003:CMF

[AW03] A. Atieg and G. A. Watson. A class of methods for fitting a curve or surface to data by minimizing the sum of squares of orthogonal distances. Journal of Computational and Applied Mathematics, 158 (2):277–296, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004485.

Ayachour:2003:FIG

[Aya03]  E. H. Ayachour.  A fast

implementation for GMRES method. Journal of Computational and Applied Mathematics, 159(2):269–283, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300534X.

Ying:2003:FDM

[aYtZ03] Lung an Ying and Xin ting Zhang. Finite difference method for detonation waves. Journal of Computational and Applied Mathematics, 159(1):185–193, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005582.

Addou:2004:SUO

[AZ04] A. Addou and J. Zahi. On the solvability of a unilateral obstacle problem on the boundary. Journal of Computational and Applied Mathematics, 162(1):193–199, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007295.

Aggoune:2004:HOP

[AZPF04] Wassila Aggoune, Hamid Zahrouni, and Michel Potier-Ferry. High-order prediction– correction algorithms for unilateral contact problems. Journal of Computational and Applied Mathematics, 168 (1–2):1–9, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703009580.

Barakat:2000:CMO

[BA00] H. M. Barakat and Y. H. Abdelkader. Computing the moments of order statistics from nonidentically distributed Weibull variables. Journal of Computational and Applied Mathematics, 117 (1):85–90, May 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704279900343X.

Backhouse:2001:RES

[Bac01] N. B. Backhouse. Resonant equations and special functions. Journal of Computational and Applied Mathematics, 133(1–2):163–169, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006415.

Bacaer:2004:NAG

[Bac04]  Nicolas Bacaër.  Numeri-

cal analysis of generalised max-plus eigenvalue problems. Journal of Computational and Applied Mathematics, 163(1):79–89, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007921.

Badr:2001:IDE

[Bad01] A. A. Badr. Integro-differential equation with Cauchy kernel. Journal of Computational and Applied Mathematics, 134 (1–2):191–199, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005367.

Baines:2001:MFE

[Bai01] M. J. Baines. Moving finite element, least squares, and finite volume approximations of steady and time-dependent PDEs in multidimensions. Journal of Computational and Applied Mathematics, 128(1–2):363–381, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005197.

Bai:2003:CAM

[Bai03a] Zhong-Zhi Bai. On the convergence of additive and multiplicative splitting iterations for systems of linear equations. Journal of Computational and Applied Mathematics, 154(1):195–214, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008221.

Bai:2003:CPN

[Bai03b] Zhong-Zhi Bai. On the convergence of parallel nonstationary multisplitting iteration methods. Journal of Computational and Applied Mathematics, 159(1):1–11, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005557.

Baker:2000:PNT

[Bak00a] Christopher T. H. Baker. A perspective on the numerical treatment of Volterra equations. Journal of Computational and Applied Mathematics, 125(1–2):217–249, December 15, 2000. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042700004702.

Baker:2000:RDE

[Bak00b] Christopher T. H. Baker. Retarded differential equations. Journal of Computational and Applied Mathematics, 125 (1–2):309–335, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004763.

Baker:2003:SSP

[Bak03] George A. Baker, Jr. Some structural properties of two counter-examples to the Baker– Gammel–Wills conjecture. Journal of Computational and Applied Mathematics, 161 (2):371–391, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007064.

Banas:2004:SCD

[Ban04] L’ubomír Banas. Solution of convection–diffusion equation by the method of characteristics. Journal of Computational and Applied Mathematics, 168(1–2):31–39, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009610.

Barrio:2002:REB

[Bar02]  Roberto Barrio.  Round-

ing error bounds for the Clenshaw and Forsythe algorithms for the evaluation of orthogonal polynomial series. Journal of Computational and Applied Mathematics, 138(2):185–204, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100382X.

Barone:2003:OPR

[Bar03]  Piero Barone.  Orthog-

onal polynomials, random matrices and the numerical inversion of the Laplace transform of positive functions. Journal of Computational and Applied Mathematics, 155(2):307–330, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008713.

Barkatou:2004:SIF

[Bar04]  M. A. Barkatou.  On

super-irreducible forms of linear differential systems with rational function coefficients. Journal of Computational and Applied Mathematics, 162(1):1–15, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007209.

Batra:2004:PNO

[Bat04] Prashant Batra. A property of the nearly optimal rootbound. Journal of Computational and Applied Mathematics, 167(2):489–491, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009397.

Bavinck:2000:DOH

[Bav00a]  H. Bavinck.  Differential

operators having Sobolevtype Gegenbauer polynomials as eigenfunctions. Journal of Computational and Applied Mathematics, 118 (1–2):23–42, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270000279X.

Bavinck:2000:NKD

[Bav00b] H. Bavinck. A note on the Koekoeks’ differential equation for generalized Jacobi polynomials. Journal of Computational and Applied Mathematics, 115(1–2):87–92, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001806.

Bavinck:2001:DOH

[Bav01]  H. Bavinck.  Differential

operators having Sobolevtype Laguerre polynomials as eigenfunctions: new developments. Journal of Computational and Applied Mathematics, 133(1–2):183–193, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006427.

Bavinck:2003:DOH

[Bav03] H. Bavinck. Differential operators having Sobolev-type Jacobi polynomials as eigenfunctions. Journal of Computational and Applied Mathematics, 151(2):271–295, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008105.

Berezansky:2000:OLE

[BB00] Leonid Berezansky and Elena Braverman. On oscillation of a logistic equation with several delays. Journal of Computational and Applied Mathematics, 113(1–2):255–265, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002605.

Baltensperger:2001:LRC

[BB01] Richard Baltensperger and Jean-Paul Berrut. The linear rational collocation method. Journal of Computational and Applied Mathematics, 134 (1–2):243–258, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005525.

Bennewitz:2002:LPC

[BB02] C. Bennewitz and B. M.

Brown. A limit point criterion with applications to nonselfadjoint equations. Journal of Computational and Applied Mathematics, 148 (1):257–265, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005848.

Berezansky:2003:LOTb

[BB03a] Leonid Berezansky and Elena Braverman. Linearized oscillation theory for a nonlinear delay impulsive equation. Journal of Computational and Applied Mathematics, 161(2):477–495, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007386.

Berezansky:2003:LOTa

[BB03b] Leonid Berezansky and Elena Braverman. Linearized oscillation theory for a nonlinear nonautonomous delay differential equation. Journal of Computational and Applied Mathematics, 151 (1):119–127, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042702007410.

Bartholomew-Biggs:2000:ADA

[BBBCD00] Michael Bartholomew-Biggs, Steven Brown, Bruce Christianson, and Laurence Dixon. Automatic differentiation of algorithms. Journal of Computational and Applied Mathematics, 124(1–2):171– 190, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700004222.

Borwein:2000:CSR

[BBC00] Jonathan M. Borwein, David M. Bradley, and Richard E. Crandall. Computational strategies for the Riemann zeta function. Journal of Computational and Applied Mathematics, 121 (1–2):247–296, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003368.

Betts:2000:COV

[BBCH00] John T. Betts, Neil Biehn, Stephen L. Campbell, and William P. Huffman. Compensating for order variation in mesh refinement for direct transcription methods. Journal of Computational and Applied Mathematics, 125 (1–2):147–158, December 15, 2000. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004659.

Betts:2002:COV

[BBCH02] J. T. Betts, N. Biehn, S. L.

Campbell, and W. P. Huffman. Compensating for order variation in mesh refinement for direct transcription methods II: computational experience. Journal of Computational and Applied Mathematics, 143(2):237–261, June 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100509X.

Berres:2004:CSS

[BBK04] S. Berres, R. Bürger, and K. H.

Karlsen. Central schemes and systems of conservation laws with discontinuous coefficients modeling gravity separation of polydisperse suspensions. Journal of Computational and Applied Mathematics, 164–165(??):53–80, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004965.

Bauer:2000:NMO

[BBKS00] Irene Bauer, Hans Georg Bock, Stefan Körkel, and Johannes P. Schlöder. Numerical methods for optimum experimental design in DAE systems. Journal of Computational and Applied Mathematics, 120(1–2):1–25, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003009.

Burrage:2000:NSS

[BBM00] Kevin Burrage, Pamela Burrage, and Taketomo Mitsui. Numerical solutions of stochastic differential equations — implementation and stability issues. Journal of Computational and Applied Mathematics, 125(1– 2):171–182, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004672.

Berroukeche:2004:HBT

[BBMS04] M. Berroukeche, A. Boussari, E. K. Maschke, and B. Saramito. Hopf bifurcation for tearing instability in magnetohydrodynamics, theory and numerical results. Journal of Computational and Applied Mathematics, 170 (2):291–301, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042704000718.

Binding:2002:SLP

[BBW02] Paul A. Binding, Patrick J.

Browne, and Bruce A. Watson. Sturm–Liouville problems with boundary conditions rationally dependent on the eigenparameter, II. Journal of Computational and Applied Mathematics, 148(1):147–168, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005794.

Binding:2004:RFS

[BBW04] Paul A. Binding, Patrick J.

Browne, and Bruce A. Watson. Recovery of the m-function from spectral data for generalized Sturm–Liouville problems. Journal of Computational and Applied Mathematics, 171(1–2):73–91, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000445.

Berriochoa:2002:DPC

[BC02] E. Berriochoa and A. Cachafeiro.

Differential properties for a class of Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 146 (2):361–372, September 15, 2002. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702003692.

Berriochoa:2003:NCE

[BC03] E. Berriochoa and A. Cachafeiro. A necessary condition for the extension of Szegő’s asymptotics inside the disk in the Sobolev case. Journal of Computational and Applied Mathematics, 153(1–2):73–78, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006064.

Bellenger:2004:NAA

[BC04] Emmanuel Bellenger and Patrice Coorevits. New approach for automation of finite element analysis. Journal of Computational and Applied Mathematics, 168 (1–2):65–75, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703009646.

Biehn:2000:SCD

[BCJW00] Neil Biehn, Stephen L. Campbell, Laurent Jay, and Tracey Westbrook. Some comments on DAE theory for IRK methods and trajectory optimization. Journal of Computational and Applied Mathematics, 120(1–2):109–131, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000306X.

Berriochoa:2001:DPS

[BCM01] E. Berriochoa, A. Cachafeiro, and F. Marcellán. Differential properties for Sobolev orthogonality on the unit circle. Journal of Computational and Applied Mathematics, 133(1–2):231–239, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006452.

Birgin:2003:OPE

[BCM03] Ernesto G. Birgin, Ivan E.

Chambouleyron, and José Mario Martínez. Optimization problems in the estimation of parameters of thin films and the elimination of the influence of the substrate. Journal of Computational and Applied Mathematics, 152(1–2):35–50, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006957.

Bavinck:2000:CFD

[BD00a] H. Bavinck and H. A. Dieterman. Closed-form dynamic response of damped mass-spring cascades. Journal of Computational and Applied Mathematics, 114(2):291–303, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704279900206X.

Bultheel:2000:RAL

[BD00b] A. Bultheel and B. De Moor. Rational approximation in linear systems and control. Journal of Computational and Applied Mathematics, 121 (1–2):355–378, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003393.

Beska:2004:MOT

[BD04a] M. Beśka and K. Dziedziul. Mazur–Orlicz theorem and asymptotic formula for the error in supreme norm. Journal of Computational and Applied Mathematics, 167(1):219–225, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008987.

Boglaev:2004:UMD

[BD04b] Igor Boglaev and Vic Duoba. On an uniform multidomain decomposition method applied to a singularly perturbed problem with regular boundary layers. Journal of Computational and Applied Mathematics, 166(1):13–29, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008719.

Bultheel:2001:CBQ

[BDGV01] Adhemar Bultheel, Leyla Daruis, and Pablo González-Vera. A connection between quadrature formulas on the unit circle and the interval [-1,1]. Journal of Computational and Applied Mathematics, 132(1):1–14, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000594X.

Bracciali:2002:CSS

[BDR02] C. F. Bracciali, D. K. Dimitrov, and A. Sri Ranga. Chain sequences and symmetric generalized orthogonal polynomials. Journal of Computational and Applied Mathematics, 143(1):95–106, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100499X.

Bracciali:2003:SPS

[BdSR03] C. F. Bracciali, A. P. da Silva, and A. Sri Ranga. Szegő polynomials: some relations to L-orthogonal and orthogonal polynomials. Journal of Computational and Applied Mathematics, 153(1–2):79–88, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006052.

Balinsky:2002:SPB

[BE02a] A. A. Balinsky and W. D. Evans. On the spectral properties of the Brown–Ravenhall operator. Journal of Computational and Applied Mathematics, 148(1):239–255, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005836.

Brown:2002:ACR

[BE02b] B. M. Brown and M. S. P. Eastham. Analytic continuation and resonance-free regions for Sturm–Liouville potentials with power decay. Journal of Computational and Applied Mathematics, 148(1):49–63, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005721.

Brown:2002:LRE

[BE02c] B. M. Brown and M. S. P.

Eastham. On the location of resonances of exponentially decaying Sturm–Liouville potentials. Journal of Computational and Applied Mathematics, 139(1):163–171, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003843.

Brown:2002:P

[BE02d] B. M. Brown and W. D. Evans.

Preface. Journal of Computational and Applied Mathematics, 148(1):ix, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005691.

Brown:2002:PME

[BE02e] B. M. Brown and W. D. Evans.

Prof. Michael eastham laudatum. Journal of Computational and Applied Mathematics, 148(1):xi–xiii, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007562.

Bermejo:2003:FES

[BE03] Rodolfo Bermejo and Mofdi El Amrani. A finite element semi-Lagrangian explicit Runge– Kutta–Chebyshev method for convection dominated reaction– diffusion problems. Journal of Computational and Applied Mathematics, 154 (1):27–61, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270200746X.

Bennewitz:2004:IAM

[BE04a] C. Bennewitz and W. D. Evans. Inequalities associated with magnetic fields. Journal of Computational and Applied Mathematics, 171(1–2):59–72, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000482.

Brown:2004:EHR

[BE04b] B. M. Brown and M. S. P. Eastham. Extended Hurwitz results for hypergeometric functions arising in spectral theory. Journal of Computational and Applied Mathematics, 171(1–2):113–121, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000457.

Brown:2004:P

[BE04c] B. M. Brown and W. D. Evans. Preface. Journal of Computational and Applied Mathematics, 171(1–2):ix, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000421.

Beardon:2001:WTC

[Bea01] A. F. Beardon. Worpitzky’s theorem on continued fractions. Journal of Computational and Applied Mathematics, 131(1–2):143–148, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003186.

Beckermann:2001:CJM

[Bec01] Bernhard Beckermann. Complex Jacobi matrices. Journal of Computational and Applied Mathematics, 127 (1–2):17–65, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004921.

Brown:2004:L

[BEE04] B. M. Brown, M. S. P. Eastham, and W. D. Evans. Laudatum. Journal of Computational and Applied Mathematics, 171(1–2):xi–xiii, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042704000433.

Behncke:2002:SDO

[Beh02] Horst Behncke. The spectrum of differential operators with almost constant coefficients II. Journal of Computational and Applied Mathematics, 148(1):287–305, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005861.

Belmehdi:2001:GGO

[Bel01]  S. Belmehdi.  Generalized

Gegenbauer orthogonal polynomials. Journal of Computational and Applied Mathematics, 133(1–2):195–205, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006439.

Ben-El-Mechaiekh:2000:SMA

[BEM00a] Hichem Ben-El-Mechaiekh. Spaces and maps approximation and fixed points. Journal of Computational and Applied Mathematics, 113 (1–2):283–308, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002629.

Brown:2000:RAC

[BEM00b] B. M. Brown, M. S. P. Eastham, and D. K. R. McCormack. Resonances and analytic continuation for exponentially decaying Sturm–Liouville potentials. Journal of Computational and Applied Mathematics, 116(1):181–193, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003246.

Benney:2002:NIT

[Ben02] D. J. Benney. Nonlinear instability theories in hydrodynamics. Journal of Computational and Applied Mathematics, 144(1–2):1–5, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005489.

Betts:2000:VLT

[Bet00] John T. Betts. Very lowthrust trajectory optimization using a direct SQP method. Journal of Computational and Applied Mathematics, 120(1–2):27–40, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003010.

Bialecki:2001:OSC

[BF01] B. Bialecki and G. Fairweather. Orthogonal spline collocation methods for partial differential equations. Journal of Computational and Applied Mathematics, 128(1–2):55–82, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005094.

Behncke:2002:SPD

[BF02] Horst Behncke and Andreas Fischer. Spectral properties of a differential system. Journal of Computational and Applied Mathematics, 148(1):267–285, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200585X.

Bartel:2002:MME

[BG02] A. Bartel and M. Günther. A multirate W -method for electrical networks in statespace formulation. Journal of Computational and Applied Mathematics, 147 (2):411–425, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004764.

Bjerregaard:2003:CAS

[BG03] Pablo Alberca Bjerregaard and Cándido Martín González. Computational approach to the simplicity of f_{4}(O_{s}, −) in the characteristic two case. Journal of Computational and Applied Mathematics, 158(1):1–10, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004722.

Belforte:2000:SNP

[BGM00] G. Belforte, P. Gay, and G. Monegato. Some new properties of Chebyshev polynomials. Journal of Computational and Applied Mathematics, 117(2):175–181, May 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002715. See er-

ratum [BGM01].

Belforte:2001:ESN

[BGM01] G. Belforte, P. Gay, and G. Monegato. Erratum to: “Some new properties of Chebyshev polynomials” [J. Comput. Appl. Math. 117 (2000) 175–181]. Journal of Computational and Applied Mathematics, 131 (1–2):505, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700005926.

See [BGM00].

Breimesser:2002:VDS

[BGP02] S. V. Breimesser, James D. E. Grant, and D. B. Pearson. Value distribution and spectral theory of Schrödinger operators with L^{2}-sparse potentials. Journal of Computational and Applied Mathematics, 148(1):307–322, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005873.

Broeckx:2002:PI

[BGPW02] F. Broeckx, M. J. Goovaerts, R. Piessens, and L. Wuytack. Proceedings of ICCAM-2000. Journal of Computational and Applied Mathematics, 140(1–2):xi–xii, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005970.

Bultheel:2001:DRM

[BGVHN01a] A. Bultheel, P. González-Vera, E. Hendriksen, and O. Njåstad. Determinacy of a rational moment problem. Journal of Computational and Applied Mathematics, 133(1–2):241–252, August 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006890.

Bultheel:2001:QOR

[BGVHN01b] A. Bultheel, P. González-Vera, E. Hendriksen, and Olav Njåstad. Quadrature and orthogonal rational functions. Journal of Computational and Applied Mathematics, 127(1–2):67–91, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004933.

Bultheel:2003:ORF

[BGVHN03] A. Bultheel, P. González-Vera, E. Hendriksen, and O. Njåstad. Orthogonal rational functions and tridiagonal matrices. Journal of Computational and Applied Mathematics, 153(1–2):89–97, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006027.

Bellen:2000:NSN

[BGZ00] Alfredo Bellen, Nicola Guglielmi, and Marino Zennaro. Numerical stability of nonlinear delay differential equations of neutral type. Journal of Computational and Applied Mathematics, 125 (1–2):251–263, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004714.

Budd:2003:NAE

[BH03] C. J. Budd and A. R. Humphries. Numerical and analytical estimates of existence regions for semi-linear elliptic equations with critical Sobolev exponents in cuboid and cylindrical domains. Journal of Computational and Applied Mathematics, 151(1):59–84, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007379.

Brown:2004:RFB

[BH04] Richard C. Brown and Don B. Hinton. Relative form boundedness and compactness for a second-order differential operator. Journal of Computational and Applied Mathematics, 171(1–2):123–140, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000536.

Bhaskar:2002:CTN

[Bha02] T. Gnana Bhaskar. Comparison theorem for a nonlinear boundary value problem on time scales. Journal of Computational and Applied Mathematics, 141(1– 2):117–122, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004393.

Burrage:2004:ASB

[BHB04] P. M. Burrage, R. Herdiana, and K. Burrage. Adaptive stepsize based on control theory for stochastic differential equations. Journal of Computational and Applied Mathematics, 170 (2):317–336, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000731.

Berselli:2003:HOS

[BI03] Luigi C. Berselli and Traian Iliescu. A higher-order subfilterscale model for large eddy simulation. Journal of Computational and Applied Mathematics, 159(2):411–430, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005442.

Bujanda:2002:ARK

[BJ02] B. Bujanda and J. C. Jorge.

Additive Runge–Kutta methods for the resolution of linear parabolic problems. Journal of Computational and Applied Mathematics, 140(1–2):99–117, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004757.

Bujanda:2004:ELI

[BJ04] B. Bujanda and J. C. Jorge. Efficient linearly implicit methods for nonlinear multidimensional parabolic problems. Journal of Computational and Applied Mathematics, 164– 165(??):159–174, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300949X.

Bhuruth:2002:PIM

[BJG02] M. Bhuruth, M. K. Jain, and A. Gopaul. Preconditioned iterative methods for the nine-point approximation to the convection–diffusion equation. Journal of Computational and Applied Mathematics, 138(1):73–92, January 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003624.

Banda:2003:KDF

[BJK03] Mapundi K. Banda, Michael Junk, and Axel Klar. Kinetic derivation of a finite difference scheme for the incompressible Navier–Stokes equation. Journal of Computational and Applied Mathematics, 154(2):341–354, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008579.

Bartoszewski:2004:CIM

[BJK04a] Z. Bartoszewski, T. Jankowski, and M. Kwapisz. On the convergence of iterative methods for general differential-algebraic systems. Journal of Computational and Applied Mathematics, 169(2):393–418, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000056.

Bielecki:2004:PAW

[BJK04b] Andrzej Bielecki, Dariusz Jabloński, and Marek Kedzierski. Properties and applications of weakly nonlinear neurons. Journal of Computational and Applied Mathematics, 164–165 (??):93–106, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703008264.

Bitterlich:2002:EMS

[BK02] S. Bitterlich and P. Knabner. An efficient method for solving an inverse problem for the Richards equation. Journal of Computational and Applied Mathematics, 147(1):153–173, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004302.

Bourchtein:2004:SPV

[BKB04] A. Bourchtein, V. Kadychnikov, and L. Bourchtein. Some properties of the vertical structure matrices of the atmospheric models. Journal of Computational and Applied Mathematics, 164– 165(??):131–144, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300640X.

Balasubramanian:2003:IFL

[BKR03] R. Balasubramanian, S. Kanemitsu, and K. Ramachandra. On ideal-function-like functions. Journal of Computational and Applied Mathematics, 160(1–2):27–36, November 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006113.

Banas:2001:SIS

[BL01a] Józef Banaś and Millenia Lecko. Solvability of infinite systems of differential equations in Banach sequence spaces. Journal of Computational and Applied Mathematics, 137(2):363–375, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007081.

Beardon:2001:ASP

[BL01b] Alan Beardon and Lisa Lorentzen.

Approximants of Śleszyński– Pringsheim continued fractions. Journal of Computational and Applied Mathematics, 132(2):467–477, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004489.

Begehr:2001:ASP

[BL01c] Heinrich Begehr and Xing Li.

Approximate solution of periodic Riemann boundary value problem for analytic functions. Journal of Computational and Applied Mathematics, 134 (1–2):85–93, September 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005306.

Bowman:2004:TDG

[BL04] Douglas Bowman and James Mc Laughlin. A theorem on divergence in the general sense for continued fractions. Journal of Computational and Applied Mathematics, 172(2):363–373, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001244.

Binding:2004:ORS

[BLM04] P. Binding, H. Langer, and M. Möller. Oscillation results for Sturm–Liouville problems with an indefinite weight function. Journal of Computational and Applied Mathematics, 171(1–2):93–101, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270400055X.

Benamou:2003:GOB

[BLSS03] J.-D. Benamou, O. Lafitte, R. Sentis, and I. Solliec. A geometrical optics-based numerical method for high frequency electromagnetic fields computations near fold caustics — Part I. Journal of Computational and Applied Mathematics, 156(1):93–125, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200907X.

Benamou:2004:GOM

[BLSS04] J.-D. Benamou, O. Lafitte, I. Solliec, and R. Sentis. A geometric optics method for high-frequency electromagnetic fields computations near fold caustics — Part II. The energy. Journal of Computational and Applied Mathematics, 167(1):91–134, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300880X.

Baxley:2000:PSS

[BM00a] John V. Baxley and Jenifer Clason Martin. Positive solutions of singular nonlinear boundary value problems. Journal of Computational and Applied Mathematics, 113(1–2):381–399, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900268X.

Billups:2000:CP

[BM00b] Stephen C. Billups and Katta G.

Murty. Complementarity problems. Journal of Computational and Applied Mathematics, 124(1–2):303–318, December 1, 2000. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042700004325.

Buskens:2000:SMS

[BM00c] Christof Büskens and Helmut Maurer. SQP-methods for solving optimal control problems with control and state constraints: adjoint variables, sensitivity analysis and real-time control. Journal of Computational and Applied Mathematics, 120(1–2):85–108, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003058.

Beckett:2001:UAF

[BM01a] G. Beckett and J. A. Mackenzie. On a uniformly accurate finite difference approximation of a singularly perturbed reaction–diffusion problem using grid equidistribution. Journal of Computational and Applied Mathematics, 131(1–2):381–405, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002600.

Boros:2001:DSR

[BM01b] George Boros and Victor H.

Moll. The double square root, Jacobi polynomials and Ramanujan’s Master Theorem. Journal of Computational and Applied Mathematics, 130(1–2):337–344, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003726.

Blanes:2002:PSP

[BM02] S. Blanes and P. C. Moan.

Practical symplectic partitioned Runge–Kutta and Runge–Kutta–Nyström methods. Journal of Computational and Applied Mathematics, 142(2):313–330, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004927.

Bellavia:2003:GCE

[BM03a] Stefania Bellavia and Benedetta Morini. Global convergence enhancement of classical linesearch interior point methods for MCPs. Journal of Computational and Applied Mathematics, 151 (1):171–199, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042702007458.

Buchwald:2003:CBS

[BM03b] B. Buchwald and G. Mühlbach. Construction of B-splines for generalized spline spaces generated from local ECTsystems. Journal of Computational and Applied Mathematics, 159(2):249–267, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005338.

Berrut:2004:APS

[BM04a] Jean-Paul Berrut and Hans D. Mittelmann. Adaptive point shifts in rational approximation with optimized denominator. Journal of Computational and Applied Mathematics, 164–165(??):81–92, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004850.

Brugnano:2004:BCN

[BM04b] Luigi Brugnano and Cecilia Magherini. The BiM code for the numerical solution of ODEs. Journal of Computational and Applied Mathematics, 164– 165(??):145–158, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008227.

Buchwald:2004:RBS

[BM04c] B. Buchwald and G. Mühlbach.

On rational B-splines with prescribed poles. Journal of Computational and Applied Mathematics, 167(2):271–291, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009221.

Boros:2001:ITP

[BMN01] George Boros, Victor H. Moll, and Roopa Nalam. An integral with three parameters. Part 2. Journal of Computational and Applied Mathematics, 134 (1–2):113–126, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000532X.

Baker:2000:PNA

[BMPV00] Christopher Baker, Giovanni Monegato, John Pryce, and Guido Vanden Berghe. Preface: Numerical Analysis 2000: Vol. VI: Ordinary differential equations and integral equations. Journal of Computational and Applied Mathematics, 125 (1–2):xi–xviii, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004544.

Barrio:2003:NEL

[BMS03] R. Barrio, B. Melendo, and S. Serrano. On the numerical evaluation of linear recurrences. Journal of Computational and Applied Mathematics, 150(1):71–86, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005654.

Brown:2000:CAP

[BMZ00] B. M. Brown, D. K. R. Mc-Cormack, and A. Zettl. On a computer assisted proof of the existence of eigenvalues below the essential spectrum of the Sturm–Liouville problem. Journal of Computational and Applied Mathematics, 125 (1–2):385–393, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004817.

Benammou:2002:FEM

[BO02] Samir Benammou and Khaled Omrani. A finite element method for the Sivashinsky equation. Journal of Computational and Applied Mathematics, 142(2):419–431, May 15, 2002. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003703.

Baratella:2004:NAN

[BO04a] Paola Baratella and Annamaria Palamara Orsi. A new approach to the numerical solution of weakly singular Volterra integral equations. Journal of Computational and Applied Mathematics, 163(2):401–418, February 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007635.

Bielecki:2004:SPA

[BO04b] Andrzej Bielecki and Jerzy Ombach. Shadowing property in analysis of neural networks dynamics. Journal of Computational and Applied Mathematics, 164– 165(??):107–115, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004862.

Boglaev:2001:DDS

[Bog01] Igor Boglaev. Domain decomposition for a singularly perturbed parabolic problem with a convection-dominated term. Journal of Computational and Applied Mathematics, 134 (1–2):283–299, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005550.

Bogatyrev:2003:FPM

[Bog03] A. Bogatyrev. Fibers of periods map are cells? Journal of Computational and Applied Mathematics, 153(1– 2):547–548, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702008877.

Boglaev:2004:MIA

[Bog04a] Igor Boglaev. Monotone iterative algorithms for a nonlinear singularly perturbed parabolic problem. Journal of Computational and Applied Mathematics, 172(2):313–335, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001219.

Boglaev:2004:MIM

[Bog04b]  Igor Boglaev.  On mono-

tone iterative methods for a nonlinear singularly perturbed reaction–diffusion problem. Journal of Computational and Applied Mathematics, 162(2):445–466, January 15, 2004. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007593.

Borzi:2003:MMP

[Bor03]  Alfio Borzì.  Multigrid

methods for parabolic distributed optimal control problems. Journal of Computational and Applied Mathematics, 157(2):365–382, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004175.

Boulmakoul:2004:GPF

[Bou04]  A. Boulmakoul.  General-

ized path-finding algorithms on semirings and the fuzzy shortest path problem. Journal of Computational and Applied Mathematics, 162(1): 263–272, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703007118.

Berg:2001:CMF

[BP01a] Christian Berg and Henrik L. Pedersen. A completely monotone function related to the gamma function. Journal of Computational and Applied Mathematics, 133(1–2):219–230, August 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006440.

Budd:2001:GIS

[BP01b] C. J. Budd and M. D. Piggott. The geometric integration of scale-invariant ordinary and partial differential equations. Journal of Computational and Applied Mathematics, 128(1–2):399–422, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005215.

Bacchelli:2004:FWT

[BP04] Silvia Bacchelli and Serena Papi. Filtered wavelet thresholding methods. Journal of Computational and Applied Mathematics, 164–165 (??):39–52, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703008239.

Baker:2002:DAE

[BPT02] C. T. H. Baker, C. A. H. Paul, and H. Tian. Differential algebraic equations with after-effect. Journal of Computational and Applied Mathematics, 140(1–2):63–80, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701006008.

Balasubramanian:2002:HFF

[BPV02] R. Balasubramanian, S. Ponnusamy, and M. Vuorinen. On hypergeometric functions and function spaces. Journal of Computational and Applied Mathematics, 139 (2):299–322, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004174.

Brown:2002:LBM

[BPW02] B. M. Brown, R. A. Peacock, and R. Weikard. A local Borg–Marchenko theorem for complex potentials. Journal of Computational and Applied Mathematics, 148(1):115–131, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005770.

Barral:2000:NAV

[BQ00] P. Barral and P. Quintela. Numerical analysis of a viscoplastic problem with contact condition taking place in an aluminium casting. Journal of Computational and Applied Mathematics, 115(1–2):63–86, March 1, 2000. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003003.

Bocharov:2000:NMB

[BR00] Gennadii A. Bocharov and Fathalla A. Rihan. Numerical modelling in biosciences using delay differential equations. Journal of Computational and Applied Mathematics, 125 (1–2):183–199, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004684.

Bultheel:2001:P

[BR01] Adhemar Bultheel and Dirk Roose. Preface. Journal of Computational and Applied Mathematics, 132(1): ix–x, July 1, 2001. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042700005938.

Brown:2002:CEF

[BR02] B. M. Brown and W. Reichel. Computing eigenvalues and fucik-spectrum of the radially symmetric p-Laplacian. Journal of Computational and Applied Mathematics, 148 (1):183–211, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005812.

Brezinski:2000:CAD

[Bre00a] C. Brezinski. Convergence acceleration during the 20th century. Journal of Computational and Applied Mathematics, 122(1–2):1–21, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003605.

Brezinski:2000:NAV

[Bre00b]  C. Brezinski.  Numerical

analysis 2000 vol. II: Interpolation and extrapolation. Journal of Computational and Applied Mathematics, 122(1–2):ix–xi, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003526.

Brigadnov:2004:DML

[Bri04] Igor A. Brigadnov. Duality method for limit analysis of dielectrics in powerful electric fields. Journal of Computational and Applied Mathematics, 168(1–2):87–94, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300966X.

Banas:2000:CUS

[BRS00] Józef Banaś, Juan R. Rodriguez, and Kishin Sadarangani. On a class of Urysohn– Stieltjes quadratic integral equations and their applications. Journal of Computational and Applied Mathematics, 113(1–2):35–50, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002423.

Banas:2003:SNI

[BRS03a] J. Banas, J. Rocha, and K. B. Sadarangani. Solvability of a nonlinear integral equation of Volterra type. Journal of Computational and Applied Mathematics, 157(1):31–48, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300373X.

Bermudez:2003:FEA

[BRS03b] A. Bermúdez, R. Rodríguez, and D. Santamarina. Finite element approximation of a displacement formulation for time-domain elastoacoustic vibrations. Journal of Computational and Applied Mathematics, 152(1–2):17–34, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042702006945.

Brugnano:2000:BBB

[Bru00] Luigi Brugnano. Blended block BVMS (B_{3}VMs): a family of economical implicit methods for ODES. Journal of Computational and Applied Mathematics, 116(1):41–62, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002800.

Brezinski:2000:MPA

[BRZS00] C. Brezinski, M. Redivo-Zaglia, and H. Sadok. The matrix and polynomial approaches to Lanczos-type algorithms. Journal of Computational and Applied Mathematics, 123(1–2):241– 260, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700003976.

Bartels:2000:RSR

[BS00a] Richard H. Bartels and Faramarz F. Samavati. Reversing subdivision rules: local linear conditions and observations on inner products. Journal of Computational and Applied Mathematics, 119(1–2):29–67, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003708.

Becken:2000:ACG

[BS00b] W. Becken and P. Schmelcher. The analytic continuation of the Gaussian hypergeometric function _{2}F_{1}(a, b; c; z) for arbitrary parameters. Journal of Computational and Applied Mathematics, 126 (1–2):449–478, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002673.

Berselli:2000:NSD

[BS00c] Luigi C. Berselli and Fausto Saleri. New substructuring domain decomposition methods for advection–diffusion equations. Journal of Computational and Applied Mathematics, 116(2):201–220, April 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003179.

Borovykh:2000:RCB

[BS00d] N. Borovykh and M. N. Spijker. Resolvent conditions and bounds on the powers of matrices, with relevance to numerical stability of initial value problems. Journal of Computational and Applied Mathematics, 125(1–2):41–56, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000457X.

Binz:2002:ITL

[BS02a] Ernst Binz and Walter Schempp.

Information technology: the Lie groups defining the filter banks of the compact disc. Journal of Computational and Applied Mathematics, 144(1–2):85–103, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005520.

Borovykh:2002:BPS

[BS02b] N. Borovykh and M. N.

Spijker. Bounding partial sums of Fourier series in weighted L^{2}-norms, with applications to matrix analysis. Journal of Computational and Applied Mathematics, 147(2):349–368, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004417.

Baruah:2003:SGT

[BS03] Nayandeep Deka Baruah and Nipen Saikia. Some general theorems on the explicit evaluations of Ramanujan’s cubic continued fraction. Journal of Computational and Applied Mathematics, 160(1–2):37–51, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006125.

Burenkov:2004:III

[BS04] V. I. Burenkov and A. Senouci. On integral inequalities involving differences. Journal of Computational and Applied Mathematics, 171 (1–2):141–149, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000469.

Banda:2004:CIL

[BSKP04] M. K. Banda, M. Seaïd, A. Klar, and L. Pareschi. Compressible and incompressible limits for hyperbolic systems with relaxation. Journal of Computational and Applied Mathematics, 168(1–2):41–52, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009622.

Beals:2003:CFI

[BSS03] Richard Beals, David H. Sattinger, and Jacek Szmigielski. Continued fractions and integrable systems. Journal of Computational and Applied Mathematics, 153 (1–2):47–60, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702006076.

Boggs:2000:SQP

[BT00a] Paul T. Boggs and Jon W.

Tolle. Sequential quadratic programming for large-scale nonlinear optimization. Journal of Computational and Applied Mathematics, 124 (1–2):123–137, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004295.

Bridson:2000:RAI

[BT00b] Robert Bridson and Wei-Pai Tang. Refining an approximate inverse. Journal of Computational and Applied Mathematics, 123 (1–2):293–306, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000399X.

Burrage:2001:CEM

[BT01] Kevin Burrage and Tianhai Tian. The composite Euler method for stiff stochastic differential equations. Journal of Computational and Applied Mathematics, 131(1–2):407–426, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002594.

Boyadjiev:2002:NAM

[BTFY02] T. L. Boyadjiev, M. D. Todorov, P. P. Fiziev, and S. S. Yazadjiev. A numerical algorithm for modelling of bosonfermion stars in dilatonic gravity. Journal of Computational and Applied Mathematics, 145(1):113–131, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005398.

Bazdidi-Tehrani:2004:NAS

[BTSHK04] F. Bazdidi-Tehrani, A. Shahmir, and A. Haghparast-Kashani. Numerical analysis of a single row of coolant jets injected into a heated crossflow. Journal of Computational and Applied Mathematics, 168(1–2):53–63, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009634.

Buckwar:2000:INA

[Buc00] Evelyn Buckwar. Introduction to the numerical analysis of stochastic delay differential equations. Journal of Computational and Applied Mathematics, 125 (1–2):297–307, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004751.

Buhring:2000:SOL

[Büh00] Wolfgang Bühring. Secondorder linear differential equations with two irregular singular points of rank three: the characteristic exponent. Journal of Computational and Applied Mathematics, 118 (1–2):43–69, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700002818.

Butcher:2000:NMO

[But00] J. C. Butcher. Numerical methods for ordinary differential equations in the 20th century. Journal of Computational and Applied Mathematics, 125(1–2):1–29, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004556.

Blom:2000:CIM

[BV00] J. G. Blom and J. G. Verwer. A comparison of integration methods for atmospheric transport-chemistry problems. Journal of Computational and Applied Mathematics, 126 (1–2):381–396, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003660.

Benito:2002:PTL

[BV02] Manuel Benito and Juan L. Varona. Pythagorean triangles with legs less than n. Journal of Computational and Applied Mathematics, 143(1):117–126, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004964.

Botchev:2003:NAM

[BV03] M. A. Botchev and J. G. Verwer. A new approximate matrix factorization for implicit time integration in air pollution modeling. Journal of Computational and Applied Mathematics, 157(2):309–327, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S037704270300414X.

Botchev:2001:PNI

[BvdV01] Mike A. Botchev and Henk A.

van der Vorst. A parallel nearly implicit time-stepping scheme. Journal of Computational and Applied Mathematics, 137(2):229–243, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003582.

Bultheel:2004:OBF

[BVV04] A. Bultheel, M. Van Barel, and P. Van Gucht. Orthogonal basis functions in discrete leastsquares rational approximation. Journal of Computational and Applied Mathematics, 164–165(??):175–194, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004977.

Bai:2002:CNM

[BW02] Zhong-Zhi Bai and Chuan-Long Wang. On the convergence of nonstationary multisplitting two-stage iteration methods for Hermitian positive definite linear systems. Journal of Computational and Applied Mathematics, 138(2):287–296, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003764.

Butler:2003:LAB

[BW03] Ronald W. Butler and Andrew T. A. Wood. Laplace approximation for Bessel functions of matrix argument. Journal of Computational and Applied Mathematics, 155(2):359–382, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008749.

Bors:2004:SNE

[BW04a] D. Bors and S. Walczak. Stability of nonlinear elliptic systems with distributed parameters and variable boundary data. Journal of Computational and Applied Mathematics, 164–165(??):117– 130, March 1, 2004. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042703008355.

Breen:2004:ABL

[BW04b] Sinéad Breen and Alastair Wood. On the asymptotic behaviour of Laplacetype multiple integral solutions of linear differential equations. Journal of Computational and Applied Mathematics, 171(1–2):103–112, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000652.

Brenner:2002:FMI

[BX02] Marty Brenner and Yuesheng Xu. A factorization method for identification of Volterra systems. Journal of Computational and Applied Mathematics, 144(1–2):105–117, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005532.

Berndt:2003:TPP

[BYY03] Bruce C. Berndt, Ae Ja Yee, and Jinhee Yi. Theorems on partitions from a page in Ramanujan’s lost notebook. Journal of Computational and Applied Mathematics, 160(1–2):53–68, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006137.

Boutyour:2004:ANM

[BZPFB04] E. H. Boutyour, H. Zahrouni, M. Potier-Ferry, and M. Boudi. Asymptotic-numerical method for buckling analysis of shell structures with large rotations. Journal of Computational and Applied Mathematics, 168(1–2):77–85, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009658.

Brezinski:2002:RFO

[BZS02] C. Brezinski, M. Redivo Zaglia, and H. Sadok. A review of formal orthogonality in Lanczos-based methods. Journal of Computational and Applied Mathematics, 140(1–2):81–98, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004125.

Carey:2004:SAA

[CACK04] Graham F. Carey, M. Anderson, B. Carnes, and B. Kirk. Some aspects of adaptive grid technology related to boundary and interior layers. Journal of Computational and Applied Mathematics, 166(1):55–86, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008689.

Cai:2002:EGM

[Cai02] Donghan Cai. An economic growth model with endogenous fertility: multiple growth paths, poverty trap and bifurcation. Journal of Computational and Applied Mathematics, 144(1–2):119–130, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005544.

Campos:2001:SSE

[Cam01] L. M. B. C. Campos. On some solutions of the extended confluent hypergeometric differential equation. Journal of Computational and Applied Mathematics, 137 (1):177–200, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007068.

Cao:2002:CIM

[Cao02] Zhi-Hao Cao. On the convergence of iterative methods for solving singular linear systems. Journal of Computational and Applied Mathematics, 145(1):1–9, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005313.

Cap:2004:IEP

[Cap04] Ferdinand F. Cap. An inhomogeneous eigenvalue problem. Journal of Computational and Applied Mathematics, 167(1):243–249, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007969.

Carfi:2003:DOS

[Car03] David Carfì. Dirac-orthogonality in the space of tempered distributions. Journal of Computational and Applied Mathematics, 153(1–2):99–107, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006349.

Cash:2000:MEB

[Cas00] J. R. Cash. Modified extended backward differentiation formulae for the numerical solution of stiff initial value problems in ODEs and DAEs. Journal of Computational and Applied Mathematics, 125 (1–2):117–130, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004635.

Casas:2004:NIM

[Cas04] Fernando Casas. Numerical integration methods for the double-bracket flow. Journal of Computational and Applied Mathematics, 166(2): 477–495, April 15, 2004.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703007714.

Cervantes:2000:SED

[CB00] Arturo M. Cervantes and Lorenz T. Biegler. A stable elemental decomposition for dynamic process optimization. Journal of Computational and Applied Mathematics, 120(1–2):41–57, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003022.

Capobianco:2003:QCP

[CC03a] M. R. Capobianco and G. Criscuolo.

On quadrature for Cauchy principal value integrals of oscillatory functions. Journal of Computational and Applied Mathematics, 156 (2):471–486, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703003881.

Chuli:2003:WEE

[CC03b] Fu Chuli and Qiu Chunyu.

Wavelet and error estimation of surface heat flux. Journal of Computational and Applied Mathematics, 150(1):143–155, January 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200657X.

Cakoni:2002:LSM

[CCH02] Fioralba Cakoni, David Colton, and Houssem Haddar. The linear sampling method for anisotropic media. Journal of Computational and Applied Mathematics, 146 (2):285–299, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003618.

Cao:2003:MAE

[CCL03] Li-Qun Cao, Jun-Zhi Cui, and Jian-Lan Luo. Multiscale asymptotic expansion and a post-processing algorithm for second-order elliptic problems with highly oscillatory coefficients over general convex domains. Journal of Computational and Applied Mathematics, 157(1):1–29, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003728.

Chien:2001:TBR

[CCM01] C.-S. Chien, S.-L. Chang, and Z. Mei. Tracing the buckling of a rectangular plate with the block GMRES method.

Journal of Computational and Applied Mathematics, 136 (1–2):199–218, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006117.

Coleman:2000:MCM

[CD00] John P. Coleman and Suzanne C.

Duxbury. Mixed collocation methods for y′′ = f(x, y). Journal of Computational and Applied Mathematics, 126 (1–2):47–75, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003404.

Coroian:2001:CLP

[CD01] Dan I. Coroian and Peter Dragnev. Constrained Leja points and the numerical solution of the constrained energy problem. Journal of Computational and Applied Mathematics, 131(1–2):427–444, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002582.

Chamayou:2003:RDE

[CD03a] J. F. Chamayou and J. L.

Dunau. Random difference equations: an asymptotical result. Journal of Computational and Applied Mathematics, 154(1):183–193, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200821X.

Collino:2003:IES

[CD03b] Francis Collino and Bruno Despres. Integral equations via saddle point problems for timeharmonic Maxwell’s equations. Journal of Computational and Applied Mathematics, 150(1):157–192, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006581.

Costa:2001:NIS

[CDN01] David G. Costa, Zhonghai Ding, and John M. Neuberger. A numerical investigation of sign-changing solutions to superlinear elliptic equations on symmetric domains. Journal of Computational and Applied Mathematics, 131(1–2):299–319, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002661.

Calvo:2003:FEN

[CDV03] N. Calvo, J. Durany, and C. Vázquez. Finite elements numerical solution of a coupled profile-velocity-temperature shallow ice sheet approximation model. Journal of Computational and Applied Mathematics, 158(1): 31–41, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703004667.

Cawood:2000:ADC

[CELM00] M. E. Cawood, V. J. Ervin, W. J. Layton, and J. M. Maubach. Adaptive defect correction methods for convection dominated, convection diffusion problems. Journal of Computational and Applied Mathematics, 116 (1):1–21, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799002782.

Cermak:2002:CVA

[Čer02] Jan Čermák. A change of variables in the asymptotic theory of differential equations with unbounded delays. Journal of Computational and Applied Mathematics, 143(1):81–93, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005003.

Calkin:2002:CCC

[CES02] Neil J. Calkin, Jonathan D. Edds, and Douglas R. Shier. Cancellation in cyclic consecutive systems. Journal of Computational and Applied Mathematics, 142(1):13–26, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004563.

Carlson:2000:RTE

[CF00] B. C. Carlson and James FitzSimons. Reduction theorems for elliptic integrands with the square root of two quadratic factors. Journal of Computational and Applied Mathematics, 118 (1–2):71–85, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270000282X.

Celledoni:2004:NLG

[CF04] Elena Celledoni and Simone Fiori. Neural learning by geometric integration of reduced ‘rigid-body’ equations. Journal of Computational and Applied Mathematics, 172 (2):247–269, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001189.

Chau:2003:VNAa

[CFHS03] O. Chau, J. R. Fernández, W. Han, and M. Sofonea. Variational and numerical analysis of a dynamic frictionless contact problem with adhesion. Journal of Computational and Applied Mathematics, 156(1):127–157, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702009093.

Cantero:2003:CBO

[CFMV03] M. J. Cantero, M. P. Ferrer, L. Moral, and L. Velázquez. A connection between orthogonal polynomials on the unit circle and matrix orthogonal polynomials on the real line. Journal of Computational and Applied Mathematics, 154(2):247–272, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008245.

Carey:2004:CCN

[CFSP04] Graham Carey, Neville Fowkes, Amanda Staelens, and Anand Pardhanani. A class of coupled nonlinear reaction diffusion models exhibiting fingering. Journal of Computational and Applied Mathematics, 166(1):87–99, April 1, 2004. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008690.

Chau:2003:VNAb

[CFSS03] O. Chau, J. R. Fernández, M. Shillor, and M. Sofonea. Variational and numerical analysis of a quasistatic viscoelastic contact problem with adhesion. Journal of Computational and Applied Mathematics, 159(2):431–465, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005478.

Chen:2002:EPM

[CFT02] Shyh-Huei Chen, Shu-Cherng Fang, and H.-S. Jacob Tsao. Entropic perturbation method for solving a system of linear inequalities. Journal of Computational and Applied Mathematics, 145(1): 133–149, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701005404.

Carnicer:2000:BHB

[CG00] J. M. Carnicer and M. Gasca. Bivariate Hermite–Birkhoff polynomial interpolation with asymptotic conditions. Journal of Computational and Applied Mathematics, 119 (1–2):69–79, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270000371X.

Cao:2002:PIE

[CG02] Weiming Cao and Benqi Guo.

A preconditioner with inexact element face solver for three dimensional pversion finite element methods. Journal of Computational and Applied Mathematics, 144(1–2):131–144, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005556.

Cao:2003:RZP

[CG03a] Jiansheng Cao and Robert Gardner. Restrictions on the zeros of a polynomial as a consequence of conditions on the coefficients of even powers and odd powers of the variable. Journal of Computational and Applied Mathematics, 155(1):153–162, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200897X.

Carfi:2003:SPN

[CG03b] David Carfì and Clara Germanà. Some properties of a new product in S′_{n}.

Journal of Computational and Applied Mathematics, 153(1–2):109–118, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006350.

Clark:2004:WTT

[CG04a] Steve Clark and Fritz Gesztesy. On Weyl–Titchmarsh theory for singular finite difference Hamiltonian systems. Journal of Computational and Applied Mathematics, 171 (1–2):151–184, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000500.

Clavero:2004:HFD

[CG04b] C. Clavero and J. L. Gracia. HODIE finite difference schemes on generalized Shishkin meshes. Journal of Computational and Applied Mathematics, 164– 165(??):195–206, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006538.

Chavarriga:2000:ICL

[CGG00] J. Chavarriga, J. Giné, and I. A. García. Isochronous centers of a linear center perturbed by fifth degree homogeneous polynomials. Journal of Computational and Applied Mathematics, 126 (1–2):351–368, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003647.

Cash:2002:MIR

[CGM02] J. R. Cash, M. P. Garcia, and D. R. Moore. Monoimplicit Runge–Kutta formulae for the numerical solution of second order nonlinear two-point boundary value problems. Journal of Computational and Applied Mathematics, 143(2):275–289, June 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100512X.

Castano-Garcia:2003:MHT

[CGMB03] Laura Castaño-García and Juan J. Moreno-Balcázar. A Mehler–Heine-type formula for Hermite–Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 150(1):25–35, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005526.

Chakrabarti:2000:SGH

[CH00] A. Chakrabarti and Hamsapriye. On some general hybrid transforms. Journal of Computational and Applied Mathematics, 116 (1):157–165, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042799003209.

Cabada:2002:ERD

[CH02a] A. Cabada and S. Heikkilä. Extremality results for discontinuous explicit and implicit diffusion problems. Journal of Computational and Applied Mathematics, 143 (1):69–80, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701005015.

Chyan:2002:TSB

[CH02b] Chuan Jen Chyan and Johnny Henderson. Twin solutions of boundary value problems for differential equations on measure chains. Journal of Computational and Applied Mathematics, 141(1–2):123–131, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270100440X.

Chen:2003:VMN

[CH03] Xiaojun Chen and Kouji Hashimoto. Verification methods for nonlinear equations with saddle point functions. Journal of Computational and Applied Mathematics, 159(1):13–24, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005703.

Chapko:2002:NSE

[Cha02] Roman Chapko. The numerical solution of an evolution problem of second order in time on a closed smooth boundary. Journal of Computational and Applied Mathematics, 145(2): 493–503, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702003515.

Chang:2004:NUB

[Cha04] Da-Wei Chang. A note on the upper bound of the spectral radius for SOR iteration matrix. Journal of Computational and Applied Mathematics, 167(1):251–253, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009002.

Cheikh:2001:SSL

[Che01] Youssèf Ben Cheikh. On some (n − 1)-symmetric linear functionals. Journal of Computational and Applied Mathematics, 133(1–2):207–218, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006476.

Cheikh:2003:RBH

[Che03a] Youssèf Ben Cheikh. Relations between harmonic analysis associated with two differential operators of different orders. Journal of Computational and Applied Mathematics, 153(1–2):61–71, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006301.

Chen:2003:UPC

[Che03b] Yanping Chen. Uniform pointwise convergence for a singularly perturbed problem using arc-length equidistribution. Journal of Computational and Applied Mathematics, 159(1):25–34, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005636.

Chibi:2001:DCD

[Chi01a]  Ahmed-Salah Chibi.  De-

fect correction and domain decomposition for secondorder boundary value problems. Journal of Computational and Applied Mathematics, 130(1–2):41–51, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003921.

Chihara:2001:YOP

[Chi01b] T. S. Chihara. 45 years of orthogonal polynomials: a view from the wings. Journal of Computational and Applied Mathematics, 133(1–2):13–21, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006324.

Chihara:2001:CSO

[Chi01c] T. S. Chihara. Convergent sequences of orthogonal polynomials. Journal of Computational and Applied Mathematics, 133(1–2):682, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007160.

Chihara:2001:SPI

[Chi01d] T. S. Chihara. Shell polynomials and indeterminate moment problems. Journal of Computational and Applied Mathematics, 133 (1–2):680–681, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700007159.

Chihara:2003:ABN

[Chi03] T. S. Chihara. An analog of the Blumenthal–Nevai theorem for unbounded intervals. Journal of Computational and Applied Mathematics, 153(1–2):535–536, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200883X.

Chouikha:2003:NTE

[Cho03] A. Raouf Chouikha. Note on trigonometric expansions of theta functions. Journal of Computational and Applied Mathematics, 153(1– 2):119–125, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702006313.

Christiansen:2001:DNU

[Chr01] Søren Christiansen. Detecting non-uniqueness of solutions to biharmonic integral equations through SVD. Journal of Computational and Applied Mathematics, 134 (1–2):23–35, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004519.

Chen:2002:CNP

[CHW02] Ke Chen, Anwar Hussein, and Hai-Bin Wan. On a class of new and practical performance indexes for approximation of fold bifurcations of nonlinear power flow equations. Journal of Computational and Applied Mathematics, 140(1–2):119–141, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004022.

Conti:2000:NSM

[CJ00] C. Conti and K. Jetter. A new subdivision method for bivariate splines on the fourdirectional mesh. Journal of Computational and Applied Mathematics, 119 (1–2):81–96, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700003721.

Chollom:2003:CTS

[CJ03] J. Chollom and Z. Jackiewicz. Construction of twostep Runge–Kutta methods with large regions of absolute stability. Journal of Computational and Applied Mathematics, 157(1):125–137, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703003820.

Chen:2004:ARS

[CJ04] Guizhi Chen and Zhongxiao Jia. An analogue of the results of Saad and Stewart for harmonic Ritz vectors. Journal of Computational and Applied Mathematics, 167(2):493–498, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009403.

Clavero:2003:UCS

[CJL03] C. Clavero, J. C. Jorge, and F. Lisbona. A uniformly convergent scheme on a nonuniform mesh for convection–diffusion parabolic problems. Journal of Computational and Applied Mathematics, 154(2):415–429, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042702008610.

Cools:2001:RIC

[CK01] Ronald Cools and K. J.

Kim. Rotation invariant cubature formulas over the ndimensional unit cube. Journal of Computational and Applied Mathematics, 132 (1):15–32, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005951.

Choe:2002:FRT

[CK02a] Geon Ho Choe and Dong Han Kim. The first return time test of pseudorandom numbers. Journal of Computational and Applied Mathematics, 143(2):263–274, June 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005106.

Cvijovic:2002:IRR

[CK02b] Djurdje Cvijović and Jacek Klinowski. Integral representations of the Riemann zeta function for odd-integer arguments. Journal of Computational and Applied Mathematics, 142(2):435–439, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042702003588.

Chen:2003:NSM

[CK03] Tsu-Fen Chen and Miao-Chen Ke. Nonlinear SOR and mesh redistribution. Journal of Computational and Applied Mathematics, 152(1– 2):51–68, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702006969.

Chapko:2004:CLM

[CK04] Roman Chapko and Philipp Kügler. A comparison of the Landweber method and the Gauss–Newton method for an inverse parabolic boundary value problem. Journal of Computational and Applied Mathematics, 169(1):183–196, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703010240.

Calvetti:2002:RAE

[CKR02] Daniela Calvetti, Sun-Mi Kim, and Lothar Reichel. The restarted QR-algorithm for eigenvalue computation of structured matrices. Journal of Computational and Applied Mathematics, 149 (2):415–422, December 15, 2002. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004867.

Chawla:2000:AOC

[CL00] Sanjay Chawla and Suzanne M.

Lenhart. Application of optimal control theory to bioremediation. Journal of Computational and Applied Mathematics, 114(1):81–102, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002903.

Cheng:2001:DPR

[CL01a] Sui Sun Cheng and Yi-Zhong Lin. Detection of positive roots of a polynomial with five parameters. Journal of Computational and Applied Mathematics, 137(1):19–48, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006282.

Chien:2001:MBG

[CL01b] C.-S. Chien and Y. S. Liao.

Multiple bifurcations generated by mode interactions in a reaction–diffusion problem. Journal of Computational and Applied Mathematics, 130(1–2):345–368, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003714.

Cui:2002:NSC

[CL02] B. T. Cui and Yongqing Liu. Necessary and sufficient conditions for oscillation of Robin boundary problems of neutral functional parabolic equations. Journal of Computational and Applied Mathematics, 146 (2):387–393, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702003722.

Chan:2003:PIC

[CL03a] Heng Huat Chan and Richard P. Lewis. Partition identities and congruences associated with the Fourier coefficients of the Euler products. Journal of Computational and Applied Mathematics, 160(1–2):69–75, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006186.

Chen:2003:MIM

[CL03b] Ren-Chuen Chen and Jinn-Liang Liu. Monotone iterative methods for the adaptive finite element solution of semiconductor equations. Journal of Computational and Applied Mathematics, 159(2):341–364, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005387.

Cui:2003:OPD

[CL03c] Bao Tong Cui and Yongqing Liu. Oscillation for partial difference equations with continuous variables. Journal of Computational and Applied Mathematics, 154(2):373–391, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008592.

Cai:2004:UCD

[CL04] X. Cai and F. Liu. Uniform convergence difference schemes for singularly perturbed mixed boundary problems. Journal of Computational and Applied Mathematics, 166(1):31–54, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008707.

Clarkson:2003:PEN

[Cla03] Peter A. Clarkson. Painlevé equations — nonlinear special functions. Journal of Computational and Applied Mathematics, 153(1– 2):127–140, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702005897.

Calvo:2000:NIO

[CLMR00] M. Calvo, M. P. Laburta, J. I. Montijano, and L. Rández. On the numerical integration of orthogonal flows with Runge–Kutta methods. Journal of Computational and Applied Mathematics, 115(1–2):121–135, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704279900182X.

Cao:2002:ASA

[CLP02] Yang Cao, Shengtai Li, and Linda Petzold. Adjoint sensitivity analysis for differentialalgebraic equations: algorithms and software. Journal of Computational and Applied Mathematics, 149 (1):171–191, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005289.

Conti:2001:BIP

[CM01] Costanza Conti and Rossana Morandi. The bidimensional interproximation problem and mixed splines. Journal of Computational and Applied Mathematics, 130 (1–2):1–16, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799003945.

Cravero:2003:SPI

[CM03] Isabella Cravero and Carla Manni. Shape-preserving interpolants with high smoothness. Journal of Computational and Applied Mathematics, 157(2):383–405, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004187.

Cousteix:2004:ANS

[CM04] J. Cousteix and J. Mauss. Approximations of the Navier– Stokes equations for high Reynolds number flows past a solid wall. Journal of Computational and Applied Mathematics, 166(1):101–122, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008677.

Coffey:2003:GCA

[CMKM03] T. Coffey, R. J. McMullan, C. T. Kelley, and D. S. McRae. Globally convergent algorithms for nonsmooth nonlinear equations in computational fluid dynamics. Journal of Computational and Applied Mathematics, 152(1–2):69–81, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006970.

Calvetti:2000:TRC

[CMRS00] D. Calvetti, S. Morigi, L. Reichel, and F. Sgallari. Tikhonov regularization and the Lcurve for large discrete illposed problems. Journal of Computational and Applied Mathematics, 123 (1–2):423–446, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004143.

Calvetti:2001:IME

[CMRS01] D. Calvetti, S. Morigi, L. Reichel, and F. Sgallari. An iterative method with error estimators. Journal of Computational and Applied Mathematics, 127(1–2):93–119, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004945.

Cools:2001:CFO

[CMS01] R. Cools, I. P. Mysovskikh, and H. J. Schmid. Cubature formulae and orthogonal polynomials. Journal of Computational and Applied Mathematics, 127(1–2):121–152, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004957.

Cash:2003:HSD

[CMSV03] J. R. Cash, D. R. Moore, N. Sumarti, and M. Van Daele. A highly stable deferred correction scheme with interpolant for systems of nonlinear two-point boundary value problems. Journal of Computational and Applied Mathematics, 155(2):339–358, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008737.

Chipot:2001:NAO

[CMV01] Michel Chipot, Riccardo March, and Domenico Vitulano. Numerical analysis of oscillations in a nonconvex problem related to image selective smoothing. Journal of Computational and Applied Mathematics, 136 (1–2):123–133, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005793.

Cabada:2000:FPA

[CN00] Alberto Cabada and Juan J. Nieto. Fixed points and approximate solutions for nonlinear operator equations. Journal of Computational and Applied Mathematics, 113(1–2):17–25, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704279900240X.

Cheung:2002:BCP

[CN02] William M. Cheung and Michael K. Ng. Blockcirculant preconditioners for systems arising from discretization of the threedimensional convection–diffusion equation. Journal of Computational and Applied Mathematics, 140(1–2):143–158, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005192.

Chan:2001:RRS

[CO01] E. S. Chan and B. H. Ong. Range restricted scattered data interpolation using convex combination of cubic Bézier triangles. Journal of Computational and Applied Mathematics, 136 (1–2):135–147, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000580X.

Calvo:2002:DBE

[CO02] Miquel Calvo and Josep M.

Oller. A distance between elliptical distributions based in an embedding into the Siegel group. Journal of Computational and Applied Mathematics, 145(2):319–334, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005842.

Cockburn:2001:DDG

[Coc01]  Bernardo Cockburn.  De-

vising discontinuous Galerkin methods for non-linear hyperbolic conservation laws. Journal of Computational and Applied Mathematics, 128(1–2):187–204, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005124.

Coffey:2003:SLC

[Cof03]  Mark W. Coffey.  On

some log–cosine integrals related to ζ(3), ζ(4), and ζ(6). Journal of Computational and Applied Mathematics, 159(2):205–215, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004382.

Coffey:2004:RPR

[Cof04] Mark W. Coffey. Relations and positivity results for the derivatives of the Riemann ξ function. Journal of Computational and Applied Mathematics, 166(2):525–534, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007970.

Collins:2001:RBP

[Col01] C. B. Collins. The role of Bell polynomials in integration. Journal of Computational and Applied Mathematics, 131(1–2):195–222, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002740.

Cools:2002:AMI

[Coo02]  Ronald Cools.  Advances

in multidimensional integration. Journal of Computational and Applied Mathematics, 149(1):1–12, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005174.

Cooper:2003:CTF

[Coo03] Shaun Cooper. Cubic theta functions. Journal of Computational and Applied Mathematics, 160(1–2):77–94, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042703006149.

Colton:2000:TEP

[CP00a] David Colton and Lassi Päivärinta. Transmission eigenvalues and a problem of Hans Lewy. Journal of Computational and Applied Mathematics, 117(2):91–104, May 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003349.

Cuevas:2000:ABV

[CP00b] Claudio Cuevas and Manuel Pinto. Asymptotic behavior in Volterra difference systems with unbounded delay. Journal of Computational and Applied Mathematics, 113(1–2):217–225, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002575.

Climent:2003:IML

[CP03] Joan-Josep Climent and Carmen Perea. Iterative methods for least-square problems based on proper splittings. Journal of Computational and Applied Mathematics, 158(1):43–48, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004655.

Calvo:2004:HOV

[CP04] M. P. Calvo and A. Portillo. Are high order variable step equistage initializers better than standard starting algorithms? Journal of Computational and Applied Mathematics, 169(2):333–344, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703010355.

Calvetti:2000:IML

[CR00] D. Calvetti and L. Reichel. Iterative methods for large continuation problems. Journal of Computational and Applied Mathematics, 123 (1–2):217–240, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004052.

Cramer:2001:CTD

[CR01] Michael Cramer and Ludger Rüschendorf. Convergence of two-dimensional branching recursions. Journal of Computational and Applied Mathematics, 130(1–2):53–73, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900391X.

Covello:2003:GFM

[CR03] Paul Covello and Garry Rodrigue. A generalized front marching algorithm for the solution of the eikonal equation. Journal of Computational and Applied Mathematics, 156(2):371–388, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003601.

Cabral-Rosetti:2000:GHF

[CRSL00] Luis G. Cabral-Rosetti and Miguel A. Sanchis-Lozano. Generalized hypergeometric functions and the evaluation of scalar one-loop integrals in Feynman diagrams. Journal of Computational and Applied Mathematics, 115(1–2):93–99, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001211.

Calvetti:2000:RLI

[CRSS00] D. Calvetti, L. Reichel, F. Sgallari, and G. Spaletta. A regularizing Lanczos iteration method for underdetermined linear systems. Journal of Computational and Applied Mathematics, 115(1–2):101–120, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002988.

Cahlon:2000:SSD

[CS00a] Baruch Cahlon and Darrell Schmidt. On stability of systems of delay differential equations. Journal of Computational and Applied Mathematics, 117(2):137–158, May 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003374.

Choi:2000:CCS

[CS00b] Junesang Choi and H. M. Srivastava. Certain classes of series associated with the zeta function and multiple gamma functions. Journal of Computational and Applied Mathematics, 118(1–2):87–109, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042700003113.

Chen:2001:SSS

[CS01] Heli Chen and Bernie D.

Shizgal. A spectral solution of the Sturm–Liouville equation: comparison of classical and nonclassical basis sets. Journal of Computational and Applied Mathematics, 136(1–2):17–35, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005732.

Chen:2002:GCT

[CS02] Xiongda Chen and Jie Sun.

Global convergence of a twoparameter family of conjugate gradient methods without line search. Journal of Computational and Applied Mathematics, 146(1):37–45, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004168.

Chu:2003:OBB

[CS03a] Wen-Chang Chu and H. M.

Srivastava. Ordinary and basic bivariate hypergeometric transformations associated with the Appell and Kampé de Fériet functions. Journal of Computational and Applied Mathematics, 156(2):355–370, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702009214.

Cusimano:2003:CBB

[CS03b] Catherine Cusimano and Sonya Stanley. Circular Bernstein–Bézier spline approximation with knot removal. Journal of Computational and Applied Mathematics, 155(1):177–185, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008993.

Cahlon:2004:SCC

[CS04a] Baruch Cahlon and Darrell Schmidt. Stability criteria for certain second-order delay differential equations with mixed coefficients. Journal of Computational and Applied Mathematics, 170 (1):79–102, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000147.

Causin:2004:MHF

[CS04b] Paola Causin and Riccardo Sacco. Mixed-hybrid finite elements for the simulation of thermal oxidation in semiconductors. Journal of Computational and Applied Mathematics, 168(1–2):95–105, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009671.

Cimrak:2004:IAS

[CS04c] Ivan Cimrák and Marián Slodicka. An iterative approximation scheme for the Landau–Lifshitz–Gilbert equation. Journal of Computational and Applied Mathematics, 169(1):17–32, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300935X.

Chamoret:2004:NSP

[CSRB04] D. Chamoret, P. Saillard, A. Rassineux, and J.-M. Bergheau. New smoothing procedures in contact mechanics. Journal of Computational and Applied Mathematics, 168(1–2):107–116, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009683.

Chen:2000:NGQ

[CT00] Ming-Po Chen and Kok-Keong Tan. A noncompact generalized quasi-variational inequality. Journal of Computational and Applied Mathematics, 113(1–2):309–315, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002630.

Cui:2002:SCC

[Cui02] Mingrong Cui. A sufficient condition for the convergence of the inexact Uzawa algorithm for saddle point problems. Journal of Computational and Applied Mathematics, 139(2):189–196, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004307.

Coussement:2001:SPM

[CV01] Els Coussement and Walter Van Assche. Some properties of multiple orthogonal polynomials associated with Macdonald functions. Journal of Computational and Applied Mathematics, 133(1–2):253–261, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006488.

Coussement:2003:AMO

[CV03] Els Coussement and Walter Van Assche. Asymptotics of multiple orthogonal polynomials associated with the modified Bessel functions of the first kind. Journal of Computational and Applied Mathematics, 153(1–2):141–149, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005964.

Caliari:2004:IDA

[CVB04] M. Caliari, M. Vianello, and L. Bergamaschi. Interpolating discrete advection–diffusion propagators at Leja sequences. Journal of Computational and Applied Mathematics, 172(1):79–99, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000998.

Chan:2000:RMM

[CW00] Tony F. Chan and W. L.

Wan. Robust multigrid methods for nonsmooth coefficient elliptic linear systems. Journal of Computational and Applied Mathematics, 123 (1–2):323–352, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004118.

Chao:2004:RSV

[CW04] Hongyang Chao and Ming Wei. Rate scalable video coding technology based on flexible block wavelet. Journal of Computational and Applied Mathematics, 163(1): 91–100, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703007933.

Chan:2000:CBD

[CY00] C. Y. Chan and Jian Yang. Complete blow-up for degenerate semilinear parabolic equations. Journal of Computational and Applied Mathematics, 113(1–2):353–364, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002666.

Cheikh:2003:SDO

[CZ03] Yousséf Ben Cheikh and Ali Zaghouani. Some discrete d-orthogonal polynomial sets. Journal of Computational and Applied Mathematics, 156(2):253–263, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702009147.

Chen:2004:NTR

[CZ04] Zhongwen Chen and Xiangsun Zhang. A nonmonotone trust-region algorithm with nonmonotone penalty parameters for constrained optimization. Journal of Computational and Applied Mathematics, 172(1):7–39, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000950.

Chen:2004:ESE

[CZW04] Shihua Chen, Qin Zhang, and Changping Wang. Existence and stability of equilibria of the continuoustime Hopfield neural network. Journal of Computational and Applied Mathematics, 169(1):117–125, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009555.

Chen:2003:EPP

[CZY03] Shihua Chen, Junhua Zhang, and Todd Young. Existence of positive periodic solution for nonautonomous predator–prey system with diffusion and time delay. Journal of Computational and Applied Mathematics, 159(2):375–386, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005405.

Daalhuis:2004:HOS

[Daa04] A. B. Olde Daalhuis. On higher-order Stokes phenomena of an inhomogeneous linear ordinary differential equation. Journal of Computational and Applied Mathematics, 169(1):235–246, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703010288.

Dachraoui:2001:WBT

[Dac01]  Azza Dachraoui.  Weyl–

Bessel transforms. Journal of Computational and Applied Mathematics, 133 (1–2):263–276, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000649X.

Dachraoui:2003:WTA

[Dac03] Azza Dachraoui. Weyl transforms associated with Laguerre functions. Journal of Computational and Applied Mathematics, 153(1– 2):151–162, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702006295.

deAmo:2003:NAP

[dACCS03] E. de Amo, I. Chitescu, M. Díaz Carrillo, and N. A. Secelean. A new approximation procedure for fractals. Journal of Computational and Applied Mathematics, 151(2):355–370, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007525.

Dahmen:2001:WMP

[Dah01] Wolfgang Dahmen. Wavelet methods for PDEs — some recent developments. Journal of Computational and Applied Mathematics, 128(1– 2):133–185, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005112.

Dai:2002:BMM

[Dai02]  Wanyang Dai.  A Brow-

nian model for multiclass queueing networks with finite buffers. Journal of Computational and Applied Mathematics, 144(1–2):145–160, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005568.

DallAglio:2001:DSO

[Dal01]  Marco Dall’Aglio.  The

Dubins–Spanier optimization problem in fair division theory. Journal of Computational and Applied Mathematics, 130(1–2):17–40, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003933.

DallOsso:2004:IBS

[Dal04] Aldo Dall’Osso. An iterative back substitution algorithm for the solution of tridiagonal matrix systems with fringes. Journal of Computational and Applied Mathematics, 169(1):87–99, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009452.

Damelin:2003:MMW

[Dam03] S. B. Damelin. On the maximum modulus of weighted polynomials in the plane, a theorem of rakhmanov, mhaskar and saff revisited. Journal of Computational and Applied Mathematics, 155(2):455–459, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702009081.

deAndrade:2003:ATS

[dAMR03] Eliana X. L. de Andrade, John H. McCabe, and A. Sri Ranga. The Q-D algorithm for transforming series expansions into a corresponding continued fraction: an extension to cope with zero coefficients. Journal of Computational and Applied Mathematics, 156(2):487–497, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003893.

Daras:2001:CPT

[Dar01] Nicholas J. Daras. Composed Padé-type approximation. Journal of Computational and Applied Mathematics, 134 (1–2):95–112, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005318.

Dattoli:2000:GPO

[Dat00] G. Dattoli. Generalized polynomials, operational identities and their applications. Journal of Computational and Applied Mathematics, 118(1–2):111–123, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002831.

Davini:2002:CEA

[Dav02] Cesare Davini. Γ-convergence of external approximations in boundary value problems involving the bi-Laplacian. Journal of Computational and Applied Mathematics, 140(1–2):185–208, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005258.

Dutt:2001:SMH

[DB01] Pravir K. Dutt and Smita Bedekar. Spectral methods for hyperbolic initial boundary value problems on parallel computers. Journal of Computational and Applied Mathematics, 134 (1–2):165–190, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005355.

deBruin:2002:RSI

[dB02] Marcel G. de Bruin. Regularity of some ‘incomplete’ Pál-type interpolation problems. Journal of Computational and Applied Mathematics, 145(2):407–415, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S037704270100591X.

deBruin:2003:SIP

[dB03] Marcel G. de Bruin. On a special interpolation problem and overconvergence. Journal of Computational and Applied Mathematics, 153(1–2):163–169, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006039.

Daya:2004:MMV

[DBDPF04] El Mostafa Daya, Bouazza Braikat, Noureddine Damil, and Michel Potier-Ferry. Modelisation of modulated vibration modes of repetitive structures. Journal of Computational and Applied Mathematics, 168(1–2):117–124, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009695.

deBruin:2001:BIN

[dBS01] Marcel G. de Bruin and A. Sharma. Birkhoff interpolation on nonuniformly distributed roots of unity II. Journal of Computational and Applied Mathematics, 133(1–2):295–303, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006464.

Coimbra:2000:MMA

[dCCSR00] Maria do Carmo Coimbra, Carlos Sereno, and Alírio Rodrigues. Modelling multicomponent adsorption process by a moving finite element method. Journal of Computational and Applied Mathematics, 115(1–2):169–179, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001715.

Driver:2001:ZUP

[DD01] Kathy Driver and Peter Duren. Zeros of ultraspherical polynomials and the Hilbert–Klein formulas. Journal of Computational and Applied Mathematics, 135(2):293–301, October 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005884.

Desjardins:2004:BLI

[DDG04] Benoît Desjardins, Emmanuel Dormy, and Emmanuel Grenier. Boundary layer instability at the top of the Earth’s outer core. Journal of Computational and Applied Mathematics, 166(1):123–131, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008665.

Dai:2003:ESW

[DDGH03a] X. Dai, Y. Diao, Q. Gu, and D. Han. The existence of subspace wavelet sets. Journal of Computational and Applied Mathematics, 155(1):83–90, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008932.

Dai:2003:FWS

[DDGH03b] X. Dai, Y. Diao, Q. Gu, and D. Han. Frame wavelet sets in R^{d}. Journal of Computational and Applied Mathematics, 155(1):69–82, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008920.

DeBaets:2001:CRC

[DDN01] B. De Baets, H. De Meyer, and H. Naessens. A class of rational cardinality-based similarity measures. Journal of Computational and Applied Mathematics, 132(1):51–69, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042700005963.

Duan:2000:WRC

[DDPT00] Qi Duan, K. Djidjeli, W. G. Price, and E. H. Twizell. Weighted rational cubic spline interpolation and its application. Journal of Computational and Applied Mathematics, 117(2):121–135, May 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003362.

DeSchutter:2000:MSS

[De 00] B. De Schutter. Minimal state-space realization in linear system theory: an overview. Journal of Computational and Applied Mathematics, 121 (1–2):331–354, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003411.

DeSchepper:2001:FEA

[De 01] Hennie De Schepper. Finite element analysis of a coupling eigenvalue problem on overlapping domains. Journal of Computational and Applied Mathematics, 132(1):141–153, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042700005987.

Dongarra:2000:NLA

[DE00] Jack J. Dongarra and Victor Eijkhout. Numerical linear algebra algorithms and software. Journal of Computational and Applied Mathematics, 123 (1–2):489–514, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004003.

Dehghan:2002:SOS

[Deh02]  Mehdi Dehghan.  Second-

order schemes for a boundary value problem with Neumann’s boundary conditions. Journal of Computational and Applied Mathematics, 138(1):173–184, January 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004520.

Dieci:2004:SSL

[DEL04] L. Dieci, C. Elia, and L. Lopez. Smooth SVD on the Lorentz group with application to computation of Lyapunov exponents. Journal of Computational and Applied Mathematics, 164– 165(??):255–264, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006447.

Demetriou:2002:SDD

[Dem02]  I. C. Demetriou.  Signs

of divided differences yield least squares data fitting with constrained monotonicity or convexity. Journal of Computational and Applied Mathematics, 146 (2):179–211, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702003539.

Demetriou:2004:TPC

[Dem04] I. C. Demetriou. A theorem for piecewise convex– concave data approximation. Journal of Computational and Applied Mathematics, 164– 165(??):245–254, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008252.

Devroye:2002:LLN

[Dev02] Luc Devroye. Laws of large numbers and tail inequalities for random tries and PATRI-CIA trees. Journal of Computational and Applied Mathematics, 142(1):27–37, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004587.

Dyn:2000:SSS

[DF00] Nira Dyn and Elza Farkhi.

Spline subdivision schemes for convex compact sets. Journal of Computational and Applied Mathematics, 119(1–2):133–144, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003757.

Deng:2002:BOS

[DFC02] Jiansong Deng, Yuyu Feng, and Falai Chen. Best onesided approximation of polynomials under L^{1} norm. Journal of Computational and Applied Mathematics, 144(1–2):161–174, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100557X.

Dyn:2002:ATB

[DFI02] N. Dyn, M. S. Floater, and A. Iske. Adaptive thinning for bivariate scattered data. Journal of Computational and Applied Mathematics, 145(2):505–517, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042702003527.

Bonis:2004:NMS

[DFM04] M. C. De Bonis, C. Frammartino, and G. Mastroianni. Numerical methods for some special Fredholm integral equations on the real line. Journal of Computational and Applied Mathematics, 164– 165(??):225–243, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006526.

deFrutos:2000:EPC

[dFN00] Javier de Frutos and Julia Novo. An enhanced pseudospectral Chebyshev method for dissipative partial differential equations. Journal of Computational and Applied Mathematics, 115(1– 2):137–150, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042799001739.

Domingo-Ferrer:2004:DRA

[DFT04] Josep Domingo-Ferrer and Vicenç Torra. Disclosure risk assessment in statistical data protection. Journal of Computational and Applied Mathematics, 164– 165(??):285–293, March 1, 2004. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006435.

deGraaf:2001:EEP

[dG01]  J. de Graaf.  Evolution

equations for polynomials and rational functions which are conformal on the unit disk. Journal of Computational and Applied Mathematics, 133(1–2):305–314, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006609.

deGroen:2001:STW

[dGK01] P. P. N. de Groen and G. E.

Karadzhov. Slow travelling waves on a finite interval for burgers’-type equations. Journal of Computational and Applied Mathematics, 132(1):155–189, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005975.

Daruis:2002:GQF

[DGVM02] Leyla Daruis, Pablo González-Vera, and Francisco Marcellán. Gaussian quadrature formulae on the unit circle. Journal of Computational and Applied Mathematics, 140(1–2):159–183, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004101.

DeCoster:2000:EMP

[DH00] C. De Coster and P. Habets. Existence and multiplicity of positive solutions of the Ginzburg–Landau boundary value problem. Journal of Computational and Applied Mathematics, 113 (1–2):317–327, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002642.

Dosly:2002:NSC

[DH02] Ondrej Doslý and Stefan Hilger. A necessary and sufficient condition for oscillation of the Sturm–Liouville dynamic equation on time scales. Journal of Computational and Applied Mathematics, 141(1–2):147–158, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004423.

Dostal:2002:IFP

[DHK02] Zdenek Dostál, Jaroslav Haslinger, and Radek Kucera. Implementation of the fixed point method in contact problems with Coulomb friction based on a dual splitting type technique. Journal of Computational and Applied Mathematics, 140(1–2):245–256, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004058.

Davis:2002:ULB

[DHP02] John M. Davis, Johnny Henderson, and K. Rajendra Prasad. Upper and lower bounds for the solution of the general matrix Riccati differential equation on a time scale. Journal of Computational and Applied Mathematics, 141(1–2):133–145, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004411.

DeSchepper:2004:P

[DHV04] H. De Schepper, M. Hogge, and R. Van Keer. Preface. Journal of Computational and Applied Mathematics, 168(1–2):xi–xii, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001499.

DeGersem:2004:FCC

[DHW04] Herbert De Gersem, Kay Hameyer, and Thomas Weiland. Field-circuit coupled models in electromagnetic simulation. Journal of Computational and Applied Mathematics, 168(1–2):125–133, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009701.

Deng:2004:SCI

[DHZ04] Yuanbei Deng, Xiyan Hu, and Lei Zhang. The solvability conditions for the inverse eigenvalue problem of the symmetrizable matrices. Journal of Computational and Applied Mathematics, 163(1):101–106, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007945.

Diatta:2003:MMC

[Dia03] Jean Diatta. A mixed measure of content on the set of real numbers. Journal of Computational and Applied Mathematics, 151(1):85–105, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007380.

Diethelm:1998:EBS

[Die98] Kai Diethelm. Error bounds for spline-based quadrature methods for strongly singular integrals. Journal of Computational and Applied Mathematics, 89(2):257–261, March 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797002458. See er-

ratum [Die02].

Diethelm:2002:EEB

[Die02] Kai Diethelm. Erratum to “Error bounds for splinebased quadrature methods for strongly singular integrals” [J. Comput. Appl. Math. 89 (1998) 257]. Journal of Computational and Applied Mathematics, 142(2):449–450, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003564.  See

[Die98].

Dikshit:2003:PTI

[Dik03] H. P. Dikshit. Pál-type interpolation on nonuniformly distributed nodes on the unit circle. Journal of Computational and Applied Mathematics, 155(2):253–261, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008683.

Dimitrov:2001:C

[Dim01a] Dimitar K. Dimitrov. Comment. Journal of Computational and Applied Mathematics, 133(1–2):685–686, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700007196.

Dimitrov:2001:CCZ

[Dim01b] Dimitar K. Dimitrov. Connection coefficients and zeros of orthogonal polynomials. Journal of Computational and Applied Mathematics, 133(1–2):331–340, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006531.

Dimitrov:2003:CEZ

[Dim03] Dimitar K. Dimitrov. Convexity of the extreme zeros of Gegenbauer and Laguerre polynomials. Journal of Computational and Applied Mathematics, 153(1– 2):171–180, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702006453.

Ding:2002:WSH

[Din02] Hongming Ding. The Nwidths of spaces of holomorphic functions on bounded symmetric domains, II. Journal of Computational and Applied Mathematics, 144(1– 2):175–186, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701005581.

Ding:2003:EAS

[Din03] Xie Ping Ding. Existence and algorithm of solutions for nonlinear mixed variationallike inequalities in Banach spaces. Journal of Computational and Applied Mathematics, 157(2):419–434, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004217.

Ding:2004:IAS

[Din04] Xie Ping Ding. Iterative algorithms for solving generalized nonlinear mixed variational inequalities. Journal of Computational and Applied Mathematics, 172 (2):399–408, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001293.

Damelin:2001:NCW

[DJK01a] S. B. Damelin, H. S. Jung, and K. H. Kwon. Necessary conditions for weighted mean convergence of Lagrange interpolation for exponential weights. Journal of Computational and Applied Mathematics, 132(2):357–369, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004398.

Damelin:2001:NMC

[DJK01b] S. B. Damelin, H. S. Jung, and K. H. Kwon. A note on mean convergence of Lagrange interpolation in L_{p}, (0 < p ≤ 1). Journal of Computational and Applied Mathematics, 133(1–2):277–282, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006506.

Damelin:2001:MCH

[DJK01c] S. B. Damelin, H. S. Jung, and K. H. Kwon. On mean convergence of Hermite– Fejér and Hermite interpolation for Erdős weights. Journal of Computational and Applied Mathematics, 137(1):71–76, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006981.

Dai:2003:CAN

[DKK03] Yang Dai, Sunyoung Kim, and Masakazu Kojima. Computing all nonsingular solutions of cyclic- n polynomial using polyhedral homotopy continuation methods. Journal of Computational and Applied Mathematics, 152(1–2):83–97, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006982.

Delgado:2004:SAN

[DKL04] Vladimir Delgado, Guillaume Kasperski, and Gérard Labrosse. Spectrally accurate numerical incompressible flows with deformed boundaries: a finite volume preconditioning method. Journal of Computational and Applied Mathematics, 168(1–2):145–154, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009725.

Deift:2001:RHA

[DKM^{+}01] P. Deift, T. Kriecherbauer, K. T.-R. McLaughlin, S. Venakides, and X. Zhou. A Riemann–Hilbert approach to asymptotic questions for orthogonal polynomials. Journal of Computational and Applied Mathematics, 133(1– 2):47–63, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700006348.

Du:2004:DBL

[DKST04] N. H. Du, R. Kon, K. Sato, and Y. Takeuchi. Dynamical behavior of Lotka–Volterra competition systems: nonautonomous bistable case and the effect of telegraph noise. Journal of Computational and Applied Mathematics, 170 (2):399–422, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000871.

Ding:2000:PPP

[DL00] Xie Ping Ding and Chun Lin Lou. Perturbed proximal point algorithms for general quasi-variational-like inclusions. Journal of Computational and Applied Mathematics, 113(1–2):153–165, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002502.

Dagnino:2001:APB

[DL01a] C. Dagnino and P. Lamberti. On the approximation power of bivariate quadratic C^{1} splines. Journal of Computational and Applied Mathematics, 131(1–2):321–332, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000265X.

Driver:2001:ZHP

[DL01b] K. A. Driver and A. D.

Love. Zeros of _{3}F_{2} hypergeometric polynomials. Journal of Computational and Applied Mathematics, 131(1– 2):243–251, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700002727.

Daehlen:2000:MAT

[DLM^{+}00] M. Dæhlen, T. Lyche, K. Mørken, R. Schneider, and H-P. Seidel. Multiresolution analysis over triangles, based on quadratic Hermite interpolation. Journal of Computational and Applied Mathematics, 119(1–2):97–114, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003733.

DeMarchi:2002:IDS

[DLTS02] S. De Marchi, A. Ligun, S. Timchenko, and A. Shumeiko. An interpolant defined by subdivision and the analysis of the error. Journal of Computational and Applied Mathematics, 145(1):71–88, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005362.

Ding:2002:FAM

[DLZ02] Jiu Ding, Tien Yien Li, and Aihui Zhou. Finite approximations of Markov operators. Journal of Computational and Applied Mathematics, 147(1):137–152, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004296.

DelBuono:2002:EMB

[DM02a] N. Del Buono and C. Mastroserio. Explicit methods based on a class of four stage fourth order Runge–Kutta methods for preserving quadratic laws. Journal of Computational and Applied Mathematics, 140(1–2):231–243, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003983.

Driver:2002:QCT

[DM02b] Kathy Driver and Manfred Möller. Quadratic and cubic transformations and zeros of hypergeometric polynomials. Journal of Computational and Applied Mathematics, 142(2):411–417, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004150.

Dehesa:2001:QIE

[DMFSR01] Jesús S. Dehesa, Andrei Martínez-Finkelshtdein, and Jorge Sánchez-Ruiz. Quantum information entropies and orthogonal polynomials. Journal of Computational and Applied Mathematics, 133(1–2):23–46, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006336.

Diaz-Mendoza:2001:PAQ

[DMGVO01] C. Díaz-Mendoza, P. González-Vera, and R. Orive. Padé approximants and quadratures related to certain strong distributions. Journal of Computational and Applied Mathematics, 133(1–2):315–329, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006695.

Duran:2001:NSC

[DMN01] Mario Durán, Marcela Miguez, and Jean-Claude Nédélec. Numerical stability in the calculation of eigenfrequencies using integral equations. Journal of Computational and Applied Mathematics, 130(1–2):323–336, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003738.

deMendonca:2003:ICU

[dMPL03] Luziane Ferreira de Mendonça, Rosana Pérez, and Véra Lucia Rocha Lopes. Inverse q-Columns Updating Methods for solving nonlinear systems of equations. Journal of Computational and Applied Mathematics, 158 (2):317–337, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004515.

Dreyer:2000:MOA

[DMS00] Thomas Dreyer, Bernd Maar, and Volker Schulz. Multigrid optimization in applications. Journal of Computational and Applied Mathematics, 120(1–2):67–84, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003046.

Dmytryshyn:2004:MGF

[Dmy04] R. I. Dmytryshyn. The multidimensional generalization of g-fractions and their application. Journal of Computational and Applied Mathematics, 164–165(??):265–284, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006423.

Davydov:2000:ISD

[DN00] Oleg Davydov and Günther Nürnberger. Interpolation by C^{1} splines of degree q ≥ 4 on triangulations. Journal of Computational and Applied Mathematics, 126 (1–2):159–183, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003507.

Dai:2001:CFD

[DN01] Weizhong Dai and Raja Nassar. A compact finitedifference scheme for solving a one-dimensional heat transport equation at the microscale. Journal of Computational and Applied Mathematics, 132(2):431–441, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042700004453.

Dai:2002:USF

[DN02a] Weizhong Dai and Raja Nassar. An unconditionally stable finite difference scheme for solving a 3D heat transport equation in a sub-microscale thin film. Journal of Computational and Applied Mathematics, 145(1):247–260, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005799.

Davies:2002:SOS

[DN02b] E. B. Davies and Jiban Nath. Schrödinger operators with slowly decaying potentials. Journal of Computational and Applied Mathematics, 148(1):1–28, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005708.

Duflot:2004:FCG

[DND04] Marc Duflot and Hung Nguyen-Dang. Fatigue crack growth analysis by an enriched meshless method. Journal of Computational and Applied Mathematics, 168(1–2):155–164, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042703009749.

Dostal:2000:DBD

[DNS00] Zdenek Dostál, Francisco A.

M. Gomes Neto, and Sandra A. Santos. Duality-based domain decomposition with natural coarse-space for variational inequalities. Journal of Computational and Applied Mathematics, 126 (1–2):397–415, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003684.

Doha:2002:CIE

[Doh02] E. H. Doha. On the coefficients of integrated expansions and integrals of ultraspherical polynomials and their applications for solving differential equations. Journal of Computational and Applied Mathematics, 139 (2):275–298, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004204.

Doi:2002:MMP

[Doi02] Masao Doi. Material modeling platform. Journal of Computational and Applied Mathematics, 149(1):13–25, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005186.

Dominguez:2003:HOC

[Dom03]  Víctor Domínguez.  High-

order collocation and quadrature methods for some logarithmic kernel integral equations on open arcs. Journal of Computational and Applied Mathematics, 161 (1):145–159, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005831.

DeMatteis:2002:PP

[DP02a] A. De Matteis and S. Pagnutti. Pseudorandom permutation. Journal of Computational and Applied Mathematics, 142(2):367–375, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004253.

Ding:2002:NCG

[DP02b] Xie Ping Ding and Jong Yeoul Park. A new class of generalized nonlinear implicit quasivariational inclusions with fuzzy mappings. Journal of Computational and Applied Mathematics, 138 (2):243–257, January 15, 2002. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100379X.

DAlessio:2004:TEG

[DPM04] S. J. D. D’Alessio, J. P. Pascal, and T. B. Moodie. Thermally enhanced gravity driven flows. Journal of Computational and Applied Mathematics, 170(1):1–25, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000111.

Dragut:2002:HAB

[Dra02]  A. B. Dragut.  Heuristic

allocation based on a dynamic programming statespace representation. Journal of Computational and Applied Mathematics, 140(1– 2):257–273, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701005994.

Dragut:2004:WMB

[Dra04] A. B. Dragut. A weakly monotonic backward induction algorithm on finite bounded subsets of vector lattices. Journal of Computational and Applied Mathematics, 164– 165(??):295–306, March 1, 2004. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009488.

Dattoli:2002:NMI

[DRC02] G. Dattoli, P. E. Ricci, and C. Cesarano. A note on multi-index polynomials of Dickson type and their applications in quantum optics. Journal of Computational and Applied Mathematics, 145(2):417–424, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005921.

delRey:2003:RFO

[dRM03] A. Martín del Rey and J. Muñoz Masqué. Reducing y′′ = F (x, y, y′) to first order. Journal of Computational and Applied Mathematics, 155(2):389–404, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008762.

Dette:2002:NMV

[DS02a] Holger Dette and William J. Studden. A note on the matrix valued q-d algorithm and matrix orthogonal polynomials on [0, 1] and [0, ∞). Journal of Computational and Applied Mathematics, 148 (2):349–361, November 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005551.

Diblik:2002:ECP

[DS02b] Josef Diblík and Zdenek Svoboda. An existence criterion of positive solutions of ptype retarded functional differential equations. Journal of Computational and Applied Mathematics, 147 (2):315–331, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004399.

Didenko:2002:SAM

[DS02c] V. D. Didenko and B. Silbermann. On stability of approximation methods for the Muskhelishvili equation. Journal of Computational and Applied Mathematics, 146 (2):419–441, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200393X.

Davies:2004:USV

[DS04] Philip I. Davies and Matthew I.

Smith. Updating the singular value decomposition. Journal of Computational and Applied Mathematics, 170 (1):145–167, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000172.

DeProft:2004:DMC

[DSD04] K. De Proft, L. J. Sluys, and W. P. De Wilde. A discrete model for cyclic mode I loading. Journal of Computational and Applied Mathematics, 168(1–2):135–144, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009713.

deSilanes:2001:SAD

[dSPPT01] M. C. López de Silanes, M. C. Parra, M. Pasadas, and J. J. Torrens. Spline approximation of discontinuous multivariate functions from scattered data. Journal of Computational and Applied Mathematics, 131(1–2):281–298, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002685.

deSilanes:2000:CCP

[dSPT00] M. C. López de Silanes, M. Pasadas, and J. J. Torrens. Construction of curves with parallelism conditions.

Journal of Computational and Applied Mathematics, 115(1–2):369–382, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001181.

deSilanes:2002:VOF

[dSPT02] María Cruz López de Silanes, María Cruz Parra, and Juan José Torrens. Vertical and oblique fault detection in explicit surfaces. Journal of Computational and Applied Mathematics, 140(1–2):559–585, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100601X.

Davydov:2000:IBL

[DSS00] Oleg Davydov, Manfred Sommer, and Hans Strauss. Interpolation by bivariate linear splines. Journal of Computational and Applied Mathematics, 119(1–2):115–131, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003745.

Denis:2003:CTP

[DSS03] Remy Y. Denis, S. N. Singh, and S. P. Singh. On certain transformations of poly-basic bilateral hypergeometric series. Journal of Computational and Applied Mathematics, 160 (1–2):95–101, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006150.

Danilin:2004:IAC

[DSVB04] A. N. Danilin, D. V. Snegovski, and D. B. Volkov-Bogorodski. On implicit algorithms of continuation method with applications to dynamic systems. Journal of Computational and Applied Mathematics, 164– 165(??):207–224, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009518.

Dong:2003:TRA

[DSW^{+}03] Jichang Dong, Jianming Shi, Shouyang Wang, Yi Xue, and Shuqin Liu. A trustregion algorithm for equalityconstrained optimization via a reduced dimension approach. Journal of Computational and Applied Mathematics, 152(1–2):99–118, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006994.

Dellnitz:2002:LAZ

[DSZ02] Michael Dellnitz, Oliver Schütze, and Qinghua Zheng. Locating all the zeros of an analytic function in one complex variable. Journal of Computational and Applied Mathematics, 138(2):325–333, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003715.

Duff:2000:IHP

[Duf00] Iain S. Duff. The impact of high-performance computing in the solution of linear systems: trends and problems. Journal of Computational and Applied Mathematics, 123 (1–2):515–530, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004015.

Dular:2004:BNE

[Dul04] Patrick Dular. The benefits of nodal and edge elements coupling for discretizing global constraints in dual magnetodynamic formulations. Journal of Computational and Applied Mathematics, 168(1–2):165–178, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042703009750.

Dunkl:2003:SFG

[Dun03] Charles F. Dunkl. Special functions and generating functions associated with reflection groups. Journal of Computational and Applied Mathematics, 153(1–2):181–190, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006398.

DeSchepper:2000:NAS

[DV00] Hennie De Schepper and Roger Van Keer. On the numerical approximation of some types of nonstandard second-order eigenvalue problems for vector valued functions. Journal of Computational and Applied Mathematics, 115(1–2):151–167, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001223.

Dassios:2001:CFD

[DV01] G. Dassios and P. Vafeas. Connection formulae for differential representations in Stokes flow. Journal of Computational and Applied Mathematics, 133(1–2):283–294, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006518.

Dang:2001:CIP

[DvM01] Chuangyin Dang and Hans van Maaren. Computing an integer point of a simplex with an arbitrary starting homotopy-like simplicial algorithm. Journal of Computational and Applied Mathematics, 129(1–2):151–170, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005471.

DeBonis:2002:AHT

[DVM02] M. C. De Bonis, B. Della Vecchia, and G. Mastroianni. Approximation of the Hilbert transform on the real semiaxis using Laguerre zeros. Journal of Computational and Applied Mathematics, 140(1–2):209–229, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005295.

Diao:2004:DRD

[DWQ04] Huaian Diao, Yimin Wei, and Sanzheng Qiao. Displacement rank of the Drazin inverse. Journal of Computational and Applied Mathematics, 167(1):147–161, May 1, 2004. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008938.

Duckworth:2002:MIM

[DWZ02] W. Duckworth, N. C. Wormald, and M. Zito. Maximum induced matchings of random cubic graphs. Journal of Computational and Applied Mathematics, 142(1):39–50, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004575.

Ding:2002:NCC

[DX02] Xie Ping Ding and Fu Quan Xia. A new class of completely generalized quasi-variational inclusions in Banach spaces. Journal of Computational and Applied Mathematics, 147(2):369–383, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004430.

Dai:2000:HFE

[DZN00] Weizhong Dai, Yukong Zhang, and Raja Nassar. A hybrid finite element-alternating direction implicit method for solving parabolic differential equations on multilayers with irregular geometry. Journal of Computational and Applied Mathematics, 117(1):1–16, May 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003258.

Eiermann:2000:AAS

[EES00] Michael Eiermann, Oliver G.

Ernst, and Olaf Schneider. Analysis of acceleration strategies for restarted minimal residual methods. Journal of Computational and Applied Mathematics, 123 (1–2):261–292, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003988.

Edwards:2002:BSS

[EF02] John T. Edwards and Neville J.

Ford. Boundedness and stability of solutions to difference equations. Journal of Computational and Applied Mathematics, 140(1– 2):275–289, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004800.

Edwards:2002:NSL

[EFS02] John T. Edwards, Neville J.

Ford, and A. Charles Simpson. The numerical solution of linear multi-term fractional differential equations: systems of equations. Journal of Computational and Applied Mathematics, 148 (2):401–418, November 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005587.

Elschner:2000:NMI

[EG00] J. Elschner and I. G. Graham. Numerical methods for integral equations of Mellin type. Journal of Computational and Applied Mathematics, 125 (1–2):423–437, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004830.

Egartner:2000:WRO

[Ega00] Wolfgang Egartner. Working range optimization for turbine and compressor blading. Journal of Computational and Applied Mathematics, 120(1–2):59–65, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003034.

Eymard:2003:LBC

[EGV03] Robert Eymard, Thierry Gallouët, and Jullien Vovelle. Limit boundary conditions for finite volume approximations of some physical problems. Journal of Computational and Applied Mathematics, 161(2):349–369, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007052.

Ezquerro:2000:MSH

[EH00] J. A. Ezquerro and M. A.

Hernández. A modification of the super-Halley method under mild differentiability conditions. Journal of Computational and Applied Mathematics, 114(2):405–409, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003489.

Ezquerro:2004:HTI

[EH04] J. A. Ezquerro and M. A.

Hernández. On Halleytype iterations with free second derivative. Journal of Computational and Applied Mathematics, 170 (2):455–459, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001281.

Ehrenmark:2001:NRS

[Ehr01a] Ulf T. Ehrenmark. A note on a recent study of oscillatory integration rules. Journal of Computational and Applied Mathematics, 131(1–2):493–496, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003557.

Ehrich:2001:SFG

[Ehr01b] Sven Ehrich. Stopping functionals for Gaussian quadrature formulas. Journal of Computational and Applied Mathematics, 127(1–2): 153–171, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700004969.

Ehrich:2002:SEG

[Ehr02]  Sven Ehrich.  On strat-

ified extensions of Gauss– Laguerre and Gauss–Hermite quadrature formulas. Journal of Computational and Applied Mathematics, 140(1– 2):291–299, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701004071.

Ezquerro:2000:DSS

[EHS00] J. A. Ezquerro, M. A. Hernández, and M. A. Salanova. A discretization scheme for some conservative problems. Journal of Computational and Applied Mathematics, 115(1–2):181–192, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001156.

Elnagar:2001:NST

[EK01] Gamal N. Elnagar and Mohammad A. Kazemi. Numerical solution of timedelayed functional differential equation control systems. Journal of Computational and Applied Mathematics, 130(1–2):75–90, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003908.

Everitt:2001:OPS

[EKLW01] W. N. Everitt, K. H. Kwon, L. L. Littlejohn, and R. Wellman. Orthogonal polynomial solutions of linear ordinary differential equations. Journal of Computational and Applied Mathematics, 133(1–2):85–109, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006361.

ElBernoussi:2004:BER

[El 04] Souad El Bernoussi. Best ellipsoidal relaxation to solve a nonconvex problem. Journal of Computational and Applied Mathematics, 162(1): 183–191, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703007301.

Elbert:2001:CZB

[EL01] Árpád Elbert and Andrea Laforgia. A conjecture on the zeros of Bessel functions. Journal of Computational and Applied Mathematics, 133(1–2):683, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700007172.

Elbert:2001:SRR

[Elb01] Á. Elbert. Some recent results on the zeros of Bessel functions and orthogonal polynomials. Journal of Computational and Applied Mathematics, 133(1–2):65–83, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000635X.

Elizalde:2000:ZFF

[Eli00]  E. Elizalde.  Zeta func-

tions: formulas and applications. Journal of Computational and Applied Mathematics, 118(1–2):125–142, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002843.

Ellinas:2001:QDA

[Ell01] Demosthenes Ellinas. Quantum diffusions and Appell systems. Journal of Computational and Applied Mathematics, 133(1–2):341–353, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000652X.

Eloe:2002:MQD

[Elo02] Paul W. Eloe. The method of quasilinearization and dynamic equations on compact measure chains. Journal of Computational and Applied Mathematics, 141(1– 2):159–167, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004435.

Engelborghs:2000:NBA

[ELR00] Koen Engelborghs, Tatyana Luzyanina, and Dirk Roose.

Numerical bifurcation analysis of delay differential equations. Journal of Computational and Applied Mathematics, 125 (1–2):265–275, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004726.

Elbert:2001:CZU

[ELS01] Árpád Elbert, Andrea Laforgia, and Panayiotis Siafarikas. A conjecture on the zeros of ultraspherical polynomials. Journal of Computational and Applied Mathematics, 133(1–2):684, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007184.

Everitt:2000:LDS

[ELW00] W. N. Everitt, L. L. Littlejohn, and R. Wellman. The left-definite spectral theory for the classical Hermite differential equation. Journal of Computational and Applied Mathematics, 121 (1–2):313–330, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003356.

Everitt:2002:LPL

[ELW02] W. N. Everitt, L. L. Littlejohn, and R. Wellman. Legendre polynomials, Legendre– Stirling numbers, and the leftdefinite spectral analysis of the Legendre differential expression. Journal of Computational and Applied Mathematics, 148(1):213–238, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005824.

Everitt:2004:SOS

[ELW04] W. N. Everitt, L. L. Littlejohn, and R. Wellman. The Sobolev orthogonality and spectral analysis of the Laguerre polynomials L_{n}^{−k} for positive integers k. Journal of Computational and Applied Mathematics, 171(1–2):199–234, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000573.

Ehrhardt:2004:SDA

[EM04a] Matthias Ehrhardt and Ronald E.

Mickens. Solutions to the discrete Airy equation: application to parabolic equation calculations. Journal of Computational and Applied Mathematics, 172 (1):183–206, November 1, 2004. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001232.

El-Mikkawy:2004:FAE

[EM04b] Moawwad E. A. El-Mikkawy. A fast algorithm for evaluating n-th order tri-diagonal determinants. Journal of Computational and Applied Mathematics, 166(2):581–584, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007726.

Abd-el-Malek:2002:UGT

[eMBH02] Mina B. Abd el Malek, Nagwa A. Badran, and Hossam S. Hassan. Using group theoretic method to solve multi-dimensional diffusion equation. Journal of Computational and Applied Mathematics, 147(2):385–395, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004740.

Abd-el-Malek:2000:GTM

[eMEM00] Mina B. Abd el Malek and Samy M. A. El-Mansi. Group theoretic methods applied to Burgers’ equation. Journal of Computational and Applied Mathematics, 115(1–2):1–12, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001703.

Abd-el-Malek:2002:GTA

[eMKM02] Mina B. Abd el Malek, Magda M. Kassem, and Mohammed L. M. Meky. Group theoretic approach for solving the problem of diffusion of a drug through a thin membrane. Journal of Computational and Applied Mathematics, 140(1–2):1–11, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005167.

Abd-el-Malek:2004:SSU

[eMKM04] Mina B. Abd el Malek, Magda M. Kassem, and Mohammad L. Mekky. Similarity solutions for unsteady freeconvection flow from a continuous moving vertical surface. Journal of Computational and Applied Mathematics, 164–165(??):11–24, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006381.

Evans:2001:AIM

[EMT01] D. J. Evans, M. M. Martins, and M. E. Trigo. The AOR iterative method for new preconditioned linear systems. Journal of Computational and Applied Mathematics, 132(2):461–466, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004477.

Essaouini:2004:NMS

[ENE04] M. Essaouini, A. Nachaoui, and S. El Hajji. Numerical method for solving a class of nonlinear elliptic inverse problems. Journal of Computational and Applied Mathematics, 162(1):165–181, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007313.

Enright:2000:CNM

[Enr00] W. H. Enright. Continuous numerical methods for ODEs with defect control. Journal of Computational and Applied Mathematics, 125 (1–2):159–170, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004660.

Erbe:2002:OCS

[EP02a] Lynn Erbe and Allan Peterson. Oscillation criteria for second-order matrix dynamic equations on a time scale. Journal of Computational and Applied Mathematics, 141(1–2):169–185, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004447.

Everitt:2002:KAK

[EP02b] W. N. Everitt and Anthippi Poulkou. Kramer analytic kernels and firstorder boundary value problems. Journal of Computational and Applied Mathematics, 148(1):29–47, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200571X.

Everitt:2003:SOR

[EP03] W. N. Everitt and Anthippi Poulkou. Some observations and remarks on differential operators generated by firstorder boundary value problems. Journal of Computational and Applied Mathematics, 153(1–2):201–211, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005903.

Erbe:2000:EMN

[EPM00] Lynn Erbe, Allan Peterson, and Ronald Mathsen. Existence, multiplicity, and nonexistence of positive solutions to a differential equation on a measure chain. Journal of Computational and Applied Mathematics, 113 (1–2):365–380, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002678.

ElHajji:2004:SIL

[ERV04] Said El Hajji, Nathalie Revol, and Paul Van Dooren. Special issue on Linear Algebra and Arithmetic, Proceedings of the ALA’01 Conference held in Rabat, Morocco on May 28– 31, 2001. Journal of Computational and Applied Mathematics, 162(1):ix–x, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007428.

Escribano:2003:MIC

[EST03] C. Escribano, M. A. Sastre, and E. Torrano. Moments of infinite convolutions of symmetric Bernoulli distributions. Journal of Computational and Applied Mathematics, 153(1–2):191–199, April 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005952.

Evans:2004:MQU

[Eva04]  G. A. Evans.  Multiple

quadrature using highly oscillatory quadrature methods. Journal of Computational and Applied Mathematics, 163(1):1–13, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300788X.

Everitt:2004:PHC

[Eve04] W. N. Everitt. A personal history of the m-coefficient. Journal of Computational and Applied Mathematics, 171 (1–2):185–197, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000494.

Ewing:2001:SNM

[EW01] Richard E. Ewing and Hong Wang. A summary of numerical methods for timedependent advection-dominated partial differential equations. Journal of Computational and Applied Mathematics, 128(1–2):423–445, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005227.

Fabiano:2000:CFA

[Fab00]  R. H. Fabiano.  Coer-

cive forms associated with elastic systems with damping. Journal of Computational and Applied Mathematics, 114(1):159–171, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002940.

Fabijonas:2002:LMC

[Fab02] Bruce R. Fabijonas. Laplace’s method on a computer algebra system with an application to the real valued modified Bessel functions. Journal of Computational and Applied Mathematics, 146 (2):323–342, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702003643.

Fang:2003:CFD

[Fan03a]  Qing Fang.  Convergence

of finite difference methods for convection–diffusion problems with singular solutions. Journal of Computational and Applied Mathematics, 152(1–2):119–131, March 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007008.

Fang:2003:NCN

[Fan03b] Qing Fang. A note on the condition number of a matrix. Journal of Computational and Applied Mathematics, 157(1):231–234, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004345.

Farouki:2000:LBB

[Far00] Rida T. Farouki. Legendre– Bernstein basis transformations. Journal of Computational and Applied Mathematics, 119(1–2):145–160, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003769.

Farkas:2002:NSR

[Far02] Gyula Farkas. A numerical C^{1}- shadowing result for retarded functional differential equations. Journal of Computational and Applied Mathematics, 145(2):269–289, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042701005817.

Fasino:2002:ANF

[Fas02] Dario Fasino. Approximation of nonnegative functions by means of exponentiated trigonometric polynomials. Journal of Computational and Applied Mathematics, 140(1–2):315–329, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100406X.

Fazio:2002:SFB

[Faz02] Riccardo Fazio. A survey on free boundary identification of the truncated boundary in numerical BVPs on infinite intervals. Journal of Computational and Applied Mathematics, 140(1–2):331–344, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100526X.

Fortin:2004:SCP

[FC04] Jérôme Fortin and Patrice Coorevits. Selecting contact particles in dynamics granular mechanics systems. Journal of Computational and Applied Mathematics, 168(1–2):207–213, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009786.

Fikioris:2002:IRB

[FCP02] G. Fikioris, P. G. Cottis, and A. D. Panagopoulos. On an integral related to biaxially anisotropic media. Journal of Computational and Applied Mathematics, 146 (2):343–360, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003680.

Fahroo:2000:OAS

[FD00] Fariba Fahroo and Michael A.

Demetriou. Optimal actuator/sensor location for active noise regulator and tracking control problems. Journal of Computational and Applied Mathematics, 114 (1):137–158, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002939.

Ferdinand:2003:NSA

[Fer03] Robert R. Ferdinand. Numerical scheme approximating solution and parameters in a beam equation. Journal of Computational and Applied Mathematics, 161 (2):469–476, December 15, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007374.

Farhloul:2002:RCM

[FF02] M. Farhloul and M. Fortin. Review and complements on mixed-hybrid finite element methods for fluid flows. Journal of Computational and Applied Mathematics, 140(1–2):301–313, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005209.

Fulton:2004:AML

[FFX04] S. R. Fulton, A. S. Fokas, and C. A. Xenophontos. An analytical method for linear elliptic PDEs and its numerical implementation. Journal of Computational and Applied Mathematics, 167(2):465–483, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009336.

Fasshauer:2000:ADN

[FGJ00] Gregory E. Fasshauer, Eugene C. Gartland, Jr., and Joseph W. Jerome. Algorithms defined by Nash iteration: some implementations via multilevel collocation and smoothing. Journal of Computational and Applied Mathematics, 119(1–2):161–183, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003770.

Farrell:2004:SPC

[FHM^{+}04] P. A. Farrell, A. F. Hegarty, J. J. H. Miller, E. O’Riordan, and G. I. Shishkin. Singularly perturbed convection–diffusion problems with boundary and weak interior layers. Journal of Computational and Applied Mathematics, 166 (1):133–151, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703008653.

Farouki:2004:CCH

[FHMS04] Rida T. Farouki, Chang Yong Han, Carla Manni, and Alessandra Sestini. Characterization and construction of helical polynomial space curves. Journal of Computational and Applied Mathematics, 162(2):365–392, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007532.

Fang:2003:ICR

[FI03] Weifu Fang and Kazufumi Ito. Identification of contact regions in semiconductor transistors by level-set methods. Journal of Computational and Applied Mathematics, 159(2):399–410, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005430.

Ferrando:2001:ABW

[FK01] S. E. Ferrando and L. A. Kolasa. Averages of best wavelet basis estimates for denoising. Journal of Computational and Applied Mathematics, 136 (1–2):357–367, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006269.

Frauendiener:2004:HTF

[FK04] J. Frauendiener and C. Klein. Hyperelliptic theta-functions and spectral methods. Journal of Computational and Applied Mathematics, 167 (1):193–218, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703008975.

Fredebeul:2002:MOD

[FKM02] Ch. Fredebeul, D. Kornmaier, and M. W. Müller. Multiple order double output Runge–Kutta Fehlberg formulae: strategies for efficient application. Journal of Computational and Applied Mathematics, 144(1–2):187–196, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005593.

Foupouagnigni:2004:FFO

[FKR04] M. Foupouagnigni, W. Koepf, and A. Ronveaux. Factorization of fourth-order differential equations for perturbed classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 162(2):299–326, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300709X.

Ferreira:2003:AEL

[FL03] Chelo Ferreira and José L.

López. Asymptotic expansions of the Lauricella hypergeometric function F_{D}. Journal of Computational and Applied Mathematics, 151 (2):235–256, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008142.

Feldheim:2004:SRH

[FL04] Véronique Feldheim and Paul Lybaert. Solution of radiative heat transfer problems with the discrete transfer method applied to triangular meshes. Journal of Computational and Applied Mathematics, 168(1–2):179–190, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009762.

Frammartino:2003:NSF

[FLM03] C. Frammartino, C. Laurita, and G. Mastroianni. On the numerical solution of Fredholm integral equations on unbounded intervals. Journal of Computational and Applied Mathematics, 158 (2):355–378, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004539.

Fabijonas:2003:ACM

[FLR03] B. R. Fabijonas, D. W. Lozier, and J. M. Rappoport. Algorithms and codes for the Macdonald function: recent progress and comparisons. Journal of Computational and Applied Mathematics, 161(1):179–192, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300596X.

Fang:2001:SQS

[FLW01] Shu-Cherng Fang, Chih-Jen Lin, and Soon-Yi Wu. Solving quadratic semi-infinite programming problems by using relaxed cutting-plane scheme. Journal of Computational and Applied Mathematics, 129(1–2):89–104, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005446.

Flemisch:2004:CSV

[FMRW04] B. Flemisch, Y. Maday, F. Rapetti, and B. I. Wohlmuth. Coupling scalar and vector potentials on nonmatching grids for eddy currents in a moving conductor. Journal of Computational and Applied Mathematics, 168(1–2):191–205, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009774.

Fort:2002:AOQ

[FP02] Jean-Claude Fort and Gilles Pagès. Asymptotics of optimal quantizers for some scalar distributions. Journal of Computational and Applied Mathematics, 146 (2):253–275, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200359X.

Ferreira:2001:PEZ

[FPO01] J. A. Ferreira, F. Patrício, and F. Oliveira. A priori estimates for the zeros of interval polynomials. Journal of Computational and Applied Mathematics, 136 (1–2):271–281, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006191.

Furi:2000:FOM

[FPS00] Massimo Furi, Maria Patrizia Pera, and Marco Spadini. Forced oscillations on manifolds and multiplicity results for periodically perturbed autonomous systems. Journal of Computational and Applied Mathematics, 113(1–2):241–254, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002599.

Floater:2000:FBA

[FQR00] Michael S. Floater, Ewald G. Quak, and Martin Reimers. Filter bank algorithms for piecewise linear prewavelets on arbitrary triangulations. Journal of Computational and Applied Mathematics, 119(1–2):185–207, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003782.

Foupouagnigni:2001:FOD

[FR01] M. Foupouagnigni and A. Ronveaux. Fourth-order difference equation satisfied by the co-recursive of q-classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 133 (1–2):355–365, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006555.

Foupouagnigni:2003:DEC

[FR03] M. Foupouagnigni and A. Ronveaux. Difference equations for the co-recursive r-th associated orthogonal polynomials of the D_{q}-Laguerre–Hahn class. Journal of Computational and Applied Mathematics, 153(1–2):213–223, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006362.

Franco:2002:EPE

[Fra02] J. M. Franco. An embedded pair of exponentially fitted explicit Runge–Kutta methods. Journal of Computational and Applied Mathematics, 149(2):407–414, December 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004855.

Franco:2003:PEA

[Fra03]  J. M. Franco.  A 5(3)

pair of explicit ARKN methods for the numerical integration of perturbed oscillators. Journal of Computational and Applied Mathematics, 161(2):283–293, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007015.

Franco:2004:EFE

[Fra04] J. M. Franco. Exponentially fitted explicit Runge–Kutta– Nyström methods. Journal of Computational and Applied Mathematics, 167(1): 1–19, May 1, 2004. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042703008756.

Freund:2000:KSM

[Fre00] Roland W. Freund. Krylovsubspace methods for reducedorder modeling in circuit simulation. Journal of Computational and Applied Mathematics, 123(1–2):395– 421, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700003964.

Freund:2001:CMV

[Fre01] Roland W. Freund. Computation of matrix-valued formally orthogonal polynomials and applications. Journal of Computational and Applied Mathematics, 127 (1–2):173–199, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005057.

Freeman:2002:MDE

[Fre02]  A. J. Freeman.  Materi-

als by design and the exciting role of quantum computation/simulation. Journal of Computational and Applied Mathematics, 149(1): 27–56, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702005198.

French:2003:ACB

[Fre03] Donald A. French. Analysis of the convergence behavior of the static localized nonlinear approximation for a problem in electromagnetism. Journal of Computational and Applied Mathematics, 161(1):27–39, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004990.

Frommer:2000:AI

[FS00] Andreas Frommer and Daniel B. Szyld. On asynchronous iterations. Journal of Computational and Applied Mathematics, 123(1–2):201–216, November 1, 2000. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S037704270000409X.

Frontini:2003:MNM

[FS03] M. Frontini and E. Sormani. Modified Newton’s method with third-order convergence and multiple roots. Journal of Computational and Applied Mathematics, 156 (2):345–354, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702009202.

Ford:2003:NIU

[FT03] J. A. Ford and S. Tharmlikit.

New implicit updates in multistep quasi-Newton methods for unconstrained optimisation. Journal of Computational and Applied Mathematics, 152(1–2):133–146, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200701X.

Fang:2002:FDF

[FTY02] Qing Fang, Takuya Tsuchiya, and Tetsuro Yamamoto. Finite difference, finite element and finite volume methods applied to two-point boundary value problems. Journal of Computational and Applied Mathematics, 139(1):9–19, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003922.

Fu:2004:STF

[Fu04] Chu-Li Fu. Simplified Tikhonov and Fourier regularization methods on a general sideways parabolic equation. Journal of Computational and Applied Mathematics, 167 (2):449–463, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703009324.

Fujita:2002:CAS

[Fuj02] Hiroshi Fujita. A coherent analysis of Stokes flows under boundary conditions of friction type. Journal of Computational and Applied Mathematics, 149(1):57–69, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005204.

Furihata:2001:FDS

[Fur01]  Daisuke Furihata.  Finite-

difference schemes for nonlinear wave equation that inherit energy conservation property. Journal of Computational and Applied Mathematics, 134 (1–2):37–57, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005276.

Ford:2000:HDN

[FW00] Neville J. Ford and Volker Wulf. How do numerical methods perform for delay differential equations undergoing a Hopf bifurcation? Journal of Computational and Applied Mathematics, 125 (1–2):277–285, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004738.

Feng:2002:CSD

[FW02] B.-F Feng and G. W. Wei.

A comparison of the spectral and the discrete singular convolution schemes for the KdV-type equations. Journal of Computational and Applied Mathematics, 145(1):183–188, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100543X.

Feng:2003:APS

[FW03] Chunhua Feng and Peiguang Wang. Almost periodic solutions of forced Liénardtype equations with time delays. Journal of Computational and Applied Mathematics, 161(1):67–74, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005454.

Galantai:2000:TNM

[Gal00]  A. Galántai.  The theory

of Newton’s method. Journal of Computational and Applied Mathematics, 124 (1–2):25–44, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004350.

Ganesh:2004:FDS

[Gan04] M. Ganesh. Fully discrete spectral methods for boundary integral equations on slender spheroids. Journal of Computational and Applied Mathematics, 164– 165(??):307–322, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004886.

Gao:2001:NAN

[Gao01]  Li Gao.  Numerical al-

gorithms for nonlinear L_{p}- norm problem and its extreme case. Journal of Computational and Applied Mathematics, 129(1–2):139–150, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000546X.

Garrappa:2004:ACT

[Gar04] Roberto Garrappa. An analysis of convergence for two-stage waveform relaxation methods. Journal of Computational and Applied Mathematics, 169(2):377–392, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042704000044.

Gatteschi:2002:ABZ

[Gat02]  Luigi Gatteschi.  Asymp-

totics and bounds for the zeros of Laguerre polynomials: a survey. Journal of Computational and Applied Mathematics, 144(1–2):7–27, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005490.

Gautschi:2001:URF

[Gau01] Walter Gautschi. The use of rational functions in numerical quadrature. Journal of Computational and Applied Mathematics, 133 (1–2):111–126, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006373.

Gautschi:2002:GQA

[Gau02]  Walter Gautschi.  Gauss

quadrature approximations to hypergeometric and confluent hypergeometric functions. Journal of Computational and Applied Mathematics, 139(1):173–187, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042701003831.

Gaspar:2002:SNE

[GCL02] F. J. Gaspar, C. Clavero, and F. Lisbona. Some numerical experiments with multigrid methods on Shishkin meshes. Journal of Computational and Applied Mathematics, 138(1):21–35, January 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100365X.

Gopalakrishnan:2004:QSS

[GD04] Jayadeep Gopalakrishnan and Leszek F. Demkowicz. Quasioptimality of some spectral mixed methods. Journal of Computational and Applied Mathematics, 167 (1):163–182, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703008951.

Goovaerts:2004:CFA

[GDD04] Marc Goovaerts, Ann De Schepper, and Marc Decamps. Closed-form approximations for diffusion densities: a path integral approach. Journal of Computational and Applied Mathematics, 164– 165(??):337–364, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008240.

Gemignani:2000:CHF

[Gem00] Luca Gemignani. Computing a Hurwitz factorization of a polynomial. Journal of Computational and Applied Mathematics, 126(1– 2):369–380, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003659.

Genz:2003:FSI

[Gen03] Alan Genz. Fully symmetric interpolatory rules for multiple integrals over hyper-spherical surfaces. Journal of Computational and Applied Mathematics, 157(1):187–195, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004138.

Georg:2003:NET

[Geo03] Kurt Georg. A new exclusion test. Journal of Computational and Applied Mathematics, 152(1–2):147–160, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007021.

Gout:2001:NFE

[GG01] Christian Gout and Allal Guessab. A new family of extended Gauss quadratures with an interior interval constraint. Journal of Computational and Applied Mathematics, 131(1–2):35–53, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003241.

Gyselinck:2004:MHM

[GGDL04] J. Gyselinck, C. Geuzaine, P. Dular, and W. Legros. Multi-harmonic modelling of motional magnetic field problems using a hybrid finite element-boundary element discretisation. Journal of Computational and Applied Mathematics, 168(1–2):225–234, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009804.

Ghelardoni:2001:SCS

[GGM01] P. Ghelardoni, G. Gheri, and M. Marletta. Spectral corrections for Sturm–Liouville problems. Journal of Computational and Applied Mathematics, 132(2):443–459, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042700004465.

Ghelardoni:2004:QEP

[GGM04] P. Ghelardoni, G. Gheri, and M. Marletta. A quasiextrapolation procedure for error estimation of numerical methods in Sturm– Liouville eigenproblems. Journal of Computational and Applied Mathematics, 164– 165(??):323–335, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004898.

Guillaume:2000:MPA

[GH00] Philippe Guillaume and Alain Huard. Multivariate Padé approximation. Journal of Computational and Applied Mathematics, 121(1– 2):197–219, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000337X.

Gatica:2001:EMF

[GH01a] Gabriel N. Gatica and Norbert Heuer. An expanded mixed finite element approach via a dual-dual formulation and the minimum residual method. Journal of Computational and Applied Mathematics, 132(2):371–385, July 15, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004404.

Gottlieb:2001:SMH

[GH01b] D. Gottlieb and J. S. Hesthaven. Spectral methods for hyperbolic problems. Journal of Computational and Applied Mathematics, 128(1–2):83–131, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005100.

Guo:2003:PSI

[GH03a] Shangjiang Guo and Lihong Huang. Periodic solutions in an inhibitory two-neuron network. Journal of Computational and Applied Mathematics, 161(1):217–229, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006691.

Gyori:2003:SAS

[GH03b] István Győri and Ferenc Hartung. Stability analysis of a single neuron model with delay. Journal of Computational and Applied Mathematics, 157(1):73–92, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003765.

Ghoreishi:2004:TMN

[GH04] F. Ghoreishi and S. Mohammad Hosseini. The Tau method and a new preconditioner. Journal of Computational and Applied Mathematics, 163(2):351–379, February 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007088.

Goldberg:2003:FNM

[GHK03] Maxim J. Goldberg, Tomasz Hrycak, and Seonja Kim. A fast numerical method for evaluation of Calderón commutators. Journal of Computational and Applied Mathematics, 158 (2):473–484, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004837.

Garcia:2001:DKS

[GHM01] Antonio G. García and Miguel A.

Hernández-Medina. The discrete Kramer sampling theorem and indeterminate moment problems. Journal of Computational and Applied Mathematics, 134 (1–2):13–22, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004507.

Garcia-Huidobro:2000:EPS

[GHMY00] M. García-Huidobro, R. Manásevich, and C. S. Yarur. Existence of positive singular solutions for a class of quasilinear elliptic equations. Journal of Computational and Applied Mathematics, 113 (1–2):329–351, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002654.

Gomez:2002:SAL

[GHR02] I. Gómez, I. Higueras, and T. Roldán. Starting algorithms for low stage order RKN methods. Journal of Computational and Applied Mathematics, 140(1–2):345–367, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004794.

Grunbaum:2003:NVB

[GI03] F. Alberto Grünbaum and Plamen Iliev. A noncommutative version of the bispectral problem. Journal of Computational and Applied Mathematics, 161(1):99–118, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300548X.

Gilewicz:2001:ZFP

[Gil01]  Jacek Gilewicz.  Zeros of

Froissart polynomials. Journal of Computational and Applied Mathematics, 133 (1–2):687, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700007202.

Guoqiang:2001:ENS

[GJ01] Han Guoqiang and Wang Jiong. Extrapolation of Nystrom solution for two dimensional nonlinear Fredholm integral equations. Journal of Computational and Applied Mathematics, 134 (1–2):259–268, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005537.

Guoqiang:2000:CME

[GJHY00] Han Guoqiang, Wang Jiong, Ken Hayami, and Xu Yuesheng. Correction method and extrapolation method for singular two-point boundary value problems. Journal of Computational and Applied Mathematics, 126 (1–2):145–157, December 30, 2000. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003490.

Gill:2000:SMO

[GJL^{+}00] Philip E. Gill, Laurent O. Jay, Michael W. Leonard, Linda R. Petzold, and Vivek Sharma. An SQP method for the optimal control of large-scale dynamical systems. Journal of Computational and Applied Mathematics, 120(1–2):197–213, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003101.

Garloff:2003:LBF

[GJS03] Jürgen Garloff, Christian Jansson, and Andrew P. Smith. Lower bound functions for polynomials. Journal of Computational and Applied Mathematics, 157(1): 207–225, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703004229.

Guseinov:2002:DCS

[GK02] G. Sh. Guseinov and B. Kaymakçalan. On a disconjugacy criterion for second order dynamic equations on time scales. Journal of Computational and Applied Mathematics, 141(1–2):187–196, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004459.

Gilewicz:2003:FDP

[GK03] Jacek Gilewicz and Yuri Kryakin. Froissart doublets in Padé approximation in the case of polynomial noise. Journal of Computational and Applied Mathematics, 153(1–2):235–242, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200674X.

Galue:2000:MTG

[GKK00] L. Galué, H. G. Khajah, and S. L. Kalla. Multiplication theorems for generalized and double-index Bessel functions. Journal of Computational and Applied Mathematics, 118(1–2):143–150, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002855.

Galperin:2000:VTN

[GKMN00] E. A. Galperin, E. J. Kansa, A. Makroglou, and S. A. Nelson. Variable transformations in the numerical solution of second kind Volterra integral equations with continuous and weakly singular kernels; extensions to Fredholm integral equations. Journal of Computational and Applied Mathematics, 115(1–2):193–211, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002976.

Glaeske:2000:MBT

[GKS00] H.-J. Glaeske, Anatoly A. Kilbas, and Megumi Saigo. A modified Bessel-type integral transform and its compositions with fractional calculus operators on spaces F_{p,μ} and F′_{p,μ}. Journal of Computational and Applied Mathematics, 118(1–2):151–168, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002867.

Goodman:2000:NCA

[GL00] T. N. T. Goodman and S. L. Lee. Nonuniform cascade algorithms. Journal of Computational and Applied Mathematics, 119(1–2):223–234, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003800.

Giordano:2001:IMP

[GL01] C. Giordano and A. Laforgia.

Inequalities and monotonicity properties for the gamma function. Journal of Computational and Applied Mathematics, 133(1–2):387–396, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006592.

Grammont:2002:SSR

[GL02] Laurence Grammont and Alain Largillier. On ϵspectra and stability radii. Journal of Computational and Applied Mathematics, 147(2):453–469, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004806.

Giordano:2003:BTI

[GL03] C. Giordano and A. Laforgia.

On the Bernstein-type inequalities for ultraspherical polynomials. Journal of Computational and Applied Mathematics, 153(1–2):243–248, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005915.

Garcia:2004:AST

[GL04a] A. G. García and L. L. Littlejohn. On analytic sampling theory. Journal of Computational and Applied Mathematics, 171(1–2):235–246, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000561.

Guan:2004:SDN

[GL04b] Lutai Guan and Bin Liu. Surface design by natural splines over refined grid points. Journal of Computational and Applied Mathematics, 163(1):107–115, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007982.

Glasser:2000:HFT

[Gla00] M. L. Glasser. Hypergeometric functions and the trinomial equation. Journal of Computational and Applied Mathematics, 118(1–2):169–173, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002879.

Gorenflo:2000:WFS

[GLM00] Rudolf Gorenflo, Yuri Luchko, and Francesco Mainardi. Wright functions as scale-invariant solutions of the diffusion-wave equation. Journal of Computational and Applied Mathematics, 118(1–2):175–191, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002880.

Globisch:2003:HPU

[Glo03] Gerhard Globisch. The hierarchical preconditioning on unstructured three-dimensional grids with locally refined regions. Journal of Computational and Applied Mathematics, 150(2):265–282, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006647.

Gu:2003:EHO

[GLZ03] Yuanxian Gu, Wenyuan Liao, and Jianping Zhu. An efficient high-order algorithm for solving systems of 3-D reaction–diffusion equations. Journal of Computational and Applied Mathematics, 155(1):1–17, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008890.

Gasca:2000:ETE

[GM00a] M. Gasca and G. Mühlbach. Elimination techniques: from extrapolation to totally positive matrices and CAGD. Journal of Computational and Applied Mathematics, 122(1–2):37–50, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003563.

Gheri:2000:PSE

[GM00b] G. Gheri and P. Marzulli. Parallel shooting with error estimate for increasing the accuracy. Journal of Computational and Applied Mathematics, 115(1–2):213–227, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001764.

Greenberg:2000:NMH

[GM00c] Leon Greenberg and Marco Marletta. Numerical methods for higher order Sturm– Liouville problems. Journal of Computational and Applied Mathematics, 125 (1–2):367–383, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004805.

Gautschi:2001:P

[GMR01] Walter Gautschi, Francisco Marcellán, and Lothar Reichel. Preface. Journal of Computational and Applied Mathematics, 127(1–2):ix–xi, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004908.

Graves-Morris:2000:EAR

[GMRS00] P. R. Graves-Morris, D. E.

Roberts, and A. Salam. The epsilon algorithm and related topics. Journal of Computational and Applied Mathematics, 122(1–2):51–80, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003551.

Grossmann:2002:SRM

[GN02] C. Grossmann and A. Noack.

Smoothing and Rothe’s method for Stefan problems in enthalpy form. Journal of Computational and Applied Mathematics, 138(2):347–366, January 15, 2002. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042701003685.

Givoli:2003:HON

[GN03] Dan Givoli and Beny Neta.

High-order nonreflecting boundary conditions for the dispersive shallow water equations. Journal of Computational and Applied Mathematics, 158(1):49–60, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300462X.

Georgescu:2001:SSR

[GNPB01] Adelina Georgescu, Bogdan Nicolescu, Nicolae Popa, and Mircea Bolosteanu. On special solutions of the Reynolds equation from lubrication. Journal of Computational and Applied Mathematics, 133(1–2):367–372, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006567.

Goh:2002:GPI

[Goh02] William M. Y. Goh. On a generalization of Polya inequality and some of its statistical implications. Journal of Computational and Applied Mathematics, 142(1):51–59, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004599.

Gorbatikh:2004:MCA

[Gor04] Larissa Gorbatikh. Mathematical and computational aspects of nonuniform frictional slip modeling. Journal of Computational and Applied Mathematics, 168(1–2):215–224, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009798.

Gotz:2001:OQA

[Göt01] M. Götz. Optimal quadrature for analytic functions. Journal of Computational and Applied Mathematics, 137 (1):123–133, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007032.

Govaerts:2000:NBA

[Gov00] W. Govaerts. Numerical bifurcation analysis for ODEs. Journal of Computational and Applied Mathematics, 125 (1–2):57–68, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004581.

Gurappa:2003:NPS

[GPS03] N. Gurappa, Prasanta K.

Panigrahi, and T. Shreecharan. A new perspective on single and multi-variate differential equations. Journal of Computational and Applied Mathematics, 160 (1–2):103–112, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006162.

Gilewicz:2002:CFT

[GPTT02] Jacek Gilewicz, Maciej Pindor, Józef J. Telega, and Stanislaw Tokarzewski. Continued fractions, two-point Padé approximants and errors in the Stieltjes case. Journal of Computational and Applied Mathematics, 145(1):99–112, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005386.

Groisman:2001:ABN

[GR01] Pablo Groisman and Julio D. Rossi. Asymptotic behaviour for a numerical approximation of a parabolic problem with blowing up solutions. Journal of Computational and Applied Mathematics, 135(1):135–155, October 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005719.

Granvilliers:2004:ICC

[Gra04] Laurent Granvilliers. An interval component for continuous constraints. Journal of Computational and Applied Mathematics, 162(1):79–92, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007210.

Groeneboom:2002:HMA

[Gro02] Piet Groeneboom. Hydrodynamical methods for analyzing longest increasing subsequences. Journal of Computational and Applied Mathematics, 142(1):83–105, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004617.

Grunbaum:2001:EIZ

[Grü01] F. Alberto Grünbaum. Electrostatic interpretation for the zeros of certain polynomials and the Darboux process. Journal of Computational and Applied Mathematics, 133(1–2):397–412, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006610.

Grundy:2003:AIB

[Gru03] R. E. Grundy. The analysis of initial-boundary-value problems using Hermite interpolation. Journal of Computational and Applied Mathematics, 154(1):63–95, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008117.

Gasca:2000:HMP

[GS00] Mariano Gasca and Thomas Sauer. On the history of multivariate polynomial interpolation. Journal of Computational and Applied Mathematics, 122(1–2):23–35, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003538.

Ganesh:2001:NSN

[GS01a] M. Ganesh and O. Steinbach. The numerical solution of a nonlinear hypersingular boundary integral equation. Journal of Computational and Applied Mathematics, 131(1–2):267–280, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002697.

Gilewicz:2001:NLB

[GS01b] Jacek Gilewicz and Peter A.

Shulimanov. A new lower bound for a rational approximation on the positive real axis. Journal of Computational and Applied Mathematics, 133(1–2):383–385, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006580.

Goh:2002:LDM

[GS02] William Goh and Eric Schmutz.

Limit distribution for the maximum degree of a random recursive tree. Journal of Computational and Applied Mathematics, 142(1):61–82, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004605.

Gine:2004:INA

[GS04a] Jaume Giné and Xavier Santallusia. Implementation of a new algorithm of computation of the Poincaré–Liapunov constants. Journal of Computational and Applied Mathematics, 166(2):465–476, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007702.

Guessous:2004:MBI

[GS04b] N. Guessous and O. Souhar. Multilevel block ILU preconditioner for sparse nonsymmetric M-matrices. Journal of Computational and Applied Mathematics, 162(1): 231–246, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703007167.

Guo:2003:PSS

[GSG03] Yanping Guo, Wenrui Shan, and Weigao Ge. Positive solutions for second-order mpoint boundary value problems. Journal of Computational and Applied Mathematics, 151(2):415–424, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007392.

Guillaume:2003:RAP

[GSS03] Philippe Guillaume, Yousef Saad, and Masha Sosonkina. Rational approximation preconditioners for sparse linear systems. Journal of Computational and Applied Mathematics, 158 (2):419–442, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042703004801.

Gil:2003:CMB

[GST03] Amparo Gil, Javier Segura, and Nico M. Temme. Computation of the modified Bessel function of the third kind of imaginary orders: uniform Airy-type asymptotic expansion. Journal of Computational and Applied Mathematics, 153(1–2):225–234, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006088.

Guo:2004:TPS

[GT04] Yanping Guo and Jiwei Tian. Two positive solutions for second-order quasilinear differential equation boundary value problems with sign changing nonlinearities. Journal of Computational and Applied Mathematics, 169(2):345–357, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000020.

Gu:2004:MPT

[Gu04] Chuanqing Gu. Matrix Padétype approximant and directional matrix Padé approximant in the inner product space. Journal of Computational and Applied Mathematics, 164–165(??): 365–385, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703004874.

Gugat:2000:NMC

[Gug00]  Martin Gugat.  A New-

ton method for the computation of time-optimal boundary controls of one-dimensional vibrating systems. Journal of Computational and Applied Mathematics, 114 (1):103–119, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002915.

Golub:2000:ECC

[GvdV00] Gene H. Golub and Henk A. van der Vorst. Eigenvalue computation in the 20th century. Journal of Computational and Applied Mathematics, 123(1–2):35–65, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004131. Numeri-

cal analysis 2000, Vol. III. Lin-

ear algebra.

Georgescu:2001:CSN

[GVSJ01] Adelina Georgescu, Harry Vereecken, Holger Schwarze, and Uwe Jaekel. Classes of solutions for a nonlinear diffusion PDE. Journal of Computational and Applied Mathematics, 133(1–2):373–381, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006579.

Gallivan:2004:SEP

[GVV04] K. Gallivan, A. Vandendorpe, and P. Van Dooren. Sylvester equations and projectionbased model reduction. Journal of Computational and Applied Mathematics, 162(1): 213–229, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703007143.

Goovaerts:2004:PI

[GVW04] M. J. Goovaerts, S. Vandewalle, and L. Wuytack. Proceedings of ICCAM-2002. Journal of Computational and Applied Mathematics, 164–165(??):xi, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009543.

Gilbert:2004:SPA

[GW04] D. J. Gilbert and A. D.

Wood. Stokes’ phenomenon and the absolutely continuous spectrum of one-dimensional Schrödinger operators. Journal of Computational and Applied Mathematics, 171 (1–2):247–264, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000512.

Ge:2002:SCH

[GZ02] Lixin Ge and Jun Zhang. Symbolic computation of high order compact difference schemes for three dimensional linear elliptic partial differential equations with variable coefficients. Journal of Computational and Applied Mathematics, 143(1):9–27, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005040.

Hadjidimos:2000:SOS

[Had00]  A. Hadjidimos.  Succes-

sive overrelaxation (SOR) and related methods. Journal of Computational and Applied Mathematics, 123 (1–2):177–199, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004039.

Hagler:2001:FHG

[Hag01]  Brian A. Hagler.  d-fold

Hermite–Gauss quadrature. Journal of Computational and Applied Mathematics, 136 (1–2):53–72, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005756.

Hahner:2000:USP

[Häh00] Peter Hähner. On the uniqueness of the shape of a penetrable, anisotropic obstacle. Journal of Computational and Applied Mathematics, 116(1):167–180, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003234.

Hamina:2000:AMH

[Ham00] Martti Hamina. An approximation method for the hypersingular heat operator equation. Journal of Computational and Applied Mathematics, 115(1–2):229–243, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003015.

Hamsapriye:2001:NQR

[Ham01] Hamsapriye. A new quadrature rule based on a generalized mixed interpolation formula of exponential type. Journal of Computational and Applied Mathematics, 131 (1–2):65–88, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700003228.

Han:2003:CST

[Han03a] Bin Han. Compactly supported tight wavelet frames and orthonormal wavelets of exponential decay with a general dilation matrix. Journal of Computational and Applied Mathematics, 155 (1):43–67, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702008919.

Han:2003:PDA

[Han03b]  Deren Han.  A proximal

decomposition algorithm for variational inequality problems. Journal of Computational and Applied Mathematics, 161(1):231–244, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006708.

Hasegawa:2004:UAF

[Has04] Takemitsu Hasegawa. Uniform approximations to finite Hilbert transform and its derivative. Journal of Computational and Applied Mathematics, 163(1): 127–138, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703008008.

Hataue:2003:ADD

[Hat03] Itaru Hataue. On analogy and dissimilarity of dependence of stability on several parameters in flow simulations. Journal of Computational and Applied Mathematics, 159(1):45–53, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005661.

Hausenblas:2002:NAS

[Hau02] Erika Hausenblas. Numerical analysis of semilinear stochastic evolution equations in Banach spaces. Journal of Computational and Applied Mathematics, 147 (2):485–516, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004831.

Harris:2003:EPI

[HC03a] Paul J. Harris and Ke Chen.

On efficient preconditioners for iterative solution of a Galerkin boundary element equation for the three-dimensional exterior Helmholtz problem. Journal of Computational and Applied Mathematics, 156 (2):303–318, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702009184.

Houzeaux:2003:ISO

[HC03b] Guillaume Houzeaux and Ramon Codina. An iteration-by-subdomain overlapping Dirichlet/Robin domain decomposition method for advection– diffusion problems. Journal of Computational and Applied Mathematics, 158 (2):243–276, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004473.

Hussein:2004:FCM

[HC04] Anwar Hussein and Ke Chen. Fast computational methods for locating fold points for the power flow equations. Journal of Computational and Applied Mathematics, 164– 165(??):419–430, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006514.

Houstis:2002:GMP

[HCT^{+}02] E. N. Houstis, A. C. Catlin, P. Tsompanopoulou, D. Gottfried, G. Balakrishnan, K. Su, and J. R. Rice. GasTurbn-Lab: a multidisciplinary problem solving environment for gas turbine engine design on a network of nonhomogeneous machines. Journal of Computational and Applied Mathematics, 149(1):83– 100, December 1, 2002. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
math.utah.edu/pub/bibnet/
authors/r/rice-john-r.bib;
http://www.sciencedirect.
com/science/article/pii/
S0377042702005228.

He:2003:EPS

[He03] Zhimin He. On the existence of positive solutions of p-Laplacian difference equations. Journal of Computational and Applied Mathematics, 161(1):193–201, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006666.

Heindl:2003:BPC

[Hei03a] G. Heindl. Best possible componentwise parameter inclusions computable from a priori estimates, measurements, and bounds for the measurement errors. Journal of Computational and Applied Mathematics, 152(1–2):175–185, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007045.

Heinrichs:2003:LSS

[Hei03b] Wilhelm Heinrichs. Leastsquares spectral collocation for discontinuous and singular perturbation problems. Journal of Computational and Applied Mathematics, 157(2):329–345, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004151.

Hernandez:2001:MCK

[Her01]  M. A. Hernández.  A

modification of the classical Kantorovich conditions for Newton’s method. Journal of Computational and Applied Mathematics, 137 (1):201–205, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003934.

Herzberger:2003:EBP

[Her03]  Jürgen Herzberger.  Ex-

plicit bounds for the positive root of classes of polynomials with applications.

Journal of Computational and Applied Mathematics, 152(1–2):187–198, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007057.

He:2003:OSO

[HG03] Zhimin He and Weigao Ge.

Oscillations of second-order nonlinear impulsive ordinary differential equations. Journal of Computational and Applied Mathematics, 158 (2):397–406, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004746.

Hanhart:2004:MEF

[HGI04] Alexander L. Hanhart, Matthias K.

Gobbert, and Leighton T. Izu. A memory-efficient finite element method for systems of reaction–diffusion equations with non-smooth forcing. Journal of Computational and Applied Mathematics, 169(2):431–458, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000093.

Hardin:2003:CWP

[HH03a] Doug Hardin and Don Hong.

Construction of wavelets and prewavelets over triangulations. Journal of Computational and Applied Mathematics, 155(1):91–109, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008944.

Hong:2003:P

[HH03b] Don Hong and Tianxiao He. Preface. Journal of Computational and Applied Mathematics, 155(1):ix–x, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008889.

Henrotte:2004:CEF

[HH04] François Henrotte and Kay Hameyer. Computation of electromagnetic force densities: Maxwell stress tensor vs. virtual work principle. Journal of Computational and Applied Mathematics, 168(1–2):235–243, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009816.

Hu:2001:ACS

[HHC01] Guang-Da Hu, Guang-Di Hu, and Baruch Cahlon. Algebraic criteria for stability of linear neutral systems with a single delay. Journal of Computational and Applied Mathematics, 135(1):125–133, October 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005707.

Hounkonnou:2000:DSC

[HHR00] M. N. Hounkonnou, C. Hounga, and A. Ronveaux. Discrete semi-classical orthogonal polynomials: generalized Charlier. Journal of Computational and Applied Mathematics, 114(2):361–366, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002757.

Higham:2003:USW

[Hig03] Desmond J. Higham. Unravelling small world networks. Journal of Computational and Applied Mathematics, 158(1):61–74, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004710.

Hilger:2002:MLT

[Hil02a] Stefan Hilger. Matrix Lie theory and measure chains. Journal of Computational and Applied Mathematics, 141(1–2):197–217, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004460.

Hilscher:2002:TSV

[Hil02b]  Roman Hilscher.  A time

scales version of a Wirtingertype inequality and applications. Journal of Computational and Applied Mathematics, 141(1–2):219–226, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004472.

Hinton:2002:LLP

[Hin02] Don Hinton. On the location of the least point of the essential spectrum. Journal of Computational and Applied Mathematics, 148(1):77–89, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005745.

Hitczenko:2002:P

[Hit02a]  Pawel Hitczenko.  Pref-

ace. Journal of Computational and Applied Mathematics, 142(1):ix–x, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100454X.

Hitrik:2002:ERM

[Hit02b] Michael Hitrik. Existence of resonances in magnetic scattering. Journal of Computational and Applied Mathematics, 148(1):91–97, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005757.

Haddar:2002:STL

[HJ02] H. Haddar and P. Joly. Stability of thin layer approximation of electromagnetic waves scattering by linear and nonlinear coatings. Journal of Computational and Applied Mathematics, 143(2):201–236, June 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005088.

Hackbusch:2000:SMA

[HK00] W. Hackbusch and B. N.

Khoromskij. A sparse Hmatrix arithmetic: general complexity estimates. Journal of Computational and Applied Mathematics, 125 (1–2):479–501, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042700004866.

Hristova:2001:QBV

[HK01] S. G. Hristova and G. K. Kulev. Quasilinearization of a boundary value problem for impulsive differential equations. Journal of Computational and Applied Mathematics, 132(2):399–407, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004428.

Haslinger:2004:ANR

[HKD04] Jaroslav Haslinger, Radek Kucera, and Zdenek Dostál. An algorithm for the numerical realization of 3D contact problems with Coulomb friction. Journal of Computational and Applied Mathematics, 164– 165(??):387–408, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006484.

Hashimoto:2003:SSC

[HKT^{+}03] M. Hashimoto, S. Kanemitsu, Y. Tanigawa, M. Yoshimoto, and W.-P. Zhang. On some slowly convergent series involving the Hurwitz zeta-function. Journal of Computational and Applied Mathematics, 160 (1–2):113–123, November 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006174.

He:2002:SGI

[HL02] Yinnian He and Kaitai Li.

Stability of Galerkin and inertial algorithms with variable time step size. Journal of Computational and Applied Mathematics, 146 (2):213–230, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003540.

Ham:2004:EGC

[HL04a] YoonMee Ham and Sang-Gu Lee. Effects on global coupling of an interface problem in an activator-inhibitor system. Journal of Computational and Applied Mathematics, 162(2):393–409, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007544.

Huo:2004:PSD

[HL04b] Hai-Feng Huo and Wan-Tong Li. Periodic solution of a delayed predator– prey system with Michaelis– Menten type functional response. Journal of Computational and Applied Mathematics, 166(2):453–463, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007696.

He:2003:FEA

[HLS^{+}03] Yinnian He, Yanping Lin, Samuel S. P. Shen, Weiwei Sun, and Robert Tait. Finite element approximation for the viscoelastic fluid motion problem. Journal of Computational and Applied Mathematics, 155(2):201–222, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008646.

Huang:2004:BDR

[HLY04] Daren Huang, Feng Li, and Lihua Yang. L^{1}-based decomposition and reconstruction algorithms and W - matrices. Journal of Computational and Applied Mathematics, 163(1):139–153, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300801X.

Hoschek:2000:TBD

[HM00] Josef Hoschek and Rainer Müller. Turbine blade design by lofted B-spline surfaces. Journal of Computational and Applied Mathematics, 119(1–2):235–248, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003812.

Hoffman:2000:COC

[Hof00] Karla L. Hoffman. Combinatorial optimization: Current successes and directions for the future. Journal of Computational and Applied Mathematics, 124 (1–2):341–360, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004301.

Homeier:2000:SLT

[Hom00] Herbert H. H. Homeier. Scalar Levin-type sequence transformations. Journal of Computational and Applied Mathematics, 122(1–2):81–147, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003599.

Homeier:2003:MNM

[Hom03]  H. H. H. Homeier.  A

modified Newton method for rootfinding with cubic convergence. Journal of Computational and Applied Mathematics, 157(1):227–230, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703003911.

Homeier:2004:MNM

[Hom04] H. H. H. Homeier. A modified Newton method with cubic convergence: the multivariate case. Journal of Computational and Applied Mathematics, 169(1):161–169, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703010215.

Hayashi:2003:DNI

[HOO03] Keisuke Hayashi, Kazuei Onishi, and Yoko Ohura. Direct numerical identification of boundary values in the Laplace equation. Journal of Computational and Applied Mathematics, 152(1–2):161–174, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007033.

Hoppe:2004:AMD

[Hop04] R. H. W. Hoppe. Adaptive multigrid and domain decomposition methods in the computation of electromagnetic fields. Journal of Computational and Applied Mathematics, 168(1–2):245–254, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300983X.

Horwitz:2001:VSR

[Hor01] Alan Horwitz. A version of Simpson’s rule for multiple integrals. Journal of Computational and Applied Mathematics, 134(1–2):1–11, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004441.

Halidias:2000:ESQ

[HP00] Nikolaos Halidias and Nikolaos S. Papageorgiou. Existence of solutions for quasilinear second order differential inclusions with nonlinear boundary conditions. Journal of Computational and Applied Mathematics, 113(1–2):51–64, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002435.

Huard:2000:CAL

[HR00] Alain Huard and Vincent Robin. Continuity of approximation by least-squares multivariate Padé approximants. Journal of Computational and Applied Mathematics, 115(1–2):255–268, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001776.

Huang:2001:AMM

[HR01] Weizhang Huang and Robert D. Russell. Adaptive mesh movement — the MMPDE approach and its applications. Journal of Computational and Applied Mathematics, 128(1–2):383–398, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005203.

He:2002:DEA

[HR02a] M. X. He and P. E. Ricci. Differential equation of Appell polynomials via the factorization method. Journal of Computational and Applied Mathematics, 139 (2):231–237, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270100423X. See re-

marks [Ism03b].

Horvat:2002:TSC

[HR02b] Vilmos Horvat and Mladen Rogina. Tension spline collocation methods for singularly perturbed Volterra integrodifferential and Volterra integral equations. Journal of Computational and Applied Mathematics, 140(1– 2):381–402, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701005179.

Horak:2003:ANR

[HR03] J. Horák and W. Reichel.

Analytical and numerical results for the Fucík spectrum of the Laplacian. Journal of Computational and Applied Mathematics, 161 (2):313–338, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007039.

Hedou-Rouillier:2004:FDM

[HR04a] Véronique Hédou-Rouillier.

A finite difference method to solve the forward problem in electroencephalography (EEG). Journal of Computational and Applied Mathematics, 167(1):35–58, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300877X.

Hernandez:2004:MNM

[HR04b] M. A. Hernández and M. J.

Rubio. A modification of Newton’s method for nondifferentiable equations. Journal of Computational and Applied Mathematics, 164– 165(??):409–417, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006502.

Hernandez:2000:SLM

[HRE00] M. A. Hernández, M. J. Rubio, and J. A. Ezquerro. Secant-like methods for solving nonlinear integral equations of the Hammerstein type. Journal of Computational and Applied Mathematics, 115(1–2):245–254, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001168.

Hooman:2004:PBA

[HRK04] K. Hooman and A. A. Ranjbar-Kani. A perturbation based analysis to investigate forced convection in a porous saturated tube. Journal of Computational and Applied Mathematics, 162(2):411–419, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007556.

Hitczenko:2002:GFA

[HRS02] Pawel Hitczenko, Cecil Rousseau, and Carla D. Savage. A generating functionology approach to a problem of Wilf. Journal of Computational and Applied Mathematics, 142(1):107–114, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004629.

Hernandez:2000:MKA

[HS00] M. A. Hernández and M. A.

Salanova. Modification of the Kantorovich assumptions for semilocal convergence of the Chebyshev method. Journal of Computational and Applied Mathematics, 126 (1–2):131–143, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003477.

Han:2001:TDV

[HS01] Weimin Han and Mircea Sofonea. Time-dependent variational inequalities for viscoelastic contact problems. Journal of Computational and Applied Mathematics, 136 (1–2):369–387, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042700006270.

Hagiwara:2002:HPE

[HS02] Ichiro Hagiwara and Junichi Shinoda. The high precise/ efficient pre/postprocessor. Journal of Computational and Applied Mathematics, 149 (1):325–333, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005423.

Han:2001:VNA

[HSS01] Weimin Han, Meir Shillor, and Mircea Sofonea. Variational and numerical analysis of a quasistatic viscoelastic problem with normal compliance, friction and damage. Journal of Computational and Applied Mathematics, 137(2):377–398, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000707X.

Huybrechs:2004:NWN

[HSV04] Daan Huybrechs, Jo Simoens, and Stefan Vandewalle. A note on wave number dependence of wavelet matrix compression for integral equations with oscillatory kernel. Journal of Computational and Applied Mathematics, 172(2):233–246, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001177.

Hsiao:2000:DDM

[HSW00] G. C. Hsiao, O. Steinbach, and W. L. Wendland. Domain decomposition methods via boundary integral equations. Journal of Computational and Applied Mathematics, 125 (1–2):521–537, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000488X.

Howle:2001:EMI

[HT01] V. E. Howle and Lloyd N.

Trefethen. Eigenvalues and musical instruments. Journal of Computational and Applied Mathematics, 135 (1):23–40, October 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005604.

Hu:2001:SFV

[Hu01]  Cheng-Feng Hu.  Solving

fuzzy variational inequalities over a compact set. Journal of Computational and Applied Mathematics, 129(1–2):185–193, April 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005495.

Hu:2002:GVI

[Hu02]  Cheng-Feng Hu.  Gener-

alized variational inequalities with fuzzy relation. Journal of Computational and Applied Mathematics, 146(1):47–56, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200417X.

Hugger:2001:TLP

[Hug01] Jens Hugger. A theory for local, a posteriori, pointwise, residual-based estimation of the finite element error. Journal of Computational and Applied Mathematics, 135(2):241–292, October 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005872.

Hui:2004:UCA

[Hui04] W. H. Hui. A unified coordinates approach to computational fluid dynamics. Journal of Computational and Applied Mathematics, 163(1):15–28, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007891.

Hendrickx:2002:KPV

[HV02] Jef Hendrickx and Marc Van Barel. A Kronecker product variant of the FACR method for solving the generalized Poisson equation. Journal of Computational and Applied Mathematics, 140(1–2):369–380, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004095.

Hairer:2000:OSS

[HW00] Ernst Hairer and Gerhard Wanner. Order stars and stiff integrators. Journal of Computational and Applied Mathematics, 125 (1–2):93–105, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004611.

Han:2002:REI

[HW02] Guoqiang Han and Ruifang Wang. Richardson extrapolation of iterated discrete Galerkin solution for twodimensional Fredholm integral equations. Journal of Computational and Applied Mathematics, 139(1):

49–63, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701003909.

He:2003:ASM

[HW03] Yong He and Min Wu. Absolute stability for multiple delay general Lur’e control systems with multiple nonlinearities. Journal of Computational and Applied Mathematics, 159(2):241–248, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004576.

Wang:2003:DTP

[hWqX03] Ren hong Wang and Zhi qiang Xu. Discrete truncated powers and lattice points in rational polytope. Journal of Computational and Applied Mathematics, 159(1):149–159, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005612.

Huang:2004:AAD

[HWT04] C.-S. Huang, S. Wang, and K. L. Teo. On application of an alternating direction method to Hamilton–Jacobin–Bellman equations. Journal of Computational and Applied Mathematics, 166(1):153–166, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008641.

He:2002:PBV

[HY02] Zhimin He and Jianshe Yu. Periodic boundary value problem for first-order impulsive functional differential equations. Journal of Computational and Applied Mathematics, 138(2):205–217, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003818.

Han:2003:NOD

[HY03] Houde Han and Dongsheng Yin. A non-overlap domain decomposition method for the forward-backward heat equation. Journal of Computational and Applied Mathematics, 159(1):35–44, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005570.

He:2003:BTQ

[HZZ03] T.-X. He, R. Zhang, and Y. S.

Zhou. Boundary-type quadrature and boundary element method. Journal of Computational and Applied Mathematics, 155(1):19–41, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008907.

Ifantis:2001:SMC

[Ifa01] E. K. Ifantis. On the spectral measure of a class of orthogonal polynomials. Journal of Computational and Applied Mathematics, 133 (1–2):688–689, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700007214.

Ifantis:2003:STC

[IK03] E. K. Ifantis and C. G. Kokologiannaki. Stieltjes transforms of a class of probability measures. Journal of Computational and Applied Mathematics, 153(1–2):249–258, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006167.

Iannelli:2001:ALM

[IM01] M. Iannelli and F. A. Milner. On the approximation of the Lotka–McKendrick equation with finite life-span. Journal of Computational and Applied Mathematics, 136 (1–2):245–254, November 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006166.

Itoh:2003:IDB

[IN03] Shoji Itoh and Yusuke Namekawa.

An improvement in DS– BiCGstab(l) and its application for linear systems in lattice QCD. Journal of Computational and Applied Mathematics, 159(1):65–75, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005648.

Inglese:2003:FLA

[Ing03] Gabriele Inglese. Fourier– Legendre approximation of a probability density from discrete data. Journal of Computational and Applied Mathematics, 154(1):161–173, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008191.

Ifantis:2001:LPE

[IP01] E. K. Ifantis and P. N.

Panagopoulos. Limit points of eigenvalues of truncated tridiagonal operators. Journal of Computational and Applied Mathematics, 133 (1–2):413–422, August 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006634.

Ipsen:2000:ORT

[Ips00] Ilse C. F. Ipsen. An overview of relative sin Θ theorems for invariant subspaces of complex matrices. Journal of Computational and Applied Mathematics, 123 (1–2):131–153, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004040.

Iqbal:2000:PCM

[Iqb00] M. Iqbal. On photon correlation measurements of colloidal size distributions using Bayesian strategies. Journal of Computational and Applied Mathematics, 126 (1–2):77–89, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003416.

Iqbal:2003:DRN

[Iqb03]  M. Iqbal.  Deconvolution

and regularization for numerical solutions of incorrectly posed problems. Journal of Computational and Applied Mathematics, 151 (2):463–476, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005678.

Ixaru:2001:WEF

[IRVD01] L. Gr. Ixaru, M. Rizea, G. Vanden Berghe, and H. De Meyer. Weights of the exponential fitting multistep algorithms for first-order ODEs. Journal of Computational and Applied Mathematics, 132(1):83–93, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005999.

Ismail:2003:ATT

[IS03] Mourad E. H. Ismail and Dennis Stanton. Applications of q-Taylor theorems. Journal of Computational and Applied Mathematics, 153(1–2):259–272, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006441.

Iske:2003:PSD

[Isk03] Armin Iske. Progressive scattered data filtering. Journal of Computational and Applied Mathematics, 158 (2):297–316, September 15, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004497.

Ismail:2000:PSE

[Ism00] A. I. Ismail. Periodic solutions of equations of motion of a heavy solid applying Krylov–Bogoliubov– Mitropolski method. Journal of Computational and Applied Mathematics, 114 (2):345–359, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002745.

Ismail:2003:GTB

[Ism03a] Mourad E. H. Ismail. A generalization of a theorem of Bochner. Journal of Computational and Applied Mathematics, 159(2):319–324, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005363.

Ismail:2003:RDE

[Ism03b] Mourad E. H. Ismail. Remarks on “Differential equation of Appell polynomials … ”. Journal of Computational and Applied Mathematics, 154(1):243–245, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008798.  See

[HR02a].

Ixaru:2002:FEE

[IVD02] L. Gr. Ixaru, G. Vanden Berghe, and H. De Meyer. Frequency evaluation in exponential fitting multistep algorithms for ODEs. Journal of Computational and Applied Mathematics, 140(1–2):423–434, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004745.

Iwai:2002:GMM

[Iwa02] Toshihiro Iwai. Geometric mechanics of many-body systems. Journal of Computational and Applied Mathematics, 140(1–2):403–422, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004009.

Ixaru:2000:CMS

[Ixa00] L. Gr. Ixaru. CP methods for the Schrödinger equation. Journal of Computational and Applied Mathematics, 125 (1–2):347–357, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004787.

Ishiwata:2003:BRF

[IY03] Tetsuya Ishiwata and Shigetoshi Yazaki. On the blow-up rate for fast blowup solutions arising in an anisotropic crystalline motion. Journal of Computational and Applied Mathematics, 159(1):55–64, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005569.

Inui:2003:RIE

[IYO03] Hirokazu Inui, Katsu Yamatani, and Kohzaburo Ohnaka. A reliable identification of electric current dipoles using harmonic functions. Journal of Computational and Applied Mathematics, 157(1): 107–123, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703003790.

Iserles:2000:PAI

[IZ00] Arieh Iserles and Antonella Zanna. Preserving algebraic invariants with Runge–Kutta methods. Journal of Computational and Applied Mathematics, 125(1–2):69–81, December 15, 2000. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004593.

Ito:2001:ISS

[IZ01] Kazufumi Ito and Jun Zou.

Identification of some source densities of the distribution type. Journal of Computational and Applied Mathematics, 132(2):295–308, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003320.

Jankowski:2002:DEI

[Jan02] Tadeusz Jankowski. Differential equations with integral boundary conditions. Journal of Computational and Applied Mathematics, 147 (1):1–8, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702003710.

Jankowski:2003:ESB

[Jan03a]  Tadeusz Jankowski.  Exis-

tence of solutions of boundary value problems for differential equations with delayed arguments. Journal of Computational and Applied Mathematics, 156(1):239–252, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702009160.

Jankowski:2003:WRM

[Jan03b] Tadeusz Jankowski. Waveform relaxation methods for periodic differential-functional systems. Journal of Computational and Applied Mathematics, 156(2):457–469, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703003868.

Jorge:2004:NME

[JB04] J. C. Jorge and B. Bujanda. Numerical methods for evolutionary reaction–diffusion problems with nonlinear reaction terms. Journal of Computational and Applied Mathematics, 166(1):167–180, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300863X.

Johnston:2001:GTM

[JE01] Peter R. Johnston and David Elliott. A generalisation of Telles’ method for evaluating weakly singular boundary element integrals. Journal of Computational and Applied Mathematics, 131(1– 2):223–241, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700002739.

Johnston:2002:TES

[JE02] Peter R. Johnston and David Elliott. Transformations for evaluating singular boundary element integrals. Journal of Computational and Applied Mathematics, 146 (2):231–251, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003576.

Jeannerod:2004:MPM

[Jea04] Claude-Pierre Jeannerod. On matrix perturbations with minimal leading Jordan structure. Journal of Computational and Applied Mathematics, 162(1):113–132, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007222.

Jeon:2000:SBI

[Jeo00]  Youngmok Jeon.  Scalar

boundary integral equation formulas for the biharmonic equation — numerical experiments. Journal of Computational and Applied Mathematics, 115(1–2):269–282, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001260.

Jeon:2001:NIS

[Jeo01] Youngmok Jeon. New indirect scalar boundary integral equation formulas for the biharmonic equation. Journal of Computational and Applied Mathematics, 135 (2):313–324, October 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005902.

Jiang:2001:MMC

[JFW01] Daqing Jiang, Meng Fan, and Aying Wan. A monotone method for constructing extremal solutions to second-order periodic boundary value problems. Journal of Computational and Applied Mathematics, 136 (1–2):189–197, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006105.

Guan:2004:BSM

[jGwHsZ04] Yu jing Guan, Li wen Han, and Yun shi Zhou. M-band subdivision of multivariate Bspline. Journal of Computational and Applied Mathematics, 163(1):117–126, February 1, 2004. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007994.

Ji:2004:GCA

[Ji04] H. B. Ji. A global convergent algorithm for flows in a two-dimensional network. Journal of Computational and Applied Mathematics, 167(1):135–146, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008926.

Johnson:2000:BMR

[JKN00] Russell Johnson, Mikhail Kamenskii, and Paolo Nistri. Bifurcation and multiplicity results for periodic solutions of a damped wave equation in a thin domain. Journal of Computational and Applied Mathematics, 113 (1–2):123–139, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002484.

Jukic:2004:LSF

[JKS04] Dragan Jukić, Gordana Kralik, and Rudolf Scitovski. Leastsquares fitting Gompertz curve. Journal of Computational and Applied Mathematics, 169(2):359–375, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000032.

Jin:2004:CPW

[JkSlS04] Xiao-Qing Jin, Vai kuong Sin, and Li li Song. Circulant preconditioned WR– BVM methods for ODE systems. Journal of Computational and Applied Mathematics, 162(1):201–211, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007325.

Jagannathan:2003:ICS

[JKVV03] R. Jagannathan, S. Kanemitsu, G. Vanden Berghe, and W. Van Assche. International Conference on Special Functions and Their Applications (ICSF2002). Journal of Computational and Applied Mathematics, 160 (1–2):xi–xiv, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006071.

Jaros:2002:OPS

[JKY02] Jaroslav Jaros, Takaši Kusano, and Norio Yoshida. Oscillation properties of solutions of a class of nonlinear parabolic equations. Journal of Computational and Applied Mathematics, 146 (2):277–284, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003606.

Jou:2000:PFE

[JL00] Jang Jou and Jinn-Liang Liu. An a posteriori finite element error analysis for the Stokes equations. Journal of Computational and Applied Mathematics, 114(2):333–343, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900271X.

Junghanns:2000:CMN

[JM00] Peter Junghanns and Katja Müller. A collocation method for nonlinear Cauchy singular integral equations. Journal of Computational and Applied Mathematics, 115(1– 2):283–300, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704279900117X.

Janas:2004:STC

[JNS04] Jan Janas, Serguei Naboko, and Gunter Stolz. Spectral theory for a class of periodically perturbed unbounded Jacobi matrices: elementary methods. Journal of Computational and Applied Mathematics, 171(1–2):265–276, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000639.

Jung:2002:SMM

[JNT02] M. Jung, S. Nicaise, and J. Tabka. Some multilevel methods on graded meshes. Journal of Computational and Applied Mathematics, 138(1):151–171, January 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003545.

Jiang:2003:TSS

[JO03] Qingtang Jiang and Peter Oswald. Triangular 3subdivision schemes: the regular case. Journal of Computational and Applied Mathematics, 156(1):47–75, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702009044.

John:2002:SFP

[Joh02] Volker John. Slip with friction and penetration with resistance boundary conditions for the Navier–Stokes equations-numerical tests and aspects of the implementation. Journal of Computational and Applied Mathematics, 147(2):287–300, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004375.

Jackiewicz:2004:CHS

[JPW04] Z. Jackiewicz, H. Podhaisky, and R. Weiner. Construction of highly stable two-step W -methods for ordinary differential equations. Journal of Computational and Applied Mathematics, 167 (2):389–403, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703009282.

Jackiewicz:2002:NSP

[JR02] Z. Jackiewicz and R. A. Renaut. A note on stability of pseudospectral methods for wave propagation. Journal of Computational and Applied Mathematics, 143 (1):127–139, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004952.

Jbilou:2000:VEM

[JS00a] K. Jbilou and H. Sadok. Vector extrapolation methods. Applications and numerical comparison. Journal of Computational and Applied Mathematics, 122(1–2):149–165, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003575.

Junghanns:2000:NAO

[JS00b] Peter Junghanns and Bernd Silbermann. Numerical analysis for one-dimensional Cauchy singular integral equations. Journal of Computational and Applied Mathematics, 125 (1–2):395–421, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004829.

Jukic:2003:SLS

[JS03] D. Jukić and R. Scitovski. Solution of the least-squares problem for logistic function. Journal of Computational and Applied Mathematics, 156(1):159–177, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200910X.

Jung:2004:GIP

[JS04] Jae-Hun Jung and Bernie D.

Shizgal. Generalization of the inverse polynomial reconstruction method in the resolution of the Gibbs phenomenon. Journal of Computational and Applied Mathematics, 172 (1):131–151, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001128.

Jagger:2002:WRO

[JT02a] Christopher Jagger and Andrew Thomason. Wavelength routing in optical networks of diameter two. Journal of Computational and Applied Mathematics, 142(1):115–124, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004630.

Jiang:2002:OND

[JT02b] J. C. Jiang and X. H. Tang.

Oscillation of nonlinear delay difference equations. Journal of Computational and Applied Mathematics, 146 (2):395–404, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003898.

Jayantha:2003:USI

[JT03] Pasdunkorale A. Jayantha and Ian W. Turner. On the use of surface interpolation techniques in generalised finite volume strategies for simulating transport in highly anisotropic porous media. Journal of Computational and Applied Mathematics, 152(1–2):199–216, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007069.

Junghanns:2004:OCP

[Jun04] P. Junghanns. Optimal control for a parametrized family of nonlinear Cauchy singular integral equations. Journal of Computational and Applied Mathematics, 164– 165(??):431–454, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009506.

Joshi:2003:ECC

[JV03] C. M. Joshi and Yashoverdhan Vyas. Extensions of certain classical integrals of Erdélyi for Gauss hypergeometric functions. Journal of Computational and Applied Mathematics, 160 (1–2):125–138, November 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006198.

Jiang:2000:BVP

[JW00] Daqing Jiang and Junyu Wang. On boundary value problems for singular secondorder functional differential equations. Journal of Computational and Applied Mathematics, 116(2):231–241, April 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003143.

Jin:2004:SFE

[JW04] Jicheng Jin and Xiaonan Wu. The singular finite element method for some elliptic boundary value problem with interface. Journal of Computational and Applied Mathematics, 170 (1):197–216, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000287.

Jiang:2001:LEF

[JY01] Lishang Jiang and Xingye Yue.

Local exponentially fitted finite element schemes for singularly perturbed convection– diffusion problems. Journal of Computational and Applied Mathematics, 132 (2):277–293, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700003290.

Kaiser:2001:GIP

[Kai01] Mark J. Kaiser. Generalized intersection and projection bodies. Journal of Computational and Applied Mathematics, 130(1–2):375–384, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003204.

Kainhofer:2003:QMS

[Kai03] Reinhold Kainhofer. QMC methods for the solution of delay differential equations. Journal of Computational and Applied Mathematics, 155(2):239–252, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008671.

Kalantari:2000:GTT

[Kal00]  Bahman Kalantari.  Gen-

eralization of Taylor’s theorem and Newton’s method via a new family of determinantal interpolation formulas and its applications. Journal of Computational and Applied Mathematics, 126 (1–2):287–318, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900360X.

Kalovics:2002:CHB

[Kál02] F. Kálovics. Creating and handling box valued functions used in numerical methods. Journal of Computational and Applied Mathematics, 147(2):333–348, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004405.

Kananthai:2000:CER

[Kan00] A. Kananthai. On the convolution equation related to the Ndimensional ultra-hyperbolic operator. Journal of Computational and Applied Mathematics, 115(1–2):301–308, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001727.

Kang:2004:VMR

[Kan04] Sheon-Young Kang. Volterra method for the radial Schrödinger equation. Journal of Computational and Applied Mathematics, 169(2):277–296, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703010318.

Karlsson:2000:IHF

[Kar00a] Per W. Karlsson. Inductance and hypergeometric functions. II. Journal of Computational and Applied Mathematics, 118(1–2):215–222, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002909.

Karlsson:2000:SFD

[Kar00b] Per W. Karlsson. Some formulae for double Clausenian functions. Journal of Computational and Applied Mathematics, 118(1–2):203–213, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002892.

Karatsuba:2001:ARE

[Kar01] Ekatherina A. Karatsuba. On the asymptotic representation of the Euler gamma function by Ramanujan. Journal of Computational and Applied Mathematics, 135 (2):225–240, October 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042700005860.

Kassara:2000:FSC

[Kas00]  Khalid Kassara.  Feed-

back spreading controls for semilinear parabolic systems. Journal of Computational and Applied Mathematics, 114(1):41–54, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002873.

Kobayashi:2002:MCI

[KATS02] M. Kobayashi, M. Aono, H. Takeuchi, and H. Samukawa. Matrix computations for information retrieval and major and outlier cluster detection. Journal of Computational and Applied Mathematics, 149(1):119–129, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005241.

Kazantzis:2002:EUL

[Kaz02] Nikolaos Kazantzis. On the existence and uniqueness of locally analytic invertible solutions of a system of nonlinear functional equations. Journal of Computational and Applied Mathematics, 146 (2):301–308, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200362X.

Kim:2000:QRW

[KC00a] Philsu Kim and U. Jin Choi. A quadrature rule for weighted Cauchy integrals. Journal of Computational and Applied Mathematics, 126(1– 2):221–232, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003568.

Kim:2000:QRI

[KC00b] Philsu Kim and U. Jin Choi. A quadrature rule of interpolatory type for Cauchy integrals. Journal of Computational and Applied Mathematics, 126 (1–2):207–220, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003544.

Kim:2002:PIE

[KCB02] Do Wan Kim, Jeong Ho Chu, and Hyeong-Ohk Bae. Perturbed integral equation method on determining unknown geometry in fluid flow. Journal of Computational and Applied Mathematics, 147(1):99–120, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003953.

Kim:2002:QRU

[KCI02] Kyung Joong Kim, Ronald Cools, and L. Gr. Ixaru. Quadrature rules using first derivatives for oscillatory integrands. Journal of Computational and Applied Mathematics, 140(1–2):479–497, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004836.

Khater:2002:BTE

[KCMK02] A. H. Khater, D. K. Callebaut, W. Malfliet, and E. S. Kamel. Bäcklund transformations and exact solutions for some nonlinear evolution equations in solar magnetostatic models. Journal of Computational and Applied Mathematics, 140(1–2):435–467, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004812.

Kutluay:2004:NSB

[KED04] S. Kutluay, A. Esen, and I. Dag. Numerical solutions of the Burgers’ equation by the least-squares quadratic B-spline finite element method. Journal of Computational and Applied Mathematics, 167(1):21–33, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008768.

Kalantari:2000:NMG

[KG00] Bahman Kalantari and Jürgen Gerlach. Newton’s method and generation of a determinantal family of iteration functions. Journal of Computational and Applied Mathematics, 116(1):195–200, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003611.

Khan:2002:HOD

[Kha02]  M. A. H. Khan.  High-

order differential approximants. Journal of Computational and Applied Mathematics, 149(2):457–468, December 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005617.

Kotakemori:2002:CTI

[KHMN02] Hisashi Kotakemori, Kyouji Harada, Munenori Morimoto, and Hiroshi Niki. A comparison theorem for the iterative method with the preconditioner (I + S_{max}).

Journal of Computational and Applied Mathematics, 145(2):373–378, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100588X.

Kiryakova:2000:MMM

[Kir00] Virginia S. Kiryakova. Multiple (multiindex) Mittag-Leffler functions and relations to generalized fractional calculus. Journal of Computational and Applied Mathematics, 118(1–2):241–259, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002922.

Koekoek:2000:DEG

[KK00] J. Koekoek and R. Koekoek.

Differential equations for generalized Jacobi polynomials. Journal of Computational and Applied Mathematics, 126(1–2):1–31, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003386.

Kweon:2003:EFE

[KK03] Jae Ryong Kweon and Philsu Kim. An explicit finite element method for convectiondominated compressible viscous Stokes system with inflow boundary. Journal of Computational and Applied Mathematics, 156(2):319–343, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702009196.

Kopteva:2001:USO

[KL01a] Natalia Kopteva and Torsten Linß. Uniform second-order pointwise convergence of a central difference approximation for a quasilinear convection– diffusion problem. Journal of Computational and Applied Mathematics, 137 (2):257–267, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003533.

Kwon:2001:EEL

[KL01b] K. H. Kwon and D. W. Lee. Error estimates of Lagrange interpolation and orthonormal expansions for Freud weights. Journal of Computational and Applied Mathematics, 133(1–2):445–454, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000666X.

Karonski:2002:PTR

[KL02] Michal Karoński and Tomasz Luczak. The phase transition in a random hypergraph. Journal of Computational and Applied Mathematics, 142(1):125–135, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004642.

Kress:2003:IEM

[KL03a] Rainer Kress and Kuo-Ming Lee. Integral equation methods for scattering from an impedance crack. Journal of Computational and Applied Mathematics, 161 (1):161–177, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005867.

Kubayi:2003:QSL

[KL03b] D. G. Kubayi and D. S. Lubinsky. Quadrature sums and Lagrange interpolation for general exponential weights. Journal of Computational and Applied Mathematics, 151 (2):383–414, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007471.

Kwon:2003:OFS

[KL03c] K. H. Kwon and D. W. Lee. Orthogonal functions satisfying a second-order differential equation. Journal of Computational and Applied Mathematics, 153(1–2):283–293, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200599X.

Kelner:2004:AGA

[KL04a] V. Kelner and O. Léonard. Application of genetic algorithms to lubrication pump stacking design. Journal of Computational and Applied Mathematics, 168(1–2):255–265, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009853.

Kwon:2004:TPD

[KL04b] K. H. Kwon and D. W. Lee. Tchebycheff polynomials on a disk. Journal of Computational and Applied Mathematics, 162(2):359–363, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007416.

Kwon:2001:OLC

[KLM01] K. H. Kwon, J. H. Lee, and F. Marcellán. Orthogonality of linear combinations of two orthogonal polynomial sequences. Journal of Computational and Applied Mathematics, 137(1):109–122, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007020.

Kraus:2004:VPE

[KLT04] Margarita Kraus, Matthias Langer, and Christiane Tretter. Variational principles and eigenvalue estimates for unbounded block operator matrices and applications. Journal of Computational and Applied Mathematics, 171(1–2):311–334, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000640.

Knowles:2004:RMF

[KLY04] Ian Knowles, Tuan Le, and Aimin Yan. On the recovery of multiple flow parameters from transient head data. Journal of Computational and Applied Mathematics, 169(1):1–15, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009348.

Kim:2003:NAD

[KLyD^{+}03] Woo-Je Kim, Sungmook Lim, Seung yong Doh, Soondal Park, and Jaegeun Ahn. Numerical aspects in developing LP softwares, LPAKO and LPABO. Journal of Computational and Applied Mathematics, 152(1–2):217–228, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007070.

Khater:2002:TWS

[KMCK02] A. H. Khater, W. Malfliet, D. K. Callebaut, and E. S. Kamel. Travelling wave solutions of some classes of nonlinear evolution equations in (1 + 1) and (2 + 1) dimensions. Journal of Computational and Applied Mathematics, 140(1–2):469–477, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004824.

Kalogiratou:2003:SIN

[KMS03] Z. Kalogiratou, Th. Monovasilis, and T. E. Simos. Symplectic integrators for the numerical solution of the Schrödinger equation. Journal of Computational and Applied Mathematics, 158(1):83–92, September 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004783.

Kzaz:2003:AEL

[KMV03] M. Kzaz, F. Maach, and J. Van Iseghem. Asymptotic expansions for Laguerrelike orthogonal polynomials. Journal of Computational and Applied Mathematics, 159(2):325–339, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005375.

Kondo:2002:DSS

[KN02] Koichi Kondo and Yoshimasa Nakamura. Determinantal solutions of solvable chaotic systems. Journal of Computational and Applied Mathematics, 145(2):361–372, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005878.

Knowles:2001:PIE

[Kno01] Ian Knowles. Parameter identification for elliptic problems. Journal of Computational and Applied Mathematics, 131(1–2):175–194, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002752.

Knowles:2002:IPH

[Kno02] Ian W. Knowles. An inverse problem for Hamiltonian systems. Journal of Computational and Applied Mathematics, 148(1):99–113, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005769.

Kiepiela:2003:GPT

[KNS03] K. Kiepiela, I. Naraniecka, and J. Szynal. The Gegenbauer polynomials and typically real functions. Journal of Computational and Applied Mathematics, 153(1– 2):273–282, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702006428.

Kohno:2000:ITM

[KNSmG00] Toshiyuki Kohno, Hiroshi Niki, Hideo Sawami, and Yi ming Gao. An iterative test for H-matrix. Journal of Computational and Applied Mathematics, 115(1–2):349–355, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042799003039.

Khan:1999:CFE

[KO99] Ishtiaq Rasool Khan and Ryoji Ohba. Closed-form expressions for the finite difference approximations of first and higher derivatives based on Taylor series. Journal of Computational and Applied Mathematics, 107(2):179–193, July 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799000886. See cor-

rigendum [KO01].

Khan:2000:NFD

[KO00] Ishtiaq Rasool Khan and Ryoji Ohba. New finite difference formulas for numerical differentiation. Journal of Computational and Applied Mathematics, 126 (1–2):269–276, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003581.

Khan:2001:CCF

[KO01] Ishtiaq Rasool Khan and Ryoji Ohba. Corrigendum to: “Closed-form expressions for the finite difference approximations of first and higher derivatives based on Taylor series” [J. Comput.

Appl. Math. 107 (1999) 179– 193]. Journal of Computational and Applied Mathematics, 130(1–2):385, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002636.  See

[KO99].

Khan:2003:TSB

[KO03] Ishtiaq Rasool Khan and Ryoji Ohba. Taylor series based finite difference approximations of higher-degree derivatives. Journal of Computational and Applied Mathematics, 154(1):115–124, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008166.

Kobza:2000:IFE

[Kob00] J. Kobza. Iterative functional equation x(x(t))=f(t) with f(t) piecewise linear. Journal of Computational and Applied Mathematics, 115(1–2):331–347, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003088.

Koeber:2003:ISD

[Koe03] M. Koeber. Inclusion of solutions of Darboux problems for quasilinear hyperbolic equations. Journal of Computational and Applied Mathematics, 152(1–2):243–262, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007094.

Kohaupt:2001:DCS

[Koh01] L. Kohaupt. Differential calculus for some p-norms of the fundamental matrix with applications. Journal of Computational and Applied Mathematics, 135(1):1–21, October 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005598.

Kohaupt:2002:DCN

[Koh02]  L. Kohaupt.  Differen-

tial calculus for p-norms of complex-valued vector functions with applications. Journal of Computational and Applied Mathematics, 145(2): 425–457, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701005945.

Khan:2003:MPC

[KOH03a] Ishtiaq Rasool Khan, Ryoji Ohba, and Noriyuki Hozumi. Mathematical proof of closed form expressions for finite difference approximations based on Taylor series. Journal of Computational and Applied Mathematics, 150(2):303–309, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006672.

Kohaupt:2003:EFD

[Koh03b]  L. Kohaupt.  Extension

and further development of the differential calculus for matrix norms with applications. Journal of Computational and Applied Mathematics, 156(2):433–456, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703003856.

Kontoghiorghes:2004:CMM

[Kon04] Erricos John Kontoghiorghes. Computational methods for modifying seemingly unrelated regressions models. Journal of Computational and Applied Mathematics, 162(1): 247–261, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703007179.

Korman:2001:MAU

[Kor01]  Philip Korman.  Mono-

tone approximations of unstable solutions. Journal of Computational and Applied Mathematics, 136 (1–2):309–315, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006221.

Koto:2002:SRK

[Kot02] Toshiyuki Koto. Stability of Runge–Kutta methods for delay integro-differential equations. Journal of Computational and Applied Mathematics, 145(2):483–492, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005969.

Koto:2003:SMD

[Kot03]  Toshiyuki Koto.  Stabil-

ity of θ-methods for delay integro-differential equations. Journal of Computational and Applied Mathematics, 161(2):393–404, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300712X.

Kouatchou:2003:CTS

[Kou03] Jules Kouatchou. Comparison of time and spatial collocation methods for the heat equation. Journal of Computational and Applied Mathematics, 150(1):129–141, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006568.

Kozubowski:2000:CSG

[Koz00] Tomasz J. Kozubowski. Computer simulation of geometric stable distributions. Journal of Computational and Applied Mathematics, 116(2):221–229, April 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003180.

Kobayashi:2003:NCW

[KOZ03] K. Kobayashi, H. Okamoto, and J. Zhu. Numerical computation of water and solitary waves by the double exponential transform. Journal of Computational and Applied Mathematics, 152(1–2):229–241, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007082.

Kouibia:2000:ADV

[KP00] A. Kouibia and M. Pasadas. Approximation by discrete variational splines. Journal of Computational and Applied Mathematics, 116 (1):145–156, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799003167.

Kalantari:2001:CCF

[KP01] Bahman Kalantari and Seungyoung Park. A computational comparison of the first nine members of a determinantal family of root-finding methods. Journal of Computational and Applied Mathematics, 130(1–2):197–204, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003830.

Kouibia:2002:ACF

[KP02] A. Kouibia and M. Pasadas.

Approximation of curves by fairness splines with tangent conditions. Journal of Computational and Applied Mathematics, 142(2):357–366, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004289.

Khan:2003:LTG

[KP03a] Subuhi Khan and M. A.

Pathan. Lie-theoretic generating relations of Hermite 2D polynomials. Journal of Computational and Applied Mathematics, 160 (1–2):139–146, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006344.

Kunis:2003:FSF

[KP03b] Stefan Kunis and Daniel Potts. Fast spherical Fourier algorithms. Journal of Computational and Applied Mathematics, 161(1):75–98, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005466.

Kouibia:2004:CSD

[KP04] A. Kouibia and M. Pasadas. Construction of surfaces by discrete variational splines with parallelism conditions. Journal of Computational and Applied Mathematics, 164– 165(??):455–467, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008276.

Kiepiela:2001:MPN

[KPS01] Katarzyna Kiepiela, Monika Pietrzyk, and Jan Szynal. Meixner polynomials and nonvanishing holomorphic functions. Journal of Computational and Applied Mathematics, 133(1–2):423–428, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006646.

Kokkorakis:2002:PSE

[KR02] Gerassimos C. Kokkorakis and John A. Roumeliotis. Power series expansions for spheroidal wave functions with small arguments. Journal of Computational and Applied Mathematics, 139(1):95–127, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003879.

Krattenthaler:2003:AGH

[KR03a] C. Krattenthaler and K. Srinivasa Rao. Automatic generation of hypergeometric identities by the beta integral method. Journal of Computational and Applied Mathematics, 160(1– 2):159–173, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006290.

Krattenthaler:2003:HIG

[KR03b] C. Krattenthaler and H. Rosengren. On a hypergeometric identity of Gelfand, graev and retakh. Journal of Computational and Applied Mathematics, 160 (1–2):147–158, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006204.

Krasovsky:2003:ADZ

[Kra03] I. V. Krasovsky. Asymptotic distribution of zeros of polynomials satisfying difference equations. Journal of Computational and Applied Mathematics, 150(1):57–70, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005642.

Krandick:2004:TJR

[Kra04]  Werner Krandick.  Trees

and jumps and real roots. Journal of Computational and Applied Mathematics, 162(1):51–55, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007246.

Kilbas:2002:ARH

[KRT02] Anatoly A. Kilbas, Luis Rodríguez, and Juan J. Trujillo. Asymptotic representations for hypergeometric-Bessel type function and fractional integrals. Journal of Computational and Applied Mathematics, 149 (2):469–487, December 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005629.

Koning:2000:SPC

[KRW00] Joe Koning, Garry Rodrigue, and Dan White. Scalable preconditioned conjugate gradient inversion of vector finite element mass matrices. Journal of Computational and Applied Mathematics, 123 (1–2):307–321, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004076.

Kirk:2000:SRE

[KS00] W. A. Kirk and Linda Marie Saliga. Some results on existence and approximation in metric fixed point theory. Journal of Computational and Applied Mathematics, 113(1–2):141–152, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002496. See re-

marks [Akk02].

Karatson:2001:SPN

[KS01a] J. Karátson and P. L. Simon. On the stability properties of nonnegative solutions of semilinear problems with convex or concave nonlinearity. Journal of Computational and Applied Mathematics, 131(1–2):497–501, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700007147.

Koepf:2001:SFC

[KS01b] Wolfram Koepf and Dieter Schmersau. On a structure formula for classical qorthogonal polynomials. Journal of Computational and Applied Mathematics, 136 (1–2):99–107, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000577X.

Koulorizos:2002:LPN

[KS02a] Philippos Koulorizos and Nikolas Samaris. The Landau problem for nonvanishing functions with real coefficients. Journal of Computational and Applied Mathematics, 139(1):129–139, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042701003867.

Kubiaczyk:2002:ODP

[KS02b] I. Kubiaczyk and S. H.

Saker. Oscillation of delay parabolic differential equations with several coefficients. Journal of Computational and Applied Mathematics, 147(2):263–275, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004272.

Kalogiratou:2003:NCF

[KS03a] Z. Kalogiratou and T. E.

Simos. Newton–Cotes formulae for long-time integration. Journal of Computational and Applied Mathematics, 158(1):75–82, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004795.

Konguetsof:2003:GHS

[KS03b] A. Konguetsof and T. E.

Simos. A generator of hybrid symmetric four-step methods for the numerical solution of the Schrödinger equation. Journal of Computational and Applied Mathematics, 158(1):93–106, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004692.

Khater:2000:CSC

[KSCI00] A. H. Khater, A. B. Shamardan, D. K. Callebaut, and R. S. Ibrahim. Chebyshev spectral collocation methods for nonlinear isothermal magnetostatic atmospheres. Journal of Computational and Applied Mathematics, 115(1–2):309–329, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003064.

Kravanja:2003:PRG

[KSSV03] Peter Kravanja, Tetsuya Sakurai, Hiroshi Sugiura, and Marc Van Barel. A perturbation result for generalized eigenvalue problems and its application to error estimation in a quadrature method for computing zeros of analytic functions. Journal of Computational and Applied Mathematics, 161(2):339–347, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007040.

Kilbas:2000:CFI

[KT00] A. A. Kilbas and J. J. Trujillo. Computation of fractional integrals via functions of hypergeometric and Bessel type. Journal of Computational and Applied Mathematics, 118(1–2):223–239, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002910.

Karkkainen:2001:BBO

[KT01] T. Kärkkäinen and J. Toivanen. Building blocks for odd-even multigrid with applications to reduced systems. Journal of Computational and Applied Mathematics, 131(1–2):15–33, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003265.

Kohno:2002:FES

[KT02] H. Kohno and T. Tanahashi.

Finite element simulation of single crystal growth process using GSMAC method. Journal of Computational and Applied Mathematics, 149 (1):359–371, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005435.

Kanayama:2003:FEC

[KTIS03] H. Kanayama, D. Tagami, K. Imoto, and S. Sugimoto.

Finite element computation of magnetic field problems with the displacement current. Journal of Computational and Applied Mathematics, 159(1):77–84, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005600.

Kubayi:2001:BWL

[Kub01]  D. G. Kubayi.  Bounds

for weighted Lebesgue functions for exponential weights. Journal of Computational and Applied Mathematics, 133(1–2):429–443, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006658.

Kuijlaars:2001:AAD

[Kui01a] A. B. J. Kuijlaars. Asymptotic analysis of the density of states in random matrix models associated with a slowly decaying weight. Journal of Computational and Applied Mathematics, 133(1–2):127–140, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006385.

Kuijlaars:2001:PRS

[Kui01b] Arno Kuijlaars. Problems related to slowly decaying weights. Journal of Computational and Applied Mathematics, 133(1–2):690–691, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007226.

Kumar:2002:TPF

[Kum02]  Manoj Kumar.  A three-

point finite difference method for a class of singular twopoint boundary value problems. Journal of Computational and Applied Mathematics, 145(1):89–97, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005374.

Kravanja:2001:Z

[KV01a] P. Kravanja and P. Verlinden. On the zeros of J_{n}(z) ± iJ_{n+1}(z) and [J_{n+1}(z)]^{2} − J_{n}(z)J_{n+2}(z). Journal of Computational and Applied Mathematics, 132 (2):237–245, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700003150.

Kyriakoussis:2001:ANC

[KV01b] A. Kyriakoussis and M. G.

Vamvakari. Asymptotic normality of the coefficients of polynomials associated with the Gegenbauer ones. Journal of Computational and Applied Mathematics, 133 (1–2):455–463, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006671.

Kaltofen:2004:CSV

[KV04] Erich Kaltofen and Gilles Villard. Computing the sign or the value of the determinant of an integer matrix, a complexity survey. Journal of Computational and Applied Mathematics, 162(1):133–146, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007234.

Kostreva:2000:OCS

[KW00] M. M. Kostreva and A. L. Ward. Optimal control of a system governed by an elliptic partial differential equation. Journal of Computational and Applied Mathematics, 114(1):173–187, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002952.

Kojouharov:2003:NES

[KW03] H. Kojouharov and B. Welfert.

A nonstandard Euler scheme for y′′ + g(y)y′ + f(y)y = 0. Journal of Computational and Applied Mathematics, 151(2):335–353, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007537.

Kong:2001:PSS

[KWW01] Lingbin Kong, Shoutian Wang, and Junyu Wang. Positive solution of a singular nonlinear third-order periodic boundary value problem. Journal of Computational and Applied Mathematics, 132(2):247–253, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003253.

Kong:2004:MSL

[KWZ04] Q. Kong, H. Wu, and A. Zettl. Multiplicity of Sturm–Liouville eigenvalues. Journal of Computational and Applied Mathematics, 171 (1–2):291–309, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000664.

Kwon:2000:GHT

[KY00] K. H. Kwon and G. J. Yoon. Generalized Hahn’s theorem. Journal of Computational and Applied Mathematics, 116(2):243–262, April 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003192.

Kim:2002:CIT

[KY02] Philsu Kim and Beong In Yun. On the convergence of interpolatory-type quadrature rules for evaluating Cauchy integrals. Journal of Computational and Applied Mathematics, 149(2):381–395, December 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004818.

Knowles:2004:RTP

[KY04] Ian Knowles and Aimin Yan. On the recovery of transport parameters in groundwater modelling. Journal of Computational and Applied Mathematics, 171 (1–2):277–290, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000548.

Kanzow:2004:LMM

[KYF04] Christian Kanzow, Nobuo Yamashita, and Masao Fukushima. Levenberg–Marquardt methods with strong local convergence properties for solving nonlinear equations with convex constraints. Journal of Computational and Applied Mathematics, 172 (2):375–397, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001256.

Konno:2002:CPA

[KyGUY02] Hiroshi Konno, Jun ya Gotoh, Takeaki Uno, and Atsushi Yuki. A cutting plane algorithm for semi-definite programming problems with applications to failure discriminant analysis. Journal of Computational and Applied Mathematics, 146 (1):141–154, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004247.

Karaa:2003:ASI

[KZ03] Samir Karaa and Jun Zhang.

Analysis of stationary iterative methods for the discrete convection–diffusion equation with a 9-point compact scheme. Journal of Computational and Applied Mathematics, 154(2):447–476, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008634.

Kzaz:2000:AEF

[Kza00] M. Kzaz. Asymptotic expansion of Fourier coefficients associated to functions with low continuity. Journal of Computational and Applied Mathematics, 114(2):217–230, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001909.

Lachaud:2003:SAR

[Lac03] Gilles Lachaud. Spectral analysis and the Riemann hypothesis. Journal of Computational and Applied Mathematics, 160 (1–2):175–190, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006216.

Lagomasino:2001:GHT

[Lag01] Guillermo López Lagomasino. A generalization of Hadamard’s theorem. Journal of Computational and Applied Mathematics, 133(1–2):692–693, August 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007238.

Lai:2000:CPS

[Lai00] Ming-Jun Lai. Convex preserving scattered data interpolation using bivariate C^{1} cubic splines. Journal of Computational and Applied Mathematics, 119(1–2):249–258, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003824.

Lamba:2003:ATA

[Lam03]  H. Lamba.  An adaptive

timestepping algorithm for stochastic differential equations. Journal of Computational and Applied Mathematics, 161(2):417–430, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007155.

Langlois:2004:MAF

[Lan04]  Philippe Langlois.  More

accuracy at fixed precision. Journal of Computational and Applied Mathematics, 162(1):57–77, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042703007258.

Lara:2002:AEE

[Lar02] Joaquin F. Sánchez Lara. On the asymptotic expansion of the entropy of Gegenbauer polynomials. Journal of Computational and Applied Mathematics, 142(2):401–409, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004186.

Lasser:2002:LWC

[Las02] R. Lasser. On lag windows connected with Jacobi polynomials. Journal of Computational and Applied Mathematics, 145(2):291–300, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005829.

Laschet:2004:HFF

[Las04] Gottfried Laschet. Homogenization of the fluid flow and heat transfer in transpiration cooled multi-layer plates. Journal of Computational and Applied Mathematics, 168(1–2):277–288, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009877.

Liu:2004:NSS

[LAT04] F. Liu, V. Anh, and I. Turner.

Numerical solution of the space fractional Fokker–Planck equation. Journal of Computational and Applied Mathematics, 166(1):209–219, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008616.

Laurita:2000:CNS

[Lau00] C. Laurita. Condition numbers for singular integral equations in weighted L^{2} spaces. Journal of Computational and Applied Mathematics, 116(1):23–40, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002794.

Laurie:2001:CGT

[Lau01] Dirk P. Laurie. Computation of Gauss-type quadrature formulas. Journal of Computational and Applied Mathematics, 127(1–2):201–217, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005069.

Lawrence:2002:VDR

[Law02] Bonita A. Lawrence. A variety of differentiability results for a multi-point boundary value problem. Journal of Computational and Applied Mathematics, 141(1–2):237–248, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004496.

Lyche:2000:P

[LB00] T. Lyche and M. Buhmann. Preface. Journal of Computational and Applied Mathematics, 119(1–2):ix, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003897.

Luo:2001:OAB

[LB01] J. W. Luo and D. D. Bainov. Oscillatory and asymptotic behavior of second-order neutral difference equations with maxima. Journal of Computational and Applied Mathematics, 131(1–2):333–341, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002648.

Li:2004:HOS

[LC04] Xiang-Gui Li and C. K. Chan. High-order schemes for Hamilton–Jacobi equations on triangular meshes. Journal of Computational and Applied Mathematics, 167 (1):227–241, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703008999.

Lu:2004:GAN

[LCC04] Zhonghua Lu, Xuebin Chi, and Lansun Chen. Global attractivity of nonautonomous stagestructured population models with dispersal and harvest. Journal of Computational and Applied Mathematics, 166(2):411–425, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007672.

Liu:2004:CSS

[LCF04] Mingzhu Liu, Wanrong Cao, and Zhencheng Fan. Convergence and stability of the semi-implicit Euler method for a linear stochastic differential delay equation. Journal of Computational and Applied Mathematics, 170 (2):255–268, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270400069X.

Cheng:2000:SMF

[lCHcH00] Xiao liang Cheng, Weimin Han, and Hong ci Huang. Some mixed finite element methods for biharmonic equation. Journal of Computational and Applied Mathematics, 126 (1–2):91–109, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003428.

Lagomasino:2001:SOP

[LCI01] G. López Lagomasino, H. Pijeira Cabrera, and I. Pérez Izquierdo. Sobolev orthogonal polynomials in the complex plane. Journal of Computational and Applied Mathematics, 127(1–2):219–230, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004982.

LeRoux:2001:ACD

[LCL01] Marie-Noëlle Le Roux, Marie-Isabelle Cozic, and Raphaël Loubère. Analysis of completely discrete finite element method for a free boundary diffusion problem with absorption. Journal of Computational and Applied Mathematics, 134 (1–2):143–164, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005343.

Liang:2004:ACB

[LCZ04] Xue-Zhang Liang, Li-Hong Cui, and Jie-Lin Zhang. The application of Cayley– Bacharach Theorem to bivariate Lagrange interpolation. Journal of Computational and Applied Mathematics, 163(1):177–187, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008045.

Lima:2002:NSN

[LD02] Pedro Lima and Teresa Diogo.

Numerical solution of a nonuniquely solvable Volterra integral equation using extrapolation methods. Journal of Computational and Applied Mathematics, 140(1– 2):537–557, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004083.

Leephakpreeda:2002:NDD

[Lee02] Thananchai Leephakpreeda.

Novel determination of differentialequation solutions: universal approximation method. Journal of Computational and Applied Mathematics, 146 (2):443–457, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042702003977.

Lenferink:2002:SOS

[Len02] Wim Lenferink. A second order scheme for a timedependent, singularly perturbed convection–diffusion equation. Journal of Computational and Applied Mathematics, 143(1):49–68, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005027.

Lesnic:2002:DMF

[Les02] D. Lesnic. The decomposition method for forward and backward time-dependent problems. Journal of Computational and Applied Mathematics, 147(1):27–39, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702003886.

Letessier:2001:FOD

[Let01] Jean Letessier. Fourth-order difference equation for corecursive associated Meixner and Charlier polynomials. Journal of Computational and Applied Mathematics, 133(1–2):465–476, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042700006683.

Leugering:2000:SDO

[Leu00] G. Leugering. On the semidiscretization of optimal control problems for networks of elastic strings: global optimality systems and domain decomposition. Journal of Computational and Applied Mathematics, 120 (1–2):133–157, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003071.

Leviatan:2000:SPA

[Lev00] D. Leviatan. Shape-preserving approximation by polynomials. Journal of Computational and Applied Mathematics, 121 (1–2):73–94, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003381.

Lewanowicz:2000:RFB

[Lew00] Stanislaw Lewanowicz. Recursion formulae for basic hypergeometric functions. Journal of Computational and Applied Mathematics, 121 (1–2):297–312, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042700003344.

Lewanowicz:2003:CRC

[Lew03a] Stanislaw Lewanowicz. Construction of recurrences for the coefficients of expansions in q-classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 153(1–2):295–309, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006015.

Lewanowicz:2003:RFA

[Lew03b] Stanislaw Lewanowicz. Representations for the first associated q-classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 150 (2):311–327, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006684.

Li:2001:MBM

[LF01] Dong-Hui Li and Masao Fukushima. A modified BFGS method and its global convergence in nonconvex minimization. Journal of Computational and Applied Mathematics, 129(1–2):15–35, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005409.

Lee:2003:SSA

[LF03] Jongwoo Lee and Bengt Fornberg. A split step approach for the 3-D Maxwell’s equations. Journal of Computational and Applied Mathematics, 158 (2):485–505, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004849.

Lee:2004:SUS

[LF04] Jongwoo Lee and Bengt Fornberg. Some unconditionally stable time stepping methods for the 3D Maxwell’s equations. Journal of Computational and Applied Mathematics, 166(2):497–523, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007957.

Lu:2004:EPP

[LG04] Shiping Lu and Weigao Ge. Existence of positive periodic solutions for neutral logarithmic population model with multiple delays. Journal of Computational and Applied Mathematics, 166(2):371–383, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300757X.

Li:2003:AID

[LGL03] Lei Li, James Glimm, and Xiaolin Li. All isomorphic distinct cases for multicomponent interfaces in a block. Journal of Computational and Applied Mathematics, 152(1–2):263–276, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007100.

Liu:2003:ELB

[LH03] Huan-Wen Liu and Don Hong. An explicit local basis for C^{1} cubic spline spaces over a triangulated quadrangulation. Journal of Computational and Applied Mathematics, 155(1):187–200, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702009007.

Li:2004:PSS

[LHHW04] Zi-Cai Li, Hsin-Yun Hu, Chung-Hua Hsu, and Song Wang. Particular solutions of singularly perturbed partial differential equations with constant coefficients in rectangular domains, Part I. Convergence analysis. Journal of Computational and Applied Mathematics, 166 (1):181–208, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703008628.

Li:2003:MGS

[LHYaC03] Wei Li, Ichiro Hagiwara, Takao Yasui, and Hu awei Chen. A method of generating scratched look calligraphy characters using mathematical morphology. Journal of Computational and Applied Mathematics, 159(1):85–90, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005739.

Li:2002:IRP

[Li02]  Xin Li.  Interpolation by

ridge polynomials and its application in neural networks. Journal of Computational and Applied Mathematics, 144(1–2):197–209, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100560X.

Li:2003:CMG

[Li03a] Wen Li. The convergence of the modified Gauss–Seidel methods for consistent linear systems. Journal of Computational and Applied Mathematics, 154(1):97–105, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008129.

Li:2003:EPE

[Li03b] Xin Li. An extremal problem and an estimation of the Wronskian of certain Jacobi polynomials. Journal of Computational and Applied Mathematics, 153(1–2):311–320, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006337.

Lievens:2004:EFT

[Lie04] S. Lievens. Expansion formulas for terminating balanced _{4}F_{3}-series from the Biedenharn–Elliot identity for su(1, 1). Journal of Computational and Applied Mathematics, 169(2):419–430, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000068.

Lin:2000:QPE

[Lin00] Chin-Yuan Lin. Quasilinear parabolic equations with nonlinear boundary conditions. Journal of Computational and Applied Mathematics, 126 (1–2):339–349, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003635.

Liu:2001:CGF

[Liu01]  Xian Liu.  A class of

generalized filled functions with improved computability. Journal of Computational and Applied Mathematics, 137(1):61–69, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000697X.

Liu:2002:NSF

[Liu02a] Jijun Liu. Numerical solution of forward and backward problem for 2-D heat conduction equation. Journal of Computational and Applied Mathematics, 145(2):459–482, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005957.

Liu:2002:BSH

[Liu02b]  Xiaoyan Liu.  Bivariate

splines on hexagonal lattice for digital image processing. Journal of Computational and Applied Mathematics, 144(1–2):211–220, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005623.

Liu:2003:GEN

[LJ03] Quan Liu and Song Jiang. Global existence and numerical approximation in adiabatic shearing of incompressible fluids. Journal of Computational and Applied Mathematics, 154(2):303–317, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008270.

Lamberg:2001:NSM

[LK01] L. Lamberg and M. Kaasalainen. Numerical solution of the Minkowski problem. Journal of Computational and Applied Mathematics, 137 (2):213–227, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003600.

Liu:2000:BSN

[LL00] Kangsheng Liu and Zhuangyi Liu. Boundary stabilization of a nonhomogeneous beam with rotatory inertia at the tip. Journal of Computational and Applied Mathematics, 114(1):1–10, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002848.

Landau:2001:AEB

[LL01a] L. J. Landau and N. J.

Luswili. Asymptotic expansion of a Bessel function integral using hypergeometric functions. Journal of Computational and Applied Mathematics, 132(2):387–397, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004416. See com-

ments [Sto05].

Liu:2001:OTS

[LL01b] Shu Tang Liu and Yong Qing Liu. Oscillation theorems for second-order nonlinear partial difference equations. Journal of Computational and Applied Mathematics, 132 (2):479–482, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701003739.

Lai:2002:DTN

[LL02] H. C. Lai and J. C. Lee. On duality theorems for a nondifferentiable minimax fractional programming. Journal of Computational and Applied Mathematics, 146 (1):115–126, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004223.

Levin:2003:AEI

[LL03a] Eli Levin and D. S. Lubinsky. Asymptotics for entropy integrals associated with exponential weights. Journal of Computational and Applied Mathematics, 156(2):265–283, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702009159.

Lin:2003:ICD

[LL03b] En-Bing Lin and Yi Ling. Image compression and denoising via nonseparable wavelet approximation. Journal of Computational and Applied Mathematics, 155(1):131–152, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008968.

Liang:2001:SSK

[LLC01] Xue-Zhang Liang, Ming-Cai Liu, and Xiang-Jiu Che. Solving second kind integral equations by Galerkin methods with continuous orthogonal wavelets. Journal of Computational and Applied Mathematics, 136 (1–2):149–161, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005811.

Luo:2004:FGM

[LLG04] Xiaonan Luo, Ning Liu, and Chengying Gao. Fairing geometric modeling based on 4point interpolatory subdivision scheme. Journal of Computational and Applied Mathematics, 163(1):189–197, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008057.

Liu:2003:CSF

[LLL03] M. B. Liu, G. R. Liu, and K. Y.

Lam. Constructing smoothing functions in smoothed particle hydrodynamics with applications. Journal of Computational and Applied Mathematics, 155(2):263–284, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008695.

Luyckx:2004:CMN

[LLN04] L. Luyckx, M. Loccufier, and E. Noldus. Computational methods in nonlinear stability analysis: stability boundary calculations. Journal of Computational and Applied Mathematics, 168(1–2):289–297, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009907.

Laitinen:2003:ADD

[LLP03] E. Laitinen, A. V. Lapin, and J. Pieskä. Asynchronous domain decomposition methods for continuous casting problem. Journal of Computational and Applied Mathematics, 154(2):393–413, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008609.

Liu:2004:WMG

[LLXS04] Fengshan Liu, Yi Ling, Xianggen Xia, and Xiquan Shi. Wavelet methods for ground penetrating radar imaging. Journal of Computational and Applied Mathematics, 169(2):459–474, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270400010X.

Lazzaro:2002:RBF

[LM02a] Damiana Lazzaro and Laura B. Montefusco. Radial basis functions for the multivariate interpolation of large scattered data sets. Journal of Computational and Applied Mathematics, 140(1–2):521–536, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270100485X.

Lubinsky:2002:BMO

[LM02b] D. S. Lubinsky and H. P.

Mashele. L_{p} boundedness of (C, 1) means of orthonormal expansions for general exponential weights. Journal of Computational and Applied Mathematics, 145(2):387–405, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005908.

Lopez-Moreno:2003:AEM

[LMMD03] A.-J. López-Moreno and F.- J. Muñoz-Delgado. Asymptotic expansion of multivariate conservative linear operators. Journal of Computational and Applied Mathematics, 150(2):219–251, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006611.

Lube:2001:NDD

[LMO01] G. Lube, L. Müller, and F. C. Otto. A nonoverlapping domain decomposition method for stabilized finite element approximations of the Oseen equations. Journal of Computational and Applied Mathematics, 132(2):211–236, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003216.

Lopez:2000:NTM

[LMP00] L. Lopez, C. Mastroserio, and T. Politi. Newton-type methods for solving nonlinear equations on quadratic matrix groups. Journal of Computational and Applied Mathematics, 115(1–2):357–368, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001843.

Laforgia:2003:AEM

[LMS03] Andrea Laforgia, Martin E. Muldoon, and Panayiotos D. Siafarikas. Árpád Elbert, 1939–2001: a memorial tribute. Journal of Computational and Applied Mathematics, 153(1–2):1–8, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042702006738.

Lamberg:2001:SEG

[LMYL01] L. Lamberg, K. Muinonen, J. Ylönen, and K. Lumme. Spectral estimation of Gaussian random circles and spheres. Journal of Computational and Applied Mathematics, 136(1–2):109–121, November 1, 2001. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042700005781.

Liang:2002:SFV

[LMZ02] Shengde Liang, Xiuling Ma, and Aihui Zhou. A symmetric finite volume scheme for selfadjoint elliptic problems. Journal of Computational and Applied Mathematics, 147(1):121–136, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004284.

Lewis:2003:GIS

[LN03] D. Lewis and N. Nigam. Geometric integration on spheres and some interesting applications. Journal of Computational and Applied Mathematics, 151(1):141–170, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007434.

Levin:2002:DNM

[LNBI02] Yuri Levin, Mikhail Nediak, and Adi Ben-Israel. A direct Newton method for calculus of variations. Journal of Computational and Applied Mathematics, 139 (2):197–213, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004277.

Luo:2002:RSA

[LNLL02] Xiaonan Luo, Hui Nie, Yi Li, and Zhongxuan Luo. Recurrence surfaces on arbitrary quadrilateral mesh. Journal of Computational and Applied Mathematics, 144(1–2):221–232, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005611.

Lorentz:2000:MHI

[Lor00] R. A. Lorentz. Multivariate Hermite interpolation by algebraic polynomials: a survey. Journal of Computational and Applied Mathematics, 122(1–2):167–201, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003678.

Lorente:2003:ISL

[Lor03a] Miguel Lorente. Integrable systems on the lattice and orthogonal polynomials of discrete variable. Journal of Computational and Applied Mathematics, 153(1–2):321–330, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005927.

Lorente:2003:OPS

[Lor03b]  Miguel Lorente.  Orthog-

onal polynomials, special functions and mathematical physics. Journal of Computational and Applied Mathematics, 153(1–2):543–545, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008865.

Lotstedt:2000:CAI

[Löt00] Per Lötstedt. Convergence analysis of iterative methods by pseudodifference operators. Journal of Computational and Applied Mathematics, 115(1–2):397–417, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042799001193.

Louchard:2002:RGD

[Lou02] Guy Louchard. Runs of geometrically distributed random variables: a probabilistic analysis. Journal of Computational and Applied Mathematics, 142(1):137–153, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004654.

Love:2000:AFS

[Lov00]  E. R. Love.  Another

Fourier-style expansion in series of Legendre functions. Journal of Computational and Applied Mathematics, 118(1–2):261–282, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002934.

Li:2000:SAS

[LP00] Shengtai Li and Linda Petzold. Software and algorithms for sensitivity analysis of large-scale differential algebraic systems. Journal of Computational and Applied Mathematics, 125 (1–2):131–145, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042700004647.

Liakhovetski:2001:AEL

[LP01] G. V. Liakhovetski and R. B.

Paris. Asymptotic expansions of Laplace-type integrals. III. Journal of Computational and Applied Mathematics, 132(2):409–429, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000443X.

Lorente-Pardo:2000:CBN

[LPSSP00] J. Lorente-Pardo, P. Sablonnière, and M. C. Serrano-Pérez. On the convexity of Bézier nets of quadratic Powell–Sabin splines on 12-fold refined triangulations. Journal of Computational and Applied Mathematics, 115(1–2):383–396, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003052.

Liao:2001:SNC

[LQQ01] Li-Zhi Liao, Houduo Qi, and Liqun Qi. Solving nonlinear complementarity problems with neural networks: a reformulation method approach. Journal of Computational and Applied Mathematics, 131(1–2):343–359, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002624.

Li:2002:SCT

[LR02] Xian-Fang Li and Er-Qian Rong. Solution of a class of two-dimensional integral equations. Journal of Computational and Applied Mathematics, 145(2):335–343, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005854.

Laforgia:2003:P

[LR03] A. Laforgia and P. E. Ricci. Preface. Journal of Computational and Applied Mathematics, 153(1–2):xi, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006726.

Laouafa:2004:IAF

[LR04a] F. Laouafa and P. Royis. An iterative algorithm for finite element analysis. Journal of Computational and Applied Mathematics, 164– 165(??):469–491, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008306.

Limic:2004:NSD

[LR04b] Nedzad Limić and Mladen Rogina. Numerical solutions of Dirichlet problem for elliptic operator in divergence form with a righthand side measure. Journal of Computational and Applied Mathematics, 164– 165(??):493–516, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006496.

Lehn:2002:ASN

[LRS02] J. Lehn, A. Rößler, and O. Schein. Adaptive schemes for the numerical solution of SDEs-a comparison. Journal of Computational and Applied Mathematics, 138 (2):297–308, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003752.

Lotstedt:2004:PIS

[LRvSS04] Per Lötstedt, Alison Ramage, Lina von Sydow, and Stefan Söderberg. Preconditioned implicit solution of linear hyperbolic equations with adaptivity. Journal of Computational and Applied Mathematics, 170 (2):269–289, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000706.

Luksan:2000:VMM

[LS00a] Ladislav Luksan and Emilio Spedicato. Variable metric methods for unconstrained optimization and nonlinear least squares. Journal of Computational and Applied Mathematics, 124(1–2):61–95, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004209.

Lyche:2000:NCN

[LS00b] Tom Lyche and Karl Scherer. On the p-norm condition number of the multivariate triangular Bernstein basis. Journal of Computational and Applied Mathematics, 119(1–2):259–273, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003836.

Luo:2001:SRI

[LS01] Zhiguo Luo and Jianhua Shen. Stability results for impulsive functional differential equations with infinite delays. Journal of Computational and Applied Mathematics, 131(1–2):55–64, June 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000323X.

Lang:2002:HCS

[LS02a] A. W. Lang and D. M. Sloan.

Hermite collocation solution of near-singular problems using numerical coordinate transformations based on adaptivity. Journal of Computational and Applied Mathematics, 140(1–2):499–520, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005283.

Lott:2002:SMR

[LS02b] Jason Lott and Günter Stolz.

The spectral minimum for random displacement models. Journal of Computational and Applied Mathematics, 148(1):133–146, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005782.

Lupu:2002:ESO

[LS02c] Mircea Lupu and Ernest Scheiber. Exact solution for an optimal impermeable parachute problem. Journal of Computational and Applied Mathematics, 147 (2):277–286, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004363.

Langville:2004:KPS

[LS04a] Amy N. Langville and William J. Stewart. The Kronecker product and stochastic automata networks. Journal of Computational and Applied Mathematics, 167(2):429–447, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009312.

Li:2004:MPS

[LS04b] Wan-Tong Li and Hong-Rui Sun. Multiple positive solutions for nonlinear dynamical systems on a measure chain. Journal of Computational and Applied Mathematics, 162(2):421–430, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007568.

Liu:2004:SRK

[LSY04] M. Z. Liu, M. H. Song, and Z. W. Yang. Stability of Runge–Kutta methods in the numerical solution of equation u′(t) = au(t) + a_{0}u([t]). Journal of Computational and Applied Mathematics, 166(2):361–370, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007362.

Lewis:2000:DSM

[LTT00] Robert Michael Lewis, Virginia Torczon, and Michael W. Trosset. Direct search methods: then and now. Journal of Computational and Applied Mathematics, 124 (1–2):191–207, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004234.

Lu:2003:CMF

[Lu03]  Ya Yan Lu.  Computing

a matrix function for exponential integrators. Journal of Computational and Applied Mathematics, 161 (1):203–216, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006678.

Lui:2000:DDM

[Lui00] S. H. Lui. Domain decomposition methods for eigenvalue problems. Journal of Computational and Applied Mathematics, 117(1):17–34, May 1, 2000. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704279900326X.

Lui:2001:ACN

[Lui01]  S. H. Lui.  On acceler-

ated convergence of nonoverlapping Schwarz methods. Journal of Computational and Applied Mathematics, 130(1–2):309–321, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704279900374X.

Lui:2003:MIS

[Lui03] S.-H. Lui. On monotone iteration and Schwarz methods for nonlinear parabolic PDEs. Journal of Computational and Applied Mathematics, 161(2):449–468, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007350.

Lupianez:2001:CFO

[Lup01] Francisco Gallego Lupiañez. Continued fractions and orderpreserving homeomorphism. Journal of Computational and Applied Mathematics, 136 (1–2):255–258, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006178.

Lakshmikantham:2000:EFP

[LV00] V. Lakshmikantham and A. S.

Vatsala. Existence of fixed points of fuzzy mappings via theory of fuzzy differential equations. Journal of Computational and Applied Mathematics, 113 (1–2):195–200, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002538.

Lakshmikantham:2002:HST

[LV02] V. Lakshmikantham and A. S.

Vatsala. Hybrid systems on time scales. Journal of Computational and Applied Mathematics, 141(1– 2):227–235, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004484.

Lievens:2003:RCV

[LV03] S. Lievens and J. Van der Jeugt. Realizations of coupled vectors in the tensor product of representations of ∫⊓(∞, ∞) and ∫⊓(∈). Journal of Computational and Applied Mathematics, 160 (1–2):191–208, November 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006228.

Lahaye:2004:AMP

[LVH04] Domenico Lahaye, Stefan Vandewalle, and Kay Hameyer. An algebraic multilevel preconditioner for field-circuit coupled problems. Journal of Computational and Applied Mathematics, 168(1–2):267–275, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009865.

Lau:2001:MQS

[LW01] Karen K. Lau and Robert S. Womersley. Multistage quadratic stochastic programming. Journal of Computational and Applied Mathematics, 129(1– 2):105–138, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700005458.

Liu:2002:DQS

[LW02] G. R. Liu and T. Y. Wu. Differential quadrature solutions of eighth-order boundaryvalue differential equations. Journal of Computational and Applied Mathematics, 145(1):223–235, August 1, 2002. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005775.

Lakey:2003:PIG

[LW03a] Joseph D. Lakey and Ying Wang. On perturbations of irregular Gabor frames. Journal of Computational and Applied Mathematics, 155 (1):111–129, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702008956.

Lecko:2003:GPS

[LW03b] Adam Lecko and Agnieszka Wiśniowska. Geometric properties of subclasses of starlike functions. Journal of Computational and Applied Mathematics, 155(2):383–387, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008750.

Li:2004:MQN

[LW04] Chong-Jun Li and Ren-Hong Wang. The multivariate quartic NURBS surfaces. Journal of Computational and Applied Mathematics, 163(1):155–164, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008021.

Li:2004:COG

[LWT04] Zhengxue Li, Wei Wu, and Yulong Tian. Convergence of an online gradient method for feedforward neural networks with stochastic inputs. Journal of Computational and Applied Mathematics, 163(1):165–176, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008033.

Liu:2001:CSS

[LWW01] Hailiang Liu, Jinghua Wang, and Gerald Warnecke. Convergence of a splitting scheme applied to the Ruijgrok–Wu model of the Boltzmann equation. Journal of Computational and Applied Mathematics, 134 (1–2):343–367, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005586.

Liao:2002:SAC

[LWY02] Xiaofeng Liao, Zhongfu Wu, and Juebang Yu. Stability analyses of cellular neural networks with continuous time delay. Journal of Computational and Applied Mathematics, 143(1):29–47, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005039.

Xu:2003:NAB

[lXsQsJ03] Cheng long Xu, Xiao song Qian, and Li shang Jiang. Numerical analysis on binomial tree methods for a jump-diffusion model. Journal of Computational and Applied Mathematics, 156 (1):23–45, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702009032.

Liu:2000:PEE

[LY00] Wenbin Liu and Ningning Yan. A posteriori error estimates for some model boundary control problems. Journal of Computational and Applied Mathematics, 120 (1–2):159–173, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003083.

Lin:2001:SEP

[LY01] Lai-Jiu Lin and Zenn-Tsuen Yu. On some equilibrium problems for multimaps. Journal of Computational and Applied Mathematics, 129(1–2):171–183, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005483.

Li:2002:NEE

[LY02] Zi Cai Li and Ningning Yan. New error estimates of bicubic Hermite finite element methods for biharmonic equations. Journal of Computational and Applied Mathematics, 142(2):251–285, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004940.

Li:2002:TOR

[LYC02] Xiang-Gui Li, Xi-Jun Yu, and Guang-Nan Chen. The third-order relaxation schemes for hyperbolic conservation laws. Journal of Computational and Applied Mathematics, 138(1):93–108, January 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003612.

Li:2003:SSD

[LYF03] Zi-Cai Li, Tetsuro Yamamoto, and Qing Fang. Superconvergence of solution derivatives for the Shortley–Weller difference approximation of Poisson’s equation. Part I: smoothness problems. Journal of Computational and Applied Mathematics, 151 (2):307–333, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007549.

Lin:2001:CCI

[LZ01] E. B. Lin and X. Zhou.

Connection coefficients on an interval and wavelet solutions of Burgers equation. Journal of Computational and Applied Mathematics, 135(1):63–78, October 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005628.

Li:2002:CEP

[LZ02a] Wan-Tong Li and Shunian Zhang. Classifications and existence of positive solutions of higher order nonlinear iterative functional differential equations. Journal of Computational and Applied Mathematics, 139 (2):351–367, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004149.

Liu:2002:NBB

[LZ02b] G. S. Liu and J. Z. Zhang. A new branch and bound algorithm for solving quadratic programs with linear complementarity constraints. Journal of Computational and Applied Mathematics, 146(1): 77–87, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702004193.

Li:2004:NER

[LZ04] Ruiming Li and Dian Zhou. A note on the eigenvalue relationship for USAOR iterative method applied to p-cyclic matrices. Journal of Computational and Applied Mathematics, 169(1):213–225, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703010264.

Mallier:2004:IRS

[MA04] R. Mallier and G. Alobaidi. Interest rate swaps under CIR. Journal of Computational and Applied Mathematics, 164– 165(??):543–554, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004904.

MacLeod:2002:AEZ

[Mac02a] Allan J. MacLeod. Asymptotic expansions for the zeros of certain special functions. Journal of Computational and Applied Mathematics, 145(2):261–267, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005805.

MacLeod:2002:ECS

[Mac02b] Allan J. MacLeod. The efficient computation of some generalised exponential integrals. Journal of Computational and Applied Mathematics, 148(2):363–374, November 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005563.

MacLeod:2002:REZ

[Mac02c] Allan J. MacLeod. Rational expressions for the zeros of several special functions. Journal of Computational and Applied Mathematics, 145(1):237–246, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005787.

Maire:2003:ICA

[Mai03]  Sylvain Maire.  An iter-

ative computation of approximations on Korobov-like spaces. Journal of Computational and Applied Mathematics, 157(2):261–281, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004102.

Mittal:2001:ACR

[MAK01] R. C. Mittal and Ahmad Al-Kurdi. Application of the Cramer rule in the solution of sparse systems of linear algebraic equations. Journal of Computational and Applied Mathematics, 136(1–2):1–15, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005720.

Makino:2002:GP

[Mak02] Junichiro Makino. GRAPE project. Journal of Computational and Applied Mathematics, 149(1):131–145, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005253.

Maklakov:2004:SRE

[Mak04] Dmitri V. Maklakov. Some remarks on the exact solution for an optimal impermeable parachute problem. Journal of Computational and Applied Mathematics, 166(2):591–596, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009270.

Malengier:2004:PIS

[Mal04a]  B. Malengier.  Parame-

ter identification in stationary groundwater flow problems in drainage basins. Journal of Computational and Applied Mathematics, 168(1–2):299–307, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009919.

Malfliet:2004:TMT

[Mal04b] W. Malfliet. The tanh method:

a tool for solving certain classes of nonlinear evolution and wave equations. Journal of Computational and Applied Mathematics, 164– 165(??):529–541, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042703006459.

Manni:2000:GPF

[Man00a]  Carla Manni.  A general

parametric framework for functional tension schemes. Journal of Computational and Applied Mathematics, 119(1–2):275–300, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003848.

Mantica:2000:CJM

[Man00b] Giorgio Mantica. On computing Jacobi matrices associated with recurrent and Möbius iterated function systems. Journal of Computational and Applied Mathematics, 115(1–2):419–431, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001880.

Marcellan:2001:FTE

[MÁN01] Francisco Marcellán and Renato Álvarez-Nodarse. On the “Favard theorem” and its extensions. Journal of Computational and Applied Mathematics, 127(1–2):231–254, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042700004970.

Medem:2001:PDS

[MÁNM01] J. C. Medem, R. Álvarez-Nodarse, and F. Marcellán. On the q-polynomials: a distributional study. Journal of Computational and Applied Mathematics, 135 (2):157–196, October 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005847.

Marchand:2000:UBS

[Mar00a]  Rich Marchand.  Uniform

boundary stabilization of nonlinear spherical shells by using two controls only: analysis and numerical computations. Journal of Computational and Applied Mathematics, 114(1):189–216, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002964.

Martinez:2000:PQN

[Mar00b] José Mario Martínez. Practical quasi-Newton methods for solving nonlinear systems. Journal of Computational and Applied Mathematics, 124 (1–2):97–121, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004349.

Marek:2003:QBD

[Mar03]  Ivo Marek.  Quasi-birth-

and-death processes, levelgeometric distributions. An aggregation/disaggregation approach. Journal of Computational and Applied Mathematics, 152(1–2):277–288, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007112.

Markov:2004:QSC

[Mar04a]  Svetoslav Markov.  On

quasilinear spaces of convex bodies and intervals. Journal of Computational and Applied Mathematics, 162(1):93–112, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300726X.

Markus:2004:SLG

[Mar04b]  Lawrence Markus.  The

Sturm–Liouville group. Journal of Computational and Applied Mathematics, 171 (1–2):335–365, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042704000585.

Marletta:2004:EPE

[Mar04c]  M. Marletta.  Eigenvalue

problems on exterior domains and Dirichlet to Neumann maps. Journal of Computational and Applied Mathematics, 171(1–2):367–391, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000597.

Maset:2003:NSR

[Mas03] S. Maset. Numerical solution of retarded functional differential equations as abstract Cauchy problems. Journal of Computational and Applied Mathematics, 161 (2):259–282, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007003.

Matsuo:2003:HOS

[Mat03]  Takayasu Matsuo.  High-

order schemes for conservative or dissipative systems. Journal of Computational and Applied Mathematics, 152(1–2):305–317, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042702007136.

Mermri:2003:CRS

[MC03] B. Mermri and X. Chen. On characterizations and regularity of the solution of bilateral obstacle problems. Journal of Computational and Applied Mathematics, 152(1–2):333–345, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200715X.

Milovanovic:2005:ROS

[MC05] Gradimir V. Milovanović and Aleksandar S. Cvetković. Remarks on “Orthogonality of some sequences of the rational functions and Müntz polynomials”. Journal of Computational and Applied Mathematics, 173(2):383–388, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704002262.  See

[MDSR04].

McCarthy:2000:ODT

[McC00] C. Maeve McCarthy. The optimal design of tubular structures. Journal of Computational and Applied Mathematics, 114(1):55–66, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002885.

McCartin:2001:EFD

[McC01] Brian J. McCartin. Exponential fitting of the delayed recruitment/renewal equation. Journal of Computational and Applied Mathematics, 136 (1–2):343–356, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006257.

Mazumder:2000:CDE

[MD00] B. S. Mazumder and D. C.

Dalal. Contaminant dispersion from an elevated timedependent source. Journal of Computational and Applied Mathematics, 126(1– 2):185–205, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003532.

Masset:2004:MPS

[MD04] Luc Masset and Jean-François Debongnie. Machining processes simulation: specific finite element aspects. Journal of Computational and Applied Mathematics, 168(1– 2):309–320, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703009932.

Meijer:2002:ZSO

[MdB02] H. G. Meijer and M. G. de Bruin. Zeros of Sobolev orthogonal polynomials following from coherent pairs. Journal of Computational and Applied Mathematics, 139 (2):253–274, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004216.

Marinkovic:2004:OSS

[MDSR04] S. D. Marinković, B. Danković, M. S. Stanković, and P. M. Rajković. Orthogonality of some sequences of the rational functions and the Müntz polynomials. Journal of Computational and Applied Mathematics, 163(2):419–427, February 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007647. See re-

marks [MC05].

Meek:2002:CPC

[Mee02]  D. S. Meek.  Coaxing

a planar curve to comply. Journal of Computational and Applied Mathematics, 140(1–2):599–618, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004782.

Melenk:2002:CNH

[Mel02]  J. M. Melenk.  On con-

dition numbers in hp - FEM with Gauss–Lobattobased shape functions. Journal of Computational and Applied Mathematics, 139(1): 21–48, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701003910.

Martinez-Finkelshtein:2001:AAS

[MF01] A. Martínez-Finkelshtein. Analytic aspects of Sobolev orthogonal polynomials revisited. Journal of Computational and Applied Mathematics, 127(1–2):255–266, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004994.

Martinez-Finkelshtein:2001:AZD

[MFMGO01] Andrei Martínez-Finkelshtein, Pedro Martínez-González, and Ramón Orive. On asymptotic zero distribution of Laguerre and generalized Bessel polynomials with varying parameters. Journal of Computational and Applied Mathematics, 133(1–2):477–487, August 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006543.

Martinez-Finkelshtein:2002:WAZ

[MFMGZ02] A. Martínez-Finkelshtein, P. Martínez-González, and A. Zarzo. WKB approach to zero distribution of solutions of linear second order differential equations. Journal of Computational and Applied Mathematics, 145(1):167–182, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005428.

Mainardi:2000:MLT

[MG00] Francesco Mainardi and Rudolf Gorenflo. On Mittag-Leffler-type functions in fractional evolution processes. Journal of Computational and Applied Mathematics, 118(1– 2):283–299, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700002946.

Meddahi:2001:FDB

[MG01] Salim Meddahi and María González. A fully discrete BEM–FEM method for an exterior elasticity system in the plane. Journal of Computational and Applied Mathematics, 134 (1–2):127–141, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005331.

Moghadas:2003:MSM

[MG03] S. M. Moghadas and A. B.

Gumel. A mathematical study of a model for childhood diseases with non-permanent immunity. Journal of Computational and Applied Mathematics, 157(2):347–363, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004163.

Martin:2001:MTA

[MGL01] Pablo Martín, Amelia García, and David J. López. Modified Taylor approximation of functions with periodic behaviour. Journal of Computational and Applied Mathematics, 130(1–2):91–97, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003891.

Muroya:2003:AOC

[MIB03] Yoshiaki Muroya, Emiko Ishiwata, and Hermann Brunner. On the attainable order of collocation methods for pantograph integro-differential equations. Journal of Computational and Applied Mathematics, 152(1–2):347–366, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007161.

Migorski:2001:ESP

[Mig01]  Stanislaw Migórski.  On

existence of solutions for parabolic hemivariational inequalities. Journal of Computational and Applied Mathematics, 129(1–2):77–87, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005434.

Miller:2000:CCW

[Mil00] Allen R. Miller. On the critical case of the Weber–Schafheitlin integral and a certain generalization. Journal of Computational and Applied Mathematics, 118(1–2):301–309, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002958.

Miller:2001:MTP

[Mil01a] Allen R. Miller. The Mellin transform of a product of two hypergeometric functions. Journal of Computational and Applied Mathematics, 137(1):77–82, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006993.

Milovanovic:2001:QMN

[Mil01b] Gradimir V. Milovanović.

Quadratures with multiple nodes, power orthogonality, and moment-preserving spline approximation. Journal of Computational and Applied Mathematics, 127 (1–2):267–286, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005008.

Milaszewicz:2003:BNM

[Mil03a]  J. P. Milaszewicz.  On

Brown’s and Newton’s methods with convexity hypotheses. Journal of Computational and Applied Mathematics, 150(1):1–24, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004892.

Miller:2003:KTT

[Mil03b]  Allen R. Miller.  On a

Kummer-type transformation for the generalized hypergeometric function _{2}F_{2}. Journal of Computational and Applied Mathematics, 157(2): 507–509, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703003832.

Minamoto:2001:NEU

[Min01] Teruya Minamoto. Numerical existence and uniqueness proof for solutions of nonlinear hyperbolic equations. Journal of Computational and Applied Mathematics, 135(1):79–90, October 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000563X.

Minchev:2003:SAS

[Min03] Borislav V. Minchev. Some algorithms for solving special tridiagonal block Toeplitz linear systems. Journal of Computational and Applied Mathematics, 156(1):179–200, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702009111.

Minamoto:2004:NMG

[Min04] Teruya Minamoto. Numerical method with guaranteed accuracy of a double turning point for a radially symmetric solution of the perturbed Gelfand equation. Journal of Computational and Applied Mathematics, 169(1):151–160, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009579.

Miura:2002:AEC

[Miu02]  Robert M. Miura.  Anal-

ysis of excitable cell models. Journal of Computational and Applied Mathematics, 144(1–2):29–47, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005507.

Miyakoda:2002:AFD

[Miy02] T. Miyakoda. On an almost free damping vibration equation using N-fractional calculus. Journal of Computational and Applied Mathematics, 144(1–2):233–240, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005635.

Miyake:2003:SFM

[Miy03] Katsuya Miyake. Some families of Mordell curves associated to cubic fields. Journal of Computational and Applied Mathematics, 160 (1–2):217–231, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006241.

Mazuch:2001:NRA

[MK01] Tibor Mazúch and Jan Kozánek. New recurrent algorithm for a matrix inversion. Journal of Computational and Applied Mathematics, 136 (1–2):219–226, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006130.

Mastronardi:2001:FAS

[MKS^{+}01] Nicola Mastronardi, Daniel Kressner, Vasile Sima, Paul Van Dooren, and Sabine Van Huffel. A fast algorithm for subspace state-space system identification via exploitation of the displacement structure. Journal of Computational and Applied Mathematics, 132(1):71–81, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006002.

Meng:2003:SNN

[ML03] Fan Wei Meng and Wei Nian Li. On some new nonlinear discrete inequalities and their applications. Journal of Computational and Applied Mathematics, 158 (2):407–417, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004758.

Mlotkowski:2001:CPM

[Mlo01]  Wojciech Mlotkowski.  A

class of probability measures. Journal of Computational and Applied Mathematics, 133(1–2):694, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000724X.

Mittelmann:2000:SEC

[MM00] Hans D. Mittelmann and Helmut Maurer. Solving elliptic control problems with interior point and SQP methods: control and state constraints. Journal of Computational and Applied Mathematics, 120(1–2):175–195, August 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003095.

Melnik:2002:OAO

[MM02a] K. N. Melnik and R. V. N.

Melnik. Optimal-by-accuracy and optimal-by-order cubature formulae in interpolational classes. Journal of Computational and Applied Mathematics, 147(1):233–262, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004351.

Mori:2002:P

[MM02b] Masatake Mori and Taketomo Mitsui. Preface. Journal of Computational and Applied Mathematics, 149(1):ix–x, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005162.

Muhammad:2003:DEF

[MM03] Mayinur Muhammad and Masatake Mori. Double exponential formulas for numerical indefinite integration. Journal of Computational and Applied Mathematics, 161(2):431–448, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007180.

Morhac:2002:OMN

[MMK02] M. Morhác, V. Matousek, and J. Kliman. Optimized multidimensional nonoscillating deconvolution. Journal of Computational and Applied Mathematics, 140(1– 2):639–658, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701005210.

MacMullen:2001:SOP

[MMOS01] H. MacMullen, J. J. H.

Miller, E. O’Riordan, and G. I. Shishkin. A secondorder parameter-uniform overlapping Schwarz method for reaction–diffusion problems with boundary layers. Journal of Computational and Applied Mathematics, 130(1– 2):231–244, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799003805.

Maazouz:2004:FIM

[MMPZ04] Abderrim Maazouz, Andro Mikelic, Jérôme Pousin, and Eric Zeltz. Fluid injection model without surface tension for resins in thin molds. Journal of Computational and Applied Mathematics, 164– 165(??):517–528, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004989.

Mulansky:2000:IAC

[MN00] Bernd Mulansky and Marian Neamtu. Interpolation and approximation from convex sets. II. Infinite-dimensional interpolation. Journal of Computational and Applied Mathematics, 119(1–2):333–346, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003861.

Moret:2001:IAM

[MN01] I. Moret and P. Novati. An interpolatory approximation of the matrix exponential based on Faber polynomials. Journal of Computational and Applied Mathematics, 131(1–2):361–380, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002612.

Mohamad:2002:DTA

[MN02a] S. Mohamad and A. G. Naim. Discrete-time analogues of integrodifferential equations modelling bidirectional neural networks. Journal of Computational and Applied Mathematics, 138(1):1–20, January 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003661.

Mukaihira:2002:SFO

[MN02b] Atsushi Mukaihira and Yoshimasa Nakamura. Schur flow for orthogonal polynomials on the unit circle and its integrable discretization. Journal of Computational and Applied Mathematics, 139(1):75–94, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003880.

Mastronardi:2000:NCS

[MO00a] N. Mastronardi and D. Occorsio. The numerical computation of some integrals on the real line. Journal of Computational and Applied Mathematics, 115(1–2):433–450, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900299X.

Micchelli:2000:PML

[MO00b] Charles A. Micchelli and Peder Olsen. Penalized maximumlikelihood estimation, the Baum–Welch algorithm, diagonal balancing of symmetric matrices and applications to training acoustic data. Journal of Computational and Applied Mathematics, 119(1–2):301–331, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S037704270000385X.

Mastroianni:2001:OSN

[MO01] G. Mastroianni and D. Occorsio. Optimal systems of nodes for Lagrange interpolation on bounded intervals. A survey. Journal of Computational and Applied Mathematics, 134 (1–2):325–341, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005574.

Mastroianni:2002:NAW

[MO02] G. Mastroianni and D. Occorsio. Numerical approximation of weakly singular integrals on the half line. Journal of Computational and Applied Mathematics, 140(1–2):587–598, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004113.

Moudafi:2003:CSI

[MO03] A. Moudafi and M. Oliny. Convergence of a splitting inertial proximal method for monotone operators. Journal of Computational and Applied Mathematics, 155(2):447–454, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042702009068.

Mastroianni:2004:MSG

[MO04] G. Mastroianni and D. Occorsio. Markov–Sonin Gaussian rule for singular functions. Journal of Computational and Applied Mathematics, 169(1):197–212, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703010252.

Mohanty:2000:FOF

[Moh00] R. K. Mohanty. A fourth-order finite difference method for the general one-dimensional nonlinear biharmonic problems of first kind. Journal of Computational and Applied Mathematics, 114 (2):275–290, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002022.

Monthe:2001:SSS

[Mon01] L. A. Monthé. A study of splitting scheme for hyperbolic conservation laws with source terms. Journal of Computational and Applied Mathematics, 137(1):1–12, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042700006154.

Moodie:2002:GC

[Moo02] T. B. Moodie. Gravity currents. Journal of Computational and Applied Mathematics, 144(1–2):49–83, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005519.

Morgan:2000:PES

[Mor00]  Ronald B. Morgan.  Pre-

conditioning eigenvalues and some comparison of solvers. Journal of Computational and Applied Mathematics, 123 (1–2):101–115, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003952.

Matthews:2002:NMS

[MOS02] S. Matthews, E. O’Riordan, and G. I. Shishkin. A numerical method for a system of singularly perturbed reaction–diffusion equations. Journal of Computational and Applied Mathematics, 145(1):151–166, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005416.

Matsuura:2003:MAB

[MOTT03] Yoshinori Matsuura, Shinnosuke Oharu, Takashi Takata, and Akio Tamura. Mathematical approaches to bone reformation phenomena and numerical simulations. Journal of Computational and Applied Mathematics, 158 (1):107–119, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004709.

Mourou:2003:TSA

[Mou03] Mohamed Ali Mourou. Taylor series associated with a differential-difference operator on the real line. Journal of Computational and Applied Mathematics, 153(1–2):343–354, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005939.

Milner:2001:ESH

[MP01] Fabio Augusto Milner and Curtis Allan Patton. Existence of solutions for a host-parasite model. Journal of Computational and Applied Mathematics, 137(2):331–361, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042700007093.

Mazzia:2002:MFE

[MP02a] A. Mazzia and M. Putti. Mixed-finite element and finite volume discretization for heavy brine simulations in groundwater. Journal of Computational and Applied Mathematics, 147(1):191–213, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004338.

Mohammadi:2002:AOS

[MP02b] B. Mohammadi and O. Pironneau. Applied optimal shape design. Journal of Computational and Applied Mathematics, 149(1):193–205, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005290.

Mainardi:2003:SPP

[MP03a] Francesco Mainardi and Gianni Pagnini. Salvatore pincherle: the pioneer of the Mellin–Barnes integrals. Journal of Computational and Applied Mathematics, 153(1–2):331–342, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S037704270200609X.

Milner:2003:DMH

[MP03b] Fabio Augusto Milner and Curtis Allan Patton. A diffusion model for host-parasite interaction. Journal of Computational and Applied Mathematics, 154(2):273–302, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008269.

Michiels:2001:SPV

[MR01] W. Michiels and D. Roose.

Sensitivity to perturbations in variable structure systems. Journal of Computational and Applied Mathematics, 132(1):127–140, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006014.

Maryska:2000:SCR

[MRT00] J. Maryska, M. Rozlozník, and M. Tuma. Schur complement reduction in the mixed-hybrid approximation of Darcy’s law: rounding error analysis. Journal of Computational and Applied Mathematics, 117(2):159–173, May 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042799003441.

Matos:2004:NIT

[MRV04] José Matos, Maria João Rodrigues, and Paulo B. Vasconcelos. New implementation of the Tau method for PDEs. Journal of Computational and Applied Mathematics, 164– 165(??):555–567, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008343.

Martinez:2000:CPA

[MRVM00] A. Martínez, C. Rodríguez, and M. E. Vázquez-Méndez. A control problem arising in the process of waste water purification. Journal of Computational and Applied Mathematics, 114(1):67–79, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002897.

Michalska:2001:NBL

[MS01a] M. Michalska and J. Szynal. A new bound for the Laguerre polynomials. Journal of Computational and Applied Mathematics, 133 (1–2):489–493, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042700006701.

Mori:2001:DET

[MS01b] Masatake Mori and Masaaki Sugihara. The doubleexponential transformation in numerical analysis. Journal of Computational and Applied Mathematics, 127 (1–2):287–296, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000501X.

Milovanovic:2002:QFC

[MS02] Gradimir V. Milovanović and Miodrag M. Spalević. Quadrature formulae connected to σ-orthogonal polynomials. Journal of Computational and Applied Mathematics, 140(1–2):619–637, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004769.

Mao:2003:NSS

[MS03a] Xuerong Mao and Sotirios Sabanis. Numerical solutions of stochastic differential delay equations under local Lipschitz condition. Journal of Computational and Applied Mathematics, 151(1):215–227, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007501.

Mikula:2003:SIF

[MS03b] Karol Mikula and Fiorella Sgallari. Semi-implicit finite volume scheme for image processing in 3D cylindrical geometry. Journal of Computational and Applied Mathematics, 161(1):119–132, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005491.

Milovanovic:2004:EAS

[MS04] Gradimir V. Milovanović and Miodrag M. Spalević. Error analysis in some Gauss– Turán–Radau and Gauss– Turán–Lobatto quadratures for analytic functions. Journal of Computational and Applied Mathematics, 164– 165(??):569–586, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004953.

Mahmoud:2004:RKF

[MSI04] Magdi S. Mahmoud, Peng Shi, and Abdulla Ismail. Robust Kalman filtering for discretetime Markovian jump systems with parameter uncertainty. Journal of Computational and Applied Mathematics, 169(1):53–69, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009439.

Miller:2004:GAS

[MSKS04] J. J. H. Miller, G. I. Shishkin, B. Koren, and L. P. Shishkina. Grid approximation of a singularly perturbed boundary value problem modelling heat transfer in the case of flow over a flat plate with suction of the boundary layer. Journal of Computational and Applied Mathematics, 166(1):221–232, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008598.

Mukherjee:2004:DSO

[MSM04] S. Mukherjee, Satakshi, and R. C. Mittal. Discrete system order reduction using multipoint step response matching. Journal of Computational and Applied Mathematics, 170 (2):461–466, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001517.

Mahmoud:2002:ISR

[MT02] Hosam Mahmoud and Tatsuie Tsukiji. On the internal structure of random recursive circuits. Journal of Computational and Applied Mathematics, 142(1):155–171, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004666.

Matsuoka:2003:NCP

[MT03] K. Matsuoka and K. Tomoeda. Numerical computations to pulse splitting phenomena in the thermal waves through an absorbing medium. Journal of Computational and Applied Mathematics, 152(1–2):319–332, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007148.

Meerschaert:2004:FDA

[MT04] Mark M. Meerschaert and Charles Tadjeran. Finite difference approximations for fractional advection-dispersion flow equations. Journal of Computational and Applied Mathematics, 172(1): 65–77, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704000986.

Marinova:2003:CPV

[MTA^{+}03] Rossitza S. Marinova, Tadayasu Takahashi, Hideaki Aiso, Christo I. Christov, and Tchavdar T. Marinov. Conservation properties of vectorial operator splitting. Journal of Computational and Applied Mathematics, 152(1–2):289–303, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007124.

Muhlbach:2000:ICV

[Müh00] G. Mühlbach. Interpolation by Cauchy–Vandermonde systems and applications. Journal of Computational and Applied Mathematics, 122 (1–2):203–222, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003642.

Muldoon:2001:OP

[Mul01] Martin Muldoon. Open problems. Journal of Computational and Applied Mathematics, 133(1–2):679, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042700006877.

Muldoon:2003:OP

[Mul03] Martin Muldoon. Open problems. Journal of Computational and Applied Mathematics, 153(1–2):531, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008816.

Mutafchiev:2002:SDS

[Mut02] Ljuben R. Mutafchiev. On the size of the Durfee square of a random integer partition. Journal of Computational and Applied Mathematics, 142(1):173–184, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004678.

Maes:2004:SNP

[MVDB04] Jan Maes, Evelyne Vanraes, Paul Dierckx, and Adhemar Bultheel. On the stability of normalized Powell– Sabin B-splines. Journal of Computational and Applied Mathematics, 170 (1):181–196, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000196.

Meek:2002:PHI

[MW02] D. S. Meek and D. J. Walton.

Planar G^{2} Hermite interpolation with some fair, C-shaped curves. Journal of Computational and Applied Mathematics, 139(1):141–161, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003855.

Meek:2004:ASA

[MW04a] D. S. Meek and D. J. Walton. An arc spline approximation to a clothoid. Journal of Computational and Applied Mathematics, 170(1):59–77, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000135.

Meek:2004:NFC

[MW04b] D. S. Meek and D. J. Walton.

A note on finding clothoids. Journal of Computational and Applied Mathematics, 170 (2):433–453, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000925.

Mishra:2004:CMP

[MWL04] S. K. Mishra, S. Y. Wang, and K. K. Lai. Complex minimax programming under generalized convexity. Journal of Computational and Applied Mathematics, 167(1):59–71, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008781.

Mishra:2003:NMF

[MWLS03] S. K. Mishra, S. Y. Wang, K. K. Lai, and J. M. Shi. Nondifferentiable minimax fractional programming under generalized univexity. Journal of Computational and Applied Mathematics, 158 (2):379–395, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004552.

Mehrmann:2000:NMC

[MX00] Volker Mehrmann and Hongguo Xu. Numerical methods in control. Journal of Computational and Applied Mathematics, 123 (1–2):371–394, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003927.

Matsunaga:2000:SSW

[MY00] Nami Matsunaga and Tetsuro Yamamoto. Superconvergence of the Shortley–Weller approximation for Dirichlet problems. Journal of Computational and Applied Mathematics, 116(2):263–273, April 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003210.

Mimachi:2003:RRS

[MY03a] Katsuhisa Mimachi and Masaaki Yoshida. The reciprocity relation of the Selberg function. Journal of Computational and Applied Mathematics, 160 (1–2):209–215, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300623X.

Minchev:2003:OSV

[MY03b] Emil Minchev and Norio Yoshida. Oscillations of solutions of vector differential equations of parabolic type with functional arguments. Journal of Computational and Applied Mathematics, 151(1):107–117, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042702007409.

Marletta:2002:CCE

[MZ02] Marco Marletta and Anton Zettl. Counting and computing eigenvalues of leftdefinite Sturm–Liouville problems. Journal of Computational and Applied Mathematics, 148(1):65–75, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005733.

Nachaoui:2004:NLA

[Nac04]  Abdeljalil Nachaoui.  Nu-

merical linear algebra for reconstruction inverse problems. Journal of Computational and Applied Mathematics, 162(1):147–164, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007349.

Nagatou:2004:CAP

[Nag04] K. Nagatou. A computerassisted proof on the stability of the Kolmogorov flows of incompressible viscous fluid. Journal of Computational and Applied Mathematics, 169(1):33–44, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042703009361.

Nakamura:2001:AAA

[Nak01] Yoshimasa Nakamura. Algorithms associated with arithmetic, geometric and harmonic means and integrable systems. Journal of Computational and Applied Mathematics, 131(1–2):161–174, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003162.

Nash:2000:STN

[Nas00] Stephen G. Nash. A survey of truncated-Newton methods. Journal of Computational and Applied Mathematics, 124(1–2):45–59, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000426X.

Naylor:2002:AEK

[Nay02] D. Naylor. An asymptotic expansion of the Kontorovich– Lebedev transform of damped oscillatory functions. Journal of Computational and Applied Mathematics, 145 (1):21–30, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701005337.

Naire:2004:DIS

[NBS04] S. Naire, R. J. Braun, and S. A. Snow. A 2 + 1 dimensional insoluble surfactant model for a vertical draining free film. Journal of Computational and Applied Mathematics, 166(2):385–410, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007660.

Neher:2003:IVB

[Neh03] M. Neher. Improved validated bounds for Taylor coefficients and for Taylor remainder series. Journal of Computational and Applied Mathematics, 152(1–2):393–404, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007197.

Neininger:2002:BST

[Nei02]  Ralph Neininger.  On bi-

nary search tree recursions with monomials as toll functions. Journal of Computational and Applied Mathematics, 142(1):185–196, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270100468X.

Neumaier:2003:ECZ

[Neu03] Arnold Neumaier. Enclosing clusters of zeros of polynomials. Journal of Computational and Applied Mathematics, 156(2):389–401, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003807.

Ng:2003:CSC

[Ng03] Michael K. Ng. Circulant and skew-circulant splitting methods for Toeplitz systems. Journal of Computational and Applied Mathematics, 159(1):101–108, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005624.

Nicolet:2004:MEW

[NGGZ04] André Nicolet, Sébastien Guenneau, Christophe Geuzaine, and Frédéric Zolla. Modelling of electromagnetic waves in periodic media with finite elements. Journal of Computational and Applied Mathematics, 168(1–2):321–329, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009956.

Niki:2004:SPU

[NHMS04] Hiroshi Niki, Kyouji Harada, Munenori Morimoto, and Michio Sakakihara. The survey of preconditioners used for accelerating the rate of convergence in the Gauss–Seidel method. Journal of Computational and Applied Mathematics, 164– 165(??):587–600, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009531.

Nievergelt:2000:THL

[Nie00] Yves Nievergelt. A tutorial history of least squares with applications to astronomy and geodesy. Journal of Computational and Applied Mathematics, 121 (1–2):37–72, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003435.

Niederreiter:2003:EBQ

[Nie03] Harald Niederreiter. Error bounds for quasi-Monte Carlo integration with uniform point sets. Journal of Computational and Applied Mathematics, 150(2):283–292, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042702006659.

Nishida:2003:NMO

[Nis03a]  Kotoba Nishida.  Numeri-

cal method for Oseen’s linearized equations in threedimensional exterior domains. Journal of Computational and Applied Mathematics, 152(1–2):405–409, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007707.

Nishizawa:2003:SHS

[Nis03b] Michitomo Nishizawa. U_{r,s}(}↕_{△})- symmetry for (r, s)-hypergeometric series. Journal of Computational and Applied Mathematics, 160(1–2):233–239, November 1, 2003. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042703006253.

Natesan:2003:PUN

[NJVA03] S. Natesan, J. Jayakumar, and J. Vigo-Aguiar. Parameter uniform numerical method for singularly perturbed turning point problems exhibiting boundary layers. Journal of Computational and Applied Mathematics, 158 (1):121–134, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300476X.

Nasir:2003:MTF

[NKK03] H. M. Nasir, T. Kako, and D. Koyama. A mixedtype finite element approximation for radiation problems using fictitious domain method. Journal of Computational and Applied Mathematics, 152(1–2):377–392, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007185.

Nakata:2002:PAG

[NMST02] Kazuto Nakata, Tadashi Murase, Toshihiro Sakuma, and Toshikazu Takada. A parallel algorithm for generating molecular integrals over MO basis sets. Journal of Computational and Applied Mathematics, 149(1):351–357, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005411.

Noubary:2004:STS

[NN04] Farzad Noubary and Reza Noubary. On survival times of sport records. Journal of Computational and Applied Mathematics, 169(1): 227–234, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703010276.

Noguchi:2002:ATR

[NOI02] Hiroshi Noguchi, Masaru Ogawa, and Hiroaki Ishii. The appropriate total ranking method using DEA for multiple categorized purposes. Journal of Computational and Applied Mathematics, 146 (1):155–166, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004259.

Noor:2001:MRS

[Noo01] Muhammad Aslam Noor.

Modified resolvent splitting algorithms for general mixed variational inequalities. Journal of Computational and Applied Mathematics, 135(1): 111–124, October 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005690.

Notaris:2001:IQF

[Not01] Sotirios E. Notaris. Interpolatory quadrature formulae with Chebyshev abscissae. Journal of Computational and Applied Mathematics, 133(1–2):507–517, August 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006725.

Notaris:2003:NIQ

[Not03] Sotirios E. Notaris. New interpolatory quadrature formulae with Gegenbauer abscissae. Journal of Computational and Applied Mathematics, 161(2):295–312, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007027.

Natalini:2003:SMR

[NP03] Pierpaolo Natalini and Biagio Palumbo. Some monotonicity results on the zeros of the generalized Laguerre polynomials. Journal of Computational and Applied Mathematics, 153(1–2):355–360, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005885.

Natalini:2001:CNS

[NR01] Pierpaolo Natalini and Paolo Emilio Ricci. Computation of Newton sum rules for associated and co-recursive classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 133(1–2):495–505, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006713.

Nieto:2003:RPB

[NRL03] Juan J. Nieto and Rosana Rodríguez-López. Remarks on periodic boundary value problems for functional differential equations. Journal of Computational and Applied Mathematics, 158 (2):339–353, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004527.

Nakane:2003:GSO

[NS03] Kazuaki Nakane and Tomoko Shinohara. Global solutions to a one-dimensional hyperbolic free boundary problem which arises in peeling phenomena. Journal of Computational and Applied Mathematics, 152(1–2):367–375, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007173.

Noskov:2004:NND

[NS04a] M. V. Noskov and H. J.

Schmid. On the number of nodes in n-dimensional cubature formulae of degree 5 for integrals over the ball. Journal of Computational and Applied Mathematics, 169(2):247–254, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270301029X.

Nouisser:2004:INB

[NS04b] O. Nouisser and D. Sbibih. Interpolation by new Bsplines on a four directional mesh of the plane. Journal of Computational and Applied Mathematics, 162(1):287–298, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007192.

Natori:2003:P

[NSK03] Makoto Natori, Zhong-Ci Shi, and Takashi Kitagawa. Preface. Journal of Computational and Applied Mathematics, 159(1):ix–xi, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005740.

Noels:2004:CIE

[NSP04] L. Noels, L. Stainier, and J.- P. Ponthot. Combined implicit/explicit time-integration algorithms for the numerical simulation of sheet metal forming. Journal of Computational and Applied Mathematics, 168(1–2):331–339, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009968.

Najib:2004:PEL

[NSZ04] K. Najib, D. Sandri, and A.-M.

Zine. On a posteriori estimates for a linearized Oldroyd’s problem. Journal of Computational and Applied Mathematics, 167(2):345–361, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009257.

Norbury:2000:DCD

[NW00] John Norbury and R. Eddie Wilson. Dynamics of constrained differential delay equations. Journal of Computational and Applied Mathematics, 125(1– 2):201–215, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004696.

Nakaguchi:2003:NAS

[NY03] Etsushi Nakaguchi and Atsushi Yagi. Numerical analysis for semilinear evolution equations of parabolic type. Journal of Computational and Applied Mathematics, 159(1):91–99, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005673.

Nurnberger:2000:DBS

[NZ00a] G. Nürnberger and F. Zeilfelder. Developments in bivariate spline interpolation. Journal of Computational and Applied Mathematics, 121 (1–2):125–152, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003460.

Nurnberger:2000:ISS

[NZ00b] G. Nürnberger and F. Zeilfelder. Interpolation by spline spaces on classes of triangulations. Journal of Computational and Applied Mathematics, 119(1–2):347–376, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003873.

Oruc:2004:SFV

[OA04] Halil Oruç and Hakan K. Akmaz. Symmetric functions and the Vandermonde matrix. Journal of Computational and Applied Mathematics, 172(1):49–64, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000974.

Ogata:2003:FSM

[OASO03] Hidenori Ogata, Kaname Amano, Masaaki Sugihara, and Dai Okano. A fundamental solution method for viscous flow problems with obstacles in a periodic array. Journal of Computational and Applied Mathematics, 152(1–2):411–425, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007203.

Ohtsuka:2002:CCD

[Oht02]  Kohji Ohtsuka.  Compari-

son of criteria on the direction of crack extension. Journal of Computational and Applied Mathematics, 149 (1):335–339, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005393.

Ohtsubo:2003:VIM

[Oht03]  Yoshio Ohtsubo.  Value

iteration methods in risk minimizing stopping problems. Journal of Computational and Applied Mathematics, 152(1–2):427–439, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007215.

OLeary:2000:SBL

[O’L00] Dianne P. O’Leary. Symbiosis between linear algebra and optimization. Journal of Computational and Applied Mathematics, 123 (1–2):447–465, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004088.

Ojiro:2003:NCM

[ONU03] Keiko Ojiro, Hiroshi Niki, and Masataka Usui. A new criterion for the H-matrix property. Journal of Computational and Applied Mathematics, 150(2):293–302, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006660.

Ogreid:2001:MSR

[OO01] Odd Magne Ogreid and Per Osland. More series related to the Euler series. Journal of Computational and Applied Mathematics, 136 (1–2):389–403, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006300.

Ogreid:2002:SIS

[OO02] Odd Magne Ogreid and Per Osland. Some infinite series related to Feynman diagrams. Journal of Computational and Applied Mathematics, 140(1–2):659–671, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004873.

Okano:2003:MNC

[OOA03] Dai Okano, Hidenori Ogata, and Kaname Amano. A method of numerical conformal mapping of curved slit domains by the charge simulation method. Journal of Computational and Applied Mathematics, 152(1–2):441–450, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007227.

Okano:2003:NCM

[OOAS03] Dai Okano, Hidenori Ogata, Kaname Amano, and Masaaki Sugihara. Numerical conformal mappings of bounded multiply connected domains by the charge simulation method. Journal of Computational and Applied Mathematics, 159(1):109–117, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005727.

Ooura:2001:CET

[Oou01] Takuya Ooura. A continuous Euler transformation and its application to the Fourier transform of a slowly decaying function. Journal of Computational and Applied Mathematics, 130(1–2):259–270, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003787.

Ooura:2003:GCE

[Oou03] Takuya Ooura. A generalization of the continuous Euler transformation and its application to numerical quadrature. Journal of Computational and Applied Mathematics, 157(2):251–259, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703003789.

Oruc:2003:BPB

[OP03] Halil Oruç and George M. Phillips. q-Bernstein polynomials and Bézier curves. Journal of Computational and Applied Mathematics, 151(1):1–12, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007331.

ORiordan:2004:SPP

[OS04] E. O’Riordan and G. I.

Shishkin. Singularly perturbed parabolic problems with nonsmooth data. Journal of Computational and Applied Mathematics, 166(1):233–245, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008586.

Osada:2000:AFS

[Osa00] Naoki Osada. The E-algorithm and the Ford–Sidi algorithm. Journal of Computational and Applied Mathematics, 122 (1–2):223–230, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003666.

Oyanagi:2002:FS

[Oya02]  Yoshio Oyanagi.  Future

of supercomputing. Journal of Computational and Applied Mathematics, 149 (1):147–153, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005265.

Ozer:2004:SGP

[Öze04] Teoman Özer. On symmetry group properties and general similarity forms of the Benney equations in the Lagrangian variables. Journal of Computational and Applied Mathematics, 169(2):297–313, August 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270301032X.

Palacios:2002:KEA

[Pal02a] M. Palacios. Kepler equation and accelerated Newton method. Journal of Computational and Applied Mathematics, 138(2):335–346, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003697.

Palamides:2002:MPS

[Pal02b] P. K. Palamides. Monotone positive solutions for singular boundary value problems. Journal of Computational and Applied Mathematics, 146 (2):405–418, September 15, 2002. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003916.

Panario:2002:ASS

[Pan02]  Daniel Panario.  Asymp-

totics of subtracted singularities for generating functions with small singularities. Journal of Computational and Applied Mathematics, 142(1):197–209, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004708.

Pao:2001:NSR

[Pao01] C. V. Pao. Numerical solutions of reaction–diffusion equations with nonlocal boundary conditions. Journal of Computational and Applied Mathematics, 136(1– 2):227–243, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006142.

Park:2000:AVN

[Par00a]  Sehie Park.  Acyclic ver-

sions of the von Neumann and Nash equilibrium theorems. Journal of Computational and Applied Mathematics, 113(1–2):83–91, January 1, 2000. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002460.

Parlett:2000:TR

[Par00b] Beresford N. Parlett. For tridiagonals T replace T with LDL^{t}. Journal of Computational and Applied Mathematics, 123(1–2):117– 130, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700003940.

Numerical analysis 2000, Vol.

III. Linear algebra.

Park:2001:NDD

[Par01] Ju-Hyun Park. A new delaydependent criterion for neutral systems with multiple delays. Journal of Computational and Applied Mathematics, 136 (1–2):177–184, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005835.

Paris:2002:EBU

[Par02a] R. B. Paris. Error bounds for the uniform asymptotic expansion of the incomplete gamma function. Journal of Computational and Applied Mathematics, 147(1):215–231, October 1, 2002. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200434X.

Paris:2002:UAE

[Par02b]  R. B. Paris.  A uniform

asymptotic expansion for the incomplete gamma function. Journal of Computational and Applied Mathematics, 148 (2):323–339, November 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005538.

Paris:2003:AEG

[Par03a] R. B. Paris. The asymptotic expansion of a generalised incomplete gamma function. Journal of Computational and Applied Mathematics, 151(2):297–306, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008099.

Paris:2003:ACB

[Par03b]  R. B. Paris.  On the

asymptotic connection between two exponential sums. Journal of Computational and Applied Mathematics, 157(2):297–308, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004126.

Park:2003:GCS

[Par03c] Ju H. Park. Guaranteed cost stabilization of neutral differential systems with parametric uncertainty. Journal of Computational and Applied Mathematics, 151 (2):371–382, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007513.

Paris:2004:UHEa

[Par04a] R. B. Paris. On the use of Hadamard expansions in hyperasymptotic evaluation of Laplace-type integrals. I: real variable. Journal of Computational and Applied Mathematics, 167(2):293–319, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009245.

Paris:2004:UHEb

[Par04b]  R. B. Paris.  On the

use of Hadamard expansions in hyperasymptotic evaluation of Laplace-type integrals: II. Complex variable. Journal of Computational and Applied Mathematics, 167(2):321–343, June 1, 2004. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009415.

Pasquetti:2004:IMR

[Pas04] R. Pasquetti. On inverse methods for the resolution of the Gibbs phenomenon. Journal of Computational and Applied Mathematics, 170 (2):303–315, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270400072X.

Paul:2000:DES

[Pau00] Christopher A. H. Paul.

Designing efficient software for solving delay differential equations. Journal of Computational and Applied Mathematics, 125(1– 2):287–295, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000474X.

Peiguang:2000:OPE

[PC00] Wang Peiguang and Feng Chunhua. Oscillation of parabolic equations of neutral type. Journal of Computational and Applied Mathematics, 126(1–2):111– 120, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042799003453.

Pantale:2004:DOO

[PCR04] O. Pantalé, S. Caperaa, and R. Rakotomalala. Development of an objectoriented finite element program: application to metalforming and impact simulations. Journal of Computational and Applied Mathematics, 168(1–2):341–351, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009981.

Pathak:2003:CDB

[PD03] R. S. Pathak and M. M. Dixit. Continuous and discrete Bessel wavelet transforms. Journal of Computational and Applied Mathematics, 160 (1–2):241–250, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006265.

Pluymers:2004:AEW

[PDVS04] B. Pluymers, W. Desmet, D. Vandepitte, and P. Sas. Application of an efficient wave-based prediction technique for the analysis of vibro-acoustic radiation problems. Journal of Computational and Applied Mathematics, 168(1–2):353–364, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703010008.

Pedersen:2003:DGF

[Ped03] Henrik L. Pedersen. The double gamma function and related Pick functions. Journal of Computational and Applied Mathematics, 153(1– 2):361–369, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702006040.

Peherstorfer:2001:TLO

[Peh01] Franz Peherstorfer. On Toda lattices and orthogonal polynomials. Journal of Computational and Applied Mathematics, 133(1–2):519–534, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006737.

Peherstorfer:2003:IIP

[Peh03] Franz Peherstorfer. Inverse images of polynomial mappings and polynomials orthogonal on them. Journal of Computational and Applied Mathematics, 153(1–2):371–385, April 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006283.

Pena:2000:FBS

[Peñ00a] J. M. Peña. On the foundation of bases of spline spaces. Journal of Computational and Applied Mathematics, 119(1–2):377–390, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003885.

Penski:2000:NNM

[Pen00b] Christian Penski. A new numerical method for SDEs and its application in circuit simulation. Journal of Computational and Applied Mathematics, 115(1–2):461–470, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003040.

Petersen:2001:MMU

[Pet01]  Vigdis Petersen.  Modifi-

cation of a method using Szegő polynomials in frequency analysis: the V - process. Journal of Computational and Applied Mathematics, 133(1–2):535–544, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006749.

Petersen:2002:MZF

[Pet02a] Vigdis Brevik Petersen. Multiple zeros in frequency analysis: the T (r)-process. Journal of Computational and Applied Mathematics, 147 (2):427–444, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004776.

Petkovic:2002:CSR

[Pet02b] Miodrag S. Petković. Comments on some recent methods for the simultaneous determination of polynomial zeros. Journal of Computational and Applied Mathematics, 145(2):519–524, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004326.

Petras:2002:SVI

[Pet02c] Knut Petras. Self-validating integration and approximation of piecewise analytic functions. Journal of Computational and Applied Mathematics, 145(2):345–359, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042701005866.

Petkovic:2003:LLI

[Pet03] Miodrag S. Petković. Laguerrelike inclusion method for polynomial zeros. Journal of Computational and Applied Mathematics, 152(1– 2):451–465, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702007239.

Petrova:2004:CFS

[Pet04] Guergana Petrova. Cubature formulae for spheres, simplices and balls. Journal of Computational and Applied Mathematics, 162(2):483–496, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007611.

Patricio:2003:ILP

[PFO03] F. Patrício, J. A. Ferreira, and F. Oliveira. On the interval Legendre polynomials. Journal of Computational and Applied Mathematics, 154(1):215–227, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008233.

Pan:2002:DSM

[PGG02] Tsorng-Whay Pan, Roland Glowinski, and Giovanni P. Galdi. Direct simulation of the motion of a settling ellipsoid in Newtonian fluid. Journal of Computational and Applied Mathematics, 149(1):71–82, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005216. See ad-

dendum [PGG03].

Pan:2003:ADS

[PGG03] Tsorng-Whay Pan, Roland Glowinski, and Giovanni P. Galdi. Addendum to: “Direct simulation of the motion of a settling ellipsoid in Newtonian fluid”: [J. Comput. Appl. Math. 149 (2002) 71–82]. Journal of Computational and Applied Mathematics, 156(2):499, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003947.  See

[PGG02].

Petkovic:2001:PES

[PH01] Miodrag S. Petković and Dorde Herceg. Point estimation of simultaneous methods for solving polynomial equations: a survey. Journal of Computational and Applied Mathematics, 136 (1–2):283–307, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006208.

Piessens:2000:CIT

[Pie00]  Robert Piessens.  Com-

puting integral transforms and solving integral equations using Chebyshev polynomial approximations. Journal of Computational and Applied Mathematics, 121 (1–2):113–124, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003496.

Park:2003:ESC

[PJ03] Ju H. Park and Ho Y. Jung. On the exponential stability of a class of nonlinear systems including delayed perturbations. Journal of Computational and Applied Mathematics, 159(2):467–471, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005508.

Pathan:2003:MMP

[PKA03] M. A. Pathan, M. Kamarujjama, and M. Khursheed Alam. On multiindices and multivariables presentation of the Voigt functions. Journal of Computational and Applied Mathematics, 160 (1–2):251–257, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006277.

Poole:2000:RPS

[PN00] George Poole and Larry Neal.

The Rook’s pivoting strategy. Journal of Computational and Applied Mathematics, 123 (1–2):353–369, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004064.

Plagianakos:2001:LCP

[PNV01] V. P. Plagianakos, N. K.

Nousis, and M. N. Vrahatis. Locating and computing in parallel all the simple roots of special functions using PVM. Journal of Computational and Applied Mathematics, 133(1–2):545–554, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006750.

Pommeret:2001:TRR

[Pom01] Denys Pommeret. K terms recurrence relations and polynomial variance functions of the K th degree. Journal of Computational and Applied Mathematics, 133 (1–2):555–565, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006762.

Potthast:2000:SER

[Pot00]  Roland Potthast.  Stabil-

ity estimates and reconstructions in inverse acoustic scattering using singular sources. Journal of Computational and Applied Mathematics, 114(2):247–274, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002010.

Potzsche:2002:CRI

[Pöt02] Christian Pötzsche. Chain rule and invariance principle on measure chains. Journal of Computational and Applied Mathematics, 141(1–2):249–254, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004502.

Pardo:2000:URD

[PP00a] M. C. Pardo and J. A. Pardo. Use of Rényi’s divergence to test for the equality of the coefficients of variation. Journal of Computational and Applied Mathematics, 116(1):93–104, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900312X.

Parks:2000:EEA

[PP00b] Harold R. Parks and Jon T.

Pitts. Energy estimates for area minimizing hypersurfaces with arbitrary boundaries. Journal of Computational and Applied Mathematics, 115(1–2):451–460, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001831.

Plemmons:2000:SCP

[PP00c] Robert J. Plemmons and Victor P. Pauca. Some computational problems arising in adaptive optics imaging systems. Journal of Computational and Applied Mathematics, 123(1–2):467– 487, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700004106.

Panholzer:2002:BST

[PP02a] Alois Panholzer and Helmut Prodinger. Binary search tree recursions with harmonic toll functions. Journal of Computational and Applied Mathematics, 142(1):211–225, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004691.

Perez:2002:AND

[PP02b] T. Pérez and J. A. Pardo. Asymptotic normality for the K_{ϕ}-divergence goodness-of-fit tests. Journal of Computational and Applied Mathematics, 145(2):301–317, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005830.

Piazza:2002:UBC

[PP02c] G. Piazza and T. Politi. An upper bound for the condition number of a matrix in spectral norm. Journal of Computational and Applied Mathematics, 143(1):141–144, June 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702003965.

Prastaro:2000:GAG

[PR00] Agostino Prástaro and Themistocles M. Rassias. A geometric approach of the generalized d’Alembert equation. Journal of Computational and Applied Mathematics, 113(1–2):93–122, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002472.

Plaza:2002:ATM

[PR02] Angel Plaza and María-Cecilia Rivara. On the adjacencies of triangular meshes based on skeleton-regular partitions. Journal of Computational and Applied Mathematics, 140(1–2):673–693, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004848.

Petkovic:2004:GCS

[PR04] M. S. Petković and L. Rancić.

On the guaranteed convergence of the square-root iteration method. Journal of Computational and Applied Mathematics, 170 (1):169–179, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000184.

Precup:2000:DCM

[Pre00a] Radu Precup. Discrete continuation method for boundary value problems on bounded sets in Banach spaces. Journal of Computational and Applied Mathematics, 113 (1–2):267–281, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002617.

Prevost:2000:DAU

[Pré00b]  M. Prévost.  Diophan-

tine approximations using Padé approximations. Journal of Computational and Applied Mathematics, 122 (1–2):231–250, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003654.

Pop:2004:MFE

[PRK04] I. S. Pop, F. Radu, and P. Knabner. Mixed finite elements for the Richards’ equation: linearization procedure. Journal of Computational and Applied Mathematics, 168(1–2):365–373, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270301001X.

Pardalos:2000:RDT

[PRT00] Panos M. Pardalos, H. Edwin Romeijn, and Hoang Tuy. Recent developments and trends in global optimization. Journal of Computational and Applied Mathematics, 124 (1–2):209–228, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004258.

Peherstorfer:2001:OEP

[PS01] Franz Peherstorfer and Robert Steinbauer. Orthogonal and L_{q}-extremal polynomials on inverse images of polynomial mappings. Journal of Computational and Applied Mathematics, 127 (1–2):297–315, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005021.

Pennacchio:2002:EAS

[PS02] Micol Pennacchio and Valeria Simoncini. Efficient algebraic solution of reaction–diffusion systems for the cardiac excitation process. Journal of Computational and Applied Mathematics, 145(1):49–70, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005350.

Pelekanos:2003:ISP

[PS03a] George Pelekanos and Vassilios Sevroglou. Inverse scattering by penetrable objects in two-dimensional elastodynamics. Journal of Computational and Applied Mathematics, 151(1):129–140, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007422.

Psihoyios:2003:TFP

[PS03b] G. Psihoyios and T. E. Simos. Trigonometrically fitted predictor–corrector methods for IVPs with oscillating solutions. Journal of Computational and Applied Mathematics, 158 (1):135–144, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004813.

Pandey:2004:CFD

[PS04] R. K. Pandey and Arvind K. Singh. On the convergence of a finite difference method for a class of singular boundary value problems arising in physiology. Journal of Computational and Applied Mathematics, 166(2):553–564, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008732.

Psarrakos:2003:DTH

[Psa03] Panayiotis J. Psarrakos. Definite triples of Hermitian matrices and matrix polynomials. Journal of Computational and Applied Mathematics, 151(1):39–58, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007367.

Petropoulou:2003:CZB

[PSS03] Eugenia N. Petropoulou, Panayiotis D. Siafarikas, and Ioannis D. Stabolas. On the common zeros of Bessel functions. Journal of Computational and Applied Mathematics, 153(1–2):387–393, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006416.

Pu:2002:GCI

[PT02] Dingguo Pu and Weiwen Tian. Globally convergent inexact generalized Newton’s methods for nonsmooth equations. Journal of Computational and Applied Mathematics, 138(1):37–49, January 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003648.

Pu:2003:RDA

[PT03] Dingguo Pu and Weiwen Tian. The revised DFP algorithm without exact line search. Journal of Computational and Applied Mathematics, 154(2):319–339, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008567.

Potra:2000:IPM

[PW00] Florian A. Potra and Stephen J. Wright. Interior-point methods. Journal of Computational and Applied Mathematics, 124 (1–2):281–302, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004337.

Petersen:2002:CTD

[PW02] Vigdis Petersen and Haakon Waadeland. A case of Toeplitz determinants and theta functions in frequency analysis. Journal of Computational and Applied Mathematics, 139(2):239–251, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042701004228.

Qian:2004:RSR

[QH04] Xiaoyuan Qian and Xuegang Huang. Reconstruction of surfaces of revolution with partial sampling. Journal of Computational and Applied Mathematics, 163(1):211–217, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008070.

Jin:2004:PWL

[qJkSlS04] Xiao qing Jin, Vai kuong Sin, and Li li Song. Preconditioned WR–LMF-based method for ODE systems. Journal of Computational and Applied Mathematics, 162(2):431–444, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007581.

Qiu:2001:SRI

[QM01] Lin Qiu and Taketomo Mitsui. Stability of the Radau IA and Lobatto IIIC methods for neutral delay differential system. Journal of Computational and Applied Mathematics, 137(2):279–292, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700007123.

Qiu:2000:NSS

[QST00] Y. Qiu, D. M. Sloan, and T. Tang. Numerical solution of a singularly perturbed two-point boundary value problem using equidistribution: analysis of convergence. Journal of Computational and Applied Mathematics, 116(1):121–143, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003155.

Qi:2002:PDA

[QSZ02] Liqun Qi, Defeng Sun, and Guanglu Zhou. A primaldual algorithm for minimizing a sum of Euclidean norms. Journal of Computational and Applied Mathematics, 138(1):127–150, January 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003570.

Quintanilla:2002:ESO

[Qui02] R. Quintanilla. Exponential stability for one-dimensional problem of swelling porous elastic soils with fluid saturation. Journal of Computational and Applied Mathematics, 145(2):525–533, August 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004429.

Qi:2002:P

[QYZ02] Liqun Qi, Xiaoqi Yang, and Jian-Zhong Zhang. Preface. Journal of Computational and Applied Mathematics, 146(1):ix, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004120.

Qi:2001:P

[QZ01] Liqun Qi and Jianzhong Zhang. Preface. Journal of Computational and Applied Mathematics, 129 (1–2):ix, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005380.

See erratum [QZ03].

Qi:2003:EPS

[QZ03] Liqun Qi and Jianzhong Zhang. Erratum to the preface of the special issue on Nonlinear Programming and Variational Inequalities [J. Comput. Appl. Math. 129 (2001) ix]. Journal of Computational and Applied Mathematics, 150(1):217, January 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006556.  See

[QZ01].

Rababah:2003:DDR

[Rab03] Abedallah Rababah. Distance for degree raising and reduction of triangular Bézier surfaces. Journal of Computational and Applied Mathematics, 158(2):233–241, September 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042703004321.

Rachunkova:2000:ETP

[Rac00]  Irena Rachunková.  Exis-

tence of two positive solutions of a singular nonlinear periodic boundary value problem. Journal of Computational and Applied Mathematics, 113(1–2):27–34, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002411.

Radoux:2000:AFP

[Rad00] Christian Radoux. Addition formulas for polynomials built on classical combinatorial sequences. Journal of Computational and Applied Mathematics, 115(1–2):471–477, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900120X.

Ramanathan:2003:SPE

[Ram03]  R. Ramanathan.  Some

properties of an extended Wigner function. Journal of Computational and Applied Mathematics, 160 (1–2):259–263, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006289.

Rannacher:2001:AGF

[Ran01]  R. Rannacher.  Adaptive

Galerkin finite element methods for partial differential equations. Journal of Computational and Applied Mathematics, 128(1–2):205–233, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042700005136.

Randrianarivony:2004:AFE

[Ran04] Maharavo Randrianarivony.

Anisotropic finite elements for the Stokes problem: a posteriori error estimator and adaptive mesh. Journal of Computational and Applied Mathematics, 169(2): 255–275, August 15, 2004.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703010306.

Rathsfeld:2000:QMP

[Rat00]  A. Rathsfeld.  Quadra-

ture methods for 2D and 3D problems. Journal of Computational and Applied Mathematics, 125(1– 2):439–460, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004842.

Rivero:2004:GDS

[RCZV04] Carlos Rivero, Ángela Castillo, Pedro J. Zufiria, and Teófilo Valdés. Global dynamics of a system governing an algorithm for regression with censored and noncensored data under general errors. Journal of Computational and Applied Mathematics, 166(2):535–551, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008720.

Renouf:2004:PVN

[RDA04] Mathieu Renouf, Frédéric Dubois, and Pierre Alart. A parallel version of the non smooth contact dynamics algorithm applied to the simulation of granular media. Journal of Computational and Applied Mathematics, 168(1–2):375–382, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703010021.

Ruas:2004:MFF

[RdAR04] V. Ruas, J. H. Carneiro de Araujo, and M. A. M. Silva Ramos. Multifield finite element methods with discontinuous pressures for axisymmetric incompressible flow. Journal of Computational and Applied Mathematics, 168(1–2):393–402, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009828.

Remiche:2000:EDI

[Rem00]  Marie-Ange Remiche.  On

the exact distribution of the isotropic planar point processes of phase type. Journal of Computational and Applied Mathematics, 116 (1):77–91, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799002824.

Remesikova:2004:SCD

[Rem04] M. Remesíková. Solution of convection–diffusion problems with nonequilibrium adsorption. Journal of Computational and Applied Mathematics, 169(1):101–116, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009464.

Rheinboldt:2000:NCM

[Rhe00] Werner C. Rheinboldt. Numerical continuation methods: a perspective. Journal of Computational and Applied Mathematics, 124 (1–2):229–244, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004283.

Rhee:2002:DRN

[Rhe02] WanSoo T. Rhee. Disjoint random needles. Journal of Computational and Applied Mathematics, 142(1):227–234, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270100471X.

Ricceri:2000:GVP

[Ric00] Biagio Ricceri. A general variational principle and some of its applications. Journal of Computational and Applied Mathematics, 113 (1–2):401–410, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002691.

Rihan:2003:SAD

[Rih03] Fathalla A. Rihan. Sensitivity analysis for dynamic systems with time-lags. Journal of Computational and Applied Mathematics, 151 (2):445–462, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006593.

Rivoal:2003:SIL

[Riv03]  Tanguy Rivoal.  Selberg’s

integral and linear forms in zeta values. Journal of Computational and Applied Mathematics, 160(1– 2):265–270, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006307.

Rixen:2004:DCB

[Rix04] Daniel J. Rixen. A dual Craig–Bampton method for dynamic substructuring. Journal of Computational and Applied Mathematics, 168(1– 2):383–391, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703010045.

Rathie:2004:AMN

[RKS04] Arjun K. Rathie, Yong Sup Kim, and Hyeong Kee Song. Another method for a new twoterm relation for the hypergeometric function _{3}F_{2} due to Exton. Journal of Computational and Applied Mathematics, 167(2):485–487, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009385.

Rolania:2001:AOP

[RLS01] D. Barrios Rolanía, G. López Lagomasino, and E. B. Saff. Asymptotics of orthogonal polynomials inside the unit circle and Szegő–Padé approximants. Journal of Computational and Applied Mathematics, 133(1–2):171–181, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006786.

Ruffing:2003:DLO

[RLZ03] Andreas Ruffing, Julian Lorenz, and Konstantin Ziegler. Difference ladder operators for a harmonic Schrödinger oscillator using unitary linear lattices. Journal of Computational and Applied Mathematics, 153(1–2):395–410, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006131.

Ryoo:2003:NVS

[RN03] Cheon Seoung Ryoo and Mitsuhiro T. Nakao. Numerical verification of solutions for obstacle problems. Journal of Computational and Applied Mathematics, 161 (2):405–416, December 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007131.

Reverter:2003:CRD

[RO03] F. Reverter and J. M. Oller.

Computing the Rao distance for gamma distributions. Journal of Computational and Applied Mathematics, 157(1):155–167, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300387X.

Ronveaux:2001:ARC

[Ron01] André Ronveaux. Asymptotics for recurrence coefficients in the generalized Meixner case. Journal of Computational and Applied Mathematics, 133(1–2):695–696, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700007251.

Rossler:2004:RKM

[Röß04] Andreas Rößler. Runge–Kutta methods for Stratonovich stochastic differential equation systems with commutative noise. Journal of Computational and Applied Mathematics, 164–165(??): 613–627, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270300829X.

Rebillard:2001:EMP

[RR01] L. Rebillard and A. Ronveaux. Expansion of multivariate polynomials in products of univariate orthogonal polynomials: discrete case. Journal of Computational and Applied Mathematics, 133(1–2):567–578, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006774.

Rodriguez:2000:TRM

[RRWT00] José F. Rodríguez, John E. Renaud, Brett A. Wujek, and Ravindra V. Tappeta.

Trust region model management in multidisciplinary design optimization. Journal of Computational and Applied Mathematics, 124 (1–2):139–154, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004246.

Robbe:2000:DTL

[RS00] M. Robbé and M. Sadkane. Discrete-time Lyapunov stability of large matrices. Journal of Computational and Applied Mathematics, 115(1–2):479–494, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001855.

Rundell:2004:NTI

[RS04] William Rundell and Paul Sacks. Numerical technique for the inverse resonance problem. Journal of Computational and Applied Mathematics, 170 (2):337–347, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000743.

Repin:2004:PEE

[RSS04] Sergey Repin, Stefan Sauter, and Anton Smolianski. A posteriori error estimation for the Poisson equation with mixed Dirichlet/Neumann boundary conditions. Journal of Computational and Applied Mathematics, 164–165 (??):601–612, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703004916.

Reitzinger:2003:AMC

[RSvR03] Stefan Reitzinger, Ute Schreiber, and Ursula van Rienen. Algebraic multigrid for complex symmetric matrices and applications. Journal of Computational and Applied Mathematics, 155(2):405–421, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008774.

Rui:2002:SMF

[Rui02] Hongxing Rui. Symmetric modified finite volume element methods for self-adjoint elliptic and parabolic problems. Journal of Computational and Applied Mathematics, 146 (2):373–386, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003709.

Rump:2003:TMB

[Rum03] Siegfried M. Rump. Ten methods to bound multiple roots of polynomials. Journal of Computational and Applied Mathematics, 156(2):403–432, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003819.

Rost:2000:RIR

[RV00] Karla Rost and Zdenek Vavrín.

Rational interpolation and recursive solution of Löwner– Vandermonde systems of equations. Journal of Computational and Applied Mathematics, 114(2):319–331, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002708.

Riaza:2002:DIO

[RZ02] Ricardo Riaza and Pedro J.

Zufiria. Discretization of implicit ODEs for singular root-finding problems. Journal of Computational and Applied Mathematics, 140(1–2):695–712, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004046.

Roos:2003:SDM

[RZ03] Hans-Görg Roos and Helena Zarin. The streamlinediffusion method for a convection– diffusion problem with a point source. Journal of Computational and Applied Mathematics, 150(1):109–128, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200568X.

Rouillier:2004:EIP

[RZ04] Fabrice Rouillier and Paul Zimmermann. Efficient isolation of polynomial’s real roots. Journal of Computational and Applied Mathematics, 162(1):33–50, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007271.

Ronveaux:2000:COP

[RZAG00] A. Ronveaux, A. Zarzo, I. Area, and E. Godoy. Classical orthogonal polynomials: dependence of parameters. Journal of Computational and Applied Mathematics, 121 (1–2):95–112, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003502.

Sallam:2000:QSI

[SA00] S. Sallam and M. Naim Anwar. Quintic C^{2}-spline integration methods for solving second-order ordinary initial value problems. Journal of Computational and Applied Mathematics, 115(1– 2):495–502, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799001740.

Schonauer:2001:HWS

[SA01] Willi Schönauer and Torsten Adolph. How we solve PDEs. Journal of Computational and Applied Mathematics, 131(1–2):473–492, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002557.

Schonauer:2003:FHW

[SA03] Willi Schönauer and Torsten Adolph. FDEM: how we make the FDM more flexible than the FEM. Journal of Computational and Applied Mathematics, 158 (1):157–167, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004618.

Sadek:2001:OPC

[SAA01] I. Sadek, M. Abukhaled, and T. Abualrub. Optimal pointwise control for a parallel system of Euler–Bernoulli beams. Journal of Computational and Applied Mathematics, 137(1):83–95, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700007007.

Sadkane:2001:NNM

[Sad01]  M. Sadkane.  A note on

normal matrices. Journal of Computational and Applied Mathematics, 136 (1–2):185–187, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006099.

Smaoui:2004:MDN

[SAE04] Nejib Smaoui and Suad Al-Enezi. Modelling the dynamics of nonlinear partial differential equations using neural networks. Journal of Computational and Applied Mathematics, 170(1):27–58, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000123.

Saito:2004:HSA

[Sai04] Norikazu Saito. A holomorphic semigroup approach to the lumped mass finite element method. Journal of Computational and Applied Mathematics, 169(1):71–85, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009440.

Samaris:2001:EKC

[Sam01a]  N. Samaris.  An exten-

sion of Krzyz’s conjecture. Journal of Computational and Applied Mathematics, 137(1):13–18, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000621X.

Samoletov:2001:SCF

[Sam01b] A. A. Samoletov. A sum containing factorials. Journal of Computational and Applied Mathematics, 131(1– 2):503–504, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005653.

Sanchez:2000:NND

[San00] L. Sanchez. A note on a nonautonomous O.D.E. related to the Fisher equation. Journal of Computational and Applied Mathematics, 113(1–2):201–209, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900254X.

Santos:2003:ESM

[San03] Manuel S. Santos. Estimation by simulation of monotone dynamical systems. Journal of Computational and Applied Mathematics, 158 (1):145–156, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004631.

Sargent:2000:OC

[Sar00]  R. W. H. Sargent.  Op-

timal control. Journal of Computational and Applied Mathematics, 124(1– 2):361–371, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004180.

Sembera:2001:NGM

[SB01] Jan Sembera and Michal Benes. Nonlinear Galerkin method for reaction–diffusion systems admitting invariant regions. Journal of Computational and Applied Mathematics, 136(1–2):163– 176, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005823.

Song:2004:QBN

[SB04] Yihong Song and Christopher T. H. Baker. Qualitative behaviour of numerical approximations to Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 172 (1):101–115, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001001.

Shi:2003:OGC

[SBSA03] Peng Shi, El-Kébir Boukas, Yan Shi, and Ramesh K. Agarwal. Optimal guaranteed cost control of uncertain discrete time-delay systems. Journal of Computational and Applied Mathematics, 157(2):435–451, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004333.

Shampine:2000:IVP

[SC00] L. F. Shampine and Robert M.

Corless. Initial value problems for ODEs in problem solving environments. Journal of Computational and Applied Mathematics, 125 (1–2):31–40, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004568.

Smiley:2001:AFA

[SC01] Michael W. Smiley and Changbum Chun. An algorithm for finding all solutions of a nonlinear system. Journal of Computational and Applied Mathematics, 137 (2):293–315, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007111.

Smirnov:2004:QKM

[SC04] Yuri F. Smirnov and Corrado Campigotto. The quantum q-Krawtchouk and q-Meixner polynomials and their related D-functions for the quantum groups SU_{q}(2) and SU_{q}(1, 1). Journal of Computational and Applied Mathematics, 164– 165(??):643–660, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006411.

Schaback:2000:UTR

[Sch00a] Robert Schaback. A unified theory of radial basis functions: Native Hilbert spaces for radial basis functions II. Journal of Computational and Applied Mathematics, 121 (1–2):165–177, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003459.

Schropp:2000:CFI

[Sch00b]  Johannes Schropp.  Con-

serving first integrals under discretization with variable step size integration procedures. Journal of Computational and Applied Mathematics, 115(1–2):503–517, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001788.

Schoutens:2001:ASL

[Sch01]  Wim Schoutens.  An ap-

plication in stochastics of the Laguerre-type polynomials. Journal of Computational and Applied Mathematics, 133(1–2):593–600, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006804.

Schmidt:2002:EGP

[Sch02a] Karl Michael Schmidt. Eigenvalues in gaps of perturbed periodic Dirac operators: numerical evidence. Journal of Computational and Applied Mathematics, 148 (1):169–181, November 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005800.

Schroder:2002:SFB

[Sch02b]  Peter Schröder.  Subdivi-

sion as a fundamental building block of digital geometry processing algorithms. Journal of Computational and Applied Mathematics, 149 (1):207–219, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005307.

Schafer:2003:ABV

[Sch03a] Uwe Schäfer. Aspects for a block version of the interval Cholesky algorithm. Journal of Computational and Applied Mathematics, 152(1–2):481–491, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007252.

Schiefermayr:2003:RWS

[Sch03b] Klaus Schiefermayr. Random walks with similar transition probabilities. Journal of Computational and Applied Mathematics, 153(1– 2):423–432, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702006404.

Schlosser:2003:NSR

[Sch03c] Michael Schlosser. A nonterminating _{8}ϕ_{7} summation for the root system C_{r}. Journal of Computational and Applied Mathematics, 160 (1–2):283–296, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006320.

Schropp:2003:OMD

[Sch03d] Johannes Schropp. One- and multistep discretizations of index 2 differential algebraic systems and their use in optimization. Journal of Computational and Applied Mathematics, 150(2):375–396, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006714.

Schittkowski:2004:DFP

[Sch04a] K. Schittkowski. Data fitting in partial differential algebraic equations: some academic and industrial applications. Journal of Computational and Applied Mathematics, 163(1):29–57, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007908.

Schmidt:2004:IDE

[Sch04b] Karl Michael Schmidt. On the integrability of differential equations by quadratures according to maximovic. Journal of Computational and Applied Mathematics, 171(1–2):393–410, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000615.

Schulz:2004:SSA

[Sch04c] Volker Schulz. Simultaneous solution approaches for large optimization problems. Journal of Computational and Applied Mathematics, 164– 165(??):629–641, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008318.

Sherali:2000:ESA

[SD00] Hanif D. Sherali and Patrick J.

Driscoll. Evolution and state-of-the-art in integer programming. Journal of Computational and Applied Mathematics, 124(1–2):319– 340, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700004313.

Shariff:2003:MAL

[SD03] M. H. B. M. Shariff and J. R.

Dormand. A modified augmented Lagrangian method for a class of constrained problems. Journal of Computational and Applied Mathematics, 151(2):257–270, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008130.

Sergeant:2004:MFC

[SDMV04] P. Sergeant, L. Dupré, J. Melkebeek, and L. Vandenbossche. Magnetic field computation for optimized shielding of induction heaters. Journal of Computational and Applied Mathematics, 168(1–2):437–446, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009890.

Stortelder:2001:FEF

[SdSP01] Walter J. H. Stortelder, Jacques J. B. de Swart, and János D. Pintér. Finding elliptic Fekete points sets: two numerical solution approaches. Journal of Computational and Applied Mathematics, 130(1–2):205–216, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003829.

Sedjelmaci:2004:MRG

[Sed04] Sidi Mohammed Sedjelmaci. A modular reduction for GCD computation. Journal of Computational and Applied Mathematics, 162(1):17–31, January 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007283.

Segura:2003:ZTP

[Seg03] Javier Segura. On the zeros and turning points of special functions. Journal of Computational and Applied Mathematics, 153(1– 2):433–440, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702006143.

Silvester:2001:EPL

[SEKW01] David Silvester, Howard Elman, David Kay, and Andrew Wathen. Efficient preconditioning of the linearized Navier–Stokes equations for incompressible flow. Journal of Computational and Applied Mathematics, 128(1–2):261–279, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000515X.

Senhadji:2001:CNS

[Sen01] M. N. Senhadji. On condition numbers of the Shanks transformation. Journal of Computational and Applied Mathematics, 135(1):41–61, October 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005616.

Sesma:2004:TSR

[Ses04] Javier Sesma. The Temme’s sum rule for confluent hypergeometric functions revisited. Journal of Computational and Applied Mathematics, 163(2):429–431, February 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007659.

Seydou:2001:PIS

[Sey01] F. Seydou. Profile inversion in scattering theory: the TE case. Journal of Computational and Applied Mathematics, 137(1):49–60, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006294.

Seaid:2004:ENM

[SFK^{+}04] M. Seaïd, M. Frank, A. Klar, R. Pinnau, and G. Thömmes. Efficient numerical methods for radiation in gas turbines. Journal of Computational and Applied Mathematics, 170 (1):217–239, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000676.

Sabariego:2004:AFM

[SGG^{+}04] R. V. Sabariego, J. Gyselinck, C. Geuzaine, P. Dular, and W. Legros. Application of the fast multipole method to hybrid finite element-boundary element models. Journal of Computational and Applied Mathematics, 168(1–2):403–412, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042703009841.

Shapira:2001:AMM

[Sha01] Yair Shapira. Algebraic multilevel method with application to the Maxwell equations. Journal of Computational and Applied Mathematics, 137(1):207–211, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005118.

Shi:2002:NFE

[Shi02]  Zhong-Ci Shi.  Noncon-

forming finite element methods. Journal of Computational and Applied Mathematics, 149(1):221–225, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005319.

Shishkin:2004:DAS

[Shi04a] G. I. Shishkin. Discrete approximations of solutions and derivatives for a singularly perturbed parabolic convection– diffusion equation. Journal of Computational and Applied Mathematics, 166 (1):247–266, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042703008562.

Shishkin:2004:LAM

[Shi04b] G. I. Shishkin. Limitations of adaptive mesh refinement techniques for singularly perturbed problems with a moving interior layer. Journal of Computational and Applied Mathematics, 166(1):267–280, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008550.

Sun:2002:GCN

[SHS02] Wenyu Sun, Jiye Han, and Jie Sun. Global convergence of nonmonotone descent methods for unconstrained optimization problems. Journal of Computational and Applied Mathematics, 146(1):89–98, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200420X.

Siafarikas:2001:FMO

[Sia01a]  P. D. Siafarikas.  Find-

ing a measure of orthogonality. Journal of Computational and Applied Mathematics, 133(1–2):697, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/

S0377042700007263. See com-

ments [Van01].

Siafarikas:2001:P

[Sia01b] Panayiotis D. Siafarikas. Preface. Journal of Computational and Applied Mathematics, 133(1–2):xiii–xiv, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006889.

Sidi:2000:GRE

[Sid00] Avram Sidi. The generalized Richardson extrapolation process GREP^{(1)} and computation of derivatives of limits of sequences with applications to the d^{(1)}-transformation. Journal of Computational and Applied Mathematics, 122 (1–2):251–273, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003629.

Siegmund:2002:SND

[Sie02] Stefan Siegmund. A spectral notion for dynamic equations on time scales. Journal of Computational and Applied Mathematics, 141(1–2):255–265, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042701004514.

Simon:2000:SER

[Sim00] Alice (Chaljub) Simon. Some existence results for quasilinear elliptic problems via the fixed-point theorem of tarski. Journal of Computational and Applied Mathematics, 113(1–2):211–216, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002563.

Shimasaki:2002:FLE

[SIM02] Masaaki Shimasaki, Takeshi Iwashita, and Takeshi Mifune. Fast linear equation solvers in high performance electromagnetic field analysis. Journal of Computational and Applied Mathematics, 149(1):341–349, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200540X.

Simon:2004:AFT

[Sim04] Barry Simon. Analogs of the m-function in the theory of orthogonal polynomials on the unit circle. Journal of Computational and Applied Mathematics, 171 (1–2):411–424, October 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000627.

Shizgal:2003:TRG

[SJ03a] Bernie D. Shizgal and Jae-Hun Jung. Towards the resolution of the Gibbs phenomena. Journal of Computational and Applied Mathematics, 161(1):41–65, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005004.

Sridhar:2003:AZF

[SJ03b] R. Sridhar and R. Jagannathan. On the q-analogues of the Zassenhaus formula for disentangling exponential operators. Journal of Computational and Applied Mathematics, 160 (1–2):297–305, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006332.

Sun:2001:NSM

[SK01] Fangyu Sun and Peter Kosmol. A new simultaneous method of fourth order for finding complex zeros in circular interval arithmetic. Journal of Computational and Applied Mathematics, 130(1–2):293–307, May 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003751.

Seaid:2004:FLC

[SKD04] M. Seaïd, A. Klar, and B. Dubroca. Flux limiters in the coupling of radiation and hydrodynamic models. Journal of Computational and Applied Mathematics, 168(1–2):425–435, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009737.

Sakurai:2003:EAT

[SKSV03] Tetsuya Sakurai, Peter Kravanja, Hiroshi Sugiura, and Marc Van Barel. An error analysis of two related quadrature methods for computing zeros of analytic functions. Journal of Computational and Applied Mathematics, 152(1–2):467–480, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007240.

Song:2003:CRI

[SL03] Baorui Song and Hua Lian. Convergence of the rational interpolants of Stieltjes functions. Journal of Computational and Applied Mathematics, 159(1):129–135, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005715.

Slavova:2000:ASM

[Sla00]  Angela Slavova.  Applica-

tions of some mathematical methods in the analysis of cellular neural networks. Journal of Computational and Applied Mathematics, 114(2):387–404, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002770.

Sloan:2000:Q

[Slo00]  Ian H. Sloan.  Qualoca-

tion. Journal of Computational and Applied Mathematics, 125 (1–2):461–478, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004854.

Slodicka:2004:ASN

[Slo04] Marián Slodicka. An approximation scheme for a nonlinear degenerate parabolic equation with a second-order differential Volterra operator. Journal of Computational and Applied Mathematics, 168(1–2):447–458, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009920.

Smith:2004:RSS

[SM04a] Chris J. Smith and Roland Mallier. Removing the separation singularity in a barotropic ocean with bottom friction. Journal of Computational and Applied Mathematics, 166(1):281–290, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008604.

Sugihara:2004:RDS

[SM04b] Masaaki Sugihara and Takayasu Matsuo. Recent developments of the Sinc numerical methods. Journal of Computational and Applied Mathematics, 164– 165(??):673–689, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008380.

Smith:2003:SEE

[Smi03]  H. V. Smith.  Some er-

ror expansions for certain Gaussian quadrature rules. Journal of Computational and Applied Mathematics, 155(2):331–337, June 15, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008725.

Song:2001:SBS

[Son01]  Yongzhong Song.  Semi-

convergence of block SOR method for singular linear systems with p-cyclic matrices. Journal of Computational and Applied Mathematics, 130(1–2):217–229, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003817.

Song:2002:GVE

[Son02] Wen Song. On generalized vector equilibrium problems. Journal of Computational and Applied Mathematics, 146 (1):167–177, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004260.

Sun:2001:SVI

[SQ01] Defeng Sun and Liqun Qi.

Solving variational inequality problems via smoothingnonsmooth reformulations. Journal of Computational and Applied Mathematics, 129 (1–2):37–62, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700005410.

Slimane:2004:TLP

[SR04] Leila Slimane and Yves Renard. The treatment of the locking phenomenon for a general class of variational inequalities. Journal of Computational and Applied Mathematics, 170 (1):121–143, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000160.

Sanchez-Ruiz:2000:EIO

[SRD00] Jorge Sánchez-Ruiz and Jesús S. Dehesa. Entropic integrals of orthogonal hypergeometric polynomials with general supports. Journal of Computational and Applied Mathematics, 118(1–2):311–322, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000296X.

Sanchez-Ruiz:2001:SCL

[SRD01] Jorge Sánchez-Ruiz and Jesús S. Dehesa. Some connection and linearization problems for polynomials in and beyond the Askey scheme. Journal of Computational and Applied Mathematics, 133 (1–2):579–591, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006798.

Srivastava:2002:RSC

[Sri02] H. M. Srivastava. Remarks on a sum containing factorials. Journal of Computational and Applied Mathematics, 142(2):441–444, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003552.

Sanchez-Ruiz:2003:ESV

[SRLAD03] J. Sánchez-Ruiz, P. López-Artés, and J. S. Dehesa. Expansions in series of varying Laguerre polynomials and some applications to molecular potentials. Journal of Computational and Applied Mathematics, 153(1– 2):411–421, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702006155.

Steinebach:2004:MCR

[SRRS04] G. Steinebach, S. Rademacher, P. Rentrop, and M. Schulz. Mechanisms of coupling in river flow simulation systems. Journal of Computational and Applied Mathematics, 168(1–2):459–470, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009944.

Stys:2000:HOF

[SS00] Krystyna Styś and Tadeusz Styś. A higher-order finite difference method for solving a system of integrodifferential equations. Journal of Computational and Applied Mathematics, 126 (1–2):33–46, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003398.

Stamova:2001:LRM

[SS01] Ivanka M. Stamova and Gani T. Stamov. Lyapunov– Razumikhin method for impulsive functional differential equations and applications to the population dynamics. Journal of Computational and Applied Mathematics, 130(1–2):163–171, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003854.

Sakurai:2002:ICI

[SS02a] Tetsuya Sakurai and Hiroshi Sugiura. Improvement of convergence of an iterative method for finding polynomial factors of analytic functions. Journal of Computational and Applied Mathematics, 140(1–2):713–725, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005246.

Schulz:2002:TDM

[SS02b] Martin Schulz and Gerd Steinebach. Two-dimensional modelling of the river Rhine. Journal of Computational and Applied Mathematics, 145 (1):11–20, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701005325.

Strauss:2002:HWS

[SS02c] Daniel J. Strauss and Gabriele Steidl. Hybrid wavelet-support vector classification of waveforms. Journal of Computational and Applied Mathematics, 148(2):375–400, November 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005575.

Sahai:2003:IR

[SS03a] Vivek Sahai and Shalini Srivastava. On irreducible p, q-representations of }↕(∈). Journal of Computational and Applied Mathematics, 160 (1–2):271–281, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006319.

Sakurai:2003:PMG

[SS03b] Tetsuya Sakurai and Hiroshi Sugiura. A projection method for generalized eigenvalue problems using numerical integration. Journal of Computational and Applied Mathematics, 159(1):119–128, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300565X.

Singh:2003:LRS

[SS03c] Arindama Singh and S. Sheela. Lardy’s regularization of a singularly perturbed elliptic PDE. Journal of Computational and Applied Mathematics, 151(1):25–37, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007355.

Suzuki:2003:NIE

[SS03d] Toshio Suzuki and Tomohiro Suzuki. Numerical integration error method for zeros of analytic functions. Journal of Computational and Applied Mathematics, 152(1–2):493–505, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007264.

Shi:2004:GRA

[SS04] Zhen-Jun Shi and Jie Shen.

A gradient-related algorithm with inexact line searches. Journal of Computational and Applied Mathematics, 170 (2):349–370, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704000858.

Sima:2004:HPN

[SSV04] Vasile Sima, Diana Maria Sima, and Sabine Van Huffel. High-performance numerical algorithms and software for subspace-based linear multivariable system identification. Journal of Computational and Applied Mathematics, 170 (2):371–397, September 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270400086X.

Sameh:2000:TMM

[ST00a] Ahmed Sameh and Zhanye Tong. The trace minimization method for the symmetric generalized eigenvalue problem. Journal of Computational and Applied Mathematics, 123 (1–2):155–175, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003915.

Srivastava:2000:CCR

[ST00b] H. M. Srivastava and Hirofumi Tsumura. A certain class of rapidly convergent series representations for ζ(2n + 1). Journal of Computational and Applied Mathematics, 118(1–2):323–335, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003125.

Sirotkin:2002:PSM

[ST02] V. Sirotkin and P. Tarvainen. Parallel Schwarz methods for convection-dominated semilinear diffusion problems. Journal of Computational and Applied Mathematics, 145(1):189–211, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005751.

Stanton:2003:SPS

[Sta03] Dennis Stanton. Some positive sums. Journal of Computational and Applied Mathematics, 153(1–2):537, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008841.

Stenger:2000:SSN

[Ste00a]  Frank Stenger.  Summary

of Sinc numerical methods. Journal of Computational and Applied Mathematics, 121 (1–2):379–420, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003484. Numer-

ical analysis in the 20th century, Vol. I, Approximation theory.

Stephan:2000:MMV

[Ste00b] Ernst P. Stephan. Multilevel methods for the h-, p-, and hp-versions of the boundary element method. Journal of Computational and Applied Mathematics, 125 (1–2):503–519, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004878.

Steele:2002:BSA

[Ste02]  J. Michael Steele.  The

Bohnenblust–Spitzer algorithm and its applications. Journal of Computational and Applied Mathematics, 142 (1):235–249, May 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004721.

Stoyanova:2001:ICF

[Sto01] Srebra B. Stoyanova. Invariant cubature formulae of the ninth degree of accuracy for the hyperoctahedron. Journal of Computational and Applied Mathematics, 137(1):135–143, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007044.

Stoyanov:2005:CAE

[Sto05] B. J. Stoyanov. Comments on “Asymptotic expansion of a Bessel function integral using hypergeometric functions” by L. J. Landau and N. J. Luswili. Journal of Computational and Applied Mathematics, 176(2):259–262, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704003310.  See

[LL01a].

Strohmer:2000:NAN

[Str00] Thomas Strohmer. Numerical analysis of the non-uniform sampling problem. Journal of Computational and Applied Mathematics, 122 (1–2):297–316, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003617.

Stuben:2001:RAM

[Stü01]  K. Stüben.  A review of

algebraic multigrid. Journal of Computational and Applied Mathematics, 128(1– 2):281–309, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005161.

Singh:2000:GFP

[STW00] S. P. Singh, E. Tarafdar, and B. Watson. A generalized fixed point theorem and equilibrium point of an abstract economy. Journal of Computational and Applied Mathematics, 113(1–2):65–71, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002447.

Stys:2004:HAF

[Sty04] Tadeusz Stys. A higher accuracy finite difference method for an elliptic equation of order four. Journal of Computational and Applied Mathematics, 164–165(??): 661–672, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270300832X.

Surla:2004:UAS

[SU04] Katarina Surla and Zorica Uzelac. A uniformly accurate spline collocation method for a normalized flux. Journal of Computational and Applied Mathematics, 166(1):291–305, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008549.

Suarez:2001:ASO

[Suá01]  C. Suárez.  About semi-

classical orthogonal polynomials of the class (2,2). Journal of Computational and Applied Mathematics, 131(1–2):457–472, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002569.

Suarez:2003:ODR

[Suá03]  C. Suárez.  On the or-

thogonality of the derivative of the reciprocal sequence. Journal of Computational and Applied Mathematics, 157(1):139–154, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003844.

Sugihara:2002:DET

[Sug02a]  Masaaki Sugihara.  Dou-

ble exponential transformation in the sinc-collocation method for two-point boundary value problems. Journal of Computational and Applied Mathematics, 149 (1):239–250, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005332.

Sugita:2002:RNI

[Sug02b] Hiroshi Sugita. Robust numerical integration and pairwise independent random variables. Journal of Computational and Applied Mathematics, 139(1):1–8, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003946.

Sun:2001:SDA

[Sun01] J. Sun. Solving the discrete l_{p}-approximation problem by a method of centers. Journal of Computational and Applied Mathematics, 129(1–2):63–76, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005422.

Suri:2001:FEM

[Sur01]  Manil Suri.  The p and

hp finite element method for problems on thin domains. Journal of Computational and Applied Mathematics, 128(1–2):235–260, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005148.

Sorokin:2000:MHP

[SV00] Vladimir Sorokin and Jeannette Van Iseghem. Matrix Hermite–Padé problem and dynamical systems. Journal of Computational and Applied Mathematics, 122 (1–2):275–295, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000354X.

Stergioulas:2004:BAR

[SV04] L. K. Stergioulas and A. Vourdas. The Bargmann analytic representation in signal analysis. Journal of Computational and Applied Mathematics, 167(1):183–192, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008963.

Simos:2003:P

[SVA03] T. E. Simos and J. Vigo-Aguiar. Preface. Journal of Computational and Applied Mathematics, 158 (1):ix, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270300459X.

Saad:2000:ISL

[SvdV00] Yousef Saad and Henk A.

van der Vorst. Iterative solution of linear systems in the 20th century. Journal of Computational and Applied Mathematics, 123(1–2):1–33, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000412X.

Stergioulas:2001:OBG

[SVV01] L. K. Stergioulas, V. S. Vassiliadis, and A. Vourdas. Optimal bases of Gaussians in a Hilbert space: applications in mathematical signal analysis. Journal of Computational and Applied Mathematics, 133(1–2):601–609, August 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006853.

Shaw:2000:AST

[SW00a] Simon Shaw and J. R. Whiteman. Adaptive space-time finite element solution for Volterra equations arising in viscoelasticity problems. Journal of Computational and Applied Mathematics, 125 (1–2):337–345, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004775.

Si:2000:ASS

[SW00b] Jian-Guo Si and Xin-Ping Wang. Analytic solutions of a second-order iterative functional differential equation. Journal of Computational and Applied Mathematics, 126 (1–2):277–285, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003593.

Sofronov:2001:ELF

[SW01] I. L. Sofronov and W. L. Wendland. Exact linear farfield conditions for threedimensional aerodynamic stationary transonic flows. Journal of Computational and Applied Mathematics, 136 (1–2):317–335, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006233.

Shi:2002:SSB

[SW02a] Xiquan Shi and Renhong Wang. Spline space and its Bsplines on an n + 1 direction mesh in R^{n}. Journal of Computational and Applied Mathematics, 144(1–2):241–250, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005647.

Sloan:2002:GAS

[SW02b] Ian H. Sloan and Robert S.

Womersley. Good approximation on the sphere, with application to geodesy and the scattering of sound. Journal of Computational and Applied Mathematics, 149 (1):227–237, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005320.

Su:2003:UDT

[SW03] H. Su and C. A. Woodham. On the uncontrollable damped triple inverted pendulum. Journal of Computational and Applied Mathematics, 151(2):425–443, February 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006635.

Swann:2000:UCD

[Swa00] Howard Swann. On using the cell discretization algorithm for mixed-boundary value problems and domain decomposition. Journal of Computational and Applied Mathematics, 115(1–2):519–534, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001867.

Swann:2002:ASN

[Swa02] H. Swann. On approximating the solution of the nonstationary Stokes equations using the cell discretization algorithm. Journal of Computational and Applied Mathematics, 140(1–2):727–750, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005234.

Swann:2004:SHD

[Swa04] Howard Swann. On solenoidal high-degree polynomial approximations to solutions of the stationary Navier– Stokes equations. Journal of Computational and Applied Mathematics, 164– 165(??):691–706, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300952X.

Saucez:2004:MLS

[SWSZ04] P. Saucez, A. Vande Wouwer, W. E. Schiesser, and P. Zegeling. Method of lines study of nonlinear dispersive waves. Journal of Computational and Applied Mathematics, 168(1–2):413–423, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009889.

Spedicato:2000:AAL

[SXZ00] Emilio Spedicato, Zunquan Xia, and Liwei Zhang. ABS algorithms for linear equations and optimization. Journal of Computational and Applied Mathematics, 124 (1–2):155–170, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004192.

Yang:2001:DDC

[sY01] Xin she Yang. Density-driven compactional flow in porous media. Journal of Computational and Applied Mathematics, 130(1–2):245–257, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003799.

Syam:2003:CSI

[Sya03] Muhammed I. Syam. Cubic spline interpolation predictors over implicitly defined curves. Journal of Computational and Applied Mathematics, 157(2):283–295, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004114.

Syam:2004:FAR

[Sya04] Muhammed I. Syam. Finding all real zeros of polynomial systems using multiresultant. Journal of Computational and Applied Mathematics, 167(2):417–428, June 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009300.

Shi:2002:CCB

[SYW02] Xiquan Shi, Piqiang Yu, and Tianjun Wang. G^{1} continuous conditions of biquartic B-spline surfaces. Journal of Computational and Applied Mathematics, 144(1–2):251–262, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005659.

Saad:2001:EML

[SZ01] Yousef Saad and Jun Zhang.

Enhanced multi-level block ILU preconditioning strategies for general sparse linear systems. Journal of Computational and Applied Mathematics, 130(1–2):99–118, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900388X.

Slavova:2003:CMS

[SZ03a] Angela Slavova and Pietro Zecca. CNN model for studying dynamics and travelling wave solutions of FitzHugh– Nagumo equation. Journal of Computational and Applied Mathematics, 151(1): 13–24, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702007343.

Sun:2003:ICN

[SZ03b] Jitao Sun and Yinping Zhang.

Impulsive control of a nuclear spin generator. Journal of Computational and Applied Mathematics, 157(1): 235–242, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042703004540.

Si:2004:ASC

[SZ04] Jianguo Si and Weinian Zhang. Analytic solutions of a class of iterative functional differential equations. Journal of Computational and Applied Mathematics, 162 (2):467–481, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300760X.

Szafraniec:2001:MIM

[Sza01] Franciszek Hugon Szafraniec. On matrix integration of matrix polynomials. Journal of Computational and Applied Mathematics, 133 (1–2):611–621, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006841.

Szafnicki:2002:UAD

[Sza02] Boleslaw Szafnicki. A unified approach for degree reduction of polynomials in the Bernstein basis. Part I: Real polynomials. Journal of Computational and Applied Mathematics, 142(2):287–312, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004939.

Szynal:2003:LPK

[Szy03] Jan Szynal. Laguerre polynomials and the krzyz conjecture. Journal of Computational and Applied Mathematics, 153(1–2):539–541, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008853.

Steinberg:2001:NSS

[SZZ01] Stanly L. Steinberg, Janaína P.

Zingano, and Paulo R. Zingano. On nilpotent singular systems. Journal of Computational and Applied Mathematics, 137(1):97–107, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007019.

Tocino:2002:RKM

[TA02] A. Tocino and R. Ardanuy.

Runge–Kutta methods for numerical solution of stochastic differential equations. Journal of Computational and Applied Mathematics, 138 (2):219–241, January 15, 2002. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003806.

Tsuji:2003:RPS

[TA03] Kumiko Tsuji and Miki Aoyagi. Recognition of planar shapes by a semi-differential invariant for moving triangles. Journal of Computational and Applied Mathematics, 152(1–2):533–544, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007288.

Takano:2003:IDN

[Tak03]  Katsuo Takano.  On infi-

nite divisibility of normed product of Cauchy densities. Journal of Computational and Applied Mathematics, 150(2):253–263, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006623.

Tammeraid:2003:SRA

[Tam03] Ivar Tammeraid. Several remarks on acceleration of convergence using generalized linear methods of summability. Journal of Computational and Applied Mathematics, 159(2):365–373, October 15, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005399.

Tang:2002:CRN

[Tan02] Hui-Chin Tang. Combined random number generator via the generalized Chinese remainder theorem. Journal of Computational and Applied Mathematics, 142 (2):377–388, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701004241.

Taseli:2000:ANB

[Tas00a] H. Taseli. Accurate numerical bounds for the spectral points of singular Sturm–Liouville problems over −∞ < x < ∞. Journal of Computational and Applied Mathematics, 115(1–2):535–546, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003027.

Taswell:2000:CSS

[Tas00b]  Carl Taswell.  Constraint-

selected and search-optimized families of Daubechies wavelet filters computable by spectral factorization. Journal of Computational and Applied Mathematics, 121 (1–2):179–195, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000340X.

Taseli:2004:ANB

[Tas04] H. Taseli. Accurate numerical bounds for the spectral points of singular Sturm–Liouville problems over 0 < x < ∞. Journal of Computational and Applied Mathematics, 164– 165(??):707–722, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006460.

Taylor:2000:SPH

[Tay00] Stephen W. Taylor. A smoothing property of a hyperbolic system and boundary controllability. Journal of Computational and Applied Mathematics, 114(1):23–40, January 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002861.

Tang:2001:OCL

[TC01] Xian-Hua Tang and Sui Sun Cheng. An oscillation criterion for linear difference equations with oscillating coefficients. Journal of Computational and Applied Mathematics, 132(2):319–329, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004362.

Tellez:2000:DPD

[Tél00] Manuel Aguirre Téllez. The distributional product of Dirac’s delta in a hypercone. Journal of Computational and Applied Mathematics, 115(1– 2):13–21, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799001247.

Temme:2000:NAA

[Tem00] Nico M. Temme. Numerical and asymptotic aspects of parabolic cylinder functions. Journal of Computational and Applied Mathematics, 121 (1–2):221–246, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003472.

Temme:2003:LPC

[Tem03] Nico M. Temme. Large parameter cases of the Gauss hypergeometric function. Journal of Computational and Applied Mathematics, 153(1– 2):441–462, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702006271.

Terwilliger:2003:ILP

[Ter03]  Paul Terwilliger.  Intro-

duction to Leonard pairs. Journal of Computational and Applied Mathematics, 153(1–2):463–475, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006003.

Thomee:2001:FDF

[Tho01] Vidar Thomée. From finite differences to finite elements: a short history of numerical analysis of partial differential equations. Journal of Computational and Applied Mathematics, 128(1–2):1–54, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005070.

Thukral:2002:IIF

[Thu02]  R. Thukral.  Introduction

to the improved functional epsilon algorithm. Journal of Computational and Applied Mathematics, 147 (1):9–25, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702003734.

Tidriri:2002:AHF

[Tid02] M. D. Tidriri. Analysis of the hybrid finite element/ finite volume methods for linear hyperbolic and convectiondominated convection–diffusion problems. Journal of Computational and Applied Mathematics, 139(2):323–350, February 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042701004162.

Tidriri:2003:EEH

[Tid03] M. D. Tidriri. Error estimates for the hybrid finite element/finite volume methods for linear hyperbolic and convection-dominated problems. Journal of Computational and Applied Mathematics, 156(1):77–92, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702009056.

Tan:2004:NLM

[TJ04] Jieqing Tan and Ping Jiang.

A Neville-like method via continued fractions. Journal of Computational and Applied Mathematics, 163 (1):219–232, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042703008082.

Temme:2001:ASH

[TL01] N. M. Temme and J. L. López. The Askey scheme for hypergeometric orthogonal polynomials viewed from asymptotic analysis. Journal of Computational and Applied Mathematics, 133(1–2):623–633, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000683X.

To:2004:PES

[TL04] K. N. To and C. C. Lim. Perturbation to enhance support vector machines for classification. Journal of Computational and Applied Mathematics, 163(1):233–239, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008094.

Tong:2002:IMS

[TLQ02] Xiaojiao Tong, Dong-Hui Li, and Liqun Qi. An iterative method for solving semismooth equations. Journal of Computational and Applied Mathematics, 146(1):1–10, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042702004132.

Thiran:2003:GAM

[TML03] J. P. Thiran, M. Matelart, and B. Le Bailly. On the generalized ADI method for the matrix equation X − AXB = C. Journal of Computational and Applied Mathematics, 156(2):285–302, July 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702009172.

Toyonaga:2002:VNC

[TNW02] K. Toyonaga, M. T. Nakao, and Y. Watanabe. Verified numerical computations for multiple and nearly multiple eigenvalues of elliptic operators. Journal of Computational and Applied Mathematics, 147(1):175–190, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004314.

Terbeche:2003:FSD

[TO03] Mekki Terbeche and Broderick O. Oluyede. Fixed and sequential designs for estimation in the exponential family with comparisons and applications to binomial proportions using beta priors. Journal of Computational and Applied Mathematics, 154(1):175–182, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008208.

Tomantschger:2002:SSC

[Tom02] K. W. Tomantschger. Series solutions of coupled differential equations with one regular singular point. Journal of Computational and Applied Mathematics, 140(1–2):773–783, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005982.

Torre:2003:LRC

[Tor03] A. Torre. Linear and radial canonical transforms of fractional order. Journal of Computational and Applied Mathematics, 153(1– 2):477–486, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702006374.

Tuwankotta:2003:GNI

[TQ03] J. M. Tuwankotta and G. R. W. Quispel. Geometric numerical integration applied to the elastic pendulum at higher-order resonance. Journal of Computational and Applied Mathematics, 154(1):229–242, May 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008257.

Tamizhmani:2003:SFS

[TRTG03] K. M. Tamizhmani, A. Ramani, T. Tamizhmani, and B. Grammaticos. Special functions as solutions to discrete Painlevé equations. Journal of Computational and Applied Mathematics, 160 (1–2):307–313, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006356.

Truhar:2004:EAD

[Tru04] Ninoslav Truhar. An efficient algorithm for damper optimization for linear vibrating systems using Lyapunov equation. Journal of Computational and Applied Mathematics, 172(1):169–182, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001165.

Tsitouras:2002:ORK

[TS02] Ch. Tsitouras and T. E. Simos.

Optimized Runge–Kutta pairs for problems with oscillating solutions. Journal of Computational and Applied Mathematics, 147(2):397–409, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004752.

Tselyaev:2004:GNM

[Tse04] V. I. Tselyaev. A generalized Numerov method for linear second-order differential equations involving a first derivative term. Journal of Computational and Applied Mathematics, 170 (1):103–120, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000159.

Tsuchiya:2001:FEA

[Tsu01] Takuya Tsuchiya. Finite element analysis for parametrized nonlinear equations around turning points. Journal of Computational and Applied Mathematics, 132(2): 255–276, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042700003289.

Tan:2002:CSM

[TT02] Jieqing Tan and Shuo Tang. Composite schemes for multivariate blending rational interpolation. Journal of Computational and Applied Mathematics, 144(1–2):263–275, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005660.

Tocino:2003:NIT

[TVA03] A. Tocino and J. Vigo-Aguiar.

New Itô–Taylor expansions. Journal of Computational and Applied Mathematics, 158 (1):169–185, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004643.

Theuns:2004:VIM

[TVV04] E. Theuns, J. Vierendeels, and P. Vandevelde. Validation of the integral model for the pyrolysis of charring materials with a moving grid. Journal of Computational and Applied Mathematics, 168(1–2):471–479, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300997X.

Tai:2004:ISC

[TW04] Chiew-Lan Tai and Guo-Jin Wang. Interpolation with slackness and continuity control and convexitypreservation using singular blending. Journal of Computational and Applied Mathematics, 172(2):337–361, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001220.

Tiesinga:2002:BAI

[TWV02] G. Tiesinga, F. W. Wubs, and A. E. P. Veldman. Bifurcation analysis of incompressible flow in a driven cavity by the Newton–Picard method. Journal of Computational and Applied Mathematics, 140(1–2):751–772, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005155.

Tsuchiya:2003:YPA

[TYI03] Takuya Tsuchiya, Kazuki Yoshida, and Sae Ishioka. Yamamoto’s principle and its applications to precise finite element error analysis. Journal of Computational and Applied Mathematics, 152(1–2):507–532, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007276.

Tang:2004:VOE

[TYZL04] Chen Tang, Haiqing Yan, Hao Zhang, and WenRun Li. The various order explicit multistep exponential fitting for systems of ordinary differential equations. Journal of Computational and Applied Mathematics, 169(1):171–182, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703010227.

Ushijima:2003:EEF

[UC03a] Teruo Ushijima and Fumihiro Chiba. Error estimates for a fundamental solution method applied to reduced wave problems in a domain exterior to a disc. Journal of Computational and Applied Mathematics, 159(1):137–148, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005594.

Ushijima:2003:FSM

[UC03b] Teruo Ushijima and Fumihiro Chiba. A fundamental solution method for the reduced wave problem in a domain exterior to a disc. Journal of Computational and Applied Mathematics, 152(1–2):545–557, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200729X.

Uesaka:2004:ONP

[Ues04] Hiroshi Uesaka. Oscillation or nonoscillation property for semilinear wave equations. Journal of Computational and Applied Mathematics, 164– 165(??):723–730, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004928.

Usman:2000:ASS

[UH00] Anila Usman and George Hall. Alternative stepsize strategies for Adams predictor–corrector codes. Journal of Computational and Applied Mathematics, 116(1):105–120, April 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003131.

Ushijima:2004:CCA

[UY04] Takeo K. Ushijima and Shigetoshi Yazaki. Convergence of a crystalline approximation for an area-preserving motion. Journal of Computational and Applied Mathematics, 166(2):427–452, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703007684.

Vrahatis:2000:CGU

[VALM00] M. N. Vrahatis, G. S. Androulakis, J. N. Lambrinos, and G. D. Magoulas. A class of gradient unconstrained minimization algorithms with adaptive stepsize. Journal of Computational and Applied Mathematics, 114(2):367–386, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002769.

VanderJeugt:2000:HSR

[Van00a] J. Van der Jeugt. Hypergeometric series related to the 9-j coefficient of ∫⊓(∞, ∞). Journal of Computational and Applied Mathematics, 118(1–2):337–351, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002971.

VanLoan:2000:UKP

[Van00b] Charles F. Van Loan. The ubiquitous Kronecker product. Journal of Computational and Applied Mathematics, 123 (1–2):85–100, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003939.

VanAssche:2001:CFM

[Van01] Walter Van Assche. Comments: [“Finding a measure of orthogonality” [J. Comput. Appl. Math. 133(1–2) 697–697 (2001)]. Journal of Computational and Applied Mathematics, 133(1–2):698–699, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700007275.  See

[Sia01a].

VanderJeugt:2002:TFD

[Van02] J. Van der Jeugt. Transformation formula for a double Clausenian hypergeometric series, its q-analogue, and its invariance group. Journal of Computational and Applied Mathematics, 139(1):65–73, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003892.

Vigo-Aguiar:2003:DCE

[VAR03] Jesús Vigo-Aguiar and Higinio Ramos. Dissipative Chebyshev exponential-fitted methods for numerical solution of secondorder differential equations. Journal of Computational and Applied Mathematics, 158 (1):187–211, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004734.

Vigo-Aguiar:2003:IHP

[VASF03] M. I. Vigo-Aguiar, M. E.

Sansaturio, and J. M. Ferrándiz. Integrability of Hamiltonians with polynomial potentials. Journal of Computational and Applied Mathematics, 158 (1):213–224, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004679.

VanGucht:2001:BEF

[VB01] P. Van Gucht and A. Bultheelf.

Bernstein equiconvergence and Fejér-type theorems for general rational Fourier series. Journal of Computational and Applied Mathematics, 133(1–2):635–645, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006622.

VanDeun:2003:ORF

[VB03] J. Van Deun and A. Bultheel.

Orthogonal rational functions and quadrature on an interval. Journal of Computational and Applied Mathematics, 153(1–2):487–495, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005988.

VanDeun:2004:IAO

[VB04] J. Van Deun and A. Bultheel. An interpolation algorithm for orthogonal rational functions. Journal of Computational and Applied Mathematics, 164– 165(??):749–762, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270300493X.

Vanlaere:2004:NNA

[VBL^{+}04] W. Vanlaere, P. Buffel, G. Lagae, R. Van Impe, and J. Belis. Neural networks for assessing the failure load of a construction. Journal of Computational and Applied Mathematics, 168(1–2):501–508, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703010082.

VanAssche:2001:SCM

[VC01] Walter Van Assche and Els Coussement. Some classical multiple orthogonal polynomials. Journal of Computational and Applied Mathematics, 127(1–2):317–347, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005033.

vanDoorn:2003:BDP

[vD03a] Erik A. van Doorn. Birthdeath processes and associated polynomials. Journal of Computational and Applied Mathematics, 153(1– 2):497–506, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702005940.

VandenBerghe:2003:SIB

[VD03b] G. Vanden Berghe and H. De Meyer. A special identity between three _{2}F_{1}(a, b; c; 4) hypergeometric series. Journal of Computational and Applied Mathematics, 160 (1–2):315–321, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006368.

vanderHouwen:2000:NTI

[vdH00] P. J. van der Houwen. Note on the time integration of 3D advection-reaction equations. Journal of Computational and Applied Mathematics, 116(2):275–278, April 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042799003222.

vanderHouwen:2000:OSC

[vdHMS00] P. J. van der Houwen, E. Messina, and B. P. Sommeijer. Oscillatory Störmer– Cowell methods. Journal of Computational and Applied Mathematics, 115(1– 2):547–564, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704279900179X.

vanderHouwen:2001:AFT

[vdHS01] P. J. van der Houwen and B. P. Sommeijer. Approximate factorization for time-dependent partial differential equations. Journal of Computational and Applied Mathematics, 128(1–2):447–466, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005239.

vanderVorst:2002:ERI

[vdV02] Henk A. van der Vorst. Efficient and reliable iterative methods for linear systems. Journal of Computational and Applied Mathematics, 149(1):251–265, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042702005344.

VandenBerghe:2000:EFR

[VDVV00] G. Vanden Berghe, H. De Meyer, M. Van Daele, and T. Van Hecke. Exponentially fitted Runge–Kutta methods. Journal of Computational and Applied Mathematics, 125 (1–2):107–115, December 15, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004623.

Vecchio:2000:SBD

[Vec00]  Antonia Vecchio.  Stabil-

ity of backward differentiation formulas for Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 115(1–2):565–576, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001235.

Veling:2001:AES

[Vel01] E. J. M. Veling. Analytical evaluation of some Bessel function integrals related to Sonine’s finite integrals. Journal of Computational and Applied Mathematics, 134 (1–2):201–212, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005379.

Verlinden:2001:SSH

[Ver01] Pierre Verlinden. The sparse structure of high-dimensional integrands. Journal of Computational and Applied Mathematics, 132(1):33–49, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000604X.

vonGolitschek:2000:NPS

[vG00] M. von Golitschek. Norms of projectors onto spaces with Riesz bases. Journal of Computational and Applied Mathematics, 119(1–2):209–221, July 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003794.

VandenBerghe:2001:FDS

[VID01] G. Vanden Berghe, L. Gr. Ixaru, and H. De Meyer. Frequency determination and step-length control for exponentiallyfitted Runge–Kutta methods. Journal of Computational and Applied Mathematics, 132(1):95–105, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042700006026.

Vidunas:2003:CRH

[Vid03] Raimundas Vidunas. Contiguous relations of hypergeometric series. Journal of Computational and Applied Mathematics, 153(1–2):507–519, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200643X.

Villarroel:2003:SIS

[Vil03] Javier Villarroel. On solutions to Ito stochastic differential equations. Journal of Computational and Applied Mathematics, 158 (1):225–231, September 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004771.

Vierendeels:2004:MMN

[VMD04] Jan Vierendeels, Bart Merci, and Erik Dick. A multigrid method for natural convective heat transfer with large temperature differences. Journal of Computational and Applied Mathematics, 168(1–2):509–517, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042703010070.

VanCamp:2004:TFA

[VMV04] Ellen Van Camp, Nicola Mastronardi, and Marc Van Barel. Two fast algorithms for solving diagonal-plus-semiseparable linear systems. Journal of Computational and Applied Mathematics, 164– 165(??):731–747, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008379.

Vourdas:2001:QSF

[Vou01] A. Vourdas. Quantum systems with finite Hilbert space and Chebyshev polynomials. Journal of Computational and Applied Mathematics, 133(1–2):657–664, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006828.

Verwer:2001:NSM

[VS01] J. G. Verwer and B. P. Sommeijer. A numerical study of mixed parabolic-gradient systems. Journal of Computational and Applied Mathematics, 132(1):191–210, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042700006051.

Vermolen:2000:MMD

[VV00] F. J. Vermolen and C. Vuik.

A mathematical model for the dissolution of particles in multi-component alloys. Journal of Computational and Applied Mathematics, 126 (1–2):233–254, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003556.

VanHecke:2001:HOC

[VV01] T. Van Hecke and M. Van Daele. High-order convergent deferred correction schemes based on parameterized Runge–Kutta–Nyström methods for second-order boundary value problems. Journal of Computational and Applied Mathematics, 132 (1):107–125, July 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700006038.

Voorde:2004:FSR

[VVD04] John Vande Voorde, Jan Vierendeels, and Erik Dick. Flow simulations in rotary volumetric pumps and compressors with the fictitious domain method. Journal of Computational and Applied Mathematics, 168(1–2):491–499, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703010057.

VanAssche:2003:KSR

[VVV03a] W. Van Assche, G. Vanden Berghe, and J. Van der Jeugt. K. Srinivasa Rao and his work. Journal of Computational and Applied Mathematics, 160(1–2):1–8, November 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703006083.

VandenBerghe:2003:EFR

[VVV03b] G. Vanden Berghe, M. Van Daele, and H. Vande Vyver. Exponential fitted Runge– Kutta methods of collocation type: fixed or variable knot points? Journal of Computational and Applied Mathematics, 159(2):217–239, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004503.

Vaassen:2004:ATO

[VWE04] J.-M. Vaassen, P. Wautelet, and J.-A. Essers. Application of a third-order reconstruction scheme to hypersonic reacting flows using unstructured meshes and multigrid techniques. Journal of Computational and Applied Mathematics, 168(1–2):481–489, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009993.

Vinet:2001:MSK

[VYZ01] Luc Vinet, Oksana Yermolayeva, and Alexei Zhedanov. A method to study the Krall and q-Krall polynomials. Journal of Computational and Applied Mathematics, 133(1–2):647–656, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700006865.

Vinet:2001:STL

[VZ01] Luc Vinet and Alexei Zhedanov.

Spectral transformations of the Laurent biorthogonal polynomials. I. q-Appel polynomials. Journal of Computational and Applied Mathematics, 131(1–2):253–266, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002703.

Vinet:2004:CCS

[VZ04] Luc Vinet and Alexei Zhedanov.

A characterization of classical and semiclassical orthogonal polynomials from their dual polynomials. Journal of Computational and Applied Mathematics, 172(1): 41–48, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704000962.

Wong:2000:ETD

[WA00a] Patricia J. Y. Wong and Ravi P. Agarwal. Eigenvalue theorems for discrete multipoint conjugate boundary value problems. Journal of Computational and Applied Mathematics, 113 (1–2):227–240, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002587.

Wong:2000:FSS

[WA00b] Patricia J. Y. Wong and Ravi P. Agarwal. Fixedsign solutions of a system of higher order difference equations. Journal of Computational and Applied Mathematics, 113(1–2):167–181, January 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002514.

Wolberg:2002:EMF

[WA02] George Wolberg and Itzik Alfy. An energy-minimization framework for monotonic cubic spline interpolation. Journal of Computational and Applied Mathematics, 143(2):145–188, June 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005064.

Walz:2000:AEM

[Wal00] Guido Walz. Asymptotic expansions for multivariate polynomial approximation. Journal of Computational and Applied Mathematics, 122 (1–2):317–328, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003587.

Waldschmidt:2003:AVA

[Wal03] Michel Waldschmidt. Algebraic values of analytic functions. Journal of Computational and Applied Mathematics, 160(1–2):323–333, November 1, 2003. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S037704270300637X.

Wang:2000:MSA

[Wan00a] Ren-Hong Wang. Multivariate spline and algebraic geometry. Journal of Computational and Applied Mathematics, 121 (1–2):153–163, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003447.

Wang:2000:IRS

[Wan00b] Xiaoqun Wang. Improving the rejection sampling method in quasi-Monte Carlo methods. Journal of Computational and Applied Mathematics, 114(2):231–246, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001946.

Wang:2001:GIC

[Wan01a] Xiaoqun Wang. On generalized invariant cubature formulae. Journal of Computational and Applied Mathematics, 130(1–2):271–281, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003775.

Wang:2001:VRT

[Wan01b]  Xiaoqun Wang.  Variance

reduction techniques and quasi-Monte Carlo methods. Journal of Computational and Applied Mathematics, 132(2):309–318, July 15, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003319.

Wang:2004:SSI

[Wan04a] Shaoming Wang. A smooth surface interpolation to 3D triangulations. Journal of Computational and Applied Mathematics, 163(1):287–293, February 1, 2004. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S037704270300815X.

Wang:2004:MFV

[Wan04b] Tongke Wang. A mixed finite volume element method based on rectangular mesh for biharmonic equations. Journal of Computational and Applied Mathematics, 172 (1):117–130, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704001116.

Watkins:2000:LAE

[Wat00a] David S. Watkins. QR-like algorithms for eigenvalue problems. Journal of Computational and Applied Mathematics, 123(1–2):67–83, November 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004027.

Watson:2000:ANL

[Wat00b]  G. A. Watson.  Approx-

imation in normed linear spaces. Journal of Computational and Applied Mathematics, 121(1–2):1–36, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700003332.

Watson:2002:POH

[Wat02] Layne T. Watson. Probabilityone homotopies in computational science. Journal of Computational and Applied Mathematics, 140(1– 2):785–807, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701004733.

Watanabe:2003:IVF

[Wat03] Jiro Watanabe. On incompressible viscous fluid flows with slip boundary conditions. Journal of Computational and Applied Mathematics, 159(1):161–172, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703005685.

Wazwaz:2001:NSF

[Waz01] Abdul-Majid Wazwaz. The numerical solution of fifthorder boundary value problems by the decomposition method. Journal of Computational and Applied Mathematics, 136 (1–2):259–270, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000618X.

Watson:2000:NAV

[WBBF00] Layne T. Watson, Michael Bartholomew-Biggs, and John Ford. Numerical analysis 2000: Vol. IV: Optimization and nonlinear equations. Journal of Computational and Applied Mathematics, 124(1–2):ix–x, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004167.

Wang:2004:CHP

[WdZ04] Ren-Hong Wang and Cheng de Zheng. Cubic Hermite– Padé approximation to the exponential function. Journal of Computational and Applied Mathematics, 163 (1):259–268, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042703008124.

Wegmann:2001:CSC

[Weg01a] Rudolf Wegmann. Constructive solution of a certain class of Riemann–Hilbert problems on multiply connected circular regions. Journal of Computational and Applied Mathematics, 130(1–2):139–161, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003866.

Wegmann:2001:FCM

[Weg01b] Rudolf Wegmann. Fast conformal mapping of multiply connected regions. Journal of Computational and Applied Mathematics, 130(1– 2):119–138, May 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042799003878.

Weniger:2000:PPA

[Wen00] Ernst Joachim Weniger. Prediction properties of Aitken’s iterated Δ^{2} process, of Wynn’s epsilon algorithm, and of Brezinski’s iterated theta algorithm. Journal of Computational and Applied Mathematics, 122(1–2):329–356, October 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003630.

Wulf:2000:NHB

[WF00] Volker Wulf and Neville J.

Ford. Numerical Hopf bifurcation for a class of delay differential equations. Journal of Computational and Applied Mathematics, 115(1–2):601–616, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001818.

Wilders:2002:PSA

[WF02] P. Wilders and G. Fotia. A positive spatial advection scheme on unstructured meshes for tracer transport. Journal of Computational and Applied Mathematics, 140(1–2):809–821, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004861.

Windcliff:2001:SOF

[WFV01] H. Windcliff, P. A. Forsyth, and K. R. Vetzal. Shout options: a framework for pricing contracts which can be modified by the investor. Journal of Computational and Applied Mathematics, 134 (1–2):213–241, September 1, 2001. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005513.

Wei:2002:CFA

[WG02] G. W. Wei and Yun Gu. Conjugate filter approach for solving Burgers’ equation. Journal of Computational and Applied Mathematics, 149 (2):439–456, December 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004880.

Wang:2004:VVT

[WG04] Jinbo Wang and Chuanqing Gu. Vector valued Thiele– Werner-type osculatory rational interpolants. Journal of Computational and Applied Mathematics, 163 (1):241–252, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008100.

Weber:2000:NSV

[WH00] Régine Weber and Jacques Hureau. Numerical solution for various inverse problems. Journal of Computational and Applied Mathematics, 115(1–2):577–591, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001879.

Wang:2001:NCR

[WH01] Chuan-Long Wang and Ting-Zhu Huang. New convergence results for alternating methods. Journal of Computational and Applied Mathematics, 135(2):325–333, October 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005914.

Wang:2004:CRB

[WH04] Zhuan-De Wang and Ting-Zhu Huang. Comparison results between Jacobi and other iterative methods. Journal of Computational and Applied Mathematics, 169(1):45–51, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009373.

Winkler:2003:SDA

[Win03] Renate Winkler. Stochastic differential algebraic equations of index 1 and applications in circuit simulation. Journal of Computational and Applied Mathematics, 157(2):477–505, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703004369. See er-

ratum [Win04a].

Winkler:2004:ESD

[Win04a] Renate Winkler. Erratum to: “Stochastic differential algebraic equations of index 1 and applications in circuit simulation”: [J. Comput. Appl. Math. 157 (2003) 477–505]. Journal of Computational and Applied Mathematics, 163(2):433, February 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703009427.  See

[Win03].

Winkler:2004:SDA

[Win04b] Renate Winkler. Stochastic differential algebraic equations of index 1 and applications in circuit simulation. Journal of Computational and Applied Mathematics, 163(2):435–463, February 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703010239.

Woznicki:2001:NCM

[WJ01] Zbigniew I. Woźnicki and Henryk A. Jedrzejec. A new class of modified line-SOR algorithms. Journal of Computational and Applied Mathematics, 131(1–2):89–142, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003198.

Wang:2004:CAS

[WKM04] Rong Wang, Patrick Keast, and Paul Muir. A comparison of adaptive software for 1D parabolic PDEs. Journal of Computational and Applied Mathematics, 169(1):127–150, August 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703009567.

Watanabe:2003:CMA

[WKS^{+}03] Masaji Watanabe, Fusako Kawai, Masaru Shibata, Shigeo Yokoyama, and Yasuhiro Sudate. Computational method for analysis of polyethylene biodegradation. Journal of Computational and Applied Mathematics, 161(1):133–144, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270300551X.

Wang:2001:NAD

[WL01] Jin Wang and Ruxun Liu.

A new approach to design high-order schemes. Journal of Computational and Applied Mathematics, 134 (1–2):59–67, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005288.

Wang:2002:PAC

[WL02a] Ren-Hong Wang and Yi-Sheng Lai. Piecewise algebraic curve. Journal of Computational and Applied Mathematics, 144(1–2):277–289, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005672.

Wang:2002:ACE

[WL02b] Song Wang and Zi-Cai Li. An analysis of a conforming exponentially fitted finite element method for a convection–diffusion problem. Journal of Computational and Applied Mathematics, 143(2):291–310, June 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701005301.

Wang:2003:FRO

[WL03] Peiguang Wang and Xuewen Li. Further results on oscillation of a class of second-order neutral equations. Journal of Computational and Applied Mathematics, 157(2):407–418, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004199.

Wang:2004:PSP

[WL04] Lin-Lin Wang and Wan-Tong Li. Periodic solutions and permanence for a delayed nonautonomous ratio-dependent predator–prey model with Holling type functional response. Journal of Computational and Applied Mathematics, 162 (2):341–357, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007404.

Wang:2004:RRC

[WLYL04] Ruomei Wang, Yi Li, Fang You, and Xiaonan Luo. Rational recurrence curves and recurrence surfaces in multivariate B-form on some regions. Journal of Computational and Applied Mathematics, 163(1):277–285, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008148.

Walton:2001:CEP

[WM01] D. J. Walton and D. S. Meek. Curvature extrema of planar parametric polynomial cubic curves. Journal of Computational and Applied Mathematics, 134 (1–2):69–83, September 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270000529X.

Walton:2002:PTF

[WM02] D. J. Walton and D. S. Meek. Planar G^{2} transition with a fair Pythagorean hodograph quintic curve. Journal of Computational and Applied Mathematics, 138(1):109–126, January 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701003594.

Walton:2004:GPH

[WM04a] D. J. Walton and D. S. Meek. A generalisation of the Pythagorean hodograph quintic spiral. Journal of Computational and Applied Mathematics, 172(2):271–287, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001190.

Wang:2004:Pb

[WM04b] Song Wang and John J. H.

Miller. Preface. Journal of Computational and Applied Mathematics, 166(1):ix, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008501.

Walton:2003:PTC

[WMA03] D. J. Walton, D. S. Meek, and J. M. Ali. Planar G^{2} transition curves composed of cubic Bézier spiral segments. Journal of Computational and Applied Mathematics, 157(2):453–476, August 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004357.

Wesseling:2001:GMA

[WO01] P. Wesseling and C. W.

Oosterlee. Geometric multigrid with applications to computational fluid dynamics. Journal of Computational and Applied Mathematics, 128(1–2):311–334, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005173.

Wolfe:2000:IMA

[Wol00a] M. A. Wolfe. Interval mathematics, algebraic equations and optimization. Journal of Computational and Applied Mathematics, 124 (1–2):263–280, December 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004210.

Wollman:2000:PMV

[Wol00b] Stephen Wollman. The particle method for the Vlasov– Poisson system using equally spaced initial data points. Journal of Computational and Applied Mathematics, 115(1–2):593–600, March 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001259.

Wong:2002:AGB

[Won02a] Patricia J. Y. Wong. Abel– Gontscharoff boundary value problems on measure chains. Journal of Computational and Applied Mathematics, 142 (2):331–355, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701004290.

Wong:2002:OAG

[Won02b] Patricia J. Y. Wong. Optimal Abel–Gontscharoff interpolation error bounds on measure chains. Journal of Computational and Applied Mathematics, 141(1–2):267–282, April 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004526.

Wong:2004:AGI

[Won04] Patricia J. Y. Wong. Abel– Gontscharoff interpolation: continuous, discrete and time scale. Journal of Computational and Applied Mathematics, 164–165(??):763– 782, March 1, 2004. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042703004941.

Woznicki:2001:PSM

[Woź01] Zbigniew I. Woźnicki. On performance of SOR method for solving nonsymmetric linear systems. Journal of Computational and Applied Mathematics, 137(1):145–176, December 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007056.

Wang:2002:PEA

[WPS02] J. F. Wang, J. Periaux, and M. Sefrioui. Parallel evolutionary algorithms for optimization problems in aerospace engineering. Journal of Computational and Applied Mathematics, 149 (1):155–169, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005277.

Wright:2002:RMA

[Wri02] R. K. Wright. A robust method for accurately representing nonperiodic functions given Fourier coefficient information. Journal of Computational and Applied Mathematics, 140(1–2):837–848, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005180.

Wright:2004:LSA

[Wri04] R. K. Wright. Local spline approximation of discontinuous functions and location of discontinuities, given low-order Fourier coefficient information. Journal of Computational and Applied Mathematics, 164– 165(??):783–795, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703006472.

Woodham:2002:CIP

[WS02a] C. A. Woodham and H. Su.

A computational investigation of pole-zero cancellation for a double inverted pendulum. Journal of Computational and Applied Mathematics, 140(1–2):823–836, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701004770.

Wu:2002:HBN

[WS02b] Wei Wu and Yi Su. Hopf bifurcation near a double singular point with Z_{2}-symmetry and X_{0}-breaking. Journal of Computational and Applied Mathematics, 144(1–2):325–334, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005702.

Wright:2001:CLC

[WT01] Thomas G. Wright and Lloyd N.

Trefethen. Computing Lyapunov constants for random recurrences with smooth coefficients. Journal of Computational and Applied Mathematics, 132(2):331–340, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700004374.

Wang:2004:Pa

[WT04] R. H. Wang and K. L. Teo. Preface. Journal of Computational and Applied Mathematics, 163(1):xi–xii, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008215.

Wunsche:2001:HLP

[Wün01a] A. Wünsche. Hermite and Laguerre 2D polynomials. Journal of Computational and Applied Mathematics, 133 (1–2):665–678, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700006816.

Wunsche:2001:UEF

[Wün01b] A. Wünsche. An unorthodox entire function. Journal of Computational and Applied Mathematics, 133 (1–2):700–701, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700007287.

Wunsche:2003:RJP

[Wün03a]  A. Wünsche.  The role

of Jacobi polynomials in the theory of Hermite and Laguerre 2D polynomials. Journal of Computational and Applied Mathematics, 153(1–2):521–529, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702006325.

Wunsche:2003:SGG

[Wün03b] Alfred Wünsche. A sequence of generalizations of the geometric series. Journal of Computational and Applied Mathematics, 153(1–2):533–534, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008828.

Windmolders:2003:UPS

[WVDB03] Joris Windmolders, Evelyne Vanraes, Paul Dierckx, and Adhemar Bultheel. Uniform Powell–Sabin spline wavelets. Journal of Computational and Applied Mathematics, 154(1):125–142, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008178.

Wei:2000:CPK

[WW00a] Yimin Wei and Hebing Wu. Convergence properties of Krylov subspace methods for singular linear systems with arbitrary index. Journal of Computational and Applied Mathematics, 114 (2):305–318, February 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799902376.

Wei:2000:RAD

[WW00b] Yimin Wei and Hebing Wu. The representation and approximation for Drazin inverse. Journal of Computational and Applied Mathematics, 126 (1–2):417–432, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003696.

Wuytack:2000:NAC

[WW00c] Luc Wuytack and Jet Wimp. Numerical analysis in the 20th century vol. I: Approximation theory. Journal of Computational and Applied Mathematics, 121(1–2):ix–x, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700003903.

Wei:2001:PSS

[WW01] Yimin Wei and Hebing Wu.

On the perturbation and subproper splittings for the generalized inverse A_{T,S}^{(2)} of rectangular matrix A. Journal of Computational and Applied Mathematics, 137 (2):317–329, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270000710X.

Wang:2002:P

[WW02] Renhong Wang and Roderick Wong. Preface. Journal of Computational and Applied Mathematics, 144(1–2):ix, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005477.

Wang:2003:OCS

[WW03] Peiguang Wang and Yonghong Wu. Oscillation of certain second-order functional differential equations with damping. Journal of Computational and Applied Mathematics, 157(1):49–56, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003741.

Wang:2004:QII

[WW04] Ren-Hong Wang and Jing-Xin Wang. Quasi-interpolations with interpolation property. Journal of Computational and Applied Mathematics, 163(1):253–257, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008112.

Wiwatanapataphee:2004:NST

[WWA^{+}04] B. Wiwatanapataphee, Y. H. Wu, J. Archapitak, P. F. Siew, and B. Unyong. A numerical study of the turbulent flow of molten steel in a domain with a phase-change boundary. Journal of Computational and Applied Mathematics, 166(1):307–319, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008537.

Wang:2002:MWS

[WX02a] Ren-Hong Wang and Zhi-Qiang Xu. Multivariate weak spline function space. Journal of Computational and Applied Mathematics, 144(1–2):291–299, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042701005684.

Wu:2002:DCO

[WX02b] Wei Wu and Yuesheng Xu.

Deterministic convergence of an online gradient method for neural networks. Journal of Computational and Applied Mathematics, 144(1– 2):335–347, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042701005714.

Wan-Xie:2004:PIM

[WX04] Zhong Wan-Xie. On precise integration method. Journal of Computational and Applied Mathematics, 163(1): 59–78, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270300791X.

Wong:2002:IBL

[WY02] R. Wong and Heping Yang.

On an internal boundary layer problem. Journal of Computational and Applied Mathematics, 144(1–2):301–323, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005696.

Wei:2003:CNR

[WZ03] Yimin Wei and Naimin Zhang. Condition number related with generalized inverse A_{T,S}^{(2)} and constrained linear systems. Journal of Computational and Applied Mathematics, 157(1):57–72, August 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703003753.

Wang:2004:HBB

[WZ04a] Lin Wang and Xingfu Zou. Hopf bifurcation in bidirectional associative memory neural networks with delays: analysis and computation. Journal of Computational and Applied Mathematics, 167(1):73–90, May 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008793.

Wang:2004:CBT

[WZ04b] Ren-Hong Wang and Chun-Gang Zhu. Cayley–Bacharach Theorem of piecewise algebraic curves. Journal of Computational and Applied Mathematics, 163(1):269–276, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042703008136.

Wang:2003:NNS

[WZW03] Y. Wang, Y. B. Zhao, and G. W. Wei. A note on the numerical solution of high-order differential equations. Journal of Computational and Applied Mathematics, 159(2):387–398, October 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005417.

Xie:2002:EFO

[XCZ02] Zhongqiang Xie, Chi-Hou Chan, and Bo Zhang. An explicit fourth-order staggered finite-difference time-domain method for Maxwell’s equations. Journal of Computational and Applied Mathematics, 147(1):75–98, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003941.

Xiaofang:2001:SFQ

[XKC01] Ma Xiaofang, Fung Richard Ying Kit, and Xu Chengxian. Stable factorized quasi-Newton methods for nonlinear least-squares problems. Journal of Computational and Applied Mathematics, 129(1–2):1–14, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005392.

Xu:2002:CSO

[XL02] Jiuping Xu and Jun Li. A class of stochastic optimization problems with one quadratic & several linear objective functions and extended portfolio selection model. Journal of Computational and Applied Mathematics, 146(1):99–113, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004211.

Xiao:2000:ECB

[XLFC00] Aiguo Xiao, Shoufu Li, Hongyuan Fu, and Guangnan Chen. Extending convergence of BDF methods for a class of nonlinear strongly stiff problems. Journal of Computational and Applied Mathematics, 126(1– 2):121–130, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799003465.

Xu:2001:MSC

[Xu01a] Jinchao Xu. The method of subspace corrections. Journal of Computational and Applied Mathematics, 128(1– 2):335–362, March 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005185.

Xu:2001:OPC

[Xu01b] Yuan Xu. Orthogonal polynomials and cubature formulae on balls, simplices, and spheres. Journal of Computational and Applied Mathematics, 127(1–2):349–368, January 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005045.

Xu:2002:HGO

[Xu02] Peiliang Xu. A hybrid global optimization method: the one-dimensional case. Journal of Computational and Applied Mathematics, 147 (2):301–314, October 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004387.

Xu:2003:HGO

[Xu03a]  Peiliang Xu.  A hybrid

global optimization method: the multi-dimensional case. Journal of Computational and Applied Mathematics, 155(2):423–446, June 15, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008786.

Xu:2003:NSB

[Xu03b] Peiliang Xu. Numerical solution for bounding feasible point sets. Journal of Computational and Applied Mathematics, 156(1):201–219, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702009123.

Xu:2004:SFF

[Xu04] Guoliang Xu. Surface fairing and featuring by mean curvature motions. Journal of Computational and Applied Mathematics, 163 (1):295–309, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008161.

Xiu:2001:SGN

[XZ01] Naihua Xiu and Jianzhong Zhang. A smoothing Gauss– Newton method for the generalized HLCP. Journal of Computational and Applied Mathematics, 129(1– 2):195–208, April 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042700005501.

Xiu:2003:SRA

[XZ03] Naihua Xiu and Jianzhong Zhang. Some recent advances in projection-type methods for variational inequalities. Journal of Computational and Applied Mathematics, 152(1–2):559–585, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007306.

Zhao:2004:SPT

[xZqZqT04] Huan xi Zhao, Gong qin Zhu, and Jie qing Tan. A Śleszyński–Pringsheim theorem for vector valued continued fractions and its optimal error bounds. Journal of Computational and Applied Mathematics, 163 (1):343–350, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008203.

Zhao:2002:BTT

[xZqZX02] Huan xi Zhao, Gong qin Zhu, and Ping Xiao. A backward three-term recurrence relation for vector valued continued fractions and its applications. Journal of Computational and Applied Mathematics, 142(2):389–400, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042701004198.

Yakubovich:2000:TIT

[Yak00] S. B. Yakubovich. The Titchmarsh integral transformation by the index of a Bessel function. Journal of Computational and Applied Mathematics, 118(1–2):353–361, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700002983.

Yakubovich:2002:ITA

[Yak02] Semyon B. Yakubovich. Index transforms associated with products of Whittaker’s functions. Journal of Computational and Applied Mathematics, 148(2):419–427, November 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005599.

Yamamoto:2000:HDC

[Yam00]  Tetsuro Yamamoto.  His-

torical developments in convergence analysis for Newton’s and Newton-like methods. Journal of Computational and Applied Mathematics, 124(1–2):1–23, December 1, 2000. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700004179.

Yamamoto:2002:CCI

[Yam02] Tetsuro Yamamoto. Convergence of consistent and inconsistent finite difference schemes and an acceleration technique. Journal of Computational and Applied Mathematics, 140(1–2):849–866, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005222.

Yang:2000:IEE

[Yan00] Daoqi Yang. Improved error estimation of dynamic finite element methods for secondorder parabolic equations. Journal of Computational and Applied Mathematics, 126 (1–2):319–338, December 30, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003623.

Yang:2002:APL

[Yan02]  Suh-Yuh Yang.  Analy-

sis of piecewise linear approximations to the generalized Stokes problem in the velocity-stress-pressure formulation. Journal of Computational and Applied Mathematics, 147(1):53–73, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702003928.

Yang:2003:GCC

[Yan03]  Zheng-Hong Yang.  Gen-

eralized confluent Cauchy– Vandermonde matrices: displacement structures, inversion formulas and tangential interpolations. Journal of Computational and Applied Mathematics, 154 (2):355–371, May 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042702008580.

Yuan:2003:CPC

[YB03] Dongjin Yuan and Kevin Burrage. Convergence of the parallel chaotic waveform relaxation method for stiff systems. Journal of Computational and Applied Mathematics, 151(1):201–213, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007495.

Chen:2001:RRM

[yCjC01] Xian yao Chen and Chang jun Cheng. Reconstruction of relaxation modulus for viscoelastic medium with complete data. Journal of Computational and Applied Mathematics, 131(1–2):445–456, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700002570.

Yang:2003:NCN

[YD03] Jian Yang and Hong-Ke Du.

A note on compact normal operators. Journal of Computational and Applied Mathematics, 151(1):229–233, February 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702007483.

Ye:2003:LCM

[Ye03]  Xingde Ye.  The Legen-

dre collocation method for the Cahn–Hilliard equation. Journal of Computational and Applied Mathematics, 150(1):87–108, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005666.

Yeh:2004:SPR

[Yeh04] Li-Chin Yeh. A singularly perturbed reaction–diffusion problem with inhomogeneous environment. Journal of Computational and Applied Mathematics, 166(1):321–341, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008525.

Yano:2003:CMM

[YEWE03] Tadashi Yano, Yasuo Ezawa, Takeshi Wada, and Hiroshi Ezawa. Can Milne’s method work well for the Coulomblike potentials? Journal of Computational and Applied Mathematics, 152(1– 2):597–611, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270200732X.

Yamamura:2003:FAS

[YF03] Kiyotaka Yamamura and Tsuyoshi Fujioka. Finding all solutions of nonlinear equations using the dual simplex method. Journal of Computational and Applied Mathematics, 152(1–2):587–595, March 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702007318.

Yang:2002:DSA

[YH02] Zheng-Hong Yang and Yong-Jian Hu. Displacement structure approach to Cauchy and Cauchy–Vandermonde matrices: inversion formulas and fast algorithms. Journal of Computational and Applied Mathematics, 138 (2):259–272, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701003788.

Yamamoto:2003:EEG

[YH03] Nobito Yamamoto and Keisuke Hayakawa. Error estimation with guaranteed accuracy of finite element method in nonconvex polygonal domains. Journal of Computational and Applied Mathematics, 159(1):173–183, October 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005697.

Yun:2004:CTS

[YK04] Jae Heon Yun and Sang Wook Kim. Convergence of twostage iterative methods using incomplete factorization. Journal of Computational and Applied Mathematics, 166(2):565–580, April 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008744.

Yang:2003:NPS

[YL03] Zuodong Yang and Qishao Lu. Nonexistence of positive solutions to a quasilinear elliptic system and blowup estimates for a quasilinear reaction–diffusion system. Journal of Computational and Applied Mathematics, 150(1):37–56, January 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702005630.

Yu:2004:NSG

[YLC04] K. F. Yu, K. S. Lau, and C. K. Chan. Numerical simulation of gas-particle flow in a single-side backward-facing step flow. Journal of Computational and Applied Mathematics, 163(1):319–331, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008185.

Yang:2002:LPG

[YLD02] Yuansheng Yang, Jianhua Lin, and Yongjie Dai. Largest planar graphs and largest maximal planar graphs of diameter two. Journal of Computational and Applied Mathematics, 144(1–2):349–358, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005726.

Yamatani:2002:IME

[YOO02] Katsu Yamatani, Takashi Ohe, and Kohzaburo Ohnaka. An identification method of electric current dipoles in spherically symmetric conductor. Journal of Computational and Applied Mathematics, 143(2):189–200, June 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005076.

Yoon:2003:NSA

[Yoo03]  Jungho Yoon.  A non-

stationary approximation scheme on scattered centers in R^{d} by radial basis functions. Journal of Computational and Applied Mathematics, 155 (1):163–175, June 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702008981.

Yabe:2002:NGC

[YOT^{+}02] T. Yabe, Y. Ogata, K. Takizawa, T. Kawai, A. Segawa, and K. Sakurai. The next generation CIP as a conservative semi-Lagrangian solver for solid, liquid and gas. Journal of Computational and Applied Mathematics, 149 (1):267–277, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005356.

Yoshimura:2002:AGP

[YSNM02] Shinobu Yoshimura, Ryuji Shioya, Hirohisa Noguchi, and Tomoshi Miyamura. Advanced general-purpose computational mechanics system for large-scale analysis and design. Journal of Computational and Applied Mathematics, 149(1):279–296, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702005368.

Yamada:2002:IAM

[YTI02] Syuuji Yamada, Tetsuzo Tanino, and Masahiro Inuiguchi. An inner approximation method incorporating with a penalty function method for a reverse convex programming problem. Journal of Computational and Applied Mathematics, 146(1):57–75, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004181.

Yu:2002:CSN

[Yu02] P. Yu. Computation of the simplest normal forms with perturbation parameters based on Lie transform and rescaling. Journal of Computational and Applied Mathematics, 144(1–2):359–373, July 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005738.

Yang:2002:SLP

[YY02] Shi-Nine Yang and Yung-Jen Yang. Step length problem for trimming curve approximation in tessellating trimmed surfaces. Journal of Computational and Applied Mathematics, 140(1–2):867–876, March 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042701005271.

Yang:2002:EPF

[YYT02] Xin Min Yang, Xiao Qi Yang, and Kok Lay Teo. Explicitly B-preinvex functions. Journal of Computational and Applied Mathematics, 146(1):25–36, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702004156.

Yang:2004:LES

[YZCL04] W. B. Yang, H. Q. Zhang, C. K. Chan, and W. Y. Lin. Large eddy simulation of mixing layer. Journal of Computational and Applied Mathematics, 163(1):311–318, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008173.

Zhong:2002:ODO

[ZD02] Ping Zhong and Naiyang Deng. On a one-dimensional optimization problem derived from the efficiency analysis of Newton–PCG-like algorithms. Journal of Computational and Applied Mathematics, 146(1):11–24, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004144.

Zegeling:2004:TPA

[Zeg04]  P. A. Zegeling.  Tensor-

product adaptive grids based on coordinate transformations. Journal of Computational and Applied Mathematics, 166(1):343–360, April 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703008513.

Zhongxuan:2004:EEA

[ZFYY04] Luo Zhongxuan, Li Fengzhi, Liu Yingxi, and Li Yi. Effect of the environmental atmosphere on heat, water and gas transfer within hygroscopic fabrics. Journal of Computational and Applied Mathematics, 163(1):199–210, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008069.

Zhang:2004:PBA

[ZH04] Zeyin Zhang and Daren Huang. Parameterization of m band affine frames with pre-assigned scaling filter. Journal of Computational and Applied Mathematics, 163(1):333–342, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008197.

Zhang:2002:CPT

[Zha02] Shao-Liang Zhang. A class of product-type Krylov-subspace methods for solving nonsymmetric linear systems. Journal of Computational and Applied Mathematics, 149 (1):297–305, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704270200537X.

Zhao:2003:UAE

[Zha03] Yu-Qiu Zhao. A uniform asymptotic expansion of the single variable Bell polynomials. Journal of Computational and Applied Mathematics, 150(2):329–355, January 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702006696.

Zhedanov:2004:BLI

[Zhe04] Alexei Zhedanov. Biorthogonality of the Lagrange interpolants. Journal of Computational and Applied Mathematics, 172(1):1–6, November 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704000949.

Zhou:2001:MPA

[Zho01]  Ping Zhou.  Multivariate

Padé approximants to a meromorphic function. Journal of Computational and Applied Mathematics, 136 (1–2):37–51, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042700005744.

Zhu:2000:NPA

[Zhu00]  Detong Zhu.  Nonmono-

tonic projected algorithm with both trust region and line search for constrained optimization. Journal of Computational and Applied Mathematics, 117(1):35–60, May 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003271.

Zhu:2001:FIS

[Zhu01]  Detong Zhu.  A family

of improved secant methods via nonmonotone curvilinear paths technique for equality constrained optimization. Journal of Computational and Applied Mathematics, 136(1–2):73–97, November 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700005768.

Zhu:2003:NAS

[Zhu03a] Detong Zhu. A new affine scaling interior point algorithm for nonlinear optimization subject to linear equality and inequality constraints. Journal of Computational and Applied Mathematics, 161(1):1–25, December 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703004588.

Zhu:2003:NBT

[Zhu03b] Detong Zhu. Nonmonotonic back-tracking trust region interior point algorithm for linear constrained optimization. Journal of Computational and Applied Mathematics, 155(2):285–305, June 15, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008701.

Zegeling:2004:AMM

[ZK04] P. A. Zegeling and H. P. Kok. Adaptive moving mesh computations for reaction–diffusion systems. Journal of Computational and Applied Mathematics, 168(1–2):519–528, July 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042703010069.

Zhang:2002:CRF

[ZKO02] Jun Zhang, Jules Kouatchou, and Mohamed Othman. On cyclic reduction and finite difference schemes. Journal of Computational and Applied Mathematics, 145(1): 213–222, August 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042701005763.

Zhang:2003:HCL

[ZL03] Peng Zhang and Ru-Xun Liu. Hyperbolic conservation laws with space-dependent flux: I. Characteristics theory and Riemann problem. Journal of Computational and Applied Mathematics, 156(1):1–21, July 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702008804.

Zhang:2004:SLS

[ZL04] Zhengce Zhang and Kaitai Li. Spike-layered solutions with compact support to some singularly perturbed quasilinear elliptic problems in general smooth domains. Journal of Computational and Applied Mathematics, 162 (2):327–340, January 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703007398.

Zeng:2001:WMN

[ZLF01] Jin-Ping Zeng, Dong-Hui Li, and Masao Fukushima. Weighted max-norm estimate of additive Schwarz iteration scheme for solving linear complementarity problems. Journal of Computational and Applied Mathematics, 131(1–2):1–14, June 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042700003277.

Zou:2002:DIT

[Zou02] Xingfu Zou. Delay induced traveling wave fronts in reaction diffusion equations of KPP–Fisher type. Journal of Computational and Applied Mathematics, 146 (2):309–321, September 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702003631.

Zube:2004:NSS

[Zub04] Severinas Zube. Number systems, α-splines and refinement. Journal of Computational and Applied Mathematics, 172(2):207–231, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042704001153.

Zhang:2004:SAV

[ZV04] Chengjian Zhang and Stefan Vandewalle. Stability analysis of Volterra delayintegro-differential equations and their backward differentiation time discretization. Journal of Computational and Applied Mathematics, 164– 165(??):797–814, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703008331.

Zhang:2002:ULS

[ZW02] Zhongxin Zhang and Junyu Wang. The upper and lower solution method for a class of singular nonlinear second order three-point boundary value problems. Journal of Computational and Applied Mathematics, 147(1):41–52, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042702003904.

Zhang:2001:PNP

[ZX01] Jianzhong Zhang and Chengxian Xu. Properties and numerical performance of quasi-Newton methods with modified quasi-Newton equations. Journal of Computational and Applied Mathematics, 137 (2):269–278, December 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042700007135.

Zeng:2002:DDM

[ZZ02a] J.-P. Zeng and S. Z. Zhou.

A domain decomposition method for a kind of optimization problems. Journal of Computational and Applied Mathematics, 146 (1):127–139, September 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702004235.

Zhu:2002:MDF

[ZZ02b] Xiaolin Zhu and Gongqin Zhu. A method for directly finding the denominator values of rational interpolants. Journal of Computational and Applied Mathematics, 148(2):341–348, November 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704270200554X.

Zhao:2003:WTV

[ZZ03a] Huanxi Zhao and Gongqin Zhu. A worpitzky theorem for vector valued continued fractions. Journal of Computational and Applied Mathematics, 154(1):107–114, May 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042702008154.

Zhou:2003:NDD

[ZZ03b] Shuzi Zhou and Wuping Zhan. A new domain decomposition method for an HJB equation. Journal of Computational and Applied Mathematics, 159(1):195–204, October 1, 2003. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042703005545.